### v1

# Singleton algorithms for the Constraint Satisfaction Problem

Abstract
A natural strengthening of an algorithm for the (promise) constraint satisfaction problem is its singleton version: we first fix a constraint to some tuple from the constraint relation, then run the algorithm, and remove the tuple from the constraint if the answer is negative. We characterize the power of the singleton versions of standard universal algorithms for the (promise) CSP over a fixed template in terms of the existence of a minion homomorphism. Using the Hales-Jewett theorem, we show that for finite relational structures this minion condition is equivalent to the existence of polymorphisms with certain symmetries, called palette block symmetric polymorphisms. By proving the existence of such polymorphisms we establish that the singleton version of the BLP+AIP algorithm solves all tractable CSPs over domains of size at most 7. Finally, by providing concrete CSP templates, we illustrate the limitations of linear programming, the power of the singleton versions, and the elegance of the palette block symmetric polymorphisms.
1 Introduction
1.1 Constraint Satisfaction Problem
The Constraint Satisfaction Problem (CSP) is the problem of deciding whether there is an assignment to a set of variables subject to some specified constraints. Formally, the Constraint Satisfaction Problem is defined as a triple $\langle X,D,\mathbf{C}\rangle$, where
  • $X$ is a finite set of variables,
  • $D$ is a mapping that assigns a set $D_{x}$ to every variable $x\in X$, called the domain of $x$;
  • $\mathbf{C}$ is a set of constraints $C=R(x_{1},\dots,x_{m})$, where $m\in\mathbb{N}$, $x_{1},\dots,x_{m}\in X$, and $R\subseteq D_{x_{1}}\times\dots\times D_{x_{m}}$, called constraint relation.
By $\operatorname{FiniteCSP}$ we denote the decision problem whose input is a triple $\langle X,D,\mathbf{C}\rangle$, where all the domains are finite and constraint relations are given by listing the tuples.
In general the problem $\operatorname{FiniteCSP}$ is NP-complete, and to get tractable cases we usually restrict the set of allowed constraint. For a finite relational structure $\mathbb{A}$, also called the constraint language or template, by $\operatorname{CSP}(\mathbb{A})$ we denote the restricted version of $\operatorname{FiniteCSP}$ where each domain is $A$ and the constraint relations are from $\mathbb{A}$. Alternatively, the input $\langle X,D,\mathbf{C}\rangle$ of $\operatorname{CSP}(\mathbb{A})$ can be viewed as a relational structure $\mathbb{X}$ in the same signature $\sigma$ as $\mathbb{A}$ on the domain $X$ such that $(x_{1},\dots,x_{m})\in R^{\mathbb{X}}$ for $R\in\sigma$ if and only if $R(x_{1},\dots,x_{m})$ is a constraint from $\mathbf{C}$. Then $\operatorname{CSP}(\mathbb{A})$ is equivalent to checking whether there exists a homomorphism $\mathbb{X}\to\mathbb{A}$. Nevertheless, we prefer not to fix the target structure $\mathbb{A}$ and work with a triple $\langle X,D,\mathbf{C}\rangle$ as standard algorithms for the CSP change the constraint relations and even domains of the variables. We have the following characterization of the complexity of $\operatorname{CSP}(\mathbb{A})$ known as the CSP Dichotomy Conjecture.
Theorem 1.1 ([16, 15, 6, 7]).
Suppose $\mathbb{A}$ is a finite relational structure. Then $\operatorname{CSP}(\mathbb{A})$ is solvable in polynomial time if there exists a WNU polymorphism of $\mathbb{A}$; $\operatorname{CSP}(\mathbb{A})$ is NP-complete otherwise.
The NP-hardness in the above theorem follows from [4, 5] and [13]. The essential part of each proof of the CSP Dichotomy Conjecture is a polynomial algorithm that works for all tractable cases. Both polynomial algorithms [16, 15, 6, 7] we know are not universal in the sense that they work only for a fixed structure $\mathbb{A}$ and one of the main open questions in the area is the existence of a universal algorithm.
A natural generalization of the CSP is the Promise Constraint Satisfaction Problem (PCSP). For two relational structures $\mathbb{A}$ and $\mathbb{B}$ in the same signature such that $\mathbb{A}\to\mathbb{B}$ by $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ we denote the following decision problem. Given a structure $\mathbb{X}$ in the same signature. Distinguish between the case when $\mathbb{X}\to\mathbb{A}$ and the case when $\mathbb{X}\not\to\mathbb{B}$. Thus, we have an additional promise about the input that one of these two conditions holds. The complexity of $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ is widely open even for relational structures $\mathbb{A}$ and $\mathbb{B}$ on a 2-element set [10, 2, 1].
1.2 Universal algorithms
We would like to find an algorithm that is universal not only in the sense that it works for any constraint language $\mathbb{A}$, but also in the following sense. An algorithm $\mathfrak{A}$ is called universal if:
  1. (1)
    $\mathfrak{A}$ returns “Yes” or “No” on any instance $\mathcal{I}$ of $\operatorname{FiniteCSP}$;
  2. (2)
    $\mathfrak{A}$ runs in polynomial time;
  3. (3)
    if $\mathfrak{A}$ returns “No” then $\mathcal{I}$ has no solutions.
Thus, the No-answer of a universal algorithm is always correct, while the Yes-answer may be wrong. For example, the trivial algorithm that always returns “Yes” satisfies the above conditions. Since $\operatorname{FiniteCSP}$ is NP-complete, we cannot expect to find a universal algorithm that always returns the correct answer. Our aim, therefore, is to find a universal algorithm that works correctly on the largest possible class of inputs from $\operatorname{FiniteCSP}$. This class can be specified either by restricting the constraint language or by imposing an additional promise on the input. We envision the following major steps toward the most powerful universal algorithm.
  • CSP over a fixed template. The first goal is an algorithm that works correctly on all instances of $\operatorname{CSP}(\mathbb{A})$, whenever $\operatorname{CSP}(\mathbb{A})$ is solvable in polynomial time (by an non-universal algorithm).
  • Perfect Matching Problem. In addition, we expect the algorithm to solve the perfect matching problem correctly.
  • PCSP over a fixed template. The ultimate goal is an algorithm that solves all instances of $\operatorname{PCSP}(\mathbb{A};\mathbb{B})$ that are solvable in polynomial time. That is, the algorithm should return “No” on any instance $\mathbb{X}$ of $\operatorname{CSP}(\mathbb{A})$ whenever $\mathbb{X}\not\to\mathbb{B}$.
In the paper we consider classic universal algorithms such as the arc-consistency algorithm, Basic Linear Programming (BLP), Affine Integer Programming (AIP), and their combinations (see Section 3).
1.3 Main results
As shown in [2], polymorphisms $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ not only characterize the complexity of $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ but also determine which algorithms can solve it. For instance, the arc-consistency algorithm solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains totally symmetric functions of all arities. Similarly, BLP and AIP are characterized by symmetric and alternating polymorphisms, respectively.
Following the ideas from [2], we characterize the power of the singleton versions of such algorithms by the existence of a polymorphism in $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ that is symmetric enough. The first main result of the paper is presented in Table 1. For every algorithm or combination of algorithms from Section 3 we provide a criterion for the algorithm to solve $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$. This criterion is formulated in terms of palette block symmetric polymorphisms, which are formally defined in Section 3.4. To explain the idea of such polymorphisms, let us consider the singleton version of the BLP. As shown in Table 1, this algorithm solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\underbrace{\overline{m},\dots,\overline{m}}_{n})$-palette block symmetric function for all $m$ and $n$. Such a function has arity $m\cdot n$, with its arguments divided into $n$ blocks of $m$ variables each. We require the function to be symmetric within blocks (permutations of variables inside a block do not change the result), but only on tuples that additionally satisfy the palette property. A tuple is a palette tuple if any element that appears in it must completely fill one of the overlined blocks. For example, $(\overline{3},\overline{3},\overline{3})$-palette block symmetric function $f$ must satisfy the identity
$f(\underbrace{x,x,x},\underbrace{x,y,x},\underbrace{y,y,y})=f(\underbrace{x,x,x},\underbrace{y,x,x},\underbrace{y,y,y})$
since the first and third blocks are overlined and completely filled with $x$ and $y$, respectively.
Among the possible combinations of algorithms, the singleton version of the BLP+AIP algorithm turned out to be very powerful. The second main result of the paper is the following theorem.
Theorem 1.2.
Suppose $\mathbb{A}$ is a finite relational structure having a WNU polymorphism such that $|\operatorname{pr}_{i}(R^{\mathbb{A}})|<8$ for every $R$ and $i\in[\operatorname{ar}(R)]$. Then $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ solves $\operatorname{CSP}(\mathbb{A})$.
Table 1: Characterization of the algorithms
Algorithms $\mathfrak{A}$ $\mathfrak{A}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if for all $m,\ell,n\in\mathbb{N}$ the set $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains
ArcCons $m$-ary totally symmetric function
$\mathrm{BLP}$ $m$-ary symmetric function
$\mathrm{AIP}$ $(2\ell+1)$-ary alternating function
$\mbox{{ArcCons}}\wedge\mathrm{AIP}$ $(m,2n+1)$-block function whose first block is totally symmetric and the second block is alternating
$\mathrm{BLP}\wedge\mathrm{AIP}$ $(m,2\ell+1)$-block function whose first block is symmetric and second block is alternating
$\mbox{{ArcCons}}+\mathrm{AIP}$ $(2\ell+1)$-ary weakly alternating function
$\mathrm{BLP}+\mathrm{AIP}$ $(m+1,m)$-block symmetric function
$\mathrm{Singl}\mbox{{ArcCons}}$ $\mathrm{CSingl}\mbox{{ArcCons}}$ $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n})$-palette block totally symmetric function
$\mathrm{Singl}\mathrm{BLP}$ $\mathrm{CSingl}\mathrm{BLP}$ $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n})$-palette block symmetric function
$\mathrm{CSingl}\mathrm{AIP}$ $(\underbrace{\overline{2\ell+1},\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n})$-palette block alternating function
$\mathrm{CSingl}(\mbox{{ArcCons}}\wedge\mathrm{AIP})$ $(\underbrace{\overline{m,2\ell+1},\overline{m,2\ell+1},\dots,\overline{m,2\ell+1}}_{2n})$-palette block function whose odd blocks are totally symmetric and even blocks are alternating
$\mathrm{Singl}(\mbox{{ArcCons}}+\mathrm{AIP})$ $\mathrm{CSingl}(\mbox{{ArcCons}}+\mathrm{AIP})$ $(\underbrace{\overline{2\ell+1},\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n})$-palette block weakly alternating function
$\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ $\mathrm{CSingl}(\mathrm{BLP}+\mathrm{AIP})$ $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette block symmetric function
$\mathrm{CSingl}(\mathrm{BLP}\wedge\mathrm{AIP})$ $(\underbrace{\overline{m,2\ell+1},\overline{m,2\ell+1},\dots,\overline{m,2\ell+1}}_{2n})$-palette block function whose odd blocks are symmetric and even blocks are alternating
$\mathrm{Singl}\mathrm{BLP}+\mathrm{AIP}$ (CLAP) $\mathrm{CSingl}\mathrm{BLP}+\mathrm{AIP}$ $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n},2\ell+1)$-palette block function whose first $n$ blocks are symmetric and last block is alternating
$\mathrm{Singl}\mathrm{AIP}+\mbox{{ArcCons}}$ $\mathrm{CSingl}\mathrm{AIP}+\mbox{{ArcCons}}$ $(\underbrace{\overline{2\ell+1},\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n},m)$-palette block function whose first $n$ blocks are alternating and last block is totally symmetric
$\mathrm{CSingl}\mathrm{AIP}+\mathrm{BLP}$ $(\underbrace{\overline{2\ell+1},\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n},m)$-palette block function whose first $n$ blocks are alternating and last block is symmetric
$\mathrm{Singl}\mbox{{ArcCons}}+\mathrm{BLP}$ $\mathrm{CSingl}\mbox{{ArcCons}}+\mathrm{BLP}$ $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n},\ell)$-palette block function whose first $n$ blocks are totally symmetric and last block is symmetric.
$\mathrm{Singl}\mbox{{ArcCons}}+\mathrm{AIP}$ $\mathrm{CSingl}\mbox{{ArcCons}}+\mathrm{AIP}$ $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n},2\ell+1)$-palette block function whose first $n$ blocks are totally symmetric and last block is alternating.
Thus, up to domain 7 (even in the multi-sorted case), all tractable CSPs can be solved by this very simple algorithm. Unfortunately, as noted by Kompatscher, the structure arising from the dihedral group $\mathbb{D}_{4}$ does not admit palette block symmetric polymorphisms, which implies that $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ does not solve it.
As a byproduct, we also show that the two versions of the singleton algorithms — fixing a tuple of a constraint or a value of a variable — have the same power for most of the algorithms.
In Section 4 we define several CSP templates to illustrate the limitations of linear programming algorithms such as BLP, AIP, and $\mathrm{BLP}+\mathrm{AIP}$. By providing concrete unsatisfiable instances on which $\mathrm{BLP}+\mathrm{AIP}$ incorrectly returns “yes”, we demonstrate that the failure of the algorithm arises because the corresponding template is, in a certain sense, disconnected. We then show that the singleton version overcomes this difficulty, as fixing an element allows it to solve each connected part separately. Finally, we show that the template arising from the dihedral group $\mathbb{D}_{4}$ cannot be solved even by the singleton version of $\mathrm{BLP}+\mathrm{AIP}$. Thus, we identify another fundamental reason why linear programming algorithms fail. Notice that in this section we not only present instances that fool the algorithms but also demonstrate the characterization of the singleton versions of the algorithms by providing concrete palette block symmetric polymorphisms.
1.4 Structure of the paper
The paper is organized as follows. In Section 2 we give necessary definitions. Section 3 introduces classic universal algorithms and their combinations, defines palette block symmetric functions, and formulates criteria for these algorithms to solve $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$. In Section 4 we discuss the limitations and weaknesses of the classic universal algorithms such as BLP, AIP, and BLP+AIP, and demonstrate why the singleton versions are more powerful. Additionally, we show that the template coming from the dihedral group $\mathbb{D}_{4}$ cannot be solved by the singleton version of $\mathrm{BLP}+\mathrm{AIP}$. In Section 5 we prove the relational characterization of the dihedral group $\mathbb{D}_{4}$ and show that it does not admit palette block symmetric polymorphisms. In Section 6 we prove the characterization of the singleton versions of the algorithms in terms of existence of palette block symmetric polymorphisms. Finally, in Section 7 we prove that every finite template on a domain of size at most 7 that admits a WNU polymorphism also admits a palette block symmetric polymorphism, and therefore can be solved by the singleton version of $\mathrm{BLP}+\mathrm{AIP}$.
2 Preliminaries
For $n\in\mathbb{N}$ we denote $n]=\{1,2,\dots,n\}$. We usually denote tuples $\mathbf{a}\in A^{n}$ in bold and refer to their components as $(\mathbf{a}^{1},\dots,\mathbf{a}^{n})$. For a permutation $\sigma\colon[n]\to[n]$, we denote by $\mathbf{a}^{\sigma}$ the tuple $(\mathbf{a}^{\sigma(1)},\dots,\mathbf{a}^{\sigma(n)})$.
A signature is a set of relational symbols $R_{1},\dots,R_{s}$, each with respective arity $\operatorname{ar}(R_{i})\in\mathbb{N}$. A $\sigma$-structure $\mathbb{A}$ consists of a domain $A$ and, for each $R_{i}\in\sigma$, a relation $R_{i}^{\mathbb{A}}\subseteq A^{\operatorname{ar}(R_{i})}$.
For two $\sigma$-structures $\mathbb{A}$ and $\mathbb{B}$, a homomorphism from $\mathbb{A}$ to $\mathbb{B}$ is a mapping $\phi\colon A\to B$ such that for each symbol $R\in\sigma$ and each tuple $\mathbf{a}=(a_{1},\dots,a_{\operatorname{ar}(R)})\in R^{\mathbb{A}}$, we have $\phi(\mathbf{a}):=(\phi(a_{1}),\dots,\phi(a_{\operatorname{ar}(R)}))\in R^{\mathbb{B}}$. If such a homomorphism exists, we write $\mathbb{A}\to\mathbb{B}$; otherwise, we write $\mathbb{A}\not\to\mathbb{B}$.
We say that a pair of finite structures $(\mathbb{A},\mathbb{B})$ is a PCSP template if they have the same signature and there exists a homomorphism $\mathbb{A}\to\mathbb{B}$.
Let $\mathcal{O}_{A,B}^{(n)}$ denote the set of all mappings $A^{n}\to B$, called $n$-ary functions, and let $\mathcal{O}_{A,B}=\bigcup\limits_{n=1}^{\infty}\mathcal{O}_{A,B}^{(n)}$. If $A=B$, we simply write $\mathcal{O}_{A}$ and $\mathcal{O}_{A}^{(n)}$, and we usually use the word operation instead of function.
To distinguish between equations and identities, we use the symbol $\approx$ to emphasize that the equality holds for all evaluations of the variables. For instance, an operation $f\in\mathcal{O}_{A}$ is called idempotent if $f(x,\dots,x)\approx x$, which means that $f(a,\dots,a)=a$ for all $a\in A$. A function $w$ is called a weak near unanimity (WNU) function if it satisfies the identity:
$w(y,x,\dots,x)\approx w(x,y,x,\dots,x)\approx\dots\approx w(x,\dots,x,y).$
A function $f\colon A^{n}\to B$ is called a polymorphism from a $\sigma$-structure $\mathbb{A}$ to $\sigma$-structure $\mathbb{B}$ if, for each symbol $R\in\sigma$ and tuples $\mathbf{a}_{1},\dots,\mathbf{a}_{n}\in R^{\mathbb{A}}$, we have $f(\mathbf{a}_{1},\dots,\mathbf{a}_{n}):=(f(\mathbf{a}_{1}^{1},\dots,\mathbf{a}_{n}^{1}),\dots,f(\mathbf{a}_{1}^{\operatorname{ar}(R)},\dots,\mathbf{a}_{n}^{\operatorname{ar}(R)}))\in R^{\mathbb{B}}$. By $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ we denote the set of all polymorphisms from $\mathbb{A}$ to $\mathbb{B}$. When $\mathbb{A}=\mathbb{B}$, we write simply $\operatorname{Pol}(\mathbb{A})$.
A clone on a set $A$ is a set of operations $\mathcal{C}\subseteq\mathcal{O}_{A}$ such that:
  1. 1.
    $\mathcal{C}$ contains all projection operations $\pi_{i}^{n}(x_{1},\dots,x_{n})=x_{i}$ for all $n\geq 1$ and $i\in[n]$;
  2. 2.
    $\mathcal{C}$ is closed under composition, i.e., if $f\in\mathcal{C}$ is $k$-ary and $g_{1},\dots,g_{k}\in\mathcal{C}$ are $n$-ary, then the $n$-ary operation
    $h(x_{1},\dots,x_{n})=f\bigl(g_{1}(x_{1},\dots,x_{n}),\dots,g_{k}(x_{1},\dots,x_{n})\bigr)$
    also belongs to $\mathcal{C}$.
For a set $N$, we denote by $\mathrm{id}_{N}\colon N\to N$ the identity map. For a relation $R\subseteq A^{n}$ and $S\subseteq[n]$ by $\operatorname{pr}_{S}(R)$ we denote the projection of $R$ onto the coordinates indexed by $S$.
3 Universal algorithms
3.1 Arc-consistency algorithm
The most basic algorithm for the CSP is to enforce some local consistency. Considering the lowest level of consistency we get the arc-consistency algorithm.
Suppose $\mathcal{I}=\langle X,D,\mathbf{C}\rangle$ is an instance of $\operatorname{FiniteCSP}$. The arc-consistency algorithm is implemented by the following pseudocode. By $\mbox{{ReduceDomain}}(\mathcal{I},x,B)$ we denote the function that returns a new instance obtained from $\mathcal{I}$ by reducing the domain of the variable $x$ to $B$ and removing the corresponding tuples from the constraints containing $x$.
1:function ArcCons($\mathcal{I}$)
2:  repeat
3:   for each constraint $C=R(x_{1},\dots,x_{s})\in\mathcal{I}$ do
4:     for $i\in\{1,2,\dots,s\}$ do
5:      if $\operatorname{pr}_{i}(R)\neq D_{x_{i}}$ then
6:        $B:=\operatorname{pr}_{i}(R)$
7:        if $B=\varnothing$ then
8:         return “No”
9:        $\mathcal{I}:=\mbox{{ReduceDomain}}(\mathcal{I},x_{i},B)$
10:  until nothing changed
11:  return “Yes”
Observation 3.1.
The algorithm $\mathrm{ArcCons}$ returns “Yes” on a CSP instance $\mathcal{I}$ if and only if there exists a mapping $D^{\prime}$ that assigns a nonempty set $D_{x}^{\prime}\subseteq D_{x}$ to every variable $x$ such that for any constraint $R(x_{1},\dots,x_{s})$ of $\mathcal{I}$ and any $j\in[s]$ we have $\operatorname{pr}_{j}(R\cap(D_{x_{1}}^{\prime}\times\dots\times D_{x_{s}}^{\prime}))=D_{x_{j}}^{\prime}$.
The power of the arc-consistency algorithm has been completely characterized for both $\operatorname{CSP}(\mathbb{A})$ and $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$. An function $f\colon A^{n}\to B$ is called totally symmetric if $f(a_{1},\dots,a_{n})=f(b_{1},\dots,b_{n})$ whenever $\{a_{1},\dots,a_{n}\}=\{b_{1},\dots,b_{n}\}$.
Theorem 3.1 (Theorem 7.4 in [2]).
Let $(\mathbb{A},\mathbb{B})$ be a PCSP template. Then $\mathrm{ArcCons}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains totally symmetric functions of all arities.
3.2 Linear relaxations
The next algorithms come from the idea of applying the algorithms for solving linear equations (such as Gaussian elimination) to arbitrary CSPs over finite domains. Suppose $\mathcal{I}$ is a CSP instance given by the triple $\langle X,D,\mathbf{C}\rangle$. By $\mathcal{I}^{L}$ we denote the following system of linear equations (the meaning of + is not specified, as it is treated as a formal expression):
Variables: $C^{\mathbf{a}}$ for every constraint $C=R(x_{1},\dots,x_{m})$ in $\mathcal{I}$ and every $\mathbf{a}\in R$,
$x^{b}$ for every variable $x\in X$ and $b\in D_{x}$;
Equations: $\sum\limits_{\mathbf{a}\in R}C^{\mathbf{a}}=1$ for every constraint $C=R(x_{1},\dots,x_{m})$;
$\sum\limits_{\begin{subarray}{c}\mathbf{a}\in R\\ \mathbf{a}(i)=b\end{subarray}}C^{\mathbf{a}}=x_{i}^{b}$ for every constraint $C=R(x_{1},\dots,x_{m})$, every $i\in[m]$ and $b\in D_{x_{i}}$.
If we restrict all the variables to the set $\{0,1\}$ and interpret addition in $\mathbb{Z}$, then we obtain a one-to-one correspondence between solutions of $\mathcal{I}$ and solutions of $\mathcal{I}^{L}$ where $x=b\Leftrightarrow x^{b}=1$.
BLP Algorithm. The algorithm takes a CSP instance $\mathcal{I}$, transforms it into a system of linear equations $\mathcal{I}^{L}$, and solves it over $\mathbb{Q}\cap[0,1]$ using linear programming [11]. It returns “Yes” if a solution exists and “No” otherwise.
The power of the BLP algorithm has been completely characterized for both $\operatorname{CSP}(\mathbb{A})$ and $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$. A function $f\colon A^{n}\to B$ is called symmetric if $f(x_{1},\dots,x_{n})\approx f(x_{\sigma(1)},\dots,x_{\sigma(n)})$ for any permutation $\sigma\colon[n]\to[n]$.
Theorem 3.2 (Theorem 7.9 in [2]).
Let $(\mathbb{A},\mathbb{B})$ be a PCSP template. Then $\mathrm{BLP}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains symmetric functions of all arities.
AIP Algorithm. The algorithm takes a CSP instance $\mathcal{I}$, transforms it into a system of linear equations $\mathcal{I}^{L}$, and solves it over $\mathbb{Z}$ using linear programming [14]. It returns “Yes” if a solution exists and “No” otherwise.
The power of the AIP algorithm has been completely characterized for both $\operatorname{CSP}(\mathbb{A})$ and $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$. A function $f\colon A^{n}\to B$ is called alternating if $f(x_{1},\dots,x_{n})\approx f(x_{\sigma(1)},\dots,x_{\sigma(n)})$ for any permutation $\sigma\colon[n]\to[n]$ preserving parity, and $f(x_{1},\dots,x_{n-2},y,y)\approx f(x_{1},\dots,x_{n-2},z,z)$. Sometimes we will need a weaker version of it where the last identity is required only when the values $y$ and $z$ already appear among other variables. Formally, a function $f$ is called weakly alternating if $f(x_{1},\dots,x_{n})\approx f(x_{\sigma(1)},\dots,x_{\sigma(n)})$ for every parity-preserving permutation $\sigma\colon[n]\to[n]$, and $f(x_{1},\dots,x_{n-2},x_{i},x_{i})\approx f(x_{1},\dots,x_{n-2},x_{j},x_{j})$ for any $i,j\in[n-2]$.
Theorem 3.3 (Theorem 7.19 in [2]).
Let $(\mathbb{A},\mathbb{B})$ be a PCSP template. Then $\mathrm{AIP}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains alternating functions of all odd arities.
A solution of $\mathcal{I}^{L}$ in $\mathbb{Q}\cap[0,1]$ is called a BLP solution, and a solution of $\mathcal{I}^{L}$ in $\mathbb{Z}$ is called an AIP solution.
3.3 Combinations of the algorithms
We will use the following notations for combinations of algorithms. For universal algorithms $\mathfrak{A}$ and $\mathfrak{B}$ by $\mathfrak{A}\wedge\mathfrak{B}$ we denote the algorithm that runs both algorithms independently and return “Yes” if both return “Yes”.
[…]
Similarly, by $\mathrm{ArcCons}$$\mathfrak{A}$ we denote the algorithm that first runs the arc-consistency algorithm, returns “No” if the latter returns “No”, and otherwise applies $\mathfrak{A}$ to the reduced instance and returns its answer.
The strength of the BLP+AIP algorithm has the following nice characterization for both $\operatorname{CSP}(\mathbb{A})$ and $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ in terms of symmetric polymorphisms. A function $f:A^{k_{1}+\dots+k_{n}}\to B$ is called $(k_{1},\dots,k_{n})$-block symmetric if for any permutations $\sigma_{1},\dots,\sigma_{n}$ on $k_{1}],\dots,[k_{n}]$ respectively, we have
$f(\mathbf{x}_{1},\dots,\mathbf{x}_{n})\approx f(\mathbf{x}_{1}^{\sigma_{1}},\dots,\mathbf{x}_{n}^{\sigma_{n}}).$
Theorem 3.4 (Theorem 5.1 in [3]).
Let $(\mathbb{A},\mathbb{B})$ be a PCSP template. Then the following conditions are equivalent:
  1. (1)
    $\mathrm{BLP}+\mathrm{AIP}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$;
  2. (2)
    for any $L\in\mathbb{N}$ there exist $s\in\mathbb{N}$ and numbers $\ell_{1},\dots,\ell_{s}$ greater than $L$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\ell_{1},\dots,\ell_{s})$-block symmetric function.
  3. (3)
    for any $L\in\mathbb{N}$ the set $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(L+1,L)$-block symmetric function.
[…]
  • (f1)
    $\mathcal{I}$ has a BLP solution $\Leftrightarrow$ there exists $Q\in\mathbb{N}$ such that $\mathcal{I}$ has a $Q$-solution.
  • (f2)
    The sum of a $d_{1}$-solution and a $d_{2}$-solution is a $(d_{1}+d_{2})$-solution.
  • (f3)
    $\mathrm{BLP}+\mathrm{AIP}$ returns “Yes” on an instance $\mathcal{I}$ $\Leftrightarrow$ there exists $L\in\mathbb{N}$ such that $\mathcal{I}$ has a $d$-solution for any $d\geq L$.
(f1) follows from the fact that multiplying a BLP solution by the least common multiple $Q$ of the denominators gives a $Q$-solution, and dividing a $d$-solution by $d$ gives a BLP solution. (f2) is immediate. Let us show (f3). If $\mathrm{BLP}+\mathrm{AIP}$ returns “Yes” on $\mathcal{I}$ then there exists a BLP solution $s_{BLP}$ of $\mathcal{I}^{L}$, an AIP solution $s_{AIP}$ of $\mathcal{I}^{L}$ such that $s_{BLP}(v)>0$ whenever $s_{AIP}(v)\neq 0$. Choose a sufficiently large integer $Q$ such that $Q\cdot s_{BLP}+s_{AIP}$ is a $(Q+1)$-solution (all values are nonnegative integers). Then $Q\cdot s_{BLP}$ is a $Q$-solution. By taking linear combinations $m\cdot(Q\cdot s_{BLP}+s_{AIP})+\ell(Q\cdot s_{BLP})$, we obtain a $d$-solution for any $d\geq Q\cdot(Q+1)$. To see the opposite direction of (f3) choose a $d_{0}$-solution, with $d_{0}\geq L$, having the maximal number of nonzero values among all the $d$-solutions (for all $d\in\mathbb{N}$). Such a solution exists by (f2). Considering the difference between a $(d_{0}+1)$-solution and the $d_{0}$-solution we obtain an AIP solution, while dividing the $d_{0}$-solution by $d_{0}$ gives a BLP solution. Together, these confirm that $\mathrm{BLP}+\mathrm{AIP}$ returns “Yes” on $\mathcal{I}$.
[…]
$(x_{1},x_{2}):=(f(\mathbf{a}_{x_{1},1},\dots,\mathbf{a}_{x_{1},s}),f(\mathbf{a}_{x_{2},1},\dots,\mathbf{a}_{x_{2},s}))=f(M_{1},\dots,M_{s})\in R^{\mathbb{A}},$
where the equality follows from the $(\ell_{1},\dots,\ell_{s})$-block symmetry of $f$, and membership follows from $f\in\operatorname{Pol}(\mathbb{A},\mathbb{B})$.
Similarly, other combinations of the above algorithms (see the top part of Table 1) can be characterized, but since the proofs do not differ substantially from those in [2, 3], we omit them here.
3.4 Singleton algorithms and palette block functions
For a universal algorithm $\mathrm{Alg}$ by $\mathrm{SinglAlg}$ and $\mathrm{CSingl}\mbox{{Alg}}$ we denote the algorithms defined by the pseudocode below. Here, $\mbox{{ChangeConstraint}}(\mathcal{I},C,R)$ denotes the function that returns a new instance obtained from $\mathcal{I}$ by changing the constraint relation of the constraint $C$ to $R$. The difference between SinglAlg and CSinglAlg is that in $\mathrm{SinglAlg}$ we fix a variable to an element, whereas in $\mathrm{CSingl}\mbox{{Alg}}$ we fix a constraint to a tuple.
Remark 3.1.
It is not generally true that $\mathrm{SinglAlg}$ is weaker than $\mathrm{CSingl}\mbox{{Alg}}$, since the algorithm $\mathrm{SinglAlg}$, when fixing a variable $x$ to a value, may reduce all constraints containing $x$, while $\mathrm{CSingl}\mbox{{Alg}}$ always reduces only one constraint. Nevertheless, if $\mbox{{Alg}}\in\{\mathrm{ArcCons},\mathrm{BLP},\mathrm{ArcCons}+\mathfrak{A},\mathrm{BLP}+\mathfrak{A}\}$, then $\mathrm{SinglAlg}$ is clearly weaker than $\mathrm{CSingl}\mbox{{Alg}}$.
Open question 1.
Does there exist a PCSP template $(\mathbb{A},\mathbb{B})$ such that $\mathrm{Singl}\mathrm{AIP}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ but $\mathrm{CSingl}\mathrm{AIP}$ does not?
1:function SinglAlg($\mathcal{I}$)
2:  repeat
3:   for each variable $x$ of $\mathcal{I}$ do
4:     for $a\in D_{x}$ do
5:      if $\neg\mbox{{Alg}}(\mbox{{ReduceDomain}}(\mathcal{I},x,\{a\}$) then
6:        $B:=D_{x}\setminus\{a\}$
7:        if $B=\varnothing$ then
8:         return “No”
9:        $\mathcal{I}:=\mbox{{ReduceDomain}}(\mathcal{I},x,B)$ $\triangleright$ remove all the tuples $\triangleright$ such that $x\notin B$
10:  until nothing changed
11:  return “Yes”
1:function CSinglAlg($\mathcal{I}$)
2:  repeat
3:   for each constraint $C=R(x_{1},\dots,x_{s})\in\mathcal{I}$ do
4:     for $(a_{1},\dots,a_{s})\in R$ do
5:      if $\neg\mbox{{Alg}}(\mbox{{ChangeConstraint}}(\mathcal{I},C,\{(a_{1},\dots,a_{s})\})$ then $\triangleright$ reduce $R$ to the one-tuple relation
6:        $R^{\prime}:=R\setminus\{(a_{1},\dots,a_{s})\}$
7:        if $R^{\prime}=\varnothing$ then
8:         return “No”
9:        $\mathcal{I}:=\mbox{{ChangeConstraint}}(\mathcal{I},C,R^{\prime})$ $\triangleright$ remove the tuple $\triangleright$ from the constraint relation of $C$
10:  until nothing changed
11:  return “Yes”
For universal algorithms $\mathfrak{A}$ and $\mathfrak{B}$ we denote by $\mathrm{Singl}\mathfrak{A}+\mathfrak{B}$ the algorithm that first runs $\mathrm{Singl}\mathfrak{A}$ and returns “No” if it returns “No”; otherwise, it returns the answer of $\mathfrak{B}$ on the reduced instance $\mathcal{I}$ (see the pseudocode of SinglAlg). Similarly, by $\mathrm{CSingl}\mathfrak{A}+\mathfrak{B}$ we denote the algorithm that first runs $\mathrm{CSingl}\mathfrak{A}$ and returns “No” if it returns “No”; otherwise, it returns the answer of $\mathfrak{B}$ on the reduced instance $\mathcal{I}$ (see the pseudocode of CSinglAlg).
To characterize the power of singleton algorithms we need a generalization of symmetric functions. We call to a function $f:A^{k_{1}+\dots+k_{n}}\to B$ a $(k_{1},\dots,k_{n})$-block function to emphasize that its arguments are divided into blocks: the first $k_{1}$ arguments form the first block, the next $k_{2}$ arguments form the second block, and so on. We say that the $i$-th block of a $(k_{1},\dots,k_{n})$-block function is (totally) symmetric/(weakly) alternating if for any $\mathbf{a}_{1}\in A^{k_{1}},\dots,\mathbf{a}_{i-1}\in A^{k_{i-1}},\mathbf{a}_{i+1}\in A^{k_{i+1}},\dots,\mathbf{a}_{n}\in A^{k_{n}}$ the function $h(\mathbf{x})=f(\mathbf{a}_{1},\dots,\mathbf{a}_{i-1},\mathbf{x},\mathbf{a}_{i+1},\dots,\mathbf{a}_{n})$ is (totally) symmetric/(weakly) alternating, respectively. A function is $(k_{1},\dots,k_{n})$-block (totally) symmetric if all of its blocks are (totally) symmetric, and $(k_{1},\dots,k_{n})$-block (weakly) alternating if all of its blocks are (weakly) alternating.
We can divide a tuple $\mathbf{a}\in A^{k_{1}+\dots+k_{n}}$ into blocks and refer to each block by its position. Let $\mathcal{P}$ be a set of disjoint subsets of $[n]$. A tuple $\mathbf{a}=(\mathbf{a}_{1},\dots,\mathbf{a}_{n})\in A^{k_{1}+\dots+k_{n}}$ is called $\mathcal{P}$-$(k_{1},\dots,k_{n})$-palette if for every element $c\in A$ appearing in $\mathbf{a}$ there exists $L\in\mathcal{P}$ such that $\mathbf{a}_{\ell}=(c,c,\dots,c)$ for every $\ell\in L$. We usually use an overline to specify the set $\mathcal{P}$. For example, $(\ell,\overline{k_{1}},\dots,\overline{k_{n}})$ corresponds to $\mathcal{P}=\{\{2\},\{3\},\dots,\{n+1\}\}$ and $(\overline{k_{1},m_{1}},\dots,\overline{k_{n},m_{n}})$ corresponds to $\mathcal{P}=\{\{1,2\},\{3,4\},\dots,\{2n-1,2n\}\}$.
Similarly to block functions, we define palette block functions, where we require the block properties only on palette tuples. Formally, a function $f\colon A^{k_{1}+\dots+k_{n}}\to A$ is called $\mathcal{P}$-$(k_{1},\dots,k_{n})$-palette block function, and the $i$-th block of it is (totally) symmetric/(weakly) alternating if (1) and (2) imply (3):
  • (1)
    $\mathbf{a}$ and $\mathbf{b}$ are $\mathcal{P}$-$(k_{1},\dots,k_{n})$-palette tuples;
  • (2)
    $g(\mathbf{a})=g(\mathbf{b})$ for any $(k_{1},\dots,k_{n})$-block function $g\colon A^{k_{1}+\dots+k_{n}}\to A$ whose $i$-th block is (totally) symmetric/(weakly) alternating;
  • (3)
    $f(\mathbf{a})=f(\mathbf{b})$.
A function is $\mathcal{P}$-$(k_{1},\dots,k_{n})$-palette block (totally) symmetric/(weakly) alternating if it is a $\mathcal{P}$-$(k_{1},\dots,k_{n})$-palette block function whose blocks are (totally) symmetric/(weakly) alternating.
The power of singleton algorithms and their combinations can be characterized using palette block functions. Below we present a characterization for the algorithms $\mathrm{Singl}\mathrm{ArcCons}$ and $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$, which are proved in Section 6.3 and 6.4, respectively.
Theorem 3.5.
Let $(\mathbb{A},\mathbb{B})$ be a PCSP template. The following conditions are equivalent:
  1. (1)
    $\mathrm{SinglArcCons}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$;
  2. (2)
    $\mathrm{CSingl}\mathrm{ArcCons}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$;
  3. (3)
    $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ has an $(\overline{m_{1}},\dots,\overline{m_{n}})$-palette block totally symmetric function for every $n,m_{1},\dots,m_{n}\in\mathbb{N}$;
  4. (4)
    for every $N\in\mathbb{N}$ there exist $n,m_{1},\dots,m_{n}\geq N$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ has an $(\overline{m_{1}},\dots,\overline{m_{n}})$-palette block totally symmetric function.
Theorem 3.6.
Let $(\mathbb{A},\mathbb{B})$ be a PCSP template. The following conditions are equivalent:
  1. (1)
    $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$;
  2. (2)
    $\mathrm{CSingl}(\mathrm{BLP}+\mathrm{AIP})$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$;
  3. (3)
    $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette block symmetric function for every $\ell,n\in\mathbb{N}$;
  4. (4)
    for every $N\in\mathbb{N}$ there exist $n\geq N$, $k_{1},\dots,k_{n}\in\mathbb{N}$, and $m_{1,1},\dots,m_{1,k_{1}},\dots,m_{n,1},\dots,m_{n,k_{n}}\geq N$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\overline{m_{1,1},\dots,m_{1,k_{1}}},\overline{m_{2,1},\dots,m_{2,k_{2}}},\dots,\overline{m_{n,1},\dots,m_{n,k_{n}}})$-palette block symmetric function.
To explain the idea behind palette block symmetric functions and Theorem 3.6, let us prove the following claim.
Claim 3.1.
Suppose $(\mathbb{A},\mathbb{B})$ is a PCSP template, and for every $L\in\mathbb{N}$ there exist $n,\ell_{1},\dots,\ell_{n}\geq L$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette block symmetric function. Then $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$.
Proof.
It is sufficient to prove that if $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ returns “Yes” on an instance $\mathcal{I}$, then $\mathcal{I}$ has a solution. For every variable $x$ and every element $a$ from the obtained domain of $x$ let $\mathcal{I}_{x}^{a}$ be the instance obtained in the algorithm SinglAlg by reducing the domain of $x$ to $\{a\}$. Let $\Omega$ be the set of all such instances. Since $\mathrm{BLP}+\mathrm{AIP}$ returns “Yes” on each of these instances, the fact (f3) from Section 3.3 implies that there exists a large number $L$ such that $L>|\Omega|$ and each instance $\mathcal{I}_{i}^{a}$ has a $d$-solution for any $d\geq L$. Choose an $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette block symmetric function $f\in\operatorname{Pol}(\mathbb{A},\mathbb{B})$, where $n,\ell_{1},\dots,\ell_{n}\geq L$.
Similar to the construction in Section 3.3, for each block (say the $j$-th block) we choose an instance $\mathcal{I}_{x}^{a}$ and an $\ell_{j}$-solution $s_{j}$ of this instance. We only require every instance from $\Omega$ to be chosen for at least one block, which is possible since $n\geq L>|\Omega|$. For every variable $x$ of $\mathcal{I}$, let $\mathbf{a}_{x,j}$ be a tuple of length $\ell_{j}$ such that each element $c$ appears in this tuple exactly $s_{j}(x^{c})$ times. We claim that a (true) solution of $\mathcal{I}$ can be defined by $x:=f(\mathbf{a}_{x,1},\dots,\mathbf{a}_{x,s})$ for every variable $x$. To follow the argument from Section 3.3, it is sufficient to check that the tuple $(\mathbf{a}_{x,1},\dots,\mathbf{a}_{x,s})$ is $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette, which follows immediately from the definition of the instances in $\Omega$. ∎
As mentioned in the introduction, the algorithm $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ is very powerful and solves all tractable $\operatorname{CSP}(\mathbb{A})$ on small domains, as follows from the following algebraic result proved in Section 7.
Theorem 3.7.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$ are similar Taylor algebras, $|A_{i}|<8$ for every $i\in[s]$. Then $\mathbf{A}_{1}\times\dots\times\mathbf{A}_{s}$ has a $(\underbrace{p^{m},\dots,p^{m}}_{n})$-palette block symmetric term operation for every prime $p>8$ and every $m,n\in\mathbb{N}$.
All necessary definitions are given in Section 7. An important corollary for our purposes is the following theorem.
Theorem 1.2.
Suppose $\mathbb{A}$ is a finite relational structure having a WNU polymorphism such that $|\operatorname{pr}_{i}(R^{\mathbb{A}})|<8$ for every $R$ and $i\in[\operatorname{ar}(R)]$. Then $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ solves $\operatorname{CSP}(\mathbb{A})$.
An alternative formulation is to introduce the notion of a multi-sorted constraint language and restrict the domains by 7 instead of restricting the projections, but we chose the current formulation to avoid additional notions that are not essential in this paper.
Most other combinations of the algorithms can be characterized similarly (see Table 1). Since their proofs are almost identical to those for $\mathrm{Singl}\mathrm{ArcCons}$ and $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$, we omit them. We can also characterize the algorithm $\mathrm{Singl}\mathfrak{A}\wedge\mathfrak{B}$ for $\mathfrak{A},\mathfrak{B}\in\{\mathrm{ArcCons},\mathrm{BLP},\mathrm{AIP}\}$, but this requires a different notion and is less interesting, since the algorithms are simply run independently. For example, $\mathrm{Singl}\mathrm{BLP}\wedge\mathrm{AIP}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if for all $m,\ell,n\in\mathbb{N}$, $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains a $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n},\ell)$-block function such that the first $n$ blocks are symmetric on $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n},\ell)$-palette tuples and the last block is always alternating.
4 Limitations of linear programming
In this section, we discuss the limitations of linear programming and provide concrete templates that illustrate its weaknesses. The results are collected in Table 2, while templates are defined below.
Table 2: Examples of constraint languages
Constraint language Strongest not working Weakest working
Horn-SAT $\mathrm{AIP}$ $\mathrm{ArcCons}$
2-SAT $\mathrm{BLP}$, $\mathrm{Singl}\mathrm{AIP}$ $\mathrm{Singl}\mathrm{ArcCons}$, $\mathrm{ArcCons}+\mathrm{AIP}$
${3}\text{-}\mathrm{LIN}(p)$ $\mathrm{Singl}\mathrm{BLP}$ $\mathrm{AIP}$
$\left(\begin{smallmatrix}0&1&2&3&4\\ 1&0&3&4&2\end{smallmatrix}\right)$ $\mathrm{BLP}+\mathrm{AIP}$ $\mathrm{Singl}\mathrm{ArcCons}$, $\mathrm{Singl}\mathrm{AIP}$
$\mathbb{Z}_{2}\cup\mathbb{Z}_{3}$ $\mathrm{Singl}\mathrm{BLP}+\mathrm{AIP}$ $\mathrm{Singl}\mathrm{AIP}$
$\mathbb{Z}_{2}\leftarrow 2$ $\mathrm{Singl}\mathrm{BLP}+\mathrm{AIP}$ $\mathrm{Singl}\mathrm{AIP}$
$\mathbb{D}_{4}$ $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ ???
The first simple idea demonstrating the weakness of the $\mathrm{BLP}$ algorithm can be formulated as follows.
Claim 4.1.
Suppose $\mathbb{A}$ is a finite relational structure, for every relation $R^{\mathbb{A}}$ of an arity $m$, every $b,c\in A$, and every $i\in[m]$, we have $|\{\mathbf{a}\in R^{\mathbb{A}}\mid\mathbf{a}^{i}=b\}|=|\{\mathbf{a}\in R^{\mathbb{A}}\mid\mathbf{a}^{i}=c\}|$. Then $\mathrm{BLP}$ answers “Yes” on any instance of $\operatorname{CSP}(\mathbb{A})$.
Proof.
The uniform distribution is always a solution to $\mathcal{I}^{L}$. Specifically, it is sufficient to set $C^{\mathbf{a}}=1/|R^{\mathbb{A}}|$ for every constraint $C=R(x_{1},\dots,x_{m})$ and $\mathbf{a}\in R$, and $x^{b}=1/|A|$ for every variable $x$ and $b\in A$. ∎
Thus, $\mathrm{BLP}$ returns “Yes” whenever the uniform distribution is feasible, making it useless for solving linear equations over finite fields.
The next claim shows the main weakness of $\mathrm{AIP}$: it produces the same result even if we replace each constraint relation by its parallelogram closure. We say that a relation $R\subseteq A^{m}$ satisfies the parallelogram property if for any $I\subseteq[m]$, $\mathbf{a},\mathbf{b},\mathbf{c}\in R$, and $\mathbf{d}\in A^{m}$, we have
$\left(\operatorname{pr}_{I}(\mathbf{a})=\operatorname{pr}_{I}(\mathbf{b})\wedge\operatorname{pr}_{I}(\mathbf{c})=\operatorname{pr}_{I}(\mathbf{d})\wedge\operatorname{pr}_{[m]\setminus I}(\mathbf{a})=\operatorname{pr}_{[m]\setminus I}(\mathbf{c})\wedge\operatorname{pr}_{[m]\setminus I}(\mathbf{b})=\operatorname{pr}_{[m]\setminus I}(\mathbf{d})\right)\Longrightarrow\mathbf{d}\in R.$
For binary relations $R$ this simply means that $(b_{1},b_{2}),(b_{1},a_{2}),(a_{1},b_{2})\in R$ implies $(a_{1},a_{2})\in R$. The parallelogram closure of $R$ is the minimal relation $R^{\prime}\supseteq R$ having the parallelogram property.
Claim 4.2.
Suppose $\mathcal{I}$ is an instance of $\operatorname{FiniteCSP}$ containing a constraint $C=R(x_{1},\dots,x_{m})$, and
  • $(a_{1},\dots,a_{k},b_{k+1},\dots,b_{m}),(b_{1},\dots,b_{k},a_{k+1},\dots,a_{m}),(b_{1},\dots,b_{k},b_{k+1},\dots,b_{m})\in R$,
  • $R^{\prime}=R\cup\{(a_{1},\dots,a_{k},a_{k+1},\dots,a_{m})\}$,
  • $\mathcal{I}^{\prime}$ is obtained from $\mathcal{I}$ by replacing $R$ in $C$ by $R^{\prime}$.
Then $\mathrm{AIP}$ returns “Yes” on $\mathcal{I}$ if and only if $\mathrm{AIP}$ returns “Yes” on $\mathcal{I}^{\prime}$.
Proof.
It is sufficient to construct an AIP solution for $\mathcal{I}$ from an AIP solution for $\mathcal{I}^{\prime}$. To do so, remove the variable $C^{(a_{1},\dots,a_{k},a_{k+1},\dots,a_{m})}$ and adjust the values of only 3 variables as follows:
$C^{(a_{1},\dots,a_{k},b_{k+1},\dots,b_{m})}:=C^{(a_{1},\dots,a_{k},b_{k+1},\dots,b_{m})}+C^{(a_{1},\dots,a_{k},a_{k+1},\dots,a_{m})}$
$C^{(b_{1},\dots,b_{k},a_{k+1},\dots,a_{m})}:=C^{(b_{1},\dots,b_{k},a_{k+1},\dots,a_{m})}+C^{(a_{1},\dots,a_{k},a_{k+1},\dots,a_{m})}$
$C^{(b_{1},\dots,b_{k},b_{k+1},\dots,b_{m})}:=C^{(b_{1},\dots,b_{k},b_{k+1},\dots,b_{m})}-C^{(a_{1},\dots,a_{k},a_{k+1},\dots,a_{m})}$
Therefore, $\mathrm{AIP}$ generally cannot solve CSPs involving relations without the parallelogram property (see Sections 4.1 and 4.2). Another weakness of $\mathrm{AIP}$ is its inability to solve linear equations modulo 2 and modulo 3 simultaneously, as shown in Section 4.3.
4.1 Classic templates
Horn-3-SAT $(\{0,1\};\{0,1\}^{3}\setminus\{(1,1,0)\},\{0\},\{1\})$. Horn-3-SAT can be solved by the arc-consistency algorithm, as witnessed by the fact that the $n$-ary conjunction, which is a totally symmetric function, belongs to $\operatorname{Pol}(\text{Horn-3-SAT})$ for every $n\geq 2$. On the other hand, $\mathrm{AIP}$ cannot solve Horn-3-SAT. Indeed, the minimal relation with the parallelogram property containing $\{0,1\}^{3}\setminus\{(1,1,0)\}$ is the full relation $\{0,1\}^{3}$. By Claim 4.2, running $\mathrm{AIP}$ on Horn-3-SAT is equivalent to running $\mathrm{AIP}$ on $(\{0,1\};\{0,1\}^{3},\{0\},\{1\})$, which can only return “No” on trivial instances.
[…]
$\mathbf{n}\text{{-LIN}}\mathbf{(p)}=\{\mathbb{Z}_{p};\{a_{1}x_{1}+\dots+a_{n}x_{n}=a_{0}\mid a_{0},a_{1},\dots,a_{n}\in\mathbb{Z}_{p}\}$ consists of linear equations with $n$ variables modulo $p$. $\mathrm{AIP}$ solves it as $\operatorname{Pol}({n}\text{-}\mathrm{LIN}(p))$ includes an alternating function $x_{1}-x_{2}+x_{3}-\dots-x_{2n}+x_{2n+1}$. To see that $\mathrm{Singl}\mathrm{BLP}$ does not solve ${3}\text{-}\mathrm{LIN}(p)$, consider an instance consisting of two contradictory equations: $x_{1}+x_{2}+x_{3}=0\wedge x_{1}+x_{2}+x_{3}=1$. Fixing a variable to a concrete value and assigning a uniform distribution to the rest (as in Claim 4.1) yields a BLP solution, so $\mathrm{Singl}\mathrm{BLP}$ incorrectly returns “Yes”.
4.2 Disconnected templates
$\left(\begin{smallmatrix}\mathbf{0}&\mathbf{1}&\mathbf{2}&\mathbf{3}&\mathbf{4}\\ \mathbf{1}&\mathbf{0}&\mathbf{3}&\mathbf{4}&\mathbf{2}\end{smallmatrix}\right)=(\{0,1,2,3,4\};\{(0,1),(1,0),(2,3),(3,4),(4,2)\})$, which is a structure containing a single binary relation given by the graph of a bijection. To show that both $\mathrm{Singl}\mathrm{ArcCons}$ and $\mathrm{Singl}\mathrm{AIP}$ solve it, let us define $f\in\operatorname{Pol}\left(\begin{smallmatrix}0&1&2&3&4\\ 1&0&3&4&2\end{smallmatrix}\right)$ that is simultaneously $(\underbrace{\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n})$-palette block totally symmetric and palette block alternating. Specifically, let $f(\mathbf{x}_{1},\dots,\mathbf{x}_{n})$ return the value from the first constant block if it exists, and $\mathbf{x}_{1}^{1}$ otherwise.
Denote the only relation in this structure by $\rho$, and consider the instance $\mathcal{I}=\rho(x_{1},x_{2})\wedge\rho(x_{2},x_{3})\wedge\rho(x_{3},x_{1})\wedge\rho(x_{1},x_{4})\wedge\rho(x_{4},x_{1})$, which has no solution. By Claim 4.1, $\mathcal{I}$ nevertheless has a BLP solution in which all the tuples receive positive assignments. To show that $\mathrm{BLP}+\mathrm{AIP}$ also returns “Yes” on $\mathcal{I}$, it is sufficient to construct an AIP solution. One such solution assigns value $1$ to the tuples $(2,3),(3,4),(4,2)$ in each constraint, and value 1-1 to the tuples $(0,1),(1,0)$ in each constraint. This example illustrates that $\mathrm{AIP}$ is powerless when the instance is disconnected (here the components $\{0,1\}$ and $\{2,3,4\}$ are disconnected). This weakness becomes even more evident in the next template.
$\bm{\mathbf{Z}_{2}\cup\mathbf{Z}_{3}}=(\{0,1,0^{\prime},1^{\prime},2^{\prime}\};l_{0}^{\mathbb{Z}_{2}},l_{1}^{\mathbb{Z}_{2}},l_{0}^{\mathbb{Z}_{3}},l_{1}^{\mathbb{Z}_{3}},l_{2}^{\mathbb{Z}_{3}})$, where the relations are defined by:
$l_{a}^{\mathbb{Z}_{2}}(x_{1},x_{2},x_{3})=$ $(x_{1},x_{2},x_{3}\in\{0,1\}\wedge x_{1}+x_{2}+x_{3}=a\pmod{2})\vee x_{1},x_{2},x_{3}\in\{0^{\prime},1^{\prime},2^{\prime}\},$
$l_{b}^{\mathbb{Z}_{3}}(x_{1},x_{2},x_{3})=$ $(x_{1},x_{2},x_{3}\in\{0^{\prime},1^{\prime},2^{\prime}\}\wedge x_{1}+x_{2}+x_{3}=b\pmod{3})\vee x_{1},x_{2},x_{3}\in\mathbb{\{}0,1\}.$
Thus, we have linear equations modulo 2 on $\{0,1\}$ and independent linear equations modulo 3 on $\{0^{\prime},1^{\prime},2^{\prime}\}$. Any instance of $\operatorname{CSP}(\mathbb{Z}_{2}\cup\mathbb{Z}_{3})$ decomposes into two independent systems, one over $\mathbb{Z}_{2}$ and one over $\mathbb{Z}_{3}$. Solving the instance amounts to solving each system separately and returning “Yes” if both have solutions. Nevertheless, $\mathrm{AIP}$ fails here: it returns “Yes” on any instance of $\operatorname{CSP}(\mathbb{Z}_{2}\cup\mathbb{Z}_{3})$. The key idea comes from the following claim (the definition of a $d$-solution was given in Section 3.3).
Claim 4.3.
Suppose $\mathcal{I}$ is an instance of ${n}\text{-}\mathrm{LIN}(p)$ for a prime $p$ and $n\in\mathbb{N}$. Then $\mathcal{I}$ has a $p^{N}$-solution for every $N\geq n-1$.
Proof.
It is sufficient to set $C^{\mathbf{a}}=p^{n}/|R^{\mathbb{A}}|$ for every constraint $C=R(x_{1},\dots,x_{m})$ and every tuple $\mathbf{a}\in R$, and $x^{b}=p^{n-1}$ for every variable $x$ and every $b\in\mathbb{Z}_{p}$. ∎
Applying Claim 4.3 to the restriction of an instance of $\operatorname{CSP}(\mathbb{Z}_{2}\cup\mathbb{Z}_{3})$ to $\{0,1\}$ and to $\{0^{\prime},1^{\prime},2^{\prime}\}$, we obtain:
  • every instance has a $2^{n}$-solution for any $n\geq 2$;
  • every instance has a $3^{n}$-solution for any $n\geq 2$;
  • a linear combination of a $2^{n}$-solution and a $3^{n}$-solution yields an AIP solution.
Thus, $\mathrm{AIP}$ returns “Yes” on all instances of $\operatorname{CSP}(\mathbb{Z}_{2}\cup\mathbb{Z}_{3})$. Even $\mathrm{Singl}\mathrm{BLP}+\mathrm{AIP}$ fails, since $\mathrm{Singl}\mathrm{BLP}$ cannot reduce a system of linear equations (with more than 2 variables); therefore, running $\mathrm{Singl}\mathrm{BLP}+\mathrm{AIP}$ is not substantially different from running $\mathrm{AIP}$ alone.
In contrast, $\mathrm{Singl}\mathrm{AIP}$ solves the problem. Fixing a concrete element in the $\mathrm{Singl}$-algorithm restricts the instance to either the $\{0,1\}$-part or the $\{0^{\prime},1^{\prime},2^{\prime}\}$-part, and each part can be solved by $\mathrm{AIP}$. This is also witnessed by a $(\underbrace{\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n})$-palette block alternating function $f(\mathbf{x}_{1},\dots,\mathbf{x}_{n})$ defined as follows:
  • choose the first block $\mathbf{x}_{i}$ consisting entirely of elements from either $\{0,1\}$ or $\{0^{\prime},1^{\prime},2^{\prime}\}$;
  • if no such block exists, return $\mathbf{x}_{1}^{1}$;
  • otherwise, apply the alternating function to $\mathbf{x}_{i}$, that is, $f(\mathbf{x}_{1},\dots,\mathbf{x}_{n})=\mathbf{x}_{i}^{1}-\mathbf{x}_{i}^{2}+\mathbf{x}_{i}^{3}-\dots-\mathbf{x}_{i}^{2\ell}+\mathbf{x}_{i}^{2\ell+1}$, where ++ and - are interpreted modulo 2 if $\mathbf{x}_{i}\in\{0,1\}^{2\ell+1}$, and modulo 3 if $\mathbf{x}_{i}\in\{0^{\prime},1^{\prime},2^{\prime}\}^{2\ell+1}$.
Although this template is extremely simple, it serves as a counterexample to several algorithms that were conjectured to solve all tractable CSPs, including $\mathbb{Z}$-affine $k$-consistency, BLP+AIP, $\mathrm{BA}^{k}$, and CLAP (see [12] for details). Since hierarchies and higher levels of consistency are beyond the scope of this paper, we omit the proofs.
4.3 Implications of linear equations
The next template has a similar flavor to the previous ones, but its instances are more interesting because they combine implications with linear equations.
$\bm{\mathbf{Z}_{2}\leftarrow 2}=(\{0,1,2\};\mathcal{R}_{L}\cup\{\{0,1\},S_{0,1}^{=}\})$, where $S_{0,1}^{=}(x,y,z)=(x=y\vee z\in\{0,1\})$, and
$\mathcal{R}_{L}=\left\{(x_{1}+\dots+x_{n}=c\wedge x_{1},\dots,x_{n}\in\{0,1\})\vee(x_{1}=\dots=x_{n}=2)\mid c\in\{0,1\},n\in\{3,4\}\right\}.$
Thus, $S_{0,1}^{=}(x,y,z)$ can be read as an implication $z=2\rightarrow x=y$, and $\mathcal{R}_{L}$ consists of linear equations over $\{0,1\}$, augmented by the additional all-2 tuple $(2,2,\dots,2)$. The polymorphisms of this structure are well-understood, as it corresponds to one of the minimal Taylor clones [9]. Define the ternary operation:
[…]
Claim 4.4 ([9]).
$\operatorname{Pol}(\mathbb{Z}_{2}\leftarrow 2)=\operatorname{Clo}(f)$.
Below we define an instance $\mathcal{I}$ of $\operatorname{CSP}(\mathbb{Z}_{2}\leftarrow 2)$ that fools $\mathrm{Singl}\mathrm{BLP}+\mathrm{AIP}$. Here, each linear equation should be understood as the corresponding relation from $\mathcal{R}_{L}$, that is, a linear equation modulo 2 augmented by the tuple $(2,2,\dots,2)$.
$\begin{cases}x_{1}\in\{0,1\}\\ \hline\cr x_{1}+x_{2}+x_{3}+x_{4}=0\\ x_{3}+x_{5}+x_{6}=0\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;(L_{1})\\ x_{4}+x_{5}+x_{6}=1\\ \hline\cr x_{1}=x_{2}\vee x_{7}\in\{0,1\}\\ \hline\cr x_{7}+x_{8}+x_{9}+x_{10}=0\\ x_{9}+x_{11}+x_{12}=0\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;(L_{2})\\ x_{10}+x_{11}+x_{12}=1\\ \hline\cr x_{7}=x_{8}\vee x_{13}\in\{0,1\}\\ \hline\cr x_{13}+x_{14}+x_{15}+x_{16}=0\\ x_{13}+x_{14}+x_{17}+x_{18}=0\;\;\;\;\;\;\;\;\;\;\;\;(L_{3})\\ x_{15}+x_{16}+x_{17}+x_{18}=1\end{cases}$
We have three systems of linear equations $L_{1}$, $L_{2}$, and $L_{3}$, with implications connecting them. Each system admits the trivial fake solution where all variables are set to 2. In addition, $L_{1}$ and $L_{2}$ have nontrivial solutions on $\{0,1\}$ but $L_{3}$ does not. A natural algorithm proceeds as follows. Solve $L_{3}$ first. Since it has no solutions on $\{0,1\}$, the variable $x_{13}$ must be equal to 2. Then $x_{7}=x_{8}$, which can be added as an equation to $L_{2}$. We check that $L_{2}$ has no solution satisfying this additional equation, which implies that $x_{7}$ must be equal to 2 and $x_{1}=x_{2}$. Finally, the linear system $L_{1}$ together with $x_{1}=x_{2}$ has no solution, contradicting $x_{1}\in\{0,1\}$.
However, $\mathrm{Singl}\mathrm{BLP}+\mathrm{AIP}$ incorrectly returns “Yes” on the instance $\mathcal{I}$. $\mathrm{Singl}\mathrm{BLP}$ does not reduce any domains (except the trivial restriction $D_{x_{1}}=\{0,1\}$), since every linear equation can be satisfied by a uniform distribution. By Claim 4.2, each constraint relation can be replaced by its parallelogram closure without changing the outcome of AIP. Since the parallelogram closure of $S_{0,1}^{=}$ is the full relation, $\mathrm{AIP}$ simply ignores all $S_{0,1}^{=}$-constraints. Taking a genuine solution of $L_{1}$ over $\{0,1\}$, together with fake all-2 solutions for $L_{2}$ and $L_{3}$, yields a global a solution that violates only $S_{0,1}^{=}$-constraints, which AIP ignores. Thus $\mathrm{AIP}$ (and hence $\mathrm{Singl}\mathrm{BLP}+\mathrm{AIP}$) returns “Yes” on the unsatisfiable instance $\mathcal{I}$.
In contrast, $\mathrm{Singl}\mathrm{AIP}$ returns “No” on $\mathcal{I}$. After fixing $x_{13}$ to 0 and 11 the $\mathrm{AIP}$ algorithm returns “No”, which excludes these values from the domain of $x_{13}$ and forces $x_{13}=2$. Hence, $x_{7}=x_{8}$. After fixing $x_{7}$ to 0 or 1, $\mathrm{AIP}$ again returns “No”, reducing the domain of $x_{7}$ to $\{2\}$ and forcing $x_{1}=x_{2}$. With $x_{1}=x_{2}$, $\mathrm{AIP}$ returns “No” on the reduced instance. Thus, $\mathrm{Singl}\mathrm{AIP}$ outputs “No” on $\mathcal{I}$, as desired.
To witness that $\mathrm{Singl}\mathrm{AIP}$ correctly solves $\operatorname{CSP}(\mathbb{Z}_{2}\leftarrow 2)$ we define a $(\underbrace{\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n})$-palette block alternating function $f(\mathbf{x}_{1},\dots,\mathbf{x}_{n})$ from $\operatorname{Pol}(\mathbb{Z}_{2}\leftarrow 2)$ as follows:
  • choose the first block $\mathbf{x}_{i}$ consisting entirely of elements from $\{0,1\}$;
  • if no such block exists, return $\mathbf{x}_{1}^{1}$;
  • otherwise, apply the alternating function to $\mathbf{x}_{i}$, that is, $f(\mathbf{x}_{1},\dots,\mathbf{x}_{n})=\mathbf{x}_{i}^{1}+\mathbf{x}_{i}^{2}+\dots+\mathbf{x}_{i}^{2\ell+1}\pmod{2}$.
The crucial difference between $\mathbb{Z}_{2}\leftarrow 2$ and $\mathbb{Z}_{2}\cup\mathbb{Z}_{3}$ is that $\mathbb{Z}_{2}\cup\mathbb{Z}_{3}$ can be solved by running the $\mathrm{AIP}$ algorithm once for each variable and each value, while for $\mathbb{Z}_{2}\leftarrow 2$ we must ran $\mathrm{AIP}$ again and again, gradually reducing the instance. We believe that $\mathbb{Z}_{2}\leftarrow 2$ also cannot be solved by $\mathbb{Z}$-affine $k$-consistency or $\mathrm{BA}^{k}$, but we do not prove this formally.
4.4 Dihedral group
The last and perhaps most important example comes from the dihedral group $\mathbf{D}_{4}$, the group of symmetries of a square. This group is generated by two elements, $r$ and $s$, and is completely defined by the equalities $r^{4}=s^{2}=1$ and $rs=sr^{3}$.
We use the following representation of the dihedral group: $\mathbf{A}=(\mathbb{Z}_{2}^{3};\circ)$, where the operation is defined by $x\circ y=(x^{1}+y^{1},x^{2}+y^{2},x^{3}+y^{3}+x^{1}\cdot y^{2})$. Then $\mathbf{A}$ is isomorphic to $\mathbf{D}_{4}$, as witnessed by the mapping $(0,0,0)\to 1$, $(0,0,1)\to r^{2}$, $(1,1,0)\to r$, $(1,1,1)\to r^{3}$, $(0,1,0)\to s$, $(0,1,1)\to sr^{2}$, $(1,0,0)\to sr$, $(1,0,1)\to sr^{3}$. Although the elements of $\mathbb{Z}_{2}^{3}$ are vectors of length 3, we do not use boldface notation for them in order to distinguish them from tuples of such vectors. However, we will still refer to their components as $a^{1}$, $a^{2}$, and $a^{3}$. Let us define a relational structure $\mathbb{D}_{4}$ corresponding to the dihedral group.
$\bm{\mathbb{D}_{4}}=(\mathbb{Z}_{2}^{3};L_{3}^{1,2},L_{2,3}^{1},L_{1,3}^{2},E_{1-2},E_{2-3}^{0},E_{1-3}^{0},R,\{(0,0,0)\})$, where
$L_{3}^{1,2}(x_{1},x_{2},x_{3},x_{4})=$ $(x_{1}^{1}=x_{2}^{1}\wedge x_{1}^{2}=x_{2}^{2}\wedge x_{3}^{1}=x_{4}^{1}\wedge x_{3}^{2}=x_{4}^{2}\wedge x_{1}^{3}+x_{2}^{3}=x_{3}^{3}+x_{4}^{3}),$
$L_{2,3}^{1}(x_{1},x_{2},x_{3},x_{4})=$ $(x_{1}^{1}=x_{2}^{1}=x_{3}^{1}=x_{4}^{1}\wedge x_{1}^{2}+x_{2}^{2}=x_{3}^{2}+x_{4}^{2}\wedge x_{1}^{3}+x_{2}^{3}=x_{3}^{3}+x_{4}^{3}),$
$L_{1,3}^{2}(x_{1},x_{2},x_{3},x_{4})=$ $(x_{1}^{2}=x_{2}^{2}=x_{3}^{2}=x_{4}^{2}\wedge x_{1}^{1}+x_{2}^{1}=x_{3}^{1}+x_{4}^{1}\wedge x_{1}^{3}+x_{2}^{3}=x_{3}^{3}+x_{4}^{3}),$
$E_{1-2}(x,y)=$ $(x^{1}=y^{2}),$
$E_{2-3}^{0}(x,y)=$ $(x^{1}=y^{1}=0\wedge x^{2}=y^{3}),$
$E_{1-3}^{0}(x,y)=$ $(x^{2}=y^{2}=0\wedge x^{1}=y^{3}),$
$R(x,y)=$ $(x^{1}=y^{2}\wedge x^{2}=y^{1}\wedge x^{3}+y^{3}=x^{1}\cdot y^{1}).$
In Section 5 we prove the following lemma.
Lemma 4.1.
$\operatorname{Pol}(\mathbb{D}_{4})=\operatorname{Clo}(x\circ y)$.
Since $\operatorname{CSP}(\mathbb{D}_{4})$ is trivial — every instance admits a solution by assigning all variables to 0 — we extend $\mathbb{D}_{4}$ with all constant relations $\{\{a\}\mid a\in\mathbb{Z}_{2}^{3}\}$ and denote the resulting structure by $\mathbb{D}_{4}^{idemp}$.
Corollary 4.2.
$\operatorname{Pol}(\mathbb{D}_{4}^{idemp})$ is the set of all idempotent operations from $\operatorname{Clo}(x\circ y)$.
We now show that $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ does not solve $\operatorname{CSP}(\mathbb{D}_{4}^{idemp})$. To witness this, we construct a concrete instance of $\operatorname{CSP}(\mathbb{D}_{4}^{idemp})$ with no solution, for which $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ nevertheless returns “Yes”.
Notice that all relations in $\mathbb{D}_{4}^{idemp}$, except for $R$, are defined by conjunctions of linear equations on the components of the variables. Hence, any instance of $\operatorname{CSP}(\mathbb{D}_{4}^{idemp})$ not involving $R$ can be solved directly by the $\mathrm{AIP}$ algorithm. However, the nonlinear part in $R$ breaks $\mathrm{AIP}$. Let us build a concrete CSP instance $\mathcal{I}$:
$L_{2,3}^{1}(x_{1},x_{2},x_{3},w)\wedge L_{1,3}^{2}(y_{1},y_{2},y_{3},w)\wedge R(x_{1},y_{1})\wedge R(x_{2},y_{2})\wedge L_{3}^{1,2}(x_{3},y_{3},u,v)\wedge u=(0,0,0)\wedge v=(0,0,1).$
Claim 4.5.
$\mathcal{I}$ has no solution.
Proof.
The constraints $L_{2,3}^{1}$, $L_{1,3}^{2}$, and $L_{3}^{1,2}$ imply that $y_{3}^{1}=x_{3}^{1}=w^{1}$ and $x_{3}^{2}=y_{3}^{2}=w^{2}$ in any solution. Then $L_{2,3}^{1}$ and $L_{1,3}^{2}$ imply that $x_{1}^{2}=x_{2}^{2}$ and $y_{1}^{1}=y_{2}^{1}$, respectively. To derive a contradiction, we sum the third components of the first five constraints (equations). Substituting the values for $u$ and $v$, we obtain
$(x_{1}^{3}+x_{2}^{3}+x_{3}^{3}+w^{3})+(y_{1}^{3}+y_{2}^{3}+y_{3}^{3}+w^{3})+(x_{1}^{3}+y_{1}^{3}+x_{1}^{1}\cdot y_{1}^{1})+$
$(x_{2}^{3}+y_{2}^{3}+x_{2}^{1}\cdot y_{2}^{1})+(x_{3}^{3}+$ $y_{3}^{3}+0+1)=0.$
This simplifies to $x_{1}^{1}\cdot y_{1}^{1}+x_{2}^{1}\cdot y_{2}^{1}=1$, which is impossible because $x_{1}^{1}=x_{2}^{1}$ and $y_{1}^{1}=y_{2}^{1}$. ∎
Thus, by examining the linear equations we deduced that $x_{1}$ and $x_{2}$ must be equal in the first two components, making it impossible to satisfy the $R$ constraints. Nevertheless, even $\mathrm{Singl}\mathrm{AIP}$ returns “Yes” on the instance $\mathcal{I}$. Below we present an AIP solution of the instance, and we claim (without a proof) that a similar solution can be constructed after fixing any variable to any value.
Claim 4.6.
$\mathrm{AIP}$ returns “Yes” on $\mathcal{I}$.
Proof.
The following table shows the weights of tuples corresponding to $L_{2,3}^{1}(x_{1},x_{2},x_{3},w)$ and $L_{1,3}^{2}(y_{1},y_{2},y_{3},w)$ in an AIP solution. The weights of tuples for the remaining constraints are straightforward.
[…]
The idea behind this AIP solution is that, for AIP, having two separate constraints $x_{1}^{1}=x_{2}^{1}$ and $x_{1}^{2}=x_{2}^{2}$ is not equivalent to having a single constraint $(x_{1}^{1},x_{1}^{2})=(x_{2}^{1},x_{2}^{2})$.
For a formal proof that $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ does not solve $\operatorname{CSP}(\mathbb{D}_{4}^{idemp})$, we apply Theorem 3.6 together with the following lemma proved in Section 5 (originally proved by Michael Kompatscher).
Lemma 4.3.
Suppose $n,\ell\geq 4$. Then $\operatorname{Clo}(x\circ y)$ does not contain an idempotent $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette block symmetric operation.
Corollary 4.4.
The algorithm $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ does not solve $\operatorname{CSP}(\mathbb{D}_{4}^{idemp})$.
5 Dihedral group $\mathbf{D}_{4}$.
In this section, we prove two properties of the dihedral group $\mathbf{D}_{4}$ that were stated in Section 4.4. Recall that $\mathbb{D}_{4}=(\mathbb{Z}_{2}^{3};L_{3}^{1,2},L_{2,3}^{1},L_{1,3}^{2},E_{1-2},E_{2-3}^{0},E_{1-3}^{0},R,\{(0,0,0)\})$. In this section addition and multiplication are always modulo 2, but $x_{i}^{2}$ denotes the second component of the vector $x_{i}$, not the second power. The following claims explain the meaning of some of the relations in $\mathbb{D}_{4}$. We assume that $f$ is an $n$-ary operation on $\mathbb{Z}_{2}^{3}$.
Claim 5.1.
$f\in\operatorname{Pol}(\{L_{3}^{1,2}\})$ if and only if there exist $g_{1},g_{2},g_{3}:\mathbb{Z}_{2}^{2n}\to\mathbb{Z}_{2}$ and $a_{1},\dots,a_{n}\in\mathbb{Z}_{2}$ such that
$f^{1}(x_{1},\dots,x_{n})$ $=g_{1}(x_{1}^{1},\dots,x_{n}^{1},x_{1}^{2},\dots,x_{n}^{2}),$
$f^{2}(x_{1},\dots,x_{n})$ $=g_{2}(x_{1}^{1},\dots,x_{n}^{1},x_{1}^{2},\dots,x_{n}^{2}),$
$f^{3}(x_{1},\dots,x_{n})$ $=a_{1}x_{1}^{3}+\dots+a_{n}x_{n}^{3}+g_{3}(x_{1}^{1},\dots,x_{n}^{1},x_{1}^{2},\dots,x_{n}^{2}).$
Proof.
Let $E=\operatorname{pr}_{1,2}(L_{3}^{1,2})$. Then $E$ is an equivalence relation and it is preserved by $f$. This implies that the first two components of $f$ depend only on the first two components of the inputs, so we can choose the corresponding $g_{1}$ and $g_{2}$. Applying $f$ to the tuples $((x_{i}^{1},x_{i}^{2},x_{i}^{3}),(x_{i}^{1},x_{i}^{2},0),(0,0,x_{i}^{3}),(0,0,0))\in L_{3}^{1,2}$ for $i=1,\dots,n$ and looking at the third component, we obtain
$f^{3}(x_{1},\dots,x_{n})=f^{3}((x_{1}^{1},x_{1}^{2},0),\dots,(x_{n}^{1},x_{n}^{2},0))+f^{3}((0,0,x_{1}^{3}),\dots,(0,0,x_{n}^{3}))+f^{3}((0,0,0),\dots,(0,0,0)).$ (1)
Similarly, we derive
$f^{3}((0,0,x_{1}^{3}),\dots,(0,0,x_{n}^{3}))=$
$f^{3}((0,0,0),(0,0,x_{2}^{3}),\dots,$ $(0,0,x_{n}^{3}))+f^{3}((0,0,x_{1}^{3}),(0,0,0),\dots,(0,0,0))+f^{3}((0,0,0),\dots,(0,0,0)).$
This holds analogously when considering the $i$-th coordinate instead of the first. Choose $a_{1},\dots,a_{n},b_{1},\dots,b_{n}\in\mathbb{Z}_{2}$ so that
$f^{3}(\underbrace{(0,0,0),\dots,(0,0,0),(0,0,z)}_{i},(0,0,0),\dots,(0,0,0))=a_{i}z+b_{i}\text{\quad\quad\quad for every $i\in[n]$.}$
Using (1), we derive that $f^{3}(x_{1},\dots,x_{n})=a_{1}x_{1}^{3}+\dots+a_{n}x_{n}^{3}+g_{3}(x_{1}^{1},\dots,x_{n}^{1},x_{1}^{2},\dots,x_{n}^{2})$ for some $g_{3}$. ∎
Claim 5.2.
$f\in\operatorname{Pol}(\{E_{1-2}\})$ if and only if there exists $g:\mathbb{Z}_{2}^{n}\to Z_{2}$ such that
$f^{1}(x_{1},\dots,x_{n})$ $=g(x_{1}^{1},\dots,x_{n}^{1}),$
$f^{2}(x_{1},\dots,x_{n})$ $=g(x_{1}^{2},\dots,x_{n}^{2}).$
Proof.
Applying $f$ to tuples $((a_{i},b_{i},c_{i}),(d_{i},a_{i},e_{i})\in E_{1-2}$ for $i=1,\dots,n$, and using the fact that the result must be in $E_{1-2}$, we obtain
$f^{1}((a_{1},b_{1},c_{1}),\dots,(a_{n},b_{n},c_{n}))=f^{2}((d_{1},a_{1},e_{1}),\dots,(d_{n},a_{n},e_{n})).$
This implies that $f^{1}$ depends only on the first components, $f^{2}$ depends only on the second components, and the operations defining $f^{1}$ and $f^{2}$ are identical. ∎
Claim 5.3.
Suppose $f\in\operatorname{Pol}(\{L_{3}^{1,2},L_{2,3}^{1}\})$. Then every monomial in the polynomial representation of $g_{3}$ (see Claim 5.1) contains at most one $x_{i}^{2}$-variable (i.e, terms like $x_{1}^{2}\cdot x_{2}^{2}$ are not allowed).
Proof.
Assume the contrary. Choose monomials of the form $x_{i_{1}}^{1}\dots x_{i_{k}}^{1}x_{j_{1}}^{2}\dots x_{j_{m}}^{2}$ with minimal $m\geq 2$; among them choose one with minimal $k\geq 0$. Set $b_{i}=1$ if $i\in\{i_{1},\dots,i_{k}\}$ and $b_{i}=0$ otherwise; $c_{i}=1$ if $i\in\{j_{1},\dots,j_{m}\}$ and $c_{i}=0$ otherwise; $d_{i}=1$ if $i\in\{j_{2},\dots,j_{m}\}$ and $d_{i}=0$ otherwise; $e_{j_{1}}=1$ and $e_{i}=0$ for $i\neq j_{1}$. Applying $f$ to the tuples $((b_{i},c_{i},0),(b_{i},d_{i},0),(b_{i},e_{i},0),(b_{i},0,0))\in L_{2,3}^{1}$ for $i=1,\dots,n$, we get a tuple from $L_{2,3}^{1}$. Examining the third components of the obtained tuple, we derive:
$g_{3}(b_{1},\dots,b_{n},c_{1},\dots,c_{n})+g_{3}(b_{1},\dots,b_{n},d_{1},\dots,d_{n})+g_{3}(b_{1},\dots,b_{n},e_{1},\dots,e_{n})+g_{3}(b_{1},\dots,b_{n},0,\dots,0)=0.$
All the monomials of $g_{3}$ involving only the first components vanish, as do all linear monomials. The only monomial contributing 1 to the above sum is $x_{i_{1}}^{1}\dots x_{i_{k}}^{1}x_{j_{1}}^{2}\dots x_{j_{m}}^{2}$, which gives us a contradiction. ∎
The next claim differs from the previous one only by swapping 1 and 2, so we omit the proof.
Claim 5.4.
Suppose $f\in\operatorname{Pol}(\{L_{3}^{1,2},L_{1,3}^{2}\})$. Then every monomial in the polynomial representation of $g_{3}$ (see Claim 5.1) contains at most one $x_{i}^{1}$-variable (i.e, terms like $x_{1}^{1}\cdot x_{2}^{1}$ are not allowed).
Let us describe a set of operations $\mathcal{D}_{4}$, which we will later prove to be the clone generated by the dihedral group. Suppose $n\in\mathbb{N}$, $a:[n]\to\mathbb{Z}_{2}$, and $c:[n]\times[n]\to\mathbb{Z}_{2}$ are such that $c_{i,j}+c_{j,i}=a_{i}\cdot a_{j}$ for every $\in[n], iji\neq j$. Then we define the following $n$-ary operation $f$ on $\mathbb{Z}_{2}^{3}$:
$f^{1}(x_{1},\dots,x_{n})$ $=a_{1}x_{1}^{1}+\dots+a_{n}x_{n}^{1},$
$f^{2}(x_{1},\dots,x_{n})$ $=a_{1}x_{1}^{2}+\dots+a_{n}x_{n}^{2},$
$f^{3}(x_{1},\dots,x_{n})$ $=a_{1}x_{1}^{3}+\dots+a_{n}x_{n}^{3}+\sum\limits_{i,j\in[n]}c_{i,j}x_{i}^{1}x_{j}^{2}.$
By $\mathcal{D}_{4}$ we denote the set of all such operations $f$ for all $n$, $a$ and $c$.
Claim 5.5.
$\mathcal{D}_{4}\subseteq\operatorname{Clo}(x\circ y)$.
Proof.
Let $f$ be an $n$-ary operation from $\mathcal{D}_{4}$ defined by some mappings $a$ and $c$. Let $\{i_{1},\dots,i_{s}\}=\{i\mid a_{i}=1\}$ with $i_{1}<i_{2}<\dots<i_{s}$. Let us show that $f$ can be defined by
$f(x_{1},\dots,x_{n})=x_{i_{1}}\circ\dots\circ x_{i_{s}}\circ\prod_{i<j\colon c_{j,i}=1}(x_{i}\circ x_{j}\circ x_{i}^{-1}\circ x_{j}^{-1})\circ\prod_{i\colon c_{i,i}=1}(x_{i}\circ x_{i}).$
At the first two components the operation $\circ$ is just an addition; hence the last two products do not affect them, and the result matches the definition. We claim that the third component of $x_{i_{1}}\circ\dots\circ x_{i_{s}}$ is $x_{i_{1}}^{3}+\dots+x_{i_{s}}^{3}+\sum\limits_{j<k}x_{i_{j}}^{1}x_{i_{k}}^{2}$. This can be shown by induction on $s$, with the case $s=2$ following directly from the definition and the inductive step being straightforward. Then we check that $(x_{i}\circ x_{j}\circ x_{i}^{-1}\circ x_{j}^{-1})=(0,0,x_{i}^{1}x_{j}^{2}+x_{i}^{2}x_{j}^{1})$ and $(x_{i}\circ x_{i})=(0,0,x_{i}^{1}x_{i}^{2})$. It remains to apply the definition of $\circ$ to verify that the third component also matches the definition. ∎
Claim 5.6.
$\operatorname{Pol}(\mathbb{D}_{4})\subseteq\mathcal{D}_{4}$.
Proof.
Let $f\in\operatorname{Pol}(\mathbb{D}_{4})$. By Claim 5.1 we choose the corresponding $g_{1},g_{2},g_{3}$ and $a_{1},\dots,a_{n}$. By Claim 5.2 we can replace $g_{1}$ and $g_{2}$ with a single operation $g$. Since the relation $\{(0,0,0)\}$ is preserved by $f$, $g_{3}(0,0,\dots,0)=0$. Applying $f$ to tuples $((0,b_{1},c_{1}),(0,d_{1},b_{1})),\dots,((0,b_{n},c_{n}),(0,d_{n},b_{n}))\in E_{2-3}^{0}$, we obtain
$g(b_{1},\dots,b_{n})=a_{1}b_{1}+\dots+a_{n}b_{n}+g_{3}(0,0,\dots,0,d_{1},\dots,d_{n})$
for all $b_{1},\dots,b_{n},d_{1},\dots,d_{n}\in\mathbb{Z}_{2}$. Similarly, using the relation $E_{1-3}^{0}$, we obtain
$g(b_{1},\dots,b_{n})=a_{1}b_{1}+\dots+a_{n}b_{n}+g_{3}(d_{1},\dots,d_{n},0,\dots,0)$
for all $b_{1},\dots,b_{n},d_{1},\dots,d_{n}\in\mathbb{Z}_{2}$. These two conditions imply that $g(z_{1},\dots,z_{n})=a_{1}z_{1}+\dots+a_{n}z_{n}$ and $g_{3}$ returns 0 whenever the first $n$ or the last $n$ arguments equal 0. Therefore, the polynomial representation of $g_{3}$ cannot have a monomial $x_{i}^{1}$ or $x_{i}^{2}$. By Claims 5.3 and 5.4, the only possible monomials are of the form $x_{i}^{1}\cdot x_{j}^{2}$. Let $g_{3}(x_{1}^{1},\dots,x_{n}^{1},x_{1}^{2},\dots,x_{n}^{2})=\sum\limits_{i,j\in[n]}c_{i,j}x_{i}^{1}x_{j}^{2}$. It remains to check that the relation $R$ enforces the property $c_{i,j}+c_{j,i}=a_{i}\cdot a_{j}$. Substituting the tuple $((1,0,0),(0,1,0))$ into the $i$-th position, $((0,1,0),(1,0,0))$ into the $j$-th position, and $((0,0,0),(0,0,0))$ into the remaining positions of $f$, we obtain the tuple $((a_{i},a_{j},c_{i,j}),(a_{j},a_{i},c_{j,i}))$, which must belong to $R$. Hence, the definition of $R$ gives $a_{i}\cdot a_{j}+c_{i,j}+c_{j,i}=0$. ∎
Lemma 4.1.
$\operatorname{Pol}(\mathbb{D}_{4})=\operatorname{Clo}(x\circ y)=\mathcal{D}_{4}$.
Proof.
We prove the inclusions $\operatorname{Pol}(\mathbb{D}_{4})\subseteq\mathcal{D}_{4}\subseteq\operatorname{Clo}(x\circ y)\subseteq\operatorname{Pol}(\mathbb{D}_{4})$. The first inclusion follows from Claim 5.6, the second follows from Claim 5.5, and the last follows from the fact that $x\circ y$ preserves all the relations from $\mathbb{D}_{4}$. ∎
Lemma 4.3.
Suppose $n,\ell\geq 4$. Then $\operatorname{Clo}(x\circ y)$ does not contain an idempotent $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette block symmetric operation.
Proof.
Assume the contrary. Let $f\in\operatorname{Clo}(x\circ y)$ be such an operation. By Lemma 5.6, $f\in\mathcal{D}_{4}$, so we can choose $a:[n]\to\mathbb{Z}_{2}$ and $c:[n]\times[n]\to\mathbb{Z}_{2}$ defining $f$. Take $i,j\in[\ell+1]$ with $i\neq j$. Substitute $(1,0,0)$ for the $i$-th coordinate and into the blocks 3 and 4 (the second overlined part), substitute $(0,1,0)$ for the $j$-th coordinate and into the blocks 5 and 6 (the third overlined part), and $(0,0,0)$ for the remaining coordinates. Since $f$ is $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette block symmetric, the result must be the same for all $i$ and $j$. The first coordinate of the result is $a_{i}+a_{2\ell+2}+a_{2\ell+3}+\dots+a_{4\ell+2}$, and the third one is
$c_{i,j}+\sum\limits_{m=1}^{2\ell+1}c_{i,4\ell+2+m}+\sum\limits_{m=1}^{2\ell+1}c_{2\ell+1+m,j}+\sum\limits_{m_{1},m_{2}\in[2\ell+1]}c_{2\ell+1+m_{1},4\ell+2+m_{1}}.$
Denote $s_{i}=\sum\limits_{m=1}^{2\ell+1}c_{i,4\ell+2+m}$, $t_{j}=\sum\limits_{m=1}^{2\ell+1}c_{2\ell+1+m,j}$. Hence, $a_{1}=a_{2}=\dots=a_{\ell+1}$, and the value $c_{i,j}+s_{i}+t_{j}$ does not depend on $i$ and $j$. Since $\ell\geq 2$, there exist $i,j\in[\ell+1],i\neq j$, such that $s_{i}+t_{i}=s_{j}+t_{j}$. This implies $c_{i,j}+c_{j,i}=0$, which, by the definition of $\mathcal{D}_{4}$, means that $a_{i}\cdot a_{j}=0$, and therefore $a_{i}=a_{j}=0$. Thus, we have showed that $a_{1}=\dots=a_{\ell+1}=0$. Repeating the argument word to word, we obtain $a_{\ell+2}=\dots=a_{2\ell+1}=0$. Finally, by swapping the blocks and repeating the same argument, we prove that $a_{1}=\dots=a_{n(2l+1)}=0$. This contradicts the fact that $f$ is idempotent, since $f((1,0,0),\dots,(1,0,0))=(1,0,0)$ implies $\sum_{i=1}^{n(2\ell+1)}a_{i}=1$. ∎
6 Characterization of Singleton Algorithms
In this section we prove the characterization of the singleton algorithms in terms of palette block symmetric functions. We start with a formal definition of a minion and show how abstract minions can be used to characterize the power of algorithms. Then, in Section 6.2 we apply the Hales-Jewett theorem to show that, under certain conditions, a homomorphic image of a minion contains more symmetric objects than the original. In the last two subsections we show how to apply this result to obtain a characterization of the algorithms $\mathrm{Singl}\mathrm{ArcCons}$ and $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$, respectively.
6.1 Minions and free structure $\mathbb{F}_{\mathcal{M}}(\mathbb{A})$
To simplify understanding of the notion of an abstract minion we start with a function minion. For a function $f\colon A^{n}\to B$ and a mapping $\pi\colon[n]\to[n^{\prime}]$, the minor of $f$ given by $\pi$, denoted by $f^{\pi}$, is the function of arity $n^{\prime}$ defined by
$f^{\pi}(x_{1},\dots,x_{n^{\prime}}):=f(x_{\pi(1)},\dots,x_{\pi(n)}).$
A subset $\mathcal{M}\subseteq\mathcal{O}_{A,B}$ is called a function minion if for any $f\in\mathcal{M}\cap\mathcal{O}_{A,B}^{(n)}$ and any mapping $\pi\colon[n]\to[n^{\prime}]$, the minor $f^{\pi}$ belongs to $\mathcal{M}$. Thus, a function minion can be viewed as a weaker version of a clone, where we only allow to permute and identify variables, and add dummy variables. For instance, the sets $\mathcal{O}_{A,B}$ and $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ are function minions.
Variables (arguments) of an $n$-ary function are usually indexed by numbers $1,2,\dots,n$, which sometimes make argumentation unnecessarily cumbersome, as it requires an additional encoding. To avoid this, we consider functions $f\colon A^{N}\to B$ for finite sets $N$, denoted by $\mathcal{O}_{A,B}^{(N)}$; that is, each function is a mapping equipped with argument names from the set $N$. Then, for a function $f\colon A^{N}\to B$ and a mapping $\pi:N\to N^{\prime}$, where $N^{\prime}$ is also a finite set, the minor of $f$ given by $\pi$ is defined by $f^{\pi}(\mathbf{a})=f(\mathbf{a}\circ\pi)$ for every $\mathbf{a}\in A^{N^{\prime}}$. The advantages of this approach can already be seen in the definition of a free structure below (compare with the respective definitions in [2, 3]).
Thus, a function minion $\mathcal{M}$ is a collection of sets $\mathcal{M}^{(N)}\subseteq\mathcal{O}_{A,B}^{(N)}$ for finite sets $N$ such that for any $f\in\mathcal{M}^{(N)}$ and any mapping $\pi\colon N\to N^{\prime}$, where $N^{\prime}$ is also a finite set, the minor $f^{\pi}$ belongs to $\mathcal{M}^{(N^{\prime})}$. Then the set of all functions $\mathcal{O}_{A,B}$ can be viewed as a minion, whose $\mathcal{O}_{A,B}^{(N)}$-part consists of all mappings $A^{N}\to B$. Similarly, $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ is a minion such that $f\in\operatorname{Pol}(\mathbb{A},\mathbb{B})^{(N)}$ whenever $f\in\mathcal{O}_{A,B}^{(N)}$ and for any $R$ from the signature of $\mathbb{A}$ and any $\mathbf{a}\colon N\to R^{\mathbb{A}}$ we have $f(\mathbf{a})\in R^{\mathbb{B}}$. Thus, $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ is no different than the usual set of polymorphisms with arbitrary names chosen for the arguments.
Now let us define the notion of an abstract minion, where instead of functions we have arbitrary objects, and taking minors satisfies the same properties we have for function minions.
Minion. (Abstract) minion $\mathcal{M}$ is a collection of sets $\mathcal{M}^{(N)}$ for finite sets $N$ together with a minor map $\mathcal{M}^{(\pi)}\colon\mathcal{M}^{(N)}\to\mathcal{M}^{(N^{\prime})}$ for every mapping $\pi\colon N\to N^{\prime}$, such that $\mathcal{M}^{(\mathrm{id}_{N})}=\mathrm{id}_{\mathcal{M}^{(N)}}$ for all finite sets $N$, and $\mathcal{M}^{(\pi)}\circ\mathcal{M}^{(\pi^{\prime})}=\mathcal{M}^{(\pi\circ\pi^{\prime})}$ whenever such a composition is well defined. When a minion is clear from context, we write $f^{\pi}$ instead of $\mathcal{M}^{(\pi)}(f)$. We refer to $\mathcal{M}^{(N)}$ as to the $N$-ary part of $\mathcal{M}$.
A collection of sets $\mathcal{M}_{0}^{(N)}$ for finite sets $N$ is called a subminion of a minion $\mathcal{M}$ if $\mathcal{M}_{0}^{(N)}\subseteq\mathcal{M}^{(N)}$ for all $N$ and the minor maps $\mathcal{M}_{0}^{(\pi)}$ and $\mathcal{M}^{(\pi)}$ coincide on $\mathcal{M}_{0}$.
Minion homomorphism. Let $\mathcal{M}$ and $\mathcal{M}^{\prime}$ be minions. A minion homomorphism from $\mathcal{M}$ to $\mathcal{M}^{\prime}$ is a collection of mappings $\xi^{(N)}\colon\mathcal{M}^{(N)}\to\mathcal{M}^{\prime(N)}$ for finite sets $N$, such that they preserve taking minors; that is, $\xi^{(N^{\prime})}(\mathcal{M}^{(\pi)}(f))=\mathcal{M}^{\prime(\pi)}(\xi^{(N)}(f))$ for all finite sets $N,N^{\prime}$, all $f\in\mathcal{M}^{(N)}$, and all mappings $\pi\colon N\to N^{\prime}$. We write $\mathcal{M}_{1}\to\mathcal{M}_{2}$ to indicate the existence of a minion homomorphism from $\mathcal{M}_{1}$ to $\mathcal{M}_{2}$.
Free structure $\mathbb{F}_{\mathcal{M}}(\mathbb{A})$. For a relational structure $\mathbb{A}$ and a minion $\mathcal{M}$, the free structure $\mathbb{F}_{\mathcal{M}}(\mathbb{A})$ is the structure with domain $\mathcal{M}^{(A)}$ (potentially infinite) and with the same signature as $\mathbb{A}$. For each $k$-ary relation $R^{\mathbb{A}}$, the relation $R^{\mathbb{F}_{\mathcal{M}}(\mathbb{A})}$ consists of all tuples $(f_{1},\dots,f_{k})\in(\mathcal{M}^{(A)})^{k}$ such that there exists $g\in\mathcal{M}^{(R^{\mathbb{A}})}$ satisfying $f_{i}=g^{\pi_{i}}$ for every $i$, where $\pi_{i}\colon R^{\mathbb{A}}\to A$ is a projection onto the $i$-th coordinate.
Lemma 6.1 ([2, 3]).
Suppose $\mathbb{A}$ and $\mathbb{B}$ are finite relational structures in the same signature and $\mathcal{M}$ is a minion. Then the following conditions are equivalent:
  • (1)
    $\mathbb{X}\to\mathbb{F}_{\mathcal{M}}(\mathbb{A})$ implies $\mathbb{X}\to\mathbb{B}$ for any finite $\mathbb{X}$;
  • (2)
    $\mathbb{F}_{\mathcal{M}}(\mathbb{A})\to\mathbb{B}$;
  • (3)
    $\mathcal{M}\to\operatorname{Pol}(\mathbb{A},\mathbb{B})$.
Proof.
[…]
Corollary 6.2.
Suppose $\mathbb{A}$ and $\mathbb{B}$ are finite relational structures in the same signature, $\mathfrak{A}$ is an algorithm for $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$, $\mathcal{M}$ is a nonempty minion, and the following conditions are equivalent for any instance $\mathbb{X}$:
  • $\mathfrak{A}$ returns “Yes” on $\mathbb{X}$;
  • $\mathbb{X}\to\mathbb{F}_{\mathcal{M}}(\mathbb{A})$.
Then $\mathfrak{A}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if $\mathcal{M}\to\operatorname{Pol}(\mathbb{A},\mathbb{B})$.
Proof.
First, let us show that $\mathfrak{A}$ returns “Yes” on any Yes-instance $\mathbb{X}$. Let $\varphi\colon\mathbb{X}\to\mathbb{A}$ be a homomorphism. Since the minion $\mathcal{M}$ is nonempty, there should be a unary object $f\in\mathcal{M}^{(\{1\})}$. Then a homomorphism $\psi\colon\mathbb{X}\to\mathbb{F}_{\mathcal{M}}(\mathbb{A})$ can be defined by $\psi(x)=f^{\pi_{x}}$, where $\pi_{x}\colon\{1\}\to A$ such that $\pi_{x}(1)=\varphi(x)$. This implies that $\mathfrak{A}$ returns “Yes” on $\mathbb{X}$.
Second, let us show that the following conditions are equivalent:
  • $\mathfrak{A}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$;
  • $\mathbb{X}\to\mathbb{F}_{\mathcal{M}}(\mathbb{A})$ implies $\mathbb{X}\to\mathbb{B}$ for any finite $\mathbb{X}$.
Assume that $\mathfrak{A}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$. Then $\mathbb{X}\to\mathbb{F}_{\mathcal{M}}(\mathbb{A})$ implies that the algorithm returns “Yes” on the instance $\mathbb{X}$, which means that $\mathbb{X}\to\mathbb{A}$ or $\mathbb{X}$ is incorrect ($\mathbb{X}\to\mathbb{B}$ and $\mathbb{X}\not\to\mathbb{A}$). Hence, $\mathbb{X}\to\mathbb{F}_{\mathcal{M}}(\mathbb{A})$ implies $\mathbb{X}\to\mathbb{B}$ for any $\mathbb{X}$. For the converse, assume that $\mathbb{X}\to\mathbb{F}_{\mathcal{M}}(\mathbb{A})$ implies $\mathbb{X}\to\mathbb{B}$. Let us show that $\mathfrak{A}$ works correctly on any instance $\mathbb{X}$. For a Yes-instance, it was shown earlier. If $\mathfrak{A}$ returns “Yes” on $\mathbb{X}$, then $\mathbb{X}\to\mathbb{F}_{\mathcal{M}}(\mathbb{A})$, which implies $\mathbb{X}\to\mathbb{B}$, hence $\mathbb{X}$ is indeed a Yes-instance.
To complete the proof, it is sufficient to use the equivalence of conditions (1) and (3) in Lemma 6.1. ∎
6.2 General characterization of singleton
Theorem 6.3 (Hales-Jewett Theorem, [8]).
Suppose $A$ and $B$ are finite sets. Then there exists $n\in\mathbb{N}$ such that for any mapping $f\colon A^{n}\to B$ there exists $(v_{1},\dots,v_{n})\in(A\cup\{x\})^{n}\setminus A^{n}$, where $x$ is a variable symbol, such that the unary mapping $h(x)\approx f(v_{1},\dots,v_{n})$ is a constant.
Suppose $\circ$ is an operation defined on $\mathcal{M}^{(n)}$ for every finite set $n$. We say that an operation $\circ$ respects taking minors if for any $\pi\colon n\to N$ and any $f_{1},f_{2}\in\mathcal{M}^{(n)}$ we have $f_{1}^{\pi}\circ f_{2}^{\pi}=(f_{1}\circ f_{2})^{\pi}$.
Theorem 6.4.
Suppose $n$ and $N$ are finite sets, $\mathcal{M}_{0}$ is a minion equipped with a semigroup operation $\circ$ that respects taking minors, $\mathcal{M}\subseteq\mathcal{M}_{0}$ is a subminion, $\mathcal{N}$ is a minion, $\mathcal{N}^{(n)}$ is finite, $h\colon\mathcal{M}\to\mathcal{N}$ is a minion homomorphism, $\mathcal{A}\subseteq\mathcal{M}_{0}^{(N)}$, $I\in\mathcal{M}_{0}^{(N)}$, $\Omega$ is a set of mappings $N\to n$, $\omega\subseteq\Omega\times\Omega$, and
  1. (1)
    $U^{\sigma_{1}}=U^{\sigma_{2}}$ for any $(\sigma_{1},\sigma_{2})\in\omega$ and $U\in\mathcal{A}$;
  2. (2)
    for any $\sigma\in\Omega$ there exists $U\in\mathcal{A}$ such that $U^{\sigma}=I^{\sigma}$;
  3. (3)
    $U_{1}\circ\dots\circ U_{\ell}\in\mathcal{M}$ for any $\ell\in\mathbb{N}$ and $(U_{1},\dots,U_{\ell})\in(\mathcal{A}\cup\{I\})^{\ell}\setminus\mathcal{A}^{\ell}$.
Then there exists $M\in\mathcal{M}^{(N)}$ such that $h(M)^{\sigma_{1}}=h(M)^{\sigma_{2}}$ for any $(\sigma_{1},\sigma_{2})\in\omega$.
Proof.
Let $\mathcal{B}$ be the set of all mappings $\Omega\to(\mathcal{N}^{(n)}\cup\{*\})$. Notice that $\mathcal{B}$ is finite. Let us choose a finite subset $\mathcal{A}_{0}\subseteq\mathcal{A}$ that still satisfies condition (2). For every $L\in\mathbb{N}$ define a mapping $\xi_{L}\colon\mathcal{A}_{0}^{L}\to\mathcal{B}$ by
$\xi_{L}(U_{1},\dots,U_{L})(\sigma)=\begin{cases}h((U_{1}\circ\dots\circ U_{L})^{\sigma}),&\text{if $(U_{1}\circ\dots\circ U_{L})^{\sigma}\in\mathcal{M}$;}\\ *&\text{otherwise.}\end{cases}$
By the Hales-Jewett theorem, choose $L$ and $(T_{1},\dots,T_{L})\in(\mathcal{A}_{0}\cup\{I\})^{L}\setminus\mathcal{A}_{0}^{L}$, where $I$ replaces the variable symbol. Let us show that $M=T_{1}\circ\dots\circ T_{L}$ satisfies all required properties. By condition (3) we have $M\in\mathcal{M}$. It remains to prove that $h(M)^{\sigma_{1}}=h(M)^{\sigma_{2}}$ for any $(\sigma_{1},\sigma_{2})\in\omega$. By condition (2) there exist $A_{1},A_{2}\in\mathcal{A}_{0}$ such that $A_{1}^{\sigma_{1}}=I^{\sigma_{1}}$ and $A_{2}^{\sigma_{2}}=I^{\sigma_{2}}$. To simplify the argument, assume that $T_{1}=T_{2}=I$ and $T_{i}\neq I$ for $i\geq 3$. Then
$h(M)^{\sigma_{1}}\overset{1}{=}h(M^{\sigma_{1}})\overset{2}{=}$ $h((T_{1}\circ\dots\circ T_{L})^{\sigma_{1}})\overset{3}{=}$
$h(T_{1}^{\sigma_{1}}\circ\dots\circ T_{L}^{\sigma_{1}})\overset{4}{=}$
$h(A_{1}^{\sigma_{1}}\circ A_{1}^{\sigma_{1}}\circ T_{3}^{\sigma_{1}}\circ\dots\circ T_{L}^{\sigma_{1}})\overset{5}{=}$
$h(A_{1}^{\sigma_{2}}\circ A_{1}^{\sigma_{2}}\circ T_{3}^{\sigma_{2}}\circ\dots\circ T_{L}^{\sigma_{2}})\overset{6}{=}$
$\xi_{L}(A_{1},A_{1},T_{3},\dots,T_{L})(\sigma_{2})\overset{7}{=}$
$\xi_{L}(A_{2},A_{2},T_{3},\dots,T_{L})(\sigma_{2})\overset{8}{=}$
$h(A_{2}^{\sigma_{2}}\circ A_{2}^{\sigma_{2}}\circ T_{3}^{\sigma_{2}}\circ\dots\circ T_{L}^{\sigma_{2}})\overset{9}{=}$
$h(T_{1}^{\sigma_{2}}\circ\dots\circ T_{L}^{\sigma_{2}})\overset{10}{=}$
$h((T_{1}\circ\dots\circ T_{L})^{\sigma_{2}})\overset{11}{=}h(M^{\sigma_{2}})\overset{12}{=}h(M)^{\sigma_{2}},$
where the equalities 1 and 12 are by the definition of the minion homomorphism, 2 and 11 are by the definition, 3 and 10 are by the property of the $\circ$, 4 and 9 are by the choice of $A_{1}$ and $A_{2}$, 5 is by condition (1), 6 and 8 are by the definition of $\xi_{L}$ and the property of $\circ$, 7 is by the choice of $L$ and $T_{1},\dots,T_{L}$. ∎
6.3 Characterization of the singleton arc-consistency algorithm
Let $\mathcal{M}_{\mathrm{SArc}}$ be the minion whose N{N}-ary part consists of matrices $M\colon N\times[L]\to\{0,1\}$, where $L\in\mathbb{N}$, such that:
  1. (1)
    for every $j\in[L]$ there exists $i\in N$ with $M(i,j)=1$ (there exists 1 in every column);
  2. (2)
    if $M(i,j)=1$ for some $i$ and $j$, then there exists $\ell\in[L]$ such that $M(i,\ell)=1$ and $M(i^{\prime},\ell)=0$ for all $i^{\prime}\neq i$ (skeleton property).
Lemma 6.5.
Suppose $\mathbb{A}$ and $\mathbb{B}$ are finite relational structures in the same signature. Then the following conditions are equivalent for any instance $\mathbb{X}$ of $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$:
  • $\mathrm{CSingl}\mbox{{ArcCons}}$ returns “Yes” on $\mathbb{X}$;
  • $\mathbb{X}\to\mathbb{F}_{\mathcal{M}_{\mathrm{SArc}}}(\mathbb{A})$.
Proof.
$\Rightarrow$ Assume that $\mathrm{CSingl}\mbox{{ArcCons}}$ returns “Yes” on $\mathbb{X}$. First, notice that finding a homomorphism $\mathbb{X}\to\mathbb{F}_{\mathcal{M}_{\mathrm{SArc}}}(\mathbb{A})$ is equivalent to finding matrices from $\mathcal{M}_{\mathrm{SArc}}$ for each constraint and each variable such that they are compatible. Since the algorithm $\mathrm{CSingl}\mbox{{ArcCons}}$ returns “Yes”, for every constraint $C=R(x_{1},\dots,x_{s})$ we obtain a nonempty relation $R_{f}$ (the final version of $R$ in the pseudocode). Let $P_{1},\dots,P_{L}$ be the list of all pairs $(C,\mathbf{a})$, where $C$ is a constraint and $\mathbf{a}$ is a tuple from the corresponding $R_{f}$. Then for every $i\in[L]$, the algorithm $\mathrm{ArcCons}$ returns a nonempty subset $D_{y}^{i}$ for each variable $y$. The matrix $M_{x}\colon A\times[L]\to\{0,1\}$ corresponding to the variable $x$ is defined by $M_{x}(a,i):=\begin{cases}1,&\text{if $a\in D_{x}^{i}$}\\ 0,&\text{otherwise}\end{cases}$. Similarly, the matrix $M_{C}\colon R^{\mathbb{A}}\times[L]\to\{0,1\}$ corresponding to a constraint $C=R(x_{1},\dots,x_{s})$ is defined by
$M_{C}(\mathbf{a},i):=\begin{cases}1,&\text{if $\mathbf{a}^{j}\in D_{x_{j}}^{i}$ for every $j\in[s]$}\\ 0,&\text{otherwise}\end{cases}.$
It follows from arc-consistency that this assignment gives us a homomorphism $\mathbb{X}\to\mathbb{F}_{\mathcal{M}_{\mathrm{SArc}}}(\mathbb{A})$.
$\Leftarrow$ Assume that $h\colon\mathbb{X}\to\mathbb{F}_{\mathcal{M}_{\mathrm{SArc}}}(\mathbb{A})$ is a homomorphism. First, notice that it is sufficient to prove the claim for connected instances $\mathbb{X}$, since the algorithm $\mathrm{CSingl}\mbox{{ArcCons}}$ essentially works with disconnected parts independently. Thus, we assume that $\mathbb{X}$ is connected. Then for any two variables $x_{1},x_{2}\in X$, the matrices $h(x_{1})$ and $h(x_{2})$ have the same width $L\in\mathbb{N}$. For each constraint $C=R(x_{1},\dots,x_{s})$, let $R^{\prime}$ be the set of all tuples $\mathbf{a}\in R$ such that the $\mathbf{a}$-th row of the matrix corresponding to $C$ has a non-zero entry. Since the matrices in $\mathcal{M}_{\mathrm{SArc}}$ are skeleton, for each tuple $\mathbf{a}\in R^{\prime}$ there exists $\ell_{C,\mathbf{a}}\in[L]$ such that the $\ell_{C,\mathbf{a}}$-th column of the matrix corresponding to $C$ has 1 only in the $\mathbf{a}$-th position. Let $D_{y}^{C,\mathbf{a}}$ be the set of all $b\in D_{y}$ such that $h(y)(b,\ell_{C,\mathbf{a}})=1$. Since the sets $(D_{y}^{C,\mathbf{a}}\colon y\in X)$ form an arc-consistent reduction (see Observation 3.1), the arc-consistency algorithm must return “Yes” after we restrict the constraint $C$ to $\mathbf{a}$. Hence, $\mathrm{CSingl}\mathrm{ArcCons}$ returns “Yes”. ∎
Corollary 6.6.
Suppose $\mathbb{A}$ and $\mathbb{B}$ are finite relational structures in the same signature. Then $\mathrm{CSingl}\mathrm{ArcCons}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if there exists a minion homomorphism $\mathcal{M}_{\mathrm{SArc}}\to\operatorname{Pol}(\mathbb{A},\mathbb{B})$.
Proof.
This follows from Corollary 6.2 and Lemma 6.5. ∎
Lemma 6.7.
Suppose $n,\ell\in\mathbb{N}$, $\mathcal{N}$ is a minion whose $n$-ary part is finite, $h\colon\mathcal{M}_{SArc}\to\mathcal{N}$ is a minion homomorphism. Then $h(\mathcal{M}_{SArc})$ contains an $(\underbrace{\overline{\ell},\dots,\overline{\ell}}_{n})$-palette block totally symmetric object.
Proof.
Let us define everything we need in Theorem 6.4. Let $N=n\cdot\ell$, with $[N]$ playing the role of $N$ and $[n]$ the role of $n$. Let $\mathcal{M}=\mathcal{M}_{SArc}$, and let $\mathcal{M}_{0}$ be the set of all matrices satisfying all the properties of $\mathcal{M}_{\mathrm{SArc}}$ but property (2) (skeleton property). By $\circ$ denote the concatenation of two matrices, which is a semigroup operation that respects taking minors. For $i\in[n]$, define $\phi(i)$ to be the matrix $N]\times[1]\to\{0,1\}$ whose only column is $(\underbrace{0,\dots,0}_{(i-1)\ell},\underbrace{1,\dots,1}_{\ell},\underbrace{0,\dots,0}_{(n-i)\ell})$. Let
$\mathcal{A}=\{\phi(a_{1})\circ\dots\circ\phi(a_{N})\mid a_{1},\dots,a_{N}\in[n]\},$
and $\mathcal{I}\in\mathcal{M}$ be the identity matrix of size $N\times N$.
A mapping $\pi\colon[N]\to[n]$ can be viewed as a tuple of length $N$, so the notions of a palette tuple and of the $i$-th block can be extended naturally to such mappings. Let $\Omega$ be the set of all $(\underbrace{\overline{\ell},\dots,\overline{\ell}}_{n})$-palette mappings $N]\to[n]$. Let $\omega$ be the set of all pairs $(\sigma_{1},\sigma_{2})$ such that, for every $i\in[n]$, the $i$-th block of $\sigma_{1}$ contains the same set of elements as the $i$-th block of $\sigma_{2}$.
Let us check the required conditions (1)-(3) of Theorem 6.4. Condition (1) follows from the definition of $\phi$. For condition (2), let $\sigma\in\Omega$. For each $i$ appearing in $\sigma$ (viewed as a tuple), let $u(i)$ be the position of a block that contains only $i$ (such a block exists since $\sigma$ is palette). Then $\phi(u(i))^{\sigma}$ has $1$ at the $i$-th position and 0 elsewhere. Hence, $(\phi(u(\sigma(1)))\circ\dots\circ\phi(u(\sigma(N))))^{\sigma}=I^{\sigma}$. Condition (3) holds because adding $I$ makes the matrix skeleton. By Theorem 6.4 we obtain $M\in\mathcal{M}_{\mathrm{SArc}}^{(N)}$ such that $h(M)^{\sigma_{1}}=h(M)^{\sigma_{2}}$ for any $(\sigma_{1},\sigma_{2})\in\omega$. Thus, $h(M)$ is an $(\underbrace{\overline{\ell},\dots,\overline{\ell}}_{n})$-palette block totally symmetric object. ∎
Lemma 6.8.
Suppose $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ has $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette block totally symmetric functions for all $\ell_{1},\dots,\ell_{n}\in\mathbb{N}$. Then $\mathrm{SinglArcCons}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$.
Proof.
Notice that by identifying variables within block of an $(L_{1},\dots,L_{n})$-palette block totally symmetric (PBTS) function we obtain a PBTS function with smaller block sizes. Therefore, for all $L_{1},\dots,L_{n}\in\mathbb{N}$ we can choose an $(L_{1},\dots,L_{n})$-PBTS function $f_{L_{1},\dots,L_{n}}\in\operatorname{Pol}(\mathbb{A},\mathbb{B})$ such that identifying variables within blocks of $f_{L_{1},\dots,L_{n}}$ yields $f_{\ell_{1},\dots,\ell_{n}}$, where $\ell_{i}\leq L_{i}$ for every $i$. To simplify notation, we write $f(S_{1},\dots,S_{n})$ meaning $f_{k_{1},\dots,k_{n}}(s_{1}^{1},\dots,s_{1}^{k_{1}},\dots,s_{n}^{1},\dots,s_{n}^{k_{n}})$, where $S_{i}=\{s_{i}^{1},\dots,s_{i}^{k_{i}}\}$ for every $i\in[n]$. This notation is justified because we only apply $f_{k_{1},\dots,k_{n}}$ to palette tuples, so the order of elements in the sets $S_{i}$ does not matter.
We need to check that if $\mathrm{SinglArcCons}$ returns “Yes” on an instance $\mathcal{I}$, then the instance has a solution in $\mathbb{B}$. Let $P_{1},\dots,P_{S}$ be all pairs $(x,a)$, where $x$ is a variable of $\mathcal{I}$ and $a\in D_{x}$ at the end of the algorithm. For each $P_{j}=(x,a)$ and each variable $y$, let $D_{y}^{j}$ be the domain of $y$ after restricting $x$ to $\{a\}$ and running arc-consistency. Let us show that a solution to $\mathcal{I}^{\mathbb{B}}$ can be defined by $s(x):=f(D_{x}^{1},\dots,D_{x}^{S})$ for every variable $x$. Let $C=R(x_{1},\dots,x_{k})$ be a constraint $\mathcal{I}$. For $j\in[S]$ set $R_{j}=R^{\mathbb{A}}\cap(D^{j}_{x_{1}}\times\dots\times D^{j}_{x_{k}})$. By the assumptions on $f$, we have $f(R_{1},\dots,R_{S})=(s(x_{1}),\dots,s(x_{k}))$. Since $f$ is a polymorphism, it follows that $(s(x_{1}),\dots,s(x_{k}))\in R^{\mathbb{B}}$. Thus, $s$ is a solution of $\mathcal{I}$ in $\mathbb{B}$, completing the proof. ∎
Theorem 3.5.
Let $(\mathbb{A},\mathbb{B})$ be a PCSP template. The following conditions are equivalent:
  1. (1)
    $\mathrm{SinglArcCons}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$;
  2. (2)
    $\mathrm{CSingl}\mathrm{ArcCons}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$;
  3. (3)
    $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ has an $(\overline{m_{1}},\dots,\overline{m_{n}})$-palette block totally symmetric function for every $n,m_{1},\dots,m_{n}\in\mathbb{N}$;
  4. (4)
    for every $N\in\mathbb{N}$ there exist $n,m_{1},\dots,m_{n}\geq N$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ has an $(\overline{m_{1}},\dots,\overline{m_{n}})$-palette block totally symmetric function.
Proof.
[…]
6.4 Characterization of the singleton BLP+AIP
Let $\mathcal{M}_{\mathrm{S(BLP+AIP)}}$ be the minion whose N{N}-ary part consists of pairs of matrices $M_{\mathrm{BLP}}\colon N\times[L]\to\mathbb{Q}\cap[0,1]$ and $M_{\mathrm{AIP}}\colon N\times[L]\to\mathbb{Z}$, where $L\in\mathbb{N}$, such that:
  1. (1)
    $\sum\limits_{i\in N}M_{\mathrm{BLP}}(i,j)=1$ and $\sum\limits_{i\in N}M_{\mathrm{AIP}}(i,j)=1$ for every $j\in[L]$;
  2. (2)
    if $M_{\mathrm{AIP}}(i,j)\neq 0$ for some $i,j$, then $M_{\mathrm{BLP}}(i,j)>0$;
  3. (3)
    if $M_{\mathrm{BLP}}(i,j)>0$ for some $i$ and $j$, then there exists $\ell\in[L]$ such that $M_{\mathrm{BLP}}(i,\ell)=M_{\mathrm{AIP}}(i,\ell)=1$ and $M_{\mathrm{BLP}}(i^{\prime},\ell)=M_{\mathrm{AIP}}(i^{\prime},\ell)=0$ for all $i^{\prime}\neq i$ (skeleton property).
Lemma 6.9.
Suppose $\mathbb{A}$ and $\mathbb{B}$ are finite relational structures in the same signature. Then the following conditions are equivalent for any instance $\mathbb{X}$ of $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$:
  • $\mathrm{CSingl}(\mathrm{BLP}+\mathrm{AIP})$ returns “Yes” on $\mathbb{X}$;
  • $\mathbb{X}\to\mathbb{F}_{\mathcal{M}_{\mathrm{S(BLP+AIP)}}}(\mathbb{A})$.
Proof.
$\Rightarrow$ Assume that $\mathrm{CSingl}(\mathrm{BLP}+\mathrm{AIP})$ returns “Yes” on $\mathbb{X}$. First, notice that finding a homomorphism $\mathbb{X}\to\mathbb{F}_{\mathcal{M}_{\mathrm{S(BLP+AIP)}}}(\mathbb{A})$ is equivalent to finding pairs of matrices from $\mathcal{M}_{\mathrm{S(BLP+AIP)}}$ for each constraint and each variable so that they are compatible. Since the algorithm $\mathrm{CSingl}(\mathrm{BLP}+\mathrm{AIP})$ returns “Yes”, for every constraint $C=R(x_{1},\dots,x_{s})$ we obtain a nonempty relation $R_{f}$ (the final version of $R$ in the pseudocode). Let $P_{1},\dots,P_{L}$ be the list of all pairs $(C,\mathbf{a})$ with $C$ a constraint and $\mathbf{a}$ a tuple from the corresponding $R_{f}$. For every $i\in[L]$ the algorithm $\mathrm{BLP}+\mathrm{AIP}$ produces a BLP solution $s_{\mathrm{BLP}}^{i}$ and an AIP solution $s_{\mathrm{AIP}}^{i}$. Then, the BLP and AIP matrices corresponding to a variable $x$ can be defined by $M^{\mathrm{BLP}}_{x}(a,i)=s_{\mathrm{BLP}}^{i}(x^{a})$ and $M^{\mathrm{AIP}}_{x}(a,i)=s_{\mathrm{AIP}}^{i}(x^{a})$. Similarly, the BLP and AIP matrices corresponding to a constraint $C$ are $M^{\mathrm{BLP}}_{C}(\mathbf{a},i)=s_{\mathrm{BLP}}^{i}(C^{\mathbf{a}})$ and $M^{\mathrm{AIP}}_{C}(\mathbf{a},i)=s_{\mathrm{AIP}}^{i}(C^{\mathbf{a}})$. By definition of $\mathrm{BLP}+\mathrm{AIP}$, this assignment yields a homomorphism $\mathbb{X}\to\mathbb{F}_{\mathcal{M}_{\mathrm{S(BLP+AIP)}}}(\mathbb{A})$.
$\Leftarrow$ Assume that $h\colon\mathbb{X}\to\mathbb{F}_{\mathcal{M}_{\mathrm{S(BLP+AIP)}}}(\mathbb{A})$ is a homomorphism. First, notice that it is sufficient to prove the claim for connected instances $\mathbb{X}$ because the algorithm $\mathrm{CSingl}(\mathrm{BLP}+\mathrm{AIP})$ essentially works with disconnected parts independently. Thus, we assume that $\mathbb{X}$ is connected. Then for any two variables $x_{1},x_{2}\in X$ the matrices in $h(x_{1})$ and $h(x_{2})$ have the same width $L\in\mathbb{N}$. For each constraint $C=R(x_{1},\dots,x_{s})$, let $R^{\prime}$ be the set of all tuples $\mathbf{a}\in R$ such that the $\mathbf{a}$-th row of the BLP matrix corresponding to $C$ has a non-zero entry. Since the matrices in $\mathcal{M}_{\mathrm{S(BLP+AIP)}}$ are skeleton, for each tuple $\mathbf{a}\in R^{\prime}$ there exists $\ell_{C,\mathbf{a}}\in[L]$ such that the $\ell_{C,\mathbf{a}}$-th column of the BLP matrix corresponding to $C$ has 1 only in the $\mathbf{a}$-th position. Then BLP and AIP solutions that $\mathrm{BLP}+\mathrm{AIP}$ returns after fixing $C$ to $\mathbf{a}$ can be defined by $s_{\mathrm{BLP}}(C_{0}^{\mathbf{a}_{0}})=M^{\mathrm{BLP}}_{C_{0}}(\mathbf{a}_{0},\ell_{C,\mathbf{a}})$, $s_{\mathrm{AIP}}(C_{0}^{\mathbf{a}_{0}})=M^{\mathrm{AIP}}_{C_{0}}(\mathbf{a}_{0},\ell_{C,\mathbf{a}})$, $s_{\mathrm{BLP}}(y^{b})=M^{\mathrm{BLP}}_{y}(b,\ell_{C,\mathbf{a}})$, $s_{\mathrm{AIP}}(y^{b})=M^{\mathrm{AIP}}_{y}(b,\ell_{C,\mathbf{a}})$ for all constraints $C_{0}$, their tuples $\mathbf{a}_{0}$, variables $y$, and $b\in D_{y}$. Hence, $\mathrm{CSingl}(\mathrm{BLP}+\mathrm{AIP})$ returns “Yes” on $\mathbb{X}$. ∎
Corollary 6.10.
Suppose $\mathbb{A}$, $\mathbb{B}$ are finite relational structures in the same signature. Then $\mathrm{CSingl}(\mathrm{BLP}+\mathrm{AIP})$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if there exists a minion homomorphism $\mathcal{M}_{\mathrm{S(BLP+AIP)}}\to\operatorname{Pol}(\mathbb{A},\mathbb{B})$.
Proof.
This follows from Corollary 6.2 and Lemma 6.9. ∎
Lemma 6.11.
Suppose $n,\ell\in\mathbb{N}$, $\mathcal{N}$ is a minion whose $n$-ary part is finite, $h\colon\mathcal{M}_{\mathrm{S(BLP+AIP)}}\to\mathcal{N}$ is a minion homomorphism. Then $h(\mathcal{M}_{\mathrm{S(BLP+AIP)}})$ contains an $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette block symmetric object.
Proof.
We define all components needed to apply Theorem 6.4. Let $n\cdot(2\ell+1), [N][N$] play the role of $N$, and $[n]$ play the role of $n$. Let $\mathcal{M}=\mathcal{M}_{\mathrm{S(BLP+AIP)}}$ and $\mathcal{M}_{0}$ be the set of all pairs of matrices satisfying all properties of $\mathcal{M}_{\mathrm{S(BLP+AIP)}}$ except property (3) (skeleton property). Define $\circ$ as the concatenation of matrices, applied independently for BLP and AIP matrices. This is a semigroup operation that respects taking minors. For $i\in[n]$ define a pair of matrices $\phi(i)$. Its BLP matrix $N]\times[1]\to\mathbb{Q}\cap[0,1]$ has one column $(\underbrace{0,\dots,0}_{(i-1)(2\ell+1)},\underbrace{\frac{1}{2\ell+1},\dots,\frac{1}{2\ell+1}}_{2\ell+1},\underbrace{0,\dots,0}_{(n-i)(2\ell+1)})$. Its AIP matrix $N]\times[1]\to\mathbb{Z}$ has one column $(\underbrace{0,\dots,0}_{(i-1)(2\ell+1)},\underbrace{1,\dots,1}_{\ell+1},\underbrace{-1,\dots,-1}_{\ell},\underbrace{0,\dots,0}_{(n-i)(2\ell+1)})$. Let $\mathcal{A}=\{\phi(a_{1})\circ\dots\circ\phi(a_{N})\mid a_{1},\dots,a_{N}\in[n]\},$ and let $\mathcal{I}\in\mathcal{M}$ be the pair of identity matrix of size $N\times N$.
Mappings $\pi\colon[N]\to[n]$ can be viewed as tuples of length $N$, and the notions of a palette tuple and the $i$-th block can be extended naturally to such mappings. Let $\Omega$ be the set of all $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette mappings $N]\to[n]$. Let $\omega$ be the set of all pairs $(\sigma_{1},\sigma_{2})$ such that for every $i\in[2n]$ the $i$-th blocks of $\sigma_{1}$ and $\sigma_{2}$ differ only by a permutation of their elements.
Let us check conditions (1)-(3) of Theorem 6.4. Condition (1) follows from the definition of $\phi$. Let us prove that condition (2) is satisfied for every $\sigma\in\Omega$. For each $i$ appearing in $\sigma$ (viewed as a tuple) let $u(i)\in[n]$ be the position of an overlined block containing only $i$ (such a block exists since $\sigma$ is palette). Then both matrices of $\phi(u(i))^{\sigma}$ have $1$ at the $i$-th position and 0 elsewhere. Hence, $(\phi(u(\sigma(1)))\circ\dots\circ\phi(u(\sigma(N))))^{\sigma}=I^{\sigma}$. Condition (3) holds because adding $I$ makes the matrices skeleton. By Theorem 6.4 we can find $M\in\mathcal{M}_{\mathrm{S(BLP+AIP)}}^{(N)}$ such that $h(M)^{\sigma_{1}}=h(M)^{\sigma_{2}}$ for all $(\sigma_{1},\sigma_{2})\in\omega$. Hence, $h(M)$ is an $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette block symmetric object. ∎
Lemma 6.12.
Suppose for every $N\in\mathbb{N}$ there exist $n\geq N$, $k_{1},\dots,k_{n}\in\mathbb{N}$, and $m_{1,1},\dots,m_{1,k_{1}},\dots$, $m_{n,1},\dots,m_{n,k_{n}}\geq N$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\overline{m_{1,1},\dots,m_{1,k_{1}}},\overline{m_{2,1},\dots,m_{2,k_{2}}},\dots,\overline{m_{n,1},\dots,m_{n,k_{n}}})$-palette block symmetric function. Then $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$.
Proof.
We need to check that if $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ returns “Yes” on an instance $\mathcal{I}$, then the instance has a solution in $\mathbb{B}$. Let $P_{1},\dots,P_{S}$ be all pairs $(x,a)$ where $x$ is a variable of $\mathcal{I}$ and $a\in D_{x}$ at the end of the algorithm. For each $P_{j}=(x,a)$ there exist BLP and AIP solutions to the instance obtained by restricting the variable $x$ to $\{a\}$. Then the argument we used after the Theorem 3.4 implies that for some $N>S$, there exists a $d$-solution $s_{d}^{j}$ to this instance for every $d\geq N$.
Choose an $(\overline{m_{1,1},\dots,m_{1,k_{1}}},\overline{m_{2,1},\dots,m_{2,k_{2}}},\dots,\overline{m_{n,1},\dots,m_{n,k_{n}}})$-palette block symmetric function $f\in\operatorname{Pol}(\mathbb{A},\mathbb{B})$, where $n\geq N$, $k_{1},\dots,k_{n}\in\mathbb{N}$, and $m_{1,1},\dots,m_{1,k_{1}},\dots,m_{n,1},\dots,m_{n,k_{n}}\geq N$. For $j\in[S]$ and $i\in[k_{j}]$, let $\alpha_{x}^{j,i}$ be a tuple of length $m_{j,i}$, where each element $a\in A$ appears $s_{m_{j,i}}^{j}(x^{a})$ times. For $j\in\{S+1,S+2,\dots,n\}$ and $i\in[k_{j}]$, let $\alpha_{x}^{j,i}$ be a tuple of length $m_{j,i}$ where each element $a\in A$ appears $s_{m_{j,i}}^{1}(x^{a})$ times. Let us show that a solution to $\mathcal{I}^{\mathbb{B}}$ can be defined by $s(x):=f(\alpha_{x}^{1,1},\dots,\alpha_{x}^{1,k_{1}},\dots,\alpha_{x}^{n,1},\dots,\alpha_{x}^{n,k_{n}})$ for every variable $x$. Let $C=R(x_{1},\dots,x_{k})$ be a constraint of $\mathcal{I}$. For $j\in[S]$ and $i\in[k_{j}]$ let $M_{j,i}$ be the matrix whose columns are tuples from $R$ where each tuple $\mathbf{a}$ appears $s_{m_{j,i}}^{j}(C^{\mathbf{a}})$ times. Similarly, for $j\in\{S+1,S+2,\dots,n\}$ and $i\in[k_{j}]$ let $M_{j,i}$ be the matrix whose columns are tuples from $R$ where each tuple $\mathbf{a}$ appears $s_{m_{j,i}}^{1}(C^{\mathbf{a}})$ times. Our assumptions about $f$ imply that $f(M_{1,1},\dots,M_{1,k_{1}},\dots,M_{n,1},\dots,M_{n,k_{n}})=(s(x_{1}),\dots,s(x_{k}))$. Since $f\in\operatorname{Pol}(\mathbb{A},\mathbb{B})$, it follows that $(s(x_{1}),\dots,s(x_{k}))\in R^{\mathbb{B}}$. Thus, $s$ is a solution of $\mathcal{I}$ in $\mathbb{B}$. ∎
Theorem 3.6.
Let $(\mathbb{A},\mathbb{B})$ be a PCSP template. The following conditions are equivalent:
  1. (1)
    $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$;
  2. (2)
    $\mathrm{CSingl}(\mathrm{BLP}+\mathrm{AIP})$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$;
  3. (3)
    $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette block symmetric function for every $\ell,n\in\mathbb{N}$;
  4. (4)
    for every $N\in\mathbb{N}$ there exist $n\geq N$, $k_{1},\dots,k_{n}\in\mathbb{N}$, and $m_{1,1},\dots,m_{1,k_{1}},\dots,m_{n,1},\dots,m_{n,k_{n}}\geq N$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\overline{m_{1,1},\dots,m_{1,k_{1}}},\overline{m_{2,1},\dots,m_{2,k_{2}}},\dots,\overline{m_{n,1},\dots,m_{n,k_{n}}})$-palette block symmetric function.
Proof.
[…]
7 Palette block symmetric terms for small algebras
A proof of Theorem 3.7 will soon appear here.