### v2

# Singleton algorithms for the Constraint Satisfaction Problem

Abstract
A natural strengthening of an algorithm for the (promise) constraint satisfaction problem is its singleton version: we first fix a constraint to a tuple from the constraint relation, then run the algorithm, and remove the tuple from the constraint if the answer is negative. We characterize the power of the singleton versions of standard universal algorithms for the (promise) CSP over a fixed template in terms of the existence of a minion homomorphism. Using the Hales-Jewett theorem, we show that for finite relational structures this minion condition is equivalent to the existence of polymorphisms with certain symmetries, which we call palette symmetric polymorphisms. By proving the existence of such polymorphisms we establish that the singleton version of the BLP+AIP algorithm solves all (multi-sorted) tractable CSPs over domains of size at most 7. We further show that already for domain size 8 there exists a relational structure arising from the dihedral group $D_{4}$ that does not admit palette symmetric polymorphisms and cannot be solved by singleton BLP+AIP. By providing concrete CSP templates, we illustrate the limitations of linear programming, the power of the singleton versions, and the elegance of palette symmetric polymorphisms. Surprisingly, we also identify a concrete temporal relational structure whose CSP can be solved by a singleton algorithm after sampling, yet it does not admit palette symmetric polymorphisms. This highlights that the Hales-Jewett argument applies only to finite structures. Nevertheless, we introduce generalized palette polymorphisms for every tractable temporal relational structure and show that they yield a rounding procedure implying that the corresponding singleton algorithm solves its CSP after sampling.
1 Introduction
1.1 Constraint Satisfaction Problem
The Constraint Satisfaction Problem (CSP) is the problem of deciding whether there is an assignment to a set of variables subject to some specified constraints. Formally, the Constraint Satisfaction Problem is defined as a triple $\langle X,D,\mathbf{C}\rangle$, where
  • $X$ is a finite set of variables,
  • $D$ is a mapping that assigns a set $D_{x}$ to every variable $x\in X$, called the domain of $x$;
  • $\mathbf{C}$ is a set of constraints $C=R(x_{1},\dots,x_{m})$, where $m\in\mathbb{N}$, $x_{1},\dots,x_{m}\in X$, and $R\subseteq D_{x_{1}}\times\dots\times D_{x_{m}}$, called constraint relation.
By $\operatorname{FiniteCSP}$ we denote the decision problem whose input is a triple $\langle X,D,\mathbf{C}\rangle$, where all the domains are finite and constraint relations are given by listing the tuples.
In general the problem $\operatorname{FiniteCSP}$ is NP-complete, and to get tractable cases we usually restrict the set of allowed constraints. For a finite relational structure $\mathbb{A}$, also called the constraint language or template, by $\operatorname{CSP}(\mathbb{A})$ we denote the restricted version of $\operatorname{FiniteCSP}$ where each domain is $A$ and the constraint relations are from $\mathbb{A}$. Alternatively, the input $\langle X,D,\mathbf{C}\rangle$ of $\operatorname{CSP}(\mathbb{A})$ can be viewed as a relational structure $\mathbb{X}$ in the same signature $\sigma$ as $\mathbb{A}$ on the domain $X$ such that $(x_{1},\dots,x_{m})\in R^{\mathbb{X}}$ for $R\in\sigma$ if and only if $R(x_{1},\dots,x_{m})$ is a constraint from $\mathbf{C}$. Then $\operatorname{CSP}(\mathbb{A})$ is equivalent to checking whether there exists a homomorphism $\mathbb{X}\to\mathbb{A}$. Nevertheless, we prefer not to fix the target structure $\mathbb{A}$ and work with a triple $\langle X,D,\mathbf{C}\rangle$ as standard algorithms for the CSP change the constraint relations and even domains of the variables. We have the following characterization of the complexity of $\operatorname{CSP}(\mathbb{A})$ known as the CSP Dichotomy Conjecture.
Theorem 1.1 ([43, 42, 15, 16]).
Suppose $\mathbb{A}$ is a finite relational structure. Then $\operatorname{CSP}(\mathbb{A})$ is solvable in polynomial time if there exists a WNU polymorphism of $\mathbb{A}$; $\operatorname{CSP}(\mathbb{A})$ is NP-complete otherwise.
The NP-hardness in the above theorem follows from [18, 17] and [35]. The essential part of each proof of the CSP Dichotomy Conjecture is a polynomial algorithm that works for all tractable cases. Both polynomial algorithms [43, 42, 15, 16] we know are not universal in the sense that they work only for a fixed structure $\mathbb{A}$, and one of the main open questions in the area is the existence of a universal algorithm.
A natural generalization of the CSP is the Promise Constraint Satisfaction Problem (PCSP). For two relational structures $\mathbb{A}$ and $\mathbb{B}$ in the same signature such that $\mathbb{A}\to\mathbb{B}$ by $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ we denote the following decision problem. Given a structure $\mathbb{X}$ in the same signature. Distinguish between the case when $\mathbb{X}\to\mathbb{A}$ and the case when $\mathbb{X}\not\to\mathbb{B}$. Thus, we have an additional promise about the input that one of these two conditions holds. The complexity of $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ is widely open even for relational structures $\mathbb{A}$ and $\mathbb{B}$ on a 2-element set [31, 3, 1].
1.2 Universal algorithms and their singleton versions
We would like to find an algorithm that is universal not only in the sense that it works for any constraint language $\mathbb{A}$, but also in the following sense. An algorithm $\mathfrak{A}$ is called universal if:
  1. (1)
    $\mathfrak{A}$ returns “Yes” or “No” on any instance $\mathcal{I}$ of $\operatorname{FiniteCSP}$;
  2. (2)
    $\mathfrak{A}$ runs in polynomial time;
  3. (3)
    if $\mathfrak{A}$ returns “No” on an instance $\mathcal{I}$, then $\mathcal{I}$ has no solutions.
Thus, the No-answer of a universal algorithm is always correct, while the Yes-answer may be wrong. For example, the trivial algorithm that always returns “Yes” satisfies the above conditions. Since $\operatorname{FiniteCSP}$ is NP-complete, we cannot expect to find a universal algorithm that always returns the correct answer. Our aim, therefore, is to find a universal algorithm that works correctly on the largest possible class of inputs from $\operatorname{FiniteCSP}$. This class can be specified either by restricting the constraint language or by imposing an additional promise on the input. We envision the following major steps toward the most powerful universal algorithm.
  • CSP over a fixed template. The first goal is an algorithm that works correctly on all instances of $\operatorname{CSP}(\mathbb{A})$, whenever $\operatorname{CSP}(\mathbb{A})$ is solvable in polynomial time (by a non-universal algorithm).
  • Perfect Matching Problem. In addition, we expect the algorithm to solve the perfect matching problem correctly. This problem can be viewed as a CSP on $\{0,1\}$ over the relations 1IN2, 1IN3, 1IN4, … with the additional structural restriction that every variable appears exactly twice. We highlight this problem because it is a particularly elegant tractable CSP for which all known algorithms are non-universal and designed specifically for this problem.
  • PCSP over a fixed template. The ultimate goal is an algorithm that solves all instances of $\operatorname{PCSP}(\mathbb{A};\mathbb{B})$ that are solvable in polynomial time. That is, the algorithm should return “No” on any instance $\mathbb{X}$ of $\operatorname{CSP}(\mathbb{A})$ whenever $\mathbb{X}\not\to\mathbb{B}$.
In the paper we consider classic universal algorithms such as the arc-consistency (AC) algorithm, Basic Linear Programming (BLP), Affine Integer Programming (AIP), and their combinations (see Section 3).
The main role in this paper is played by the singleton versions of the algorithms. Such algorithms also appear in the literature under the name “cohomological” [37, 34], but we prefer the term singleton, since it appeared earlier [22] and, in our opinion, better reflects the underlying idea. For a universal algorithm $\mathfrak{A}$ by $\mathrm{CSingl}\mathfrak{A}$ we denote the algorithm that gradually reduces constraint relations in the following way. It fixes a constraint to some tuple from the constraint relation, then runs the algorithm $\mathfrak{A}$, and removes the tuple from the constraint if the answer is negative. It repeats this for every constraint and every tuple until no more tuples can be removed. If some constraint relation becomes empty, it returns “No”. Otherwise, it returns “Yes”.
A slight modification of this algorithm is $\mathrm{Singl}\mathfrak{A}$, where instead of fixing a constraint to a tuple, it fixes a variable to a value and removes the value from the domain if the answer of $\mathfrak{A}$ is negative. For most of the algorithms, $\mathrm{Singl}\mathfrak{A}$ is not stronger than $\mathrm{CSingl}\mathfrak{A}$, but this is not obvious for $\mathrm{Singl}\mathrm{AIP}$, as $\mathrm{Singl}$-algorithm changes all the constraints containing a variable $x$ when removing an element from the domain of $x$ (see Remark 3.1). The precise definitions are given in Section 3.4.
There is another natural way to strengthen a universal algorithm. Instead of arc consistency, one may consider higher levels of local consistency, such as $k$-consistency [5]. One can also combine this with AIP by first enforcing some level of consistency and then applying the AIP relaxation to partial solutions [26]. Finally, one can further strengthen such algorithms by considering the singleton (cohomological) version [37]. A major open question in the area is whether some combination of a sufficient level of consistency and a sufficient level of linear relaxation can solve all tractable CSPs. Studying the limitations of such algorithms is currently a very active topic [26, 24, 34, 20]. Only recently, the seemingly simple conjecture of Dalmau and Opršal [26] that an affine relaxation above $k$-consistency can solve all tractable CSPs was disproved [34]. Several recent papers demonstrate how such algorithms can be fooled by random instances and regular graphs [20, 21, 25, 34]. Since higher levels and hierarchies are outside the scope of the present paper, we do not define them here and instead refer the reader to the above references.
1.3 Main results
As shown in [3], the polymorphisms $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ not only characterize the complexity of $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ but also determine which algorithms can solve it. For instance, the arc-consistency algorithm solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains totally symmetric functions of all arities. Similarly, BLP and AIP are characterized by symmetric and alternating polymorphisms, respectively.
Following the ideas from [3], we characterize the power of the singleton versions of such algorithms by the existence of a polymorphism in $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ that is symmetric enough. The first main result of the paper is presented in Table 1. For every algorithm or combination of algorithms from Section 3, we provide a criterion for the algorithm to solve $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$. This criterion is formulated in terms of palette symmetric polymorphisms, which are formally defined in Section 3.4. Similar polymorphisms first appeared in [22] in the study of the power of the singleton arc-consistency algorithm. To explain the idea of such polymorphisms, let us consider the singleton version of the BLP. As shown in Table 1, this algorithm solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\underbrace{\overline{m},\dots,\overline{m}}_{n})$-palette symmetric function for all $m$ and $n$. Such a function has arity $m\cdot n$, with its arguments divided into $n$ blocks of $m$ variables each. We require the function to be symmetric within blocks (permutations of variables inside a block do not change the result), but only on tuples that additionally satisfy the palette property. A tuple is a palette tuple if any element that appears in it must completely fill one of the overlined blocks. For example, $(\overline{3},\overline{3},\overline{3})$-palette symmetric function $f$ must satisfy the identity
$f(\underbrace{x,x,x},\underbrace{x,y,x},\underbrace{y,y,y})=f(\underbrace{x,x,x},\underbrace{y,x,x},\underbrace{y,y,y})$
since the first and third blocks are overlined and completely filled with $x$ and $y$, respectively.
Among the possible combinations of algorithms, the singleton version of the BLP+AIP algorithm turned out to be very powerful. The second main result of the paper is the following theorem.
Theorem 1.2.
Suppose $\mathbb{A}$ is a finite relational structure having a WNU polymorphism such that $|\operatorname{pr}_{i}(R^{\mathbb{A}})|<8$ for every $R$ and $i\in[\operatorname{ar}(R)]$. Then $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ solves $\operatorname{CSP}(\mathbb{A})$.
Table 1: Characterization of the algorithms
Algorithms $\mathfrak{A}$ $\mathfrak{A}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if for all $m,\ell,n\in\mathbb{N}$ the set $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains
AC $m$-ary totally symmetric function
$\mathrm{BLP}$ $m$-ary symmetric function
$\mathrm{AIP}$ $(2\ell+1)$-ary alternating function
$\mbox{{AC}}\wedge\mathrm{AIP}$ $(m,2n+1)$-block function whose first block is totally symmetric and the second block is alternating
$\mathrm{BLP}\wedge\mathrm{AIP}$ $(m,2\ell+1)$-block function whose first block is symmetric and second block is alternating
$\mbox{{AC}}+\mathrm{AIP}$ $(2\ell+1)$-ary weakly alternating function
$\mathrm{BLP}+\mathrm{AIP}$ $(m+1,m)$-block symmetric function
$\mathrm{Singl}\mbox{{AC}}$ $\mathrm{CSingl}\mbox{{AC}}$ $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n})$-palette totally symmetric function
$\mathrm{Singl}\mathrm{BLP}$ $\mathrm{CSingl}\mathrm{BLP}$ $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n})$-palette symmetric function
$\mathrm{CSingl}\mathrm{AIP}$ $(\underbrace{\overline{2\ell+1},\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n})$-palette alternating function
$\mathrm{CSingl}(\mbox{{AC}}\wedge\mathrm{AIP})$ $(\underbrace{\overline{m,2\ell+1},\overline{m,2\ell+1},\dots,\overline{m,2\ell+1}}_{2n})$-palette function whose odd blocks are totally symmetric and even blocks are alternating
$\mathrm{Singl}(\mbox{{AC}}+\mathrm{AIP})$ $\mathrm{CSingl}(\mbox{{AC}}+\mathrm{AIP})$ $(\underbrace{\overline{2\ell+1},\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n})$-palette weakly alternating function
$\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ $\mathrm{CSingl}(\mathrm{BLP}+\mathrm{AIP})$ $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette symmetric function
$\mathrm{CSingl}(\mathrm{BLP}\wedge\mathrm{AIP})$ $(\underbrace{\overline{m,2\ell+1},\overline{m,2\ell+1},\dots,\overline{m,2\ell+1}}_{2n})$-palette function whose odd blocks are symmetric and even blocks are alternating
$\mathrm{Singl}\mathrm{BLP}+\mathrm{AIP}$ (CLAP) $\mathrm{CSingl}\mathrm{BLP}+\mathrm{AIP}$ $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n},2\ell+1)$-palette function whose first $n$ blocks are symmetric and last block is alternating
$\mathrm{Singl}\mathrm{AIP}+\mbox{{AC}}$ $\mathrm{CSingl}\mathrm{AIP}+\mbox{{AC}}$ $(\underbrace{\overline{2\ell+1},\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n},m)$-palette function whose first $n$ blocks are alternating and last block is totally symmetric
$\mathrm{CSingl}\mathrm{AIP}+\mathrm{BLP}$ $(\underbrace{\overline{2\ell+1},\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n},m)$-palette function whose first $n$ blocks are alternating and last block is symmetric
$\mathrm{Singl}\mbox{{AC}}+\mathrm{BLP}$ $\mathrm{CSingl}\mbox{{AC}}+\mathrm{BLP}$ $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n},\ell)$-palette function whose first $n$ blocks are totally symmetric and last block is symmetric.
$\mathrm{Singl}\mbox{{AC}}+\mathrm{AIP}$ $\mathrm{CSingl}\mbox{{AC}}+\mathrm{AIP}$ $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n},2\ell+1)$-palette function whose first $n$ blocks are totally symmetric and last block is alternating.
Thus, up to domain size 7 (even in the multi-sorted case), all tractable CSPs can be solved by this very simple algorithm. However, this is no longer true for domain size 8: as observed by Kompatscher, the structure on 8 elements arising from the dihedral group $\mathbf{D}_{4}$ does not admit palette symmetric polymorphisms, implying that $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ fails to solve it.
As a byproduct, we also show that the two versions of the singleton algorithms — fixing a tuple of a constraint ($\mathrm{CSingl}$) or a value of a variable ($\mathrm{Singl}$) — have the same power for most of the algorithms. Surprisingly, in Section 6 we discovered two structures $\mathbb{A}$ and $\mathbb{B}$, where $\mathbb{A}$ is finite and $\mathbb{B}$ is infinite, such that $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ can be solved by $\mathrm{CSingl}\mathrm{AC}$ but not by $\mathrm{Singl}\mathrm{AC}$. This emphasizes that the equivalence of $\mathrm{Singl}$ and $\mathrm{CSingl}$ for finite structures is nontrivial and the use of the Hales-Jewett theorem is essential.
In Section 4 we define several CSP templates to illustrate the limitations of linear programming algorithms such as BLP, AIP, and $\mathrm{BLP}+\mathrm{AIP}$. By providing concrete unsatisfiable instances on which $\mathrm{BLP}+\mathrm{AIP}$ incorrectly returns “yes”, we demonstrate that the failure of the algorithm arises because the corresponding template is, in a certain sense, disconnected. We then show that the singleton version overcomes this difficulty, as fixing an element allows it to solve each connected part separately. Finally, we show that the template arising from the dihedral group $\mathbf{D}_{4}$ cannot be solved even by the singleton version of $\mathrm{BLP}+\mathrm{AIP}$. Thus, we identify another fundamental reason why linear programming algorithms fail. Notice that in this section we not only present instances that fool the algorithms but also demonstrate the characterization of the singleton versions of the algorithms by providing concrete palette symmetric polymorphisms.
In Section 6, we apply our ideas to temporal CSPs, where a relational structure is called temporal if its domain is $\mathbb{Q}$ and its relations are definable by Boolean combinations of atomic formulas of the form $x<y$. The classification of the complexity of temporal CSPs has been known for a long time [9], but only recently Mottet proposed a new uniform algorithm based on universal algorithms [36]. Notice that most of the universal algorithms cannot be applied directly to infinite domain CSPs. For instance, in linear relaxations one introduces a variable for every element of the domain, which would result in a system with infinitely many variables.
One of the ways to overcome this difficulty is “sampling”: to solve an instance with $N$ variables one first restricts the domain to any $N$-element subset of $\mathbb{Q}$, then runs a universal algorithm on the restricted instance, and returns its answer. Mottet proved in [36] that every tractable temporal CSP can be solved this way using local consistency and $\mathrm{AIP}$ algorithms. The soundness of the algorithm follows from the fact that $\operatorname{PCSP}(\mathbb{A}|_{[N]},\mathbb{A})$ is solvable by the corresponding universal algorithms, where $\mathbb{A}|_{[N]}$ is the induced substructure obtained by restricting the domain to $\{1,2,\dots,N\}$.
In the paper we present an alternative proof of this result by providing concrete palette symmetric functions from $\operatorname{Pol}(\mathbb{A}|_{[N]},\mathbb{A})$, which can be used to obtain a (true) solution from the outputs of the algorithms $\mathrm{Singl}\mathrm{AC}$ and $\mathrm{CSingl}\mathrm{AIP}$. Thus, we prove the following slight strengthening of Mottet’s result, where local consistency is replaced by $\mathrm{CSingl}\mathrm{AC}$:
Theorem 1.3.
Let $\mathbb{A}$ be a temporal structure, $N\in\mathbb{N}$. Then one of the following holds:
  • $\operatorname{PCSP}(\mathbb{A}|_{[N]},\mathbb{A})$ is solvable by $\mathrm{CSingl}\mathrm{AC}\wedge\mathrm{Singl}\mathrm{AIP}$;
  • $\operatorname{CSP}(\mathbb{A})$ is NP-hard.
1.4 Structure of the paper
The paper is organized as follows. In Section 2 we give necessary definitions. Section 3 introduces classic universal algorithms and their combinations, defines palette symmetric functions, and formulates criteria for these algorithms to solve $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$. In Section 4 we discuss the limitations and weaknesses of the classic universal algorithms such as BLP, AIP, and BLP+AIP, and demonstrate why the singleton versions are more powerful. Additionally, we show that the template coming from the dihedral group $\mathbf{D}_{4}$ cannot be solved by the singleton version of $\mathrm{BLP}+\mathrm{AIP}$. In Section 5 we prove the characterization of the singleton versions of the algorithms in terms of existence of palette symmetric polymorphisms. Section 6 provides concrete palette symmetric polymorphisms for temporal relational structures and shows that all tractable temporal CSPs can be solved (after sampling) by $\mathrm{Singl}\mathrm{AIP}$ or $\mathrm{CSingl}\mathrm{AC}$. In Section 7 we prove the relational characterization of the dihedral group $\mathbf{D}_{4}$ and show that it does not admit palette symmetric polymorphisms. Finally, in Section 8 we prove that every finite template on a domain of size at most 7 that admits a WNU polymorphism also admits a palette symmetric polymorphism, and therefore can be solved by the singleton version of $\mathrm{BLP}+\mathrm{AIP}$.
2 Preliminaries
For $n\in\mathbb{N}$ we denote $n]=\{1,2,\dots,n\}$. We usually denote tuples $\mathbf{a}\in A^{n}$ in bold and refer to their components as $(\mathbf{a}^{1},\dots,\mathbf{a}^{n})$. For a permutation $\sigma\colon[n]\to[n]$, we denote by $\mathbf{a}^{\sigma}$ the tuple $(\mathbf{a}^{\sigma(1)},\dots,\mathbf{a}^{\sigma(n)})$.
A signature is a set of relational symbols $R_{1},\dots,R_{s}$, each with respective arity $\operatorname{ar}(R_{i})\in\mathbb{N}$. A $\sigma$-structure $\mathbb{A}$ consists of a domain $A$ and, for each $R_{i}\in\sigma$, a relation $R_{i}^{\mathbb{A}}\subseteq A^{\operatorname{ar}(R_{i})}$.
For two $\sigma$-structures $\mathbb{A}$ and $\mathbb{B}$, a homomorphism from $\mathbb{A}$ to $\mathbb{B}$ is a mapping $\phi\colon A\to B$ such that for each symbol $R\in\sigma$ and each tuple $\mathbf{a}=(a_{1},\dots,a_{\operatorname{ar}(R)})\in R^{\mathbb{A}}$, we have $\phi(\mathbf{a}):=(\phi(a_{1}),\dots,\phi(a_{\operatorname{ar}(R)}))\in R^{\mathbb{B}}$. If such a homomorphism exists, we write $\mathbb{A}\to\mathbb{B}$; otherwise, we write $\mathbb{A}\not\to\mathbb{B}$.
We say that a pair of finite structures $(\mathbb{A},\mathbb{B})$ is a PCSP template if they have the same signature and there exists a homomorphism $\mathbb{A}\to\mathbb{B}$.
Let $\mathcal{O}_{A,B}^{(n)}$ denote the set of all mappings $A^{n}\to B$, called $n$-ary functions, and let $\mathcal{O}_{A,B}=\bigcup\limits_{n=1}^{\infty}\mathcal{O}_{A,B}^{(n)}$. If $A=B$, we simply write $\mathcal{O}_{A}$ and $\mathcal{O}_{A}^{(n)}$, and we usually use the word operation instead of function.
To distinguish between equations and identities, we use the symbol $\approx$ to emphasize that the equality holds for all evaluations of the variables. For instance, an operation $f\in\mathcal{O}_{A}$ is called idempotent if $f(x,\dots,x)\approx x$, which means that $f(a,\dots,a)=a$ for all $a\in A$. A function $w$ is called a weak near unanimity (WNU) function if it satisfies the identity:
$w(y,x,\dots,x)\approx w(x,y,x,\dots,x)\approx\dots\approx w(x,\dots,x,y).$
A function $f\colon A^{n}\to B$ is called a polymorphism from a $\sigma$-structure $\mathbb{A}$ to a $\sigma$-structure $\mathbb{B}$ if, for each symbol $R\in\sigma$ and tuples $\mathbf{a}_{1},\dots,\mathbf{a}_{n}\in R^{\mathbb{A}}$, we have $f(\mathbf{a}_{1},\dots,\mathbf{a}_{n}):=(f(\mathbf{a}_{1}^{1},\dots,\mathbf{a}_{n}^{1}),\dots,f(\mathbf{a}_{1}^{\operatorname{ar}(R)},\dots,\mathbf{a}_{n}^{\operatorname{ar}(R)}))\in R^{\mathbb{B}}$. By $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ we denote the set of all polymorphisms from $\mathbb{A}$ to $\mathbb{B}$. When $\mathbb{A}=\mathbb{B}$, we write simply $\operatorname{Pol}(\mathbb{A})$.
A clone on a set $A$ is a set of operations $\mathcal{C}\subseteq\mathcal{O}_{A}$ such that:
  1. 1.
    $\mathcal{C}$ contains all projection operations $\pi_{i}^{n}(x_{1},\dots,x_{n})=x_{i}$ for all $n\geq 1$ and $i\in[n]$;
  2. 2.
    $\mathcal{C}$ is closed under composition, i.e., if $f\in\mathcal{C}$ is $k$-ary and $g_{1},\dots,g_{k}\in\mathcal{C}$ are $n$-ary, then the $n$-ary operation
    $h(x_{1},\dots,x_{n})=f\bigl(g_{1}(x_{1},\dots,x_{n}),\dots,g_{k}(x_{1},\dots,x_{n})\bigr)$
    also belongs to $\mathcal{C}$.
For a set of operations $F\subseteq\mathcal{O}_{A}$ by $\operatorname{Clo}(F)$ we denote the minimal clone containing $F$. For a set $N$, we denote by $\mathrm{id}_{N}\colon N\to N$ the identity map. For a relation $R\subseteq A^{n}$ and $S\subseteq[n]$ by $\operatorname{pr}_{S}(R)$ we denote the projection of $R$ onto the coordinates indexed by $S$.
3 Universal algorithms
3.1 Arc-consistency algorithm
The most basic algorithm for the CSP is to enforce some local consistency. Considering the lowest level of consistency we get the arc-consistency algorithm.
Suppose $\mathcal{I}=\langle X,D,\mathbf{C}\rangle$ is an instance of $\operatorname{FiniteCSP}$. The arc-consistency algorithm is implemented by the following pseudocode. By $\mbox{{ReduceDomain}}(\mathcal{I},x,B)$ we denote the function that returns a new instance obtained from $\mathcal{I}$ by reducing the domain of the variable $x$ to $B$ and removing the corresponding tuples from the constraints containing $x$.
1:function AC($\mathcal{I}$)
2:  repeat
3:   for each constraint $C=R(x_{1},\dots,x_{s})\in\mathcal{I}$ do
4:     for $i\in\{1,2,\dots,s\}$ do
5:      if $\operatorname{pr}_{i}(R)\neq D_{x_{i}}$ then
6:        $B:=\operatorname{pr}_{i}(R)$
7:        if $B=\varnothing$ then
8:         return “No”
9:        $\mathcal{I}:=\mbox{{ReduceDomain}}(\mathcal{I},x_{i},B)$
10:  until nothing changed
11:  return “Yes”
Observation 3.1.
The algorithm $\mathrm{AC}$ returns “Yes” on a CSP instance $\mathcal{I}$ if and only if there exists a mapping $D^{\prime}$ that assigns a nonempty set $D_{x}^{\prime}\subseteq D_{x}$ to every variable $x$ such that for any constraint $R(x_{1},\dots,x_{s})$ of $\mathcal{I}$ and any $j\in[s]$ we have $\operatorname{pr}_{j}(R\cap(D_{x_{1}}^{\prime}\times\dots\times D_{x_{s}}^{\prime}))=D_{x_{j}}^{\prime}$.
The power of the arc-consistency algorithm has been completely characterized for both $\operatorname{CSP}(\mathbb{A})$ and $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$. A function $f\colon A^{n}\to B$ is called totally symmetric if $f(a_{1},\dots,a_{n})=f(b_{1},\dots,b_{n})$ whenever $\{a_{1},\dots,a_{n}\}=\{b_{1},\dots,b_{n}\}$.
Theorem 3.1 (Theorem 7.4 in [3]).
Let $(\mathbb{A},\mathbb{B})$ be a PCSP template. Then $\mathrm{AC}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains totally symmetric functions of all arities.
One direction of this theorem can be proved as follows. Suppose $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains totally symmetric polymorphisms of all arities. First, we choose a sequence of totally symmetric polymorphisms $f_{1},f_{2},f_{n},\dots,$ such that each $f_{n}$ is of arity $n$, and identifying two variables in $f_{n}$ yields $f_{n-1}$. Then, for a set $S=\{a_{1},\dots,a_{s}\}\subseteq A$, we define $f(S):=f_{s}(a_{1},\dots,a_{s})$. If the algorithm $\mathrm{AC}$ returns “Yes” on a CSP instance $\mathcal{I}$, then it assigns a nonempty set $D_{x}^{\prime}$ to every variable $x$. We claim that a solution to $\mathcal{I}^{\mathbb{B}}$ can be defined by $s(x):=f(D_{x}^{\prime})$ for every variable $x$. Let $R(x_{1},\dots,x_{k})$ be a constraint $\mathcal{I}$. Consider $R^{\mathbb{A}}\cap(D_{x_{1}}^{\prime}\times\dots\times D_{x_{k}}^{\prime})=\{\mathbf{a}_{1},\dots,\mathbf{a}_{m}\}$. Since $f_{m}$ is a polymorphism, we have $f_{m}(\mathbf{a}_{1},\dots,\mathbf{a}_{m})\in R^{\mathbb{B}}$. By the definition of $f$, we obtain $f_{m}(\mathbf{a}_{1},\dots,\mathbf{a}_{m})=(s(x_{1}),\dots,s(x_{k}))$. Therefore, $(s(x_{1}),\dots,s(x_{k}))\in R^{\mathbb{B}}$. Thus, $s$ is a solution of $\mathcal{I}$ in $\mathbb{B}$, and $\mathrm{AC}$ solves the instance correctly.
3.2 Linear relaxations
The next algorithms arise from the idea of applying algorithms for solving linear equations (such as Gaussian elimination) to arbitrary CSPs over finite domains. Suppose $\mathcal{I}$ is a CSP instance given by the triple $\langle X,D,\mathbf{C}\rangle$. By $\mathcal{I}^{L}$ we denote the following system of linear equations (the meaning of + is not specified, as it is treated as a formal expression):
Variables: $C^{\mathbf{a}}$ for every constraint $C=R(x_{1},\dots,x_{m})$ in $\mathcal{I}$ and every $\mathbf{a}\in R$,
$x^{b}$ for every variable $x\in X$ and $b\in D_{x}$;
Equations: $\sum\limits_{\mathbf{a}\in R}C^{\mathbf{a}}=1$ for every constraint $C=R(x_{1},\dots,x_{m})$;
$\sum\limits_{\begin{subarray}{c}\mathbf{a}\in R\\ \mathbf{a}(i)=b\end{subarray}}C^{\mathbf{a}}=x_{i}^{b}$ for every constraint $C=R(x_{1},\dots,x_{m})$, every $i\in[m]$ and $b\in D_{x_{i}}$.
If we restrict all the variables to the set $\{0,1\}$ and interpret addition in $\mathbb{Z}$, then we obtain a one-to-one correspondence between solutions of $\mathcal{I}$ and solutions of $\mathcal{I}^{L}$ where $x=b\Leftrightarrow x^{b}=1$.
BLP Algorithm. The algorithm takes a CSP instance $\mathcal{I}$, transforms it into a system of linear equations $\mathcal{I}^{L}$, and solves it over $\mathbb{Q}\cap[0,1]$ using linear programming [32]. It returns “Yes” if a solution exists and “No” otherwise.
The power of the BLP algorithm has been completely characterized for both $\operatorname{CSP}(\mathbb{A})$ and $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$. A function $f\colon A^{n}\to B$ is called symmetric if $f(x_{1},\dots,x_{n})\approx f(x_{\sigma(1)},\dots,x_{\sigma(n)})$ for any permutation $\sigma\colon[n]\to[n]$.
Theorem 3.2 (Theorem 7.9 in [3]).
Let $(\mathbb{A},\mathbb{B})$ be a PCSP template. Then $\mathrm{BLP}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains symmetric functions of all arities.
AIP Algorithm. The algorithm takes a CSP instance $\mathcal{I}$, transforms it into a system of linear equations $\mathcal{I}^{L}$, and solves it over $\mathbb{Z}$ using linear programming [38]. It returns “Yes” if a solution exists and “No” otherwise.
The power of the AIP algorithm has been completely characterized for both $\operatorname{CSP}(\mathbb{A})$ and $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$. A function $f\colon A^{n}\to B$ is called alternating if $f(x_{1},\dots,x_{n})\approx f(x_{\sigma(1)},\dots,x_{\sigma(n)})$ for any permutation $\sigma\colon[n]\to[n]$ preserving parity, and $f(x_{1},\dots,x_{n-2},y,y)\approx f(x_{1},\dots,x_{n-2},z,z)$. Sometimes we will need a weaker version of it where the last identity is required only when the values $y$ and $z$ already appear among other variables. Formally, a function $f$ is called weakly alternating if $f(x_{1},\dots,x_{n})\approx f(x_{\sigma(1)},\dots,x_{\sigma(n)})$ for every parity-preserving permutation $\sigma\colon[n]\to[n]$, and $f(x_{1},\dots,x_{n-2},x_{i},x_{i})\approx f(x_{1},\dots,x_{n-2},x_{j},x_{j})$ for any $i,j\in[n-2]$.
Theorem 3.3 (Theorem 7.19 in [3]).
Let $(\mathbb{A},\mathbb{B})$ be a PCSP template. Then $\mathrm{AIP}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains alternating functions of all odd arities.
A solution of $\mathcal{I}^{L}$ in $\mathbb{Q}\cap[0,1]$ is called a BLP solution, and a solution of $\mathcal{I}^{L}$ in $\mathbb{Z}$ is called an AIP solution.
3.3 Combinations of the algorithms
We will use the following notations for combinations of algorithms. For universal algorithms $\mathfrak{A}$ and $\mathfrak{B}$ by $\mathfrak{A}\wedge\mathfrak{B}$ we denote the algorithm that runs both algorithms independently and return “Yes” if both return “Yes”.
[…]
Similarly, by $\mathrm{AC}$$\mathfrak{A}$ we denote the algorithm that first runs the arc-consistency algorithm, returns “No” if the latter returns “No”, and otherwise applies $\mathfrak{A}$ to the reduced instance and returns its answer.
The strength of the BLP+AIP algorithm has the following nice characterization for both $\operatorname{CSP}(\mathbb{A})$ and $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ in terms of symmetric polymorphisms. A function $f:A^{k_{1}+\dots+k_{n}}\to B$ is called $(k_{1},\dots,k_{n})$-block symmetric if for any permutations $\sigma_{1},\dots,\sigma_{n}$ on $k_{1}],\dots,[k_{n}]$ respectively, we have
$f(\mathbf{x}_{1},\dots,\mathbf{x}_{n})\approx f(\mathbf{x}_{1}^{\sigma_{1}},\dots,\mathbf{x}_{n}^{\sigma_{n}}).$
Theorem 3.4 (Theorem 5.1 in [14]).
Let $(\mathbb{A},\mathbb{B})$ be a PCSP template. Then the following conditions are equivalent:
  1. (1)
    $\mathrm{BLP}+\mathrm{AIP}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$;
  2. (2)
    for any $L\in\mathbb{N}$ there exist $s\in\mathbb{N}$ and numbers $\ell_{1},\dots,\ell_{s}$ greater than $L$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\ell_{1},\dots,\ell_{s})$-block symmetric function.
  3. (3)
    for any $L\in\mathbb{N}$ the set $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(L+1,L)$-block symmetric function.
[…]
  • (f1)
    $\mathcal{I}$ has a BLP solution $\Leftrightarrow$ there exists $Q\in\mathbb{N}$ such that $\mathcal{I}$ has a $Q$-solution.
  • (f2)
    The sum of a $d_{1}$-solution and a $d_{2}$-solution is a $(d_{1}+d_{2})$-solution.
  • (f3)
    $\mathrm{BLP}+\mathrm{AIP}$ returns “Yes” on an instance $\mathcal{I}$ $\Leftrightarrow$ there exists $L\in\mathbb{N}$ such that $\mathcal{I}$ has a $d$-solution for any $d\geq L$.
(f1) follows from the fact that multiplying a BLP solution by the least common multiple $Q$ of the denominators gives a $Q$-solution, and dividing a $d$-solution by $d$ gives a BLP solution. (f2) is immediate. Let us show (f3). If $\mathrm{BLP}+\mathrm{AIP}$ returns “Yes” on $\mathcal{I}$ then there exists a BLP solution $s_{BLP}$ of $\mathcal{I}^{L}$, an AIP solution $s_{AIP}$ of $\mathcal{I}^{L}$ such that $s_{BLP}(v)>0$ whenever $s_{AIP}(v)\neq 0$. Choose a sufficiently large integer $Q$ such that $Q\cdot s_{BLP}+s_{AIP}$ is a $(Q+1)$-solution (all values are nonnegative integers). Then $Q\cdot s_{BLP}$ is a $Q$-solution. By taking linear combinations $m\cdot(Q\cdot s_{BLP}+s_{AIP})+\ell(Q\cdot s_{BLP})$, we obtain a $d$-solution for any $d\geq Q\cdot(Q+1)$. To see the opposite direction of (f3) choose a $d_{0}$-solution, with $d_{0}\geq L$, having the maximal number of nonzero values among all the $d$-solutions (for all $d\in\mathbb{N}$). Such a solution exists by (f2). Considering the difference between a $(d_{0}+1)$-solution and the $d_{0}$-solution we obtain an AIP solution, while dividing the $d_{0}$-solution by $d_{0}$ gives a BLP solution. Together, these confirm that $\mathrm{BLP}+\mathrm{AIP}$ returns “Yes” on $\mathcal{I}$.
[…]
$(x_{1},x_{2}):=(f(\mathbf{a}_{x_{1},1},\dots,\mathbf{a}_{x_{1},s}),f(\mathbf{a}_{x_{2},1},\dots,\mathbf{a}_{x_{2},s}))=f(M_{1},\dots,M_{s})\in R^{\mathbb{A}},$
where the equality follows from the $(\ell_{1},\dots,\ell_{s})$-block symmetry of $f$, and membership follows from $f\in\operatorname{Pol}(\mathbb{A},\mathbb{B})$.
Similarly, other combinations of the above algorithms (see the top part of Table 1) can be characterized, but since the proofs do not differ substantially from those in [3, 14], we omit them here.
3.4 Singleton algorithms and palette functions
For a universal algorithm $\mathrm{Alg}$ by $\mathrm{SinglAlg}$ and $\mathrm{CSingl}\mbox{{Alg}}$ we denote the algorithms defined by the pseudocode below. Here, $\mbox{{ChangeConstraint}}(\mathcal{I},C,R)$ denotes the function that returns a new instance obtained from $\mathcal{I}$ by changing the constraint relation of the constraint $C$ to $R$. The difference between SinglAlg and CSinglAlg is that in $\mathrm{SinglAlg}$ we fix a variable to an element, whereas in $\mathrm{CSingl}\mbox{{Alg}}$ we fix a constraint to a tuple.
Remark 3.1.
It is not generally true that $\mathrm{SinglAlg}$ is weaker than $\mathrm{CSingl}\mbox{{Alg}}$, since the algorithm $\mathrm{SinglAlg}$, when fixing a variable $x$ to a value, may reduce all constraints containing $x$, while $\mathrm{CSingl}\mbox{{Alg}}$ always reduces only one constraint. Nevertheless, if $\mbox{{Alg}}\in\{\mathrm{AC},\mathrm{BLP},\mathrm{AC}+\mathfrak{A},\mathrm{BLP}+\mathfrak{A}\}$, then $\mathrm{SinglAlg}$ is clearly weaker than $\mathrm{CSingl}\mbox{{Alg}}$.
Open question 1.
Does there exist a PCSP template $(\mathbb{A},\mathbb{B})$ such that $\mathrm{Singl}\mathrm{AIP}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ but $\mathrm{CSingl}\mathrm{AIP}$ does not?
1:function SinglAlg($\mathcal{I}$)
2:  repeat
3:   for each variable $x$ of $\mathcal{I}$ do
4:     for $a\in D_{x}$ do
5:      if $\neg\mbox{{Alg}}(\mbox{{ReduceDomain}}(\mathcal{I},x,\{a\}$) then
6:        $B:=D_{x}\setminus\{a\}$
7:        if $B=\varnothing$ then
8:         return “No”
9:        $\mathcal{I}:=\mbox{{ReduceDomain}}(\mathcal{I},x,B)$ $\triangleright$ remove all the tuples $\triangleright$ such that $x\notin B$
10:  until nothing changed
11:  return “Yes”
1:function CSinglAlg($\mathcal{I}$)
2:  repeat
3:   for each constraint $C=R(x_{1},\dots,x_{s})\in\mathcal{I}$ do
4:     for $(a_{1},\dots,a_{s})\in R$ do
5:      if $\neg\mbox{{Alg}}(\mbox{{ChangeConstraint}}(\mathcal{I},C,\{(a_{1},\dots,a_{s})\})$ then $\triangleright$ reduce $R$ to the one-tuple relation
6:        $R^{\prime}:=R\setminus\{(a_{1},\dots,a_{s})\}$
7:        if $R^{\prime}=\varnothing$ then
8:         return “No”
9:        $\mathcal{I}:=\mbox{{ChangeConstraint}}(\mathcal{I},C,R^{\prime})$ $\triangleright$ remove the tuple $\triangleright$ from the constraint relation of $C$
10:  until nothing changed
11:  return “Yes”
For universal algorithms $\mathfrak{A}$ and $\mathfrak{B}$ we denote by $\mathrm{Singl}\mathfrak{A}+\mathfrak{B}$ the algorithm that first runs $\mathrm{Singl}\mathfrak{A}$ and returns “No” if it returns “No”; otherwise, it returns the answer of $\mathfrak{B}$ on the reduced instance $\mathcal{I}$ (see the pseudocode of SinglAlg). Similarly, by $\mathrm{CSingl}\mathfrak{A}+\mathfrak{B}$ we denote the algorithm that first runs $\mathrm{CSingl}\mathfrak{A}$ and returns “No” if it returns “No”; otherwise, it returns the answer of $\mathfrak{B}$ on the reduced instance $\mathcal{I}$ (see the pseudocode of CSinglAlg).
To characterize the power of singleton algorithms we need a generalization of symmetric functions. We call to a function $f:A^{k_{1}+\dots+k_{n}}\to B$ a $(k_{1},\dots,k_{n})$-block function to emphasize that its arguments are divided into blocks: the first $k_{1}$ arguments form the first block, the next $k_{2}$ arguments form the second block, and so on. We say that the $i$-th block of a $(k_{1},\dots,k_{n})$-block function is (totally) symmetric/(weakly) alternating if for any $\mathbf{a}_{1}\in A^{k_{1}},\dots,\mathbf{a}_{i-1}\in A^{k_{i-1}},\mathbf{a}_{i+1}\in A^{k_{i+1}},\dots,\mathbf{a}_{n}\in A^{k_{n}}$ the function $h(\mathbf{x})=f(\mathbf{a}_{1},\dots,\mathbf{a}_{i-1},\mathbf{x},\mathbf{a}_{i+1},\dots,\mathbf{a}_{n})$ is (totally) symmetric/(weakly) alternating, respectively. A function is $(k_{1},\dots,k_{n})$-block (totally) symmetric if all of its blocks are (totally) symmetric, and $(k_{1},\dots,k_{n})$-block (weakly) alternating if all of its blocks are (weakly) alternating.
We can divide a tuple $\mathbf{a}\in A^{k_{1}+\dots+k_{n}}$ into blocks and refer to each block by its position. Let $\mathcal{P}$ be a set of disjoint subsets of $[n]$. A tuple $\mathbf{a}=(\mathbf{a}_{1},\dots,\mathbf{a}_{n})\in A^{k_{1}+\dots+k_{n}}$ is called $\mathcal{P}$-$(k_{1},\dots,k_{n})$-palette if for every element $c\in A$ appearing in $\mathbf{a}$ there exists $L\in\mathcal{P}$ such that $\mathbf{a}_{\ell}=(c,c,\dots,c)$ for every $\ell\in L$. We usually use an overline to specify the set $\mathcal{P}$. For example, $(\ell,\overline{k_{1}},\dots,\overline{k_{n}})$ corresponds to $\mathcal{P}=\{\{2\},\{3\},\dots,\{n+1\}\}$ and $(\overline{k_{1},m_{1}},\dots,\overline{k_{n},m_{n}})$ corresponds to $\mathcal{P}=\{\{1,2\},\{3,4\},\dots,\{2n-1,2n\}\}$.
Similarly to block functions, we define palette functions, where we require the block properties only on palette tuples. Formally, a function $f\colon A^{k_{1}+\dots+k_{n}}\to A$ is called $\mathcal{P}$-$(k_{1},\dots,k_{n})$-palette function, and the $i$-th block of it is (totally) symmetric/(weakly) alternating if (1) and (2) imply (3):
  • (1)
    $\mathbf{a}$ and $\mathbf{b}$ are $\mathcal{P}$-$(k_{1},\dots,k_{n})$-palette tuples;
  • (2)
    $g(\mathbf{a})=g(\mathbf{b})$ for any $(k_{1},\dots,k_{n})$-block function $g\colon A^{k_{1}+\dots+k_{n}}\to A$ whose $i$-th block is (totally) symmetric/(weakly) alternating;
  • (3)
    $f(\mathbf{a})=f(\mathbf{b})$.
A function is $\mathcal{P}$-$(k_{1},\dots,k_{n})$-palette (totally) symmetric/(weakly) alternating if it is a $\mathcal{P}$-$(k_{1},\dots,k_{n})$-palette function whose blocks are (totally) symmetric/(weakly) alternating.
The power of singleton algorithms and their combinations can be characterized using palette functions. Below we present a characterization for the algorithms $\mathrm{Singl}\mathrm{AC}$ and $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$, which are proved in Section 5.3 and 5.4, respectively.
Theorem 3.5.
Let $(\mathbb{A},\mathbb{B})$ be a PCSP template. The following conditions are equivalent:
  1. (1)
    $\mathrm{SinglArcCons}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$;
  2. (2)
    $\mathrm{CSingl}\mathrm{AC}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$;
  3. (3)
    $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ has an $(\overline{m_{1}},\dots,\overline{m_{n}})$-palette totally symmetric function for every $n,m_{1},\dots,m_{n}\in\mathbb{N}$;
  4. (4)
    for every $N\in\mathbb{N}$ there exist $n,m_{1},\dots,m_{n}\geq N$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ has an $(\overline{m_{1}},\dots,\overline{m_{n}})$-palette totally symmetric function.
Theorem 3.6.
Let $(\mathbb{A},\mathbb{B})$ be a PCSP template. The following conditions are equivalent:
  1. (1)
    $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$;
  2. (2)
    $\mathrm{CSingl}(\mathrm{BLP}+\mathrm{AIP})$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$;
  3. (3)
    $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette symmetric function for every $\ell,n\in\mathbb{N}$;
  4. (4)
    for every $N\in\mathbb{N}$ there exist $n\geq N$, $k_{1},\dots,k_{n}\in\mathbb{N}$, and $m_{1,1},\dots,m_{1,k_{1}},\dots,m_{n,1},\dots,m_{n,k_{n}}\geq N$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\overline{m_{1,1},\dots,m_{1,k_{1}}},\overline{m_{2,1},\dots,m_{2,k_{2}}},\dots,\overline{m_{n,1},\dots,m_{n,k_{n}}})$-palette symmetric function.
To explain the idea behind palette symmetric functions and Theorem 3.6, let us prove the following claim.
Claim 3.1.
Suppose $(\mathbb{A},\mathbb{B})$ is a PCSP template, and for every $L\in\mathbb{N}$ there exist $n,\ell_{1},\dots,\ell_{n}\geq L$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette symmetric function. Then $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$.
Proof.
It is sufficient to prove that if $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ returns “Yes” on an instance $\mathcal{I}$, then $\mathcal{I}$ has a solution. For every variable $x$ and every element $a$ from the obtained domain of $x$ let $\mathcal{I}_{x}^{a}$ be the instance obtained in the algorithm SinglAlg by reducing the domain of $x$ to $\{a\}$. Let $\Omega$ be the set of all such instances. Since $\mathrm{BLP}+\mathrm{AIP}$ returns “Yes” on each of these instances, the fact (f3) from Section 3.3 implies that there exists a large number $L$ such that $L>|\Omega|$ and each instance $\mathcal{I}_{i}^{a}$ has a $d$-solution for any $d\geq L$. Choose an $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette symmetric function $f\in\operatorname{Pol}(\mathbb{A},\mathbb{B})$, where $n,\ell_{1},\dots,\ell_{n}\geq L$.
Similar to the construction in Section 3.3, for each block (say the $j$-th block) we choose an instance $\mathcal{I}_{x}^{a}$ and an $\ell_{j}$-solution $s_{j}$ of this instance. We only require every instance from $\Omega$ to be chosen for at least one block, which is possible since $n\geq L>|\Omega|$. For every variable $x$ of $\mathcal{I}$, let $\mathbf{a}_{x,j}$ be a tuple of length $\ell_{j}$ such that each element $c$ appears in this tuple exactly $s_{j}(x^{c})$ times. We claim that a (true) solution of $\mathcal{I}$ can be defined by $x:=f(\mathbf{a}_{x,1},\dots,\mathbf{a}_{x,s})$ for every variable $x$. To follow the argument from Section 3.3, it is sufficient to check that the tuple $(\mathbf{a}_{x,1},\dots,\mathbf{a}_{x,s})$ is $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette, which follows immediately from the definition of the instances in $\Omega$. ∎
As mentioned in the introduction, the algorithm $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ is very powerful and solves all tractable $\operatorname{CSP}(\mathbb{A})$ on small domains, as follows from the following algebraic result proved in Section 8.
Theorem 3.7.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$ are similar algebras admitting WNU term operations, $|A_{i}|<8$ for every $i\in[s]$. Then $\mathbf{A}_{1}\times\dots\times\mathbf{A}_{s}$ has a $(\underbrace{p^{m},\dots,p^{m}}_{n})$-palette symmetric term operation for every prime $p>8$ and every $m,n\in\mathbb{N}$.
All necessary definitions are given in Section 8. An important corollary for our purposes is the following theorem.
Theorem 1.2.
Suppose $\mathbb{A}$ is a finite relational structure having a WNU polymorphism such that $|\operatorname{pr}_{i}(R^{\mathbb{A}})|<8$ for every $R$ and $i\in[\operatorname{ar}(R)]$. Then $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ solves $\operatorname{CSP}(\mathbb{A})$.
An alternative formulation is to introduce the notion of a multi-sorted constraint language and restrict the domains by 7 instead of restricting the projections, but we chose the current formulation to avoid additional notions that are not essential in this paper.
Most other combinations of the algorithms can be characterized similarly (see Table 1). Since their proofs are almost identical to those for $\mathrm{Singl}\mathrm{AC}$ and $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$, we omit them. We can also characterize the algorithm $\mathrm{Singl}\mathfrak{A}\wedge\mathfrak{B}$ for $\mathfrak{A},\mathfrak{B}\in\{\mathrm{AC},\mathrm{BLP},\mathrm{AIP}\}$, but this requires a different notion and is less interesting, since the algorithms are simply run independently. For example, $\mathrm{Singl}\mathrm{BLP}\wedge\mathrm{AIP}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if for all $m,\ell,n\in\mathbb{N}$, $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains a $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n},\ell)$-block function such that the first $n$ blocks are symmetric on $(\underbrace{\overline{m},\overline{m},\dots,\overline{m}}_{n},\ell)$-palette tuples and the last block is always alternating.
4 Limitations of linear programming
In this section, we discuss the limitations of linear programming and provide concrete templates that illustrate its weaknesses. The results are collected in Table 2, while templates are defined below.
Table 2: Examples of constraint languages
Constraint language Strongest not working Weakest working
Horn-SAT $\mathrm{AIP}$ $\mathrm{AC}$
2-SAT $\mathrm{BLP}$, $\mathrm{Singl}\mathrm{AIP}$ $\mathrm{Singl}\mathrm{AC}$, $\mathrm{BLP}+\mathrm{AIP}$
${3}\text{-}\mathrm{LIN}(p)$ $\mathrm{Singl}\mathrm{BLP}$ $\mathrm{AIP}$
$\left(\begin{smallmatrix}0&1&2&3&4\\ 1&0&3&4&2\end{smallmatrix}\right)$ $\mathrm{BLP}+\mathrm{AIP}$ $\mathrm{Singl}\mathrm{AC}$, $\mathrm{Singl}\mathrm{AIP}$
$\mathbb{Z}_{2}\cup\mathbb{Z}_{3}$ $\mathrm{Singl}\mathrm{BLP}+\mathrm{AIP}$ $\mathrm{Singl}\mathrm{AIP}$
$\mathbb{Z}_{2}\leftarrow 2$ $\mathrm{Singl}\mathrm{BLP}+\mathrm{AIP}$ $\mathrm{Singl}\mathrm{AIP}$
$\mathbb{D}_{4}^{idemp}$ $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ ??? [^1] We already know that $\operatorname{CSP}(\mathbb{D}_{4}^{idemp})$ can be solved by the second level of $\mathrm{AIP}$; however higher-level algorithms are outside the scope of this paper.
The first simple idea demonstrating the weakness of the $\mathrm{BLP}$ algorithm can be formulated as follows.
Claim 4.1.
Suppose $\mathbb{A}$ is a finite relational structure, for every relation $R^{\mathbb{A}}$ of an arity $m$, every $b,c\in A$, and every $i\in[m]$, we have $|\{\mathbf{a}\in R^{\mathbb{A}}\mid\mathbf{a}^{i}=b\}|=|\{\mathbf{a}\in R^{\mathbb{A}}\mid\mathbf{a}^{i}=c\}|$. Then $\mathrm{BLP}$ answers “Yes” on any instance of $\operatorname{CSP}(\mathbb{A})$.
Proof.
The uniform distribution is always a solution to $\mathcal{I}^{L}$. Specifically, it is sufficient to set $C^{\mathbf{a}}=1/|R^{\mathbb{A}}|$ for every constraint $C=R(x_{1},\dots,x_{m})$ and $\mathbf{a}\in R$, and $x^{b}=1/|A|$ for every variable $x$ and $b\in A$. ∎
Thus, $\mathrm{BLP}$ returns “Yes” whenever the uniform distribution is feasible, making it useless for solving linear equations over finite fields.
The next claim shows the main weakness of $\mathrm{AIP}$: it produces the same result even if we replace each constraint relation by its parallelogram closure. We say that a relation $R\subseteq A^{m}$ satisfies the parallelogram property if for any $I\subseteq[m]$, $\mathbf{a},\mathbf{b},\mathbf{c}\in R$, and $\mathbf{d}\in A^{m}$, we have
$\left(\operatorname{pr}_{I}(\mathbf{a})=\operatorname{pr}_{I}(\mathbf{b})\wedge\operatorname{pr}_{I}(\mathbf{c})=\operatorname{pr}_{I}(\mathbf{d})\wedge\operatorname{pr}_{[m]\setminus I}(\mathbf{a})=\operatorname{pr}_{[m]\setminus I}(\mathbf{c})\wedge\operatorname{pr}_{[m]\setminus I}(\mathbf{b})=\operatorname{pr}_{[m]\setminus I}(\mathbf{d})\right)\Longrightarrow\mathbf{d}\in R.$
For binary relations $R$ this simply means that $(b_{1},b_{2}),(b_{1},a_{2}),(a_{1},b_{2})\in R$ implies $(a_{1},a_{2})\in R$. The parallelogram closure of $R$ is the minimal relation $R^{\prime}\supseteq R$ having the parallelogram property.
Claim 4.2.
Suppose $\mathcal{I}$ is an instance of $\operatorname{FiniteCSP}$ containing a constraint $C=R(x_{1},\dots,x_{m})$, and
  • $(a_{1},\dots,a_{k},b_{k+1},\dots,b_{m}),(b_{1},\dots,b_{k},a_{k+1},\dots,a_{m}),(b_{1},\dots,b_{k},b_{k+1},\dots,b_{m})\in R$,
  • $R^{\prime}=R\cup\{(a_{1},\dots,a_{k},a_{k+1},\dots,a_{m})\}$,
  • $\mathcal{I}^{\prime}$ is obtained from $\mathcal{I}$ by replacing $R$ in $C$ by $R^{\prime}$.
Then $\mathrm{AIP}$ returns “Yes” on $\mathcal{I}$ if and only if $\mathrm{AIP}$ returns “Yes” on $\mathcal{I}^{\prime}$.
Proof.
It is sufficient to construct an AIP solution for $\mathcal{I}$ from an AIP solution for $\mathcal{I}^{\prime}$. To do so, remove the variable $C^{(a_{1},\dots,a_{k},a_{k+1},\dots,a_{m})}$ and adjust the values of only 3 variables as follows:
$C^{(a_{1},\dots,a_{k},b_{k+1},\dots,b_{m})}:=C^{(a_{1},\dots,a_{k},b_{k+1},\dots,b_{m})}+C^{(a_{1},\dots,a_{k},a_{k+1},\dots,a_{m})}$
$C^{(b_{1},\dots,b_{k},a_{k+1},\dots,a_{m})}:=C^{(b_{1},\dots,b_{k},a_{k+1},\dots,a_{m})}+C^{(a_{1},\dots,a_{k},a_{k+1},\dots,a_{m})}$
$C^{(b_{1},\dots,b_{k},b_{k+1},\dots,b_{m})}:=C^{(b_{1},\dots,b_{k},b_{k+1},\dots,b_{m})}-C^{(a_{1},\dots,a_{k},a_{k+1},\dots,a_{m})}$
Therefore, $\mathrm{AIP}$ generally cannot solve CSPs involving relations without the parallelogram property (see Sections 4.1 and 4.2). Another weakness of $\mathrm{AIP}$ is its inability to solve linear equations modulo 2 and modulo 3 simultaneously, as shown in Section 4.3.
4.1 Classic templates
Horn-3-SAT $(\{0,1\};\{0,1\}^{3}\setminus\{(1,1,0)\},\{0\},\{1\})$. Horn-3-SAT can be solved by the arc-consistency algorithm, as witnessed by the fact that the $n$-ary conjunction, which is a totally symmetric function, belongs to $\operatorname{Pol}(\text{Horn-3-SAT})$ for every $n\geq 2$. On the other hand, $\mathrm{AIP}$ cannot solve Horn-3-SAT. Indeed, the minimal relation with the parallelogram property containing $\{0,1\}^{3}\setminus\{(1,1,0)\}$ is the full relation $\{0,1\}^{3}$. By Claim 4.2, running $\mathrm{AIP}$ on Horn-3-SAT is equivalent to running $\mathrm{AIP}$ on $(\{0,1\};\{0,1\}^{3},\{0\},\{1\})$, which can only return “No” on trivial instances.
[…]
$\mathbf{n}\text{{-LIN}}\mathbf{(p)}=\{\mathbb{Z}_{p};\{a_{1}x_{1}+\dots+a_{n}x_{n}=a_{0}\mid a_{0},a_{1},\dots,a_{n}\in\mathbb{Z}_{p}\}$ consists of linear equations with $n$ variables modulo $p$. $\mathrm{AIP}$ solves it as $\operatorname{Pol}({n}\text{-}\mathrm{LIN}(p))$ includes an alternating function $x_{1}-x_{2}+x_{3}-\dots-x_{2n}+x_{2n+1}$. To see that $\mathrm{Singl}\mathrm{BLP}$ does not solve ${3}\text{-}\mathrm{LIN}(p)$, consider an instance consisting of two contradictory equations: $x_{1}+x_{2}+x_{3}=0\wedge x_{1}+x_{2}+x_{3}=1$. Fixing a variable to a concrete value and assigning a uniform distribution to the rest (as in Claim 4.1) yields a BLP solution, so $\mathrm{Singl}\mathrm{BLP}$ incorrectly returns “Yes”.
4.2 Disconnected templates
$\left(\begin{smallmatrix}\mathbf{0}&\mathbf{1}&\mathbf{2}&\mathbf{3}&\mathbf{4}\\ \mathbf{1}&\mathbf{0}&\mathbf{3}&\mathbf{4}&\mathbf{2}\end{smallmatrix}\right)=(\{0,1,2,3,4\};\{(0,1),(1,0),(2,3),(3,4),(4,2)\})$, which is a structure containing a single binary relation given by the graph of a bijection. To show that both $\mathrm{Singl}\mathrm{AC}$ and $\mathrm{Singl}\mathrm{AIP}$ solve it, let us define $f\in\operatorname{Pol}\left(\begin{smallmatrix}0&1&2&3&4\\ 1&0&3&4&2\end{smallmatrix}\right)$ that is simultaneously $(\underbrace{\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n})$-palette totally symmetric and palette alternating. Specifically, let $f(\mathbf{x}_{1},\dots,\mathbf{x}_{n})$ return the value from the first constant block if it exists, and $\mathbf{x}_{1}^{1}$ otherwise.
Denote the only relation in this structure by $\rho$, and consider the instance $\mathcal{I}=\rho(x_{1},x_{2})\wedge\rho(x_{2},x_{3})\wedge\rho(x_{3},x_{1})\wedge\rho(x_{1},x_{4})\wedge\rho(x_{4},x_{1})$, which has no solution. By Claim 4.1, $\mathcal{I}$ nevertheless has a BLP solution in which all the tuples receive positive assignments. To show that $\mathrm{BLP}+\mathrm{AIP}$ also returns “Yes” on $\mathcal{I}$, it is sufficient to construct an AIP solution. One such solution assigns value $1$ to the tuples $(2,3),(3,4),(4,2)$ in each constraint, and value 1-1 to the tuples $(0,1),(1,0)$ in each constraint. This example illustrates that $\mathrm{AIP}$ is powerless when the instance is disconnected (here the components $\{0,1\}$ and $\{2,3,4\}$ are disconnected). This weakness becomes even more evident in the next template.
$\bm{\mathbf{Z}_{2}\cup\mathbf{Z}_{3}}=(\{0,1,0^{\prime},1^{\prime},2^{\prime}\};l_{0}^{\mathbb{Z}_{2}},l_{1}^{\mathbb{Z}_{2}},l_{0}^{\mathbb{Z}_{3}},l_{1}^{\mathbb{Z}_{3}},l_{2}^{\mathbb{Z}_{3}})$, where the relations are defined by:
$l_{a}^{\mathbb{Z}_{2}}(x_{1},x_{2},x_{3})=$ $(x_{1},x_{2},x_{3}\in\{0,1\}\wedge x_{1}+x_{2}+x_{3}=a\pmod{2})\vee x_{1},x_{2},x_{3}\in\{0^{\prime},1^{\prime},2^{\prime}\},$
$l_{b}^{\mathbb{Z}_{3}}(x_{1},x_{2},x_{3})=$ $(x_{1},x_{2},x_{3}\in\{0^{\prime},1^{\prime},2^{\prime}\}\wedge x_{1}+x_{2}+x_{3}=b\pmod{3})\vee x_{1},x_{2},x_{3}\in\mathbb{\{}0,1\}.$
Thus, we have linear equations modulo 2 on $\{0,1\}$ and independent linear equations modulo 3 on $\{0^{\prime},1^{\prime},2^{\prime}\}$. Any instance of $\operatorname{CSP}(\mathbb{Z}_{2}\cup\mathbb{Z}_{3})$ decomposes into two independent systems, one over $\mathbb{Z}_{2}$ and one over $\mathbb{Z}_{3}$. Solving the instance amounts to solving each system separately and returning “Yes” if one of them has a solution. Nevertheless, $\mathrm{AIP}$ fails here: it returns “Yes” on any instance of $\operatorname{CSP}(\mathbb{Z}_{2}\cup\mathbb{Z}_{3})$. The key idea comes from the following claim (the definition of a $d$-solution was given in Section 3.3).
Claim 4.3.
Suppose $\mathcal{I}$ is an instance of ${n}\text{-}\mathrm{LIN}(p)$ for a prime $p$ and $n\in\mathbb{N}$. Then $\mathcal{I}$ has a $p^{N}$-solution for every $N\geq n-1$.
Proof.
It is sufficient to set $C^{\mathbf{a}}=p^{n}/|R^{\mathbb{A}}|$ for every constraint $C=R(x_{1},\dots,x_{m})$ and every tuple $\mathbf{a}\in R$, and $x^{b}=p^{n-1}$ for every variable $x$ and every $b\in\mathbb{Z}_{p}$. ∎
Applying Claim 4.3 to the restriction of an instance of $\operatorname{CSP}(\mathbb{Z}_{2}\cup\mathbb{Z}_{3})$ to $\{0,1\}$ and to $\{0^{\prime},1^{\prime},2^{\prime}\}$, we obtain:
  • every instance has a $2^{n}$-solution for any $n\geq 2$;
  • every instance has a $3^{n}$-solution for any $n\geq 2$;
  • a linear combination of a $2^{n}$-solution and a $3^{n}$-solution yields an AIP solution.
Thus, $\mathrm{AIP}$ returns “Yes” on all instances of $\operatorname{CSP}(\mathbb{Z}_{2}\cup\mathbb{Z}_{3})$. Even $\mathrm{Singl}\mathrm{BLP}+\mathrm{AIP}$ fails, since $\mathrm{Singl}\mathrm{BLP}$ cannot reduce a system of linear equations (with more than 2 variables); therefore, running $\mathrm{Singl}\mathrm{BLP}+\mathrm{AIP}$ is not substantially different from running $\mathrm{AIP}$ alone.
In contrast, $\mathrm{Singl}\mathrm{AIP}$ solves the problem. Fixing a concrete element in the $\mathrm{Singl}$-algorithm restricts the instance to either the $\{0,1\}$-part or the $\{0^{\prime},1^{\prime},2^{\prime}\}$-part, and each part can be solved by $\mathrm{AIP}$. This is also witnessed by a $(\underbrace{\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n})$-palette alternating function $f(\mathbf{x}_{1},\dots,\mathbf{x}_{n})$ defined as follows:
  • choose the first block $\mathbf{x}_{i}$ consisting entirely of elements from either $\{0,1\}$ or $\{0^{\prime},1^{\prime},2^{\prime}\}$;
  • if no such block exists, return $\mathbf{x}_{1}^{1}$;
  • otherwise, apply the alternating function to $\mathbf{x}_{i}$, that is, $f(\mathbf{x}_{1},\dots,\mathbf{x}_{n})=\mathbf{x}_{i}^{1}-\mathbf{x}_{i}^{2}+\mathbf{x}_{i}^{3}-\dots-\mathbf{x}_{i}^{2\ell}+\mathbf{x}_{i}^{2\ell+1}$, where ++ and - are interpreted modulo 2 if $\mathbf{x}_{i}\in\{0,1\}^{2\ell+1}$, and modulo 3 if $\mathbf{x}_{i}\in\{0^{\prime},1^{\prime},2^{\prime}\}^{2\ell+1}$.
Although this template is extremely simple, it serves as a counterexample to several algorithms that were conjectured to solve all tractable CSPs, including $\mathbb{Z}$-affine $k$-consistency, BLP+AIP, $\mathrm{BA}^{k}$, and CLAP (see [34] for details). Since hierarchies and higher levels of consistency are beyond the scope of this paper, we omit the proofs.
4.3 Implications of linear equations
The next template has a similar flavor to the previous ones, but its instances are more interesting because they combine implications with linear equations.
$\bm{\mathbf{Z}_{2}\leftarrow 2}=(\{0,1,2\};\mathcal{R}_{L}\cup\{\{0,1\},S_{0,1}^{=}\})$, where $S_{0,1}^{=}(x,y,z)=(x=y\vee z\in\{0,1\})$, and
$\mathcal{R}_{L}=\left\{(x_{1}+\dots+x_{n}=c\wedge x_{1},\dots,x_{n}\in\{0,1\})\vee(x_{1}=\dots=x_{n}=2)\mid c\in\{0,1\},n\in\{3,4\}\right\}.$
Thus, $S_{0,1}^{=}(x,y,z)$ can be read as an implication $z=2\rightarrow x=y$, and $\mathcal{R}_{L}$ consists of linear equations over $\{0,1\}$, augmented by the additional all-2 tuple $(2,2,\dots,2)$. The polymorphisms of this structure are well-understood, as it corresponds to one of the minimal Taylor clones [29]. Define the ternary operation:
[…]
Claim 4.4 ([29]).
$\operatorname{Pol}(\mathbb{Z}_{2}\leftarrow 2)=\operatorname{Clo}(f)$.
Below we define an instance $\mathcal{I}$ of $\operatorname{CSP}(\mathbb{Z}_{2}\leftarrow 2)$ that fools $\mathrm{Singl}\mathrm{BLP}+\mathrm{AIP}$. Here, each linear equation should be understood as the corresponding relation from $\mathcal{R}_{L}$, that is, a linear equation modulo 2 augmented by the tuple $(2,2,\dots,2)$.
$\begin{cases}x_{1}\in\{0,1\}\\ \hline\cr x_{1}+x_{2}+x_{3}+x_{4}=0\\ x_{3}+x_{5}+x_{6}=0\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;(L_{1})\\ x_{4}+x_{5}+x_{6}=1\\ \hline\cr x_{1}=x_{2}\vee x_{7}\in\{0,1\}\\ \hline\cr x_{7}+x_{8}+x_{9}+x_{10}=0\\ x_{9}+x_{11}+x_{12}=0\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;(L_{2})\\ x_{10}+x_{11}+x_{12}=1\\ \hline\cr x_{7}=x_{8}\vee x_{13}\in\{0,1\}\\ \hline\cr x_{13}+x_{14}+x_{15}+x_{16}=0\\ x_{13}+x_{14}+x_{17}+x_{18}=0\;\;\;\;\;\;\;\;\;\;\;\;(L_{3})\\ x_{15}+x_{16}+x_{17}+x_{18}=1\end{cases}$
We have three systems of linear equations $L_{1}$, $L_{2}$, and $L_{3}$, with implications connecting them. Each system admits the trivial fake solution where all variables are set to 2. In addition, $L_{1}$ and $L_{2}$ have nontrivial solutions on $\{0,1\}$ but $L_{3}$ does not. A natural algorithm proceeds as follows. Solve $L_{3}$ first. Since it has no solutions on $\{0,1\}$, the variable $x_{13}$ must be equal to 2. Then $x_{7}=x_{8}$, which can be added as an equation to $L_{2}$. We check that $L_{2}$ has no solution satisfying this additional equation, which implies that $x_{7}$ must be equal to 2 and $x_{1}=x_{2}$. Finally, the linear system $L_{1}$ together with $x_{1}=x_{2}$ has no solution, contradicting $x_{1}\in\{0,1\}$.
However, $\mathrm{Singl}\mathrm{BLP}+\mathrm{AIP}$ incorrectly returns “Yes” on the instance $\mathcal{I}$. $\mathrm{Singl}\mathrm{BLP}$ does not reduce any domains (except the trivial restriction $D_{x_{1}}=\{0,1\}$), since every linear equation can be satisfied by a uniform distribution. By Claim 4.2, each constraint relation can be replaced by its parallelogram closure without changing the outcome of AIP. Since the parallelogram closure of $S_{0,1}^{=}$ is the full relation, $\mathrm{AIP}$ simply ignores all $S_{0,1}^{=}$-constraints. Taking a genuine solution of $L_{1}$ over $\{0,1\}$, together with fake all-2 solutions for $L_{2}$ and $L_{3}$, yields a global a solution that violates only $S_{0,1}^{=}$-constraints, which AIP ignores. Thus $\mathrm{AIP}$ (and hence $\mathrm{Singl}\mathrm{BLP}+\mathrm{AIP}$) returns “Yes” on the unsatisfiable instance $\mathcal{I}$.
In contrast, $\mathrm{Singl}\mathrm{AIP}$ returns “No” on $\mathcal{I}$. After fixing $x_{13}$ to 0 and 11 the $\mathrm{AIP}$ algorithm returns “No”, which excludes these values from the domain of $x_{13}$ and forces $x_{13}=2$. Hence, $x_{7}=x_{8}$. After fixing $x_{7}$ to 0 or 1, $\mathrm{AIP}$ again returns “No”, reducing the domain of $x_{7}$ to $\{2\}$ and forcing $x_{1}=x_{2}$. With $x_{1}=x_{2}$, $\mathrm{AIP}$ returns “No” on the reduced instance. Thus, $\mathrm{Singl}\mathrm{AIP}$ outputs “No” on $\mathcal{I}$, as desired.
To witness that $\mathrm{Singl}\mathrm{AIP}$ correctly solves $\operatorname{CSP}(\mathbb{Z}_{2}\leftarrow 2)$ we define a $(\underbrace{\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n})$-palette alternating function $f(\mathbf{x}_{1},\dots,\mathbf{x}_{n})$ from $\operatorname{Pol}(\mathbb{Z}_{2}\leftarrow 2)$ as follows:
  • choose the first block $\mathbf{x}_{i}$ consisting entirely of elements from $\{0,1\}$;
  • if no such block exists, return $\mathbf{x}_{1}^{1}$;
  • otherwise, apply the alternating function to $\mathbf{x}_{i}$, that is, $f(\mathbf{x}_{1},\dots,\mathbf{x}_{n})=\mathbf{x}_{i}^{1}+\mathbf{x}_{i}^{2}+\dots+\mathbf{x}_{i}^{2\ell+1}\pmod{2}$.
The crucial difference between $\mathbb{Z}_{2}\leftarrow 2$ and $\mathbb{Z}_{2}\cup\mathbb{Z}_{3}$ is that $\mathbb{Z}_{2}\cup\mathbb{Z}_{3}$ can be solved by running the $\mathrm{AIP}$ algorithm once for each variable and each value, while for $\mathbb{Z}_{2}\leftarrow 2$ we must ran $\mathrm{AIP}$ again and again, gradually reducing the instance. We believe that $\mathbb{Z}_{2}\leftarrow 2$ also cannot be solved by $\mathbb{Z}$-affine $k$-consistency or $\mathrm{BA}^{k}$, but we do not prove this formally.
4.4 Dihedral group
The last and perhaps most important example comes from the dihedral group $\mathbf{D}_{4}$, the group of symmetries of a square. This group is generated by two elements, $r$ and $s$, and is completely defined by the equalities $r^{4}=s^{2}=1$ and $rs=sr^{3}$.
We use the following representation of the dihedral group: $\mathbf{A}=(\mathbb{Z}_{2}^{3};\circ)$, where the operation is defined by $x\circ y=(x^{1}+y^{1},x^{2}+y^{2},x^{3}+y^{3}+x^{1}\cdot y^{2})$. Then $\mathbf{A}$ is isomorphic to $\mathbf{D}_{4}$, as witnessed by the mapping $(0,0,0)\to 1$, $(0,0,1)\to r^{2}$, $(1,1,0)\to r$, $(1,1,1)\to r^{3}$, $(0,1,0)\to s$, $(0,1,1)\to sr^{2}$, $(1,0,0)\to sr$, $(1,0,1)\to sr^{3}$. Although the elements of $\mathbb{Z}_{2}^{3}$ are vectors of length 3, we do not use boldface notation for them in order to distinguish them from tuples of such vectors. However, we will still refer to their components as $a^{1}$, $a^{2}$, and $a^{3}$. Let us define a relational structure $\mathbb{D}_{4}$ corresponding to the dihedral group.
$\bm{\mathbb{D}_{4}}=(\mathbb{Z}_{2}^{3};L_{3}^{1,2},L_{2,3}^{1},L_{1,3}^{2},E_{1-2},E_{2-3}^{0},E_{1-3}^{0},R,\{(0,0,0)\})$, where
$L_{3}^{1,2}(x_{1},x_{2},x_{3},x_{4})=$ $(x_{1}^{1}=x_{2}^{1}\wedge x_{1}^{2}=x_{2}^{2}\wedge x_{3}^{1}=x_{4}^{1}\wedge x_{3}^{2}=x_{4}^{2}\wedge x_{1}^{3}+x_{2}^{3}=x_{3}^{3}+x_{4}^{3}),$
$L_{2,3}^{1}(x_{1},x_{2},x_{3},x_{4})=$ $(x_{1}^{1}=x_{2}^{1}=x_{3}^{1}=x_{4}^{1}\wedge x_{1}^{2}+x_{2}^{2}=x_{3}^{2}+x_{4}^{2}\wedge x_{1}^{3}+x_{2}^{3}=x_{3}^{3}+x_{4}^{3}),$
$L_{1,3}^{2}(x_{1},x_{2},x_{3},x_{4})=$ $(x_{1}^{2}=x_{2}^{2}=x_{3}^{2}=x_{4}^{2}\wedge x_{1}^{1}+x_{2}^{1}=x_{3}^{1}+x_{4}^{1}\wedge x_{1}^{3}+x_{2}^{3}=x_{3}^{3}+x_{4}^{3}),$
$E_{1-2}(x,y)=$ $(x^{1}=y^{2}),$
$E_{2-3}^{0}(x,y)=$ $(x^{1}=y^{1}=0\wedge x^{2}=y^{3}),$
$E_{1-3}^{0}(x,y)=$ $(x^{2}=y^{2}=0\wedge x^{1}=y^{3}),$
$R(x,y)=$ $(x^{1}=y^{2}\wedge x^{2}=y^{1}\wedge x^{3}+y^{3}=x^{1}\cdot y^{1}).$
In Section 7 we prove the following lemma.
Lemma 4.1.
$\operatorname{Pol}(\mathbb{D}_{4})=\operatorname{Clo}(x\circ y)$.
Since $\operatorname{CSP}(\mathbb{D}_{4})$ is trivial — every instance admits a solution by assigning all variables to 0 — we extend $\mathbb{D}_{4}$ with all constant relations $\{\{a\}\mid a\in\mathbb{Z}_{2}^{3}\}$ and denote the resulting structure by $\mathbb{D}_{4}^{idemp}$.
Corollary 4.2.
$\operatorname{Pol}(\mathbb{D}_{4}^{idemp})$ is the set of all idempotent operations from $\operatorname{Clo}(x\circ y)$.
We now show that $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ does not solve $\operatorname{CSP}(\mathbb{D}_{4}^{idemp})$. To witness this, we construct a concrete instance of $\operatorname{CSP}(\mathbb{D}_{4}^{idemp})$ with no solution, for which $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ nevertheless returns “Yes”.
Notice that all relations in $\mathbb{D}_{4}^{idemp}$, except for $R$, are defined by conjunctions of linear equations on the components of the variables. Hence, any instance of $\operatorname{CSP}(\mathbb{D}_{4}^{idemp})$ not involving $R$ can be solved directly by the $\mathrm{AIP}$ algorithm. However, the nonlinear part in $R$ breaks $\mathrm{AIP}$. Let us build a concrete CSP instance $\mathcal{I}$:
$L_{2,3}^{1}(x_{1},x_{2},x_{3},w)\wedge L_{1,3}^{2}(y_{1},y_{2},y_{3},w)\wedge R(x_{1},y_{1})\wedge R(x_{2},y_{2})\wedge L_{3}^{1,2}(x_{3},y_{3},u,v)\wedge u=(0,0,0)\wedge v=(0,0,1).$
Claim 4.5.
$\mathcal{I}$ has no solution.
Proof.
The constraints $L_{2,3}^{1}$, $L_{1,3}^{2}$, and $L_{3}^{1,2}$ imply that $y_{3}^{1}=x_{3}^{1}=w^{1}$ and $x_{3}^{2}=y_{3}^{2}=w^{2}$ in any solution. Then $L_{2,3}^{1}$ and $L_{1,3}^{2}$ imply that $x_{1}^{2}=x_{2}^{2}$ and $y_{1}^{1}=y_{2}^{1}$, respectively. To derive a contradiction, we sum the third components of the first five constraints (equations). Substituting the values for $u$ and $v$, we obtain
$(x_{1}^{3}+x_{2}^{3}+x_{3}^{3}+w^{3})+(y_{1}^{3}+y_{2}^{3}+y_{3}^{3}+w^{3})+(x_{1}^{3}+y_{1}^{3}+x_{1}^{1}\cdot y_{1}^{1})+$
$(x_{2}^{3}+y_{2}^{3}+x_{2}^{1}\cdot y_{2}^{1})+(x_{3}^{3}+$ $y_{3}^{3}+0+1)=0.$
This simplifies to $x_{1}^{1}\cdot y_{1}^{1}+x_{2}^{1}\cdot y_{2}^{1}=1$, which is impossible because $x_{1}^{1}=x_{2}^{1}$ and $y_{1}^{1}=y_{2}^{1}$. ∎
Thus, by examining the linear equations we deduced that $x_{1}$ and $x_{2}$ must be equal in the first two components, making it impossible to satisfy the $R$ constraints. Nevertheless, even $\mathrm{Singl}\mathrm{AIP}$ returns “Yes” on the instance $\mathcal{I}$. Below we present an AIP solution of the instance, and we claim (without a proof) that a similar solution can be constructed after fixing any variable to any value.
Claim 4.6.
$\mathrm{AIP}$ returns “Yes” on $\mathcal{I}$.
Proof.
The following table shows the weights of tuples corresponding to $L_{2,3}^{1}(x_{1},x_{2},x_{3},w)$ and $L_{1,3}^{2}(y_{1},y_{2},y_{3},w)$ in an AIP solution. The weights of tuples for the remaining constraints are straightforward.
[…]
The idea behind this AIP solution is that, for AIP, having two separate constraints $x_{1}^{1}=x_{2}^{1}$ and $x_{1}^{2}=x_{2}^{2}$ is not equivalent to having a single constraint $(x_{1}^{1},x_{1}^{2})=(x_{2}^{1},x_{2}^{2})$.
For a formal proof that $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ does not solve $\operatorname{CSP}(\mathbb{D}_{4}^{idemp})$, we apply Theorem 3.6 together with the following lemma proved in Section 7 (originally proved by Michael Kompatscher).
Lemma 4.3.
Suppose $n,\ell\geq 4$. Then $\operatorname{Clo}(x\circ y)$ does not contain an idempotent $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette symmetric operation.
Corollary 4.4.
The algorithm $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ does not solve $\operatorname{CSP}(\mathbb{D}_{4}^{idemp})$.
5 Characterization of Singleton Algorithms
In this section we prove the characterization of the singleton algorithms in terms of palette symmetric functions. We start with a formal definition of a minion and show how abstract minions can be used to characterize the power of algorithms. Then, in Section 5.2 we apply the Hales-Jewett theorem to show that, under certain conditions, a homomorphic image of a minion contains more symmetric objects than the original. In the last two subsections we show how to apply this result to obtain a characterization of the algorithms $\mathrm{Singl}\mathrm{AC}$ and $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$, respectively.
5.1 Minions and free structure $\mathbb{F}_{\mathcal{M}}(\mathbb{A})$
To simplify understanding of the notion of an abstract minion we start with a function minion. For a function $f\colon A^{n}\to B$ and a mapping $\pi\colon[n]\to[n^{\prime}]$, the minor of $f$ given by $\pi$, denoted by $f^{\pi}$, is the function of arity $n^{\prime}$ defined by
$f^{\pi}(x_{1},\dots,x_{n^{\prime}}):=f(x_{\pi(1)},\dots,x_{\pi(n)}).$
A subset $\mathcal{M}\subseteq\mathcal{O}_{A,B}$ is called a function minion if for any $f\in\mathcal{M}\cap\mathcal{O}_{A,B}^{(n)}$ and any mapping $\pi\colon[n]\to[n^{\prime}]$, the minor $f^{\pi}$ belongs to $\mathcal{M}$. Thus, a function minion can be viewed as a weaker version of a clone, where we only allow to permute and identify variables, and add dummy variables. For instance, the sets $\mathcal{O}_{A,B}$ and $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ are function minions.
Variables (arguments) of an $n$-ary function are usually indexed by numbers $1,2,\dots,n$, which sometimes make argumentation unnecessarily cumbersome, as it requires an additional encoding. To avoid this, we consider functions $f\colon A^{N}\to B$ for finite sets $N$, denoted by $\mathcal{O}_{A,B}^{(N)}$; that is, each function is a mapping equipped with argument names from the set $N$. Then, for a function $f\colon A^{N}\to B$ and a mapping $\pi:N\to N^{\prime}$, where $N^{\prime}$ is also a finite set, the minor of $f$ given by $\pi$ is defined by $f^{\pi}(\mathbf{a})=f(\mathbf{a}\circ\pi)$ for every $\mathbf{a}\in A^{N^{\prime}}$. The advantages of this approach can already be seen in the definition of a free structure below (compare with the respective definitions in [3, 14]).
Thus, a function minion $\mathcal{M}$ is a collection of sets $\mathcal{M}^{(N)}\subseteq\mathcal{O}_{A,B}^{(N)}$ for finite sets $N$ such that for any $f\in\mathcal{M}^{(N)}$ and any mapping $\pi\colon N\to N^{\prime}$, where $N^{\prime}$ is also a finite set, the minor $f^{\pi}$ belongs to $\mathcal{M}^{(N^{\prime})}$. Then the set of all functions $\mathcal{O}_{A,B}$ can be viewed as a minion, whose $\mathcal{O}_{A,B}^{(N)}$-part consists of all mappings $A^{N}\to B$. Similarly, $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ is a minion such that $f\in\operatorname{Pol}(\mathbb{A},\mathbb{B})^{(N)}$ whenever $f\in\mathcal{O}_{A,B}^{(N)}$ and for any $R$ from the signature of $\mathbb{A}$ and any $\mathbf{a}\colon N\to R^{\mathbb{A}}$ we have $f(\mathbf{a})\in R^{\mathbb{B}}$. Thus, $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ is no different than the usual set of polymorphisms with arbitrary names chosen for the arguments.
Now let us define the notion of an abstract minion, where instead of functions we have arbitrary objects, and taking minors satisfies the same properties we have for function minions.
Minion. (Abstract) minion $\mathcal{M}$ is a collection of sets $\mathcal{M}^{(N)}$ for finite sets $N$ together with a minor map $\mathcal{M}^{(\pi)}\colon\mathcal{M}^{(N)}\to\mathcal{M}^{(N^{\prime})}$ for every mapping $\pi\colon N\to N^{\prime}$, such that $\mathcal{M}^{(\mathrm{id}_{N})}=\mathrm{id}_{\mathcal{M}^{(N)}}$ for all finite sets $N$, and $\mathcal{M}^{(\pi)}\circ\mathcal{M}^{(\pi^{\prime})}=\mathcal{M}^{(\pi\circ\pi^{\prime})}$ whenever such a composition is well defined. When a minion is clear from context, we write $f^{\pi}$ instead of $\mathcal{M}^{(\pi)}(f)$. We refer to $\mathcal{M}^{(N)}$ as to the $N$-ary part of $\mathcal{M}$.
A collection of sets $\mathcal{M}_{0}^{(N)}$ for finite sets $N$ is called a subminion of a minion $\mathcal{M}$ if $\mathcal{M}_{0}^{(N)}\subseteq\mathcal{M}^{(N)}$ for all $N$ and the minor maps $\mathcal{M}_{0}^{(\pi)}$ and $\mathcal{M}^{(\pi)}$ coincide on $\mathcal{M}_{0}$.
Minion homomorphism. Let $\mathcal{M}$ and $\mathcal{M}^{\prime}$ be minions. A minion homomorphism from $\mathcal{M}$ to $\mathcal{M}^{\prime}$ is a collection of mappings $\xi^{(N)}\colon\mathcal{M}^{(N)}\to\mathcal{M}^{\prime(N)}$ for finite sets $N$, such that they preserve taking minors; that is, $\xi^{(N^{\prime})}(\mathcal{M}^{(\pi)}(f))=\mathcal{M}^{\prime(\pi)}(\xi^{(N)}(f))$ for all finite sets $N,N^{\prime}$, all $f\in\mathcal{M}^{(N)}$, and all mappings $\pi\colon N\to N^{\prime}$. We write $\mathcal{M}_{1}\to\mathcal{M}_{2}$ to indicate the existence of a minion homomorphism from $\mathcal{M}_{1}$ to $\mathcal{M}_{2}$.
Free structure $\mathbb{F}_{\mathcal{M}}(\mathbb{A})$. For a relational structure $\mathbb{A}$ and a minion $\mathcal{M}$, the free structure $\mathbb{F}_{\mathcal{M}}(\mathbb{A})$ is the structure with domain $\mathcal{M}^{(A)}$ (potentially infinite) and with the same signature as $\mathbb{A}$. For each $k$-ary relation $R^{\mathbb{A}}$, the relation $R^{\mathbb{F}_{\mathcal{M}}(\mathbb{A})}$ consists of all tuples $(f_{1},\dots,f_{k})\in(\mathcal{M}^{(A)})^{k}$ such that there exists $g\in\mathcal{M}^{(R^{\mathbb{A}})}$ satisfying $f_{i}=g^{\pi_{i}}$ for every $i$, where $\pi_{i}\colon R^{\mathbb{A}}\to A$ is a projection onto the $i$-th coordinate.
Lemma 5.1 ([3, 14]).
Suppose $\mathbb{A}$ and $\mathbb{B}$ are finite relational structures in the same signature and $\mathcal{M}$ is a minion. Then the following conditions are equivalent:
  • (1)
    $\mathbb{X}\to\mathbb{F}_{\mathcal{M}}(\mathbb{A})$ implies $\mathbb{X}\to\mathbb{B}$ for any finite $\mathbb{X}$;
  • (2)
    $\mathbb{F}_{\mathcal{M}}(\mathbb{A})\to\mathbb{B}$;
  • (3)
    $\mathcal{M}\to\operatorname{Pol}(\mathbb{A},\mathbb{B})$.
Proof.
[…]
Corollary 5.2.
Suppose $\mathbb{A}$ and $\mathbb{B}$ are finite relational structures in the same signature, $\mathfrak{A}$ is an algorithm for $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$, $\mathcal{M}$ is a nonempty minion, and the following conditions are equivalent for any instance $\mathbb{X}$:
  • $\mathfrak{A}$ returns “Yes” on $\mathbb{X}$;
  • $\mathbb{X}\to\mathbb{F}_{\mathcal{M}}(\mathbb{A})$.
Then $\mathfrak{A}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if $\mathcal{M}\to\operatorname{Pol}(\mathbb{A},\mathbb{B})$.
Proof.
First, let us show that $\mathfrak{A}$ returns “Yes” on any Yes-instance $\mathbb{X}$. Let $\varphi\colon\mathbb{X}\to\mathbb{A}$ be a homomorphism. Since the minion $\mathcal{M}$ is nonempty, there should be a unary object $f\in\mathcal{M}^{(\{1\})}$. Then a homomorphism $\psi\colon\mathbb{X}\to\mathbb{F}_{\mathcal{M}}(\mathbb{A})$ can be defined by $\psi(x)=f^{\pi_{x}}$, where $\pi_{x}\colon\{1\}\to A$ such that $\pi_{x}(1)=\varphi(x)$. This implies that $\mathfrak{A}$ returns “Yes” on $\mathbb{X}$.
Second, let us show that the following conditions are equivalent:
  • $\mathfrak{A}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$;
  • $\mathbb{X}\to\mathbb{F}_{\mathcal{M}}(\mathbb{A})$ implies $\mathbb{X}\to\mathbb{B}$ for any finite $\mathbb{X}$.
Assume that $\mathfrak{A}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$. Then $\mathbb{X}\to\mathbb{F}_{\mathcal{M}}(\mathbb{A})$ implies that the algorithm returns “Yes” on the instance $\mathbb{X}$, which means that $\mathbb{X}\to\mathbb{A}$ or $\mathbb{X}$ is incorrect ($\mathbb{X}\to\mathbb{B}$ and $\mathbb{X}\not\to\mathbb{A}$). Hence, $\mathbb{X}\to\mathbb{F}_{\mathcal{M}}(\mathbb{A})$ implies $\mathbb{X}\to\mathbb{B}$ for any $\mathbb{X}$. For the converse, assume that $\mathbb{X}\to\mathbb{F}_{\mathcal{M}}(\mathbb{A})$ implies $\mathbb{X}\to\mathbb{B}$. Let us show that $\mathfrak{A}$ works correctly on any instance $\mathbb{X}$. For a Yes-instance, it was shown earlier. If $\mathfrak{A}$ returns “Yes” on $\mathbb{X}$, then $\mathbb{X}\to\mathbb{F}_{\mathcal{M}}(\mathbb{A})$, which implies $\mathbb{X}\to\mathbb{B}$, hence $\mathbb{X}$ is indeed a Yes-instance.
To complete the proof, it is sufficient to use the equivalence of conditions (1) and (3) in Lemma 5.1. ∎
5.2 General characterization of singleton
Theorem 5.3 (Hales-Jewett Theorem, [27]).
Suppose $A$ and $B$ are finite sets. Then there exists $n\in\mathbb{N}$ such that for any mapping $f\colon A^{n}\to B$ there exists $(v_{1},\dots,v_{n})\in(A\cup\{x\})^{n}\setminus A^{n}$, where $x$ is a variable symbol, such that the unary mapping $h(x)\approx f(v_{1},\dots,v_{n})$ is a constant.
Suppose $\circ$ is an operation defined on $\mathcal{M}^{(n)}$ for every finite set $n$. We say that an operation $\circ$ respects taking minors if for any $\pi\colon n\to N$ and any $f_{1},f_{2}\in\mathcal{M}^{(n)}$ we have $f_{1}^{\pi}\circ f_{2}^{\pi}=(f_{1}\circ f_{2})^{\pi}$.
Theorem 5.4.
Suppose $n$ and $N$ are finite sets, $\mathcal{M}_{0}$ is a minion equipped with a semigroup operation $\circ$ that respects taking minors, $\mathcal{M}\subseteq\mathcal{M}_{0}$ is a subminion, $\mathcal{N}$ is a minion, $\mathcal{N}^{(n)}$ is finite, $h\colon\mathcal{M}\to\mathcal{N}$ is a minion homomorphism, $\mathcal{A}\subseteq\mathcal{M}_{0}^{(N)}$, $I\in\mathcal{M}_{0}^{(N)}$, $\Omega$ is a set of mappings $N\to n$, $\omega\subseteq\Omega\times\Omega$, and
  1. (1)
    $U^{\sigma_{1}}=U^{\sigma_{2}}$ for any $(\sigma_{1},\sigma_{2})\in\omega$ and $U\in\mathcal{A}$;
  2. (2)
    for any $\sigma\in\Omega$ there exists $U\in\mathcal{A}$ such that $U^{\sigma}=I^{\sigma}$;
  3. (3)
    $U_{1}\circ\dots\circ U_{\ell}\in\mathcal{M}$ for any $\ell\in\mathbb{N}$ and $(U_{1},\dots,U_{\ell})\in(\mathcal{A}\cup\{I\})^{\ell}\setminus\mathcal{A}^{\ell}$.
Then there exists $M\in\mathcal{M}^{(N)}$ such that $h(M)^{\sigma_{1}}=h(M)^{\sigma_{2}}$ for any $(\sigma_{1},\sigma_{2})\in\omega$.
Proof.
Let $\mathcal{B}$ be the set of all mappings $\Omega\to(\mathcal{N}^{(n)}\cup\{*\})$. Notice that $\mathcal{B}$ is finite. Let us choose a finite subset $\mathcal{A}_{0}\subseteq\mathcal{A}$ that still satisfies condition (2). For every $L\in\mathbb{N}$ define a mapping $\xi_{L}\colon\mathcal{A}_{0}^{L}\to\mathcal{B}$ by
$\xi_{L}(U_{1},\dots,U_{L})(\sigma)=\begin{cases}h((U_{1}\circ\dots\circ U_{L})^{\sigma}),&\text{if $(U_{1}\circ\dots\circ U_{L})^{\sigma}\in\mathcal{M}$;}\\ *&\text{otherwise.}\end{cases}$
By the Hales-Jewett theorem, choose $L$ and $(T_{1},\dots,T_{L})\in(\mathcal{A}_{0}\cup\{I\})^{L}\setminus\mathcal{A}_{0}^{L}$, where $I$ replaces the variable symbol. Let us show that $M=T_{1}\circ\dots\circ T_{L}$ satisfies all required properties. By condition (3) we have $M\in\mathcal{M}$. It remains to prove that $h(M)^{\sigma_{1}}=h(M)^{\sigma_{2}}$ for any $(\sigma_{1},\sigma_{2})\in\omega$. By condition (2) there exist $A_{1},A_{2}\in\mathcal{A}_{0}$ such that $A_{1}^{\sigma_{1}}=I^{\sigma_{1}}$ and $A_{2}^{\sigma_{2}}=I^{\sigma_{2}}$. To simplify the argument, assume that $T_{1}=T_{2}=I$ and $T_{i}\neq I$ for $i\geq 3$. Then
$h(M)^{\sigma_{1}}\overset{1}{=}h(M^{\sigma_{1}})\overset{2}{=}$ $h((T_{1}\circ\dots\circ T_{L})^{\sigma_{1}})\overset{3}{=}$
$h(T_{1}^{\sigma_{1}}\circ\dots\circ T_{L}^{\sigma_{1}})\overset{4}{=}$
$h(A_{1}^{\sigma_{1}}\circ A_{1}^{\sigma_{1}}\circ T_{3}^{\sigma_{1}}\circ\dots\circ T_{L}^{\sigma_{1}})\overset{5}{=}$
$h(A_{1}^{\sigma_{2}}\circ A_{1}^{\sigma_{2}}\circ T_{3}^{\sigma_{2}}\circ\dots\circ T_{L}^{\sigma_{2}})\overset{6}{=}$
$\xi_{L}(A_{1},A_{1},T_{3},\dots,T_{L})(\sigma_{2})\overset{7}{=}$
$\xi_{L}(A_{2},A_{2},T_{3},\dots,T_{L})(\sigma_{2})\overset{8}{=}$
$h(A_{2}^{\sigma_{2}}\circ A_{2}^{\sigma_{2}}\circ T_{3}^{\sigma_{2}}\circ\dots\circ T_{L}^{\sigma_{2}})\overset{9}{=}$
$h(T_{1}^{\sigma_{2}}\circ\dots\circ T_{L}^{\sigma_{2}})\overset{10}{=}$
$h((T_{1}\circ\dots\circ T_{L})^{\sigma_{2}})\overset{11}{=}h(M^{\sigma_{2}})\overset{12}{=}h(M)^{\sigma_{2}},$
where the equalities 1 and 12 are by the definition of the minion homomorphism, 2 and 11 are by the definition, 3 and 10 are by the property of the $\circ$, 4 and 9 are by the choice of $A_{1}$ and $A_{2}$, 5 is by condition (1), 6 and 8 are by the definition of $\xi_{L}$ and the property of $\circ$, 7 is by the choice of $L$ and $T_{1},\dots,T_{L}$. ∎
5.3 Characterization of the singleton arc-consistency algorithm
We characterize minions for the singleton versions via matrices, following the ideas of Ciardo and Živný in [23]. Let $\mathcal{M}_{\mathrm{SArc}}$ be the minion whose N{N}-ary part consists of matrices $M\colon N\times[L]\to\{0,1\}$, where $L\in\mathbb{N}$, such that:
  1. (1)
    for every $j\in[L]$ there exists $i\in N$ with $M(i,j)=1$ (there exists 1 in every column);
  2. (2)
    if $M(i,j)=1$ for some $i$ and $j$, then there exists $\ell\in[L]$ such that $M(i,\ell)=1$ and $M(i^{\prime},\ell)=0$ for all $i^{\prime}\neq i$ (skeleton property).
Lemma 5.5.
Suppose $\mathbb{A}$ and $\mathbb{B}$ are finite relational structures in the same signature. Then the following conditions are equivalent for any instance $\mathbb{X}$ of $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$:
  • $\mathrm{CSingl}\mbox{{AC}}$ returns “Yes” on $\mathbb{X}$;
  • $\mathbb{X}\to\mathbb{F}_{\mathcal{M}_{\mathrm{SArc}}}(\mathbb{A})$.
Proof.
$\Rightarrow$ Assume that $\mathrm{CSingl}\mbox{{AC}}$ returns “Yes” on $\mathbb{X}$. First, notice that finding a homomorphism $\mathbb{X}\to\mathbb{F}_{\mathcal{M}_{\mathrm{SArc}}}(\mathbb{A})$ is equivalent to finding matrices from $\mathcal{M}_{\mathrm{SArc}}$ for each constraint and each variable such that they are compatible. Since the algorithm $\mathrm{CSingl}\mbox{{AC}}$ returns “Yes”, for every constraint $C=R(x_{1},\dots,x_{s})$ we obtain a nonempty relation $R_{f}$ (the final version of $R$ in the pseudocode). Let $P_{1},\dots,P_{L}$ be the list of all pairs $(C,\mathbf{a})$, where $C$ is a constraint and $\mathbf{a}$ is a tuple from the corresponding $R_{f}$. Then for every $i\in[L]$, the algorithm $\mathrm{AC}$ returns a nonempty subset $D_{y}^{i}$ for each variable $y$. The matrix $M_{x}\colon A\times[L]\to\{0,1\}$ corresponding to the variable $x$ is defined by $M_{x}(a,i):=\begin{cases}1,&\text{if $a\in D_{x}^{i}$}\\ 0,&\text{otherwise}\end{cases}$. Similarly, the matrix $M_{C}\colon R^{\mathbb{A}}\times[L]\to\{0,1\}$ corresponding to a constraint $C=R(x_{1},\dots,x_{s})$ is defined by
$M_{C}(\mathbf{a},i):=\begin{cases}1,&\text{if $\mathbf{a}^{j}\in D_{x_{j}}^{i}$ for every $j\in[s]$}\\ 0,&\text{otherwise}\end{cases}.$
Let us show that the matrices for constraints and for variables are consistent, that is, for any constraint $C=R(x_{1},\dots,x_{s})$ and $i\in[s]$, we have $M_{C}^{\pi_{i}}=M_{x}$, where $\pi_{i}\colon R^{\mathbb{A}}\to A$ is a projection onto the $i$-th coordinate. Since the $i$-th columns of all matrices come from the output of the arc-consistency algorithm, the $i$-th columns are consistent. Taking a minor of a matrix object is equivalent to taking the corresponding minor of each column and then gluing the results, so the matrices themselves are consistent. Hence the assignment we defined yields a homomorphism $\mathbb{X}\to\mathbb{F}_{\mathcal{M}_{\mathrm{SArc}}}(\mathbb{A})$.
$\Leftarrow$ Assume that $h\colon\mathbb{X}\to\mathbb{F}_{\mathcal{M}_{\mathrm{SArc}}}(\mathbb{A})$ is a homomorphism. First, notice that it is sufficient to prove the claim for connected instances $\mathbb{X}$, since the algorithm $\mathrm{CSingl}\mbox{{AC}}$ essentially works with disconnected parts independently. Thus, we assume that $\mathbb{X}$ is connected. Then for any two variables $x_{1},x_{2}\in X$, the matrices $h(x_{1})$ and $h(x_{2})$ have the same width $L\in\mathbb{N}$. For each constraint $C=R(x_{1},\dots,x_{s})$, let $R^{\prime}$ be the set of all tuples $\mathbf{a}\in R$ such that the $\mathbf{a}$-th row of the matrix corresponding to $C$ has a non-zero entry. Since the matrices in $\mathcal{M}_{\mathrm{SArc}}$ are skeleton, for each tuple $\mathbf{a}\in R^{\prime}$ there exists $\ell_{C,\mathbf{a}}\in[L]$ such that the $\ell_{C,\mathbf{a}}$-th column of the matrix corresponding to $C$ has 1 only in the $\mathbf{a}$-th position. Let $D_{y}^{C,\mathbf{a}}$ be the set of all $b\in D_{y}$ such that $h(y)(b,\ell_{C,\mathbf{a}})=1$. Since the sets $(D_{y}^{C,\mathbf{a}}\colon y\in X)$ form an arc-consistent reduction (see Observation 3.1), the arc-consistency algorithm must return “Yes” after we restrict the constraint $C$ to $\mathbf{a}$. Hence, $\mathrm{CSingl}\mathrm{AC}$ returns “Yes”. ∎
Corollary 5.6.
Suppose $\mathbb{A}$ and $\mathbb{B}$ are finite relational structures in the same signature. Then $\mathrm{CSingl}\mathrm{AC}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if there exists a minion homomorphism $\mathcal{M}_{\mathrm{SArc}}\to\operatorname{Pol}(\mathbb{A},\mathbb{B})$.
Proof.
This follows from Corollary 5.2 and Lemma 5.5. ∎
Lemma 5.7.
Suppose $n,\ell\in\mathbb{N}$, $\mathcal{N}$ is a minion whose $n$-ary part is finite, $h\colon\mathcal{M}_{SArc}\to\mathcal{N}$ is a minion homomorphism. Then $h(\mathcal{M}_{SArc})$ contains an $(\underbrace{\overline{\ell},\dots,\overline{\ell}}_{n})$-palette totally symmetric object.
Proof.
Let us define everything we need in Theorem 5.4. Let $N=n\cdot\ell$, with $[N]$ playing the role of $N$ and $[n]$ the role of $n$. Let $\mathcal{M}=\mathcal{M}_{SArc}$, and let $\mathcal{M}_{0}$ be the set of all matrices satisfying all the properties of $\mathcal{M}_{\mathrm{SArc}}$ but property (2) (skeleton property). By $\circ$ denote the concatenation of two matrices, which is a semigroup operation that respects taking minors. For $i\in[n]$, define $\phi(i)$ to be the matrix $N]\times[1]\to\{0,1\}$ whose only column is $(\underbrace{0,\dots,0}_{(i-1)\ell},\underbrace{1,\dots,1}_{\ell},\underbrace{0,\dots,0}_{(n-i)\ell})$. Let
$\mathcal{A}=\{\phi(a_{1})\circ\dots\circ\phi(a_{N})\mid a_{1},\dots,a_{N}\in[n]\},$
and $\mathcal{I}\in\mathcal{M}$ be the identity matrix of size $N\times N$.
A mapping $\pi\colon[N]\to[n]$ can be viewed as a tuple of length $N$, so the notions of a palette tuple and of the $i$-th block can be extended naturally to such mappings. Let $\Omega$ be the set of all $(\underbrace{\overline{\ell},\dots,\overline{\ell}}_{n})$-palette mappings $N]\to[n]$. Let $\omega$ be the set of all pairs $(\sigma_{1},\sigma_{2})$ such that, for every $i\in[n]$, the $i$-th block of $\sigma_{1}$ contains the same set of elements as the $i$-th block of $\sigma_{2}$.
Let us check the required conditions (1)-(3) of Theorem 5.4. Condition (1) follows from the definition of $\phi$. For condition (2), let $\sigma\in\Omega$. For each $i$ appearing in $\sigma$ (viewed as a tuple), let $u(i)$ be the position of a block that contains only $i$ (such a block exists since $\sigma$ is palette). Then $\phi(u(i))^{\sigma}$ has $1$ at the $i$-th position and 0 elsewhere. Hence, $(\phi(u(\sigma(1)))\circ\dots\circ\phi(u(\sigma(N))))^{\sigma}=I^{\sigma}$. Condition (3) holds because adding $I$ makes the matrix skeleton. By Theorem 5.4 we obtain $M\in\mathcal{M}_{\mathrm{SArc}}^{(N)}$ such that $h(M)^{\sigma_{1}}=h(M)^{\sigma_{2}}$ for any $(\sigma_{1},\sigma_{2})\in\omega$. Thus, $h(M)$ is an $(\underbrace{\overline{\ell},\dots,\overline{\ell}}_{n})$-palette totally symmetric object. ∎
Lemma 5.8.
Suppose $\mathbb{A}$ and $\mathbb{B}$ are relational structures in the same signature, $\mathbb{A}$ is finite, $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ has $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette totally symmetric functions for all $\ell_{1},\dots,\ell_{n}\in\mathbb{N}$. Then $\mathrm{SinglArcCons}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$.
Proof.
Notice that by identifying variables within blocks of an $(L_{1},\dots,L_{n})$-palette totally symmetric (PTS) function we obtain a PTS function with smaller block sizes. Therefore, for all $L_{1},\dots,L_{n}\in\mathbb{N}$ we can choose an $(L_{1},\dots,L_{n})$-PTS function $f_{L_{1},\dots,L_{n}}\in\operatorname{Pol}(\mathbb{A},\mathbb{B})$ such that identifying variables within blocks of $f_{L_{1},\dots,L_{n}}$ yields $f_{\ell_{1},\dots,\ell_{n}}$, where $\ell_{i}\leq L_{i}$ for every $i$. To simplify notation, we write $f(S_{1},\dots,S_{n})$ meaning $f_{k_{1},\dots,k_{n}}(s_{1}^{1},\dots,s_{1}^{k_{1}},\dots,s_{n}^{1},\dots,s_{n}^{k_{n}})$, where $S_{i}=\{s_{i}^{1},\dots,s_{i}^{k_{i}}\}$ for every $i\in[n]$. This notation is justified because we only apply $f_{k_{1},\dots,k_{n}}$ to palette tuples, so the order of elements in the sets $S_{i}$ does not matter.
We need to check that if $\mathrm{SinglArcCons}$ returns “Yes” on an instance $\mathcal{I}$, then the instance has a solution in $\mathbb{B}$. Let $P_{1},\dots,P_{S}$ be all pairs $(x,a)$, where $x$ is a variable of $\mathcal{I}$ and $a\in D_{x}$ at the end of the algorithm. For each $P_{j}=(x,a)$ and each variable $y$, let $D_{y}^{j}$ be the domain of $y$ after restricting $x$ to $\{a\}$ and running arc-consistency. Let us show that a solution to $\mathcal{I}^{\mathbb{B}}$ can be defined by $s(x):=f(D_{x}^{1},\dots,D_{x}^{S})$ for every variable $x$. Let $C=R(x_{1},\dots,x_{k})$ be a constraint $\mathcal{I}$. For $j\in[S]$ set $R_{j}=R^{\mathbb{A}}\cap(D^{j}_{x_{1}}\times\dots\times D^{j}_{x_{k}})$. By the assumptions on $f$, we have $f(R_{1},\dots,R_{S})=(s(x_{1}),\dots,s(x_{k}))$. Since $f$ is a polymorphism, it follows that $(s(x_{1}),\dots,s(x_{k}))\in R^{\mathbb{B}}$. Thus, $s$ is a solution of $\mathcal{I}$ in $\mathbb{B}$, completing the proof. ∎
Theorem 3.5.
Let $(\mathbb{A},\mathbb{B})$ be a PCSP template. The following conditions are equivalent:
  1. (1)
    $\mathrm{SinglArcCons}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$;
  2. (2)
    $\mathrm{CSingl}\mathrm{AC}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$;
  3. (3)
    $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ has an $(\overline{m_{1}},\dots,\overline{m_{n}})$-palette totally symmetric function for every $n,m_{1},\dots,m_{n}\in\mathbb{N}$;
  4. (4)
    for every $N\in\mathbb{N}$ there exist $n,m_{1},\dots,m_{n}\geq N$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ has an $(\overline{m_{1}},\dots,\overline{m_{n}})$-palette totally symmetric function.
Proof.
[…]
5.4 Characterization of the singleton BLP+AIP
Let $\mathcal{M}_{\mathrm{S(BLP+AIP)}}$ be the minion whose N{N}-ary part consists of pairs of matrices $M_{\mathrm{BLP}}\colon N\times[L]\to\mathbb{Q}\cap[0,1]$ and $M_{\mathrm{AIP}}\colon N\times[L]\to\mathbb{Z}$, where $L\in\mathbb{N}$, such that:
  1. (1)
    $\sum\limits_{i\in N}M_{\mathrm{BLP}}(i,j)=1$ and $\sum\limits_{i\in N}M_{\mathrm{AIP}}(i,j)=1$ for every $j\in[L]$;
  2. (2)
    if $M_{\mathrm{AIP}}(i,j)\neq 0$ for some $i,j$, then $M_{\mathrm{BLP}}(i,j)>0$;
  3. (3)
    if $M_{\mathrm{BLP}}(i,j)>0$ for some $i$ and $j$, then there exists $\ell\in[L]$ such that $M_{\mathrm{BLP}}(i,\ell)=M_{\mathrm{AIP}}(i,\ell)=1$ and $M_{\mathrm{BLP}}(i^{\prime},\ell)=M_{\mathrm{AIP}}(i^{\prime},\ell)=0$ for all $i^{\prime}\neq i$ (skeleton property).
Lemma 5.9.
Suppose $\mathbb{A}$ and $\mathbb{B}$ are finite relational structures in the same signature. Then the following conditions are equivalent for any instance $\mathbb{X}$ of $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$:
  • $\mathrm{CSingl}(\mathrm{BLP}+\mathrm{AIP})$ returns “Yes” on $\mathbb{X}$;
  • $\mathbb{X}\to\mathbb{F}_{\mathcal{M}_{\mathrm{S(BLP+AIP)}}}(\mathbb{A})$.
Proof.
$\Rightarrow$ Assume that $\mathrm{CSingl}(\mathrm{BLP}+\mathrm{AIP})$ returns “Yes” on $\mathbb{X}$. First, notice that finding a homomorphism $\mathbb{X}\to\mathbb{F}_{\mathcal{M}_{\mathrm{S(BLP+AIP)}}}(\mathbb{A})$ is equivalent to finding pairs of matrices from $\mathcal{M}_{\mathrm{S(BLP+AIP)}}$ for each constraint and each variable so that they are compatible. Since the algorithm $\mathrm{CSingl}(\mathrm{BLP}+\mathrm{AIP})$ returns “Yes”, for every constraint $C=R(x_{1},\dots,x_{s})$ we obtain a nonempty relation $R_{f}$ (the final version of $R$ in the pseudocode). Let $P_{1},\dots,P_{L}$ be the list of all pairs $(C,\mathbf{a})$ with $C$ a constraint and $\mathbf{a}$ a tuple from the corresponding $R_{f}$. For every $i\in[L]$ the algorithm $\mathrm{BLP}+\mathrm{AIP}$ produces a BLP solution $s_{\mathrm{BLP}}^{i}$ and an AIP solution $s_{\mathrm{AIP}}^{i}$. Then, the BLP and AIP matrices corresponding to a variable $x$ can be defined by $M^{\mathrm{BLP}}_{x}(a,i)=s_{\mathrm{BLP}}^{i}(x^{a})$ and $M^{\mathrm{AIP}}_{x}(a,i)=s_{\mathrm{AIP}}^{i}(x^{a})$. Similarly, the BLP and AIP matrices corresponding to a constraint $C$ are $M^{\mathrm{BLP}}_{C}(\mathbf{a},i)=s_{\mathrm{BLP}}^{i}(C^{\mathbf{a}})$ and $M^{\mathrm{AIP}}_{C}(\mathbf{a},i)=s_{\mathrm{AIP}}^{i}(C^{\mathbf{a}})$. By definition of $\mathrm{BLP}+\mathrm{AIP}$, this assignment yields a homomorphism $\mathbb{X}\to\mathbb{F}_{\mathcal{M}_{\mathrm{S(BLP+AIP)}}}(\mathbb{A})$.
$\Leftarrow$ Assume that $h\colon\mathbb{X}\to\mathbb{F}_{\mathcal{M}_{\mathrm{S(BLP+AIP)}}}(\mathbb{A})$ is a homomorphism. First, notice that it is sufficient to prove the claim for connected instances $\mathbb{X}$ because the algorithm $\mathrm{CSingl}(\mathrm{BLP}+\mathrm{AIP})$ essentially works with disconnected parts independently. Thus, we assume that $\mathbb{X}$ is connected. Then for any two variables $x_{1},x_{2}\in X$ the matrices in $h(x_{1})$ and $h(x_{2})$ have the same width $L\in\mathbb{N}$. For each constraint $C=R(x_{1},\dots,x_{s})$, let $R^{\prime}$ be the set of all tuples $\mathbf{a}\in R$ such that the $\mathbf{a}$-th row of the BLP matrix corresponding to $C$ has a non-zero entry. Since the matrices in $\mathcal{M}_{\mathrm{S(BLP+AIP)}}$ are skeleton, for each tuple $\mathbf{a}\in R^{\prime}$ there exists $\ell_{C,\mathbf{a}}\in[L]$ such that the $\ell_{C,\mathbf{a}}$-th column of the BLP matrix corresponding to $C$ has 1 only in the $\mathbf{a}$-th position. Then BLP and AIP solutions that $\mathrm{BLP}+\mathrm{AIP}$ returns after fixing $C$ to $\mathbf{a}$ can be defined by $s_{\mathrm{BLP}}(C_{0}^{\mathbf{a}_{0}})=M^{\mathrm{BLP}}_{C_{0}}(\mathbf{a}_{0},\ell_{C,\mathbf{a}})$, $s_{\mathrm{AIP}}(C_{0}^{\mathbf{a}_{0}})=M^{\mathrm{AIP}}_{C_{0}}(\mathbf{a}_{0},\ell_{C,\mathbf{a}})$, $s_{\mathrm{BLP}}(y^{b})=M^{\mathrm{BLP}}_{y}(b,\ell_{C,\mathbf{a}})$, $s_{\mathrm{AIP}}(y^{b})=M^{\mathrm{AIP}}_{y}(b,\ell_{C,\mathbf{a}})$ for all constraints $C_{0}$, their tuples $\mathbf{a}_{0}$, variables $y$, and $b\in D_{y}$. Hence, $\mathrm{CSingl}(\mathrm{BLP}+\mathrm{AIP})$ returns “Yes” on $\mathbb{X}$. ∎
Corollary 5.10.
Suppose $\mathbb{A}$, $\mathbb{B}$ are finite relational structures in the same signature. Then $\mathrm{CSingl}(\mathrm{BLP}+\mathrm{AIP})$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if there exists a minion homomorphism $\mathcal{M}_{\mathrm{S(BLP+AIP)}}\to\operatorname{Pol}(\mathbb{A},\mathbb{B})$.
Proof.
This follows from Corollary 5.2 and Lemma 5.9. ∎
Lemma 5.11.
Suppose $n,\ell\in\mathbb{N}$, $\mathcal{N}$ is a minion whose $n$-ary part is finite, $h\colon\mathcal{M}_{\mathrm{S(BLP+AIP)}}\to\mathcal{N}$ is a minion homomorphism. Then $h(\mathcal{M}_{\mathrm{S(BLP+AIP)}})$ contains an $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette symmetric object.
Proof.
We define all components needed to apply Theorem 5.4. Let $n\cdot(2\ell+1), [N][N$] play the role of $N$, and $[n]$ play the role of $n$. Let $\mathcal{M}=\mathcal{M}_{\mathrm{S(BLP+AIP)}}$ and $\mathcal{M}_{0}$ be the set of all pairs of matrices satisfying all properties of $\mathcal{M}_{\mathrm{S(BLP+AIP)}}$ except property (3) (skeleton property). Define $\circ$ as the concatenation of matrices, applied independently for BLP and AIP matrices. This is a semigroup operation that respects taking minors. For $i\in[n]$ define a pair of matrices $\phi(i)$. Its BLP matrix $N]\times[1]\to\mathbb{Q}\cap[0,1]$ has one column $(\underbrace{0,\dots,0}_{(i-1)(2\ell+1)},\underbrace{\frac{1}{2\ell+1},\dots,\frac{1}{2\ell+1}}_{2\ell+1},\underbrace{0,\dots,0}_{(n-i)(2\ell+1)})$. Its AIP matrix $N]\times[1]\to\mathbb{Z}$ has one column $(\underbrace{0,\dots,0}_{(i-1)(2\ell+1)},\underbrace{1,\dots,1}_{\ell+1},\underbrace{-1,\dots,-1}_{\ell},\underbrace{0,\dots,0}_{(n-i)(2\ell+1)})$. Let $\mathcal{A}=\{\phi(a_{1})\circ\dots\circ\phi(a_{N})\mid a_{1},\dots,a_{N}\in[n]\},$ and let $\mathcal{I}\in\mathcal{M}$ be the pair of identity matrix of size $N\times N$.
Mappings $\pi\colon[N]\to[n]$ can be viewed as tuples of length $N$, and the notions of a palette tuple and the $i$-th block can be extended naturally to such mappings. Let $\Omega$ be the set of all $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette mappings $N]\to[n]$. Let $\omega$ be the set of all pairs $(\sigma_{1},\sigma_{2})$ such that for every $i\in[2n]$ the $i$-th blocks of $\sigma_{1}$ and $\sigma_{2}$ differ only by a permutation of their elements.
Let us check conditions (1)-(3) of Theorem 5.4. Condition (1) follows from the definition of $\phi$. Let us prove that condition (2) is satisfied for every $\sigma\in\Omega$. For each $i$ appearing in $\sigma$ (viewed as a tuple) let $u(i)\in[n]$ be the position of an overlined block containing only $i$ (such a block exists since $\sigma$ is palette). Then both matrices of $\phi(u(i))^{\sigma}$ have $1$ at the $i$-th position and 0 elsewhere. Hence, $(\phi(u(\sigma(1)))\circ\dots\circ\phi(u(\sigma(N))))^{\sigma}=I^{\sigma}$. Condition (3) holds because adding $I$ makes the matrices skeleton. By Theorem 5.4 we can find $M\in\mathcal{M}_{\mathrm{S(BLP+AIP)}}^{(N)}$ such that $h(M)^{\sigma_{1}}=h(M)^{\sigma_{2}}$ for all $(\sigma_{1},\sigma_{2})\in\omega$. Hence, $h(M)$ is an $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette symmetric object. ∎
Lemma 5.12.
Suppose for every $N\in\mathbb{N}$ there exist $n\geq N$, $k_{1},\dots,k_{n}\in\mathbb{N}$, and $m_{1,1},\dots,m_{1,k_{1}},\dots$, $m_{n,1},\dots,m_{n,k_{n}}\geq N$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\overline{m_{1,1},\dots,m_{1,k_{1}}},\overline{m_{2,1},\dots,m_{2,k_{2}}},\dots,\overline{m_{n,1},\dots,m_{n,k_{n}}})$-palette symmetric function. Then $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$.
Proof.
We need to check that if $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ returns “Yes” on an instance $\mathcal{I}$, then the instance has a solution in $\mathbb{B}$. Let $P_{1},\dots,P_{S}$ be all pairs $(x,a)$ where $x$ is a variable of $\mathcal{I}$ and $a\in D_{x}$ at the end of the algorithm. For each $P_{j}=(x,a)$ there exist BLP and AIP solutions to the instance obtained by restricting the variable $x$ to $\{a\}$. Then the argument we used after the Theorem 3.4 implies that for some $N>S$, there exists a $d$-solution $s_{d}^{j}$ to this instance for every $d\geq N$.
Choose an $(\overline{m_{1,1},\dots,m_{1,k_{1}}},\overline{m_{2,1},\dots,m_{2,k_{2}}},\dots,\overline{m_{n,1},\dots,m_{n,k_{n}}})$-palette symmetric function $f\in\operatorname{Pol}(\mathbb{A},\mathbb{B})$, where $n\geq N$, $k_{1},\dots,k_{n}\in\mathbb{N}$, and $m_{1,1},\dots,m_{1,k_{1}},\dots,m_{n,1},\dots,m_{n,k_{n}}\geq N$. For $j\in[S]$ and $i\in[k_{j}]$, let $\alpha_{x}^{j,i}$ be a tuple of length $m_{j,i}$, where each element $a\in A$ appears $s_{m_{j,i}}^{j}(x^{a})$ times. For $j\in\{S+1,S+2,\dots,n\}$ and $i\in[k_{j}]$, let $\alpha_{x}^{j,i}$ be a tuple of length $m_{j,i}$ where each element $a\in A$ appears $s_{m_{j,i}}^{1}(x^{a})$ times. Let us show that a solution to $\mathcal{I}^{\mathbb{B}}$ can be defined by $s(x):=f(\alpha_{x}^{1,1},\dots,\alpha_{x}^{1,k_{1}},\dots,\alpha_{x}^{n,1},\dots,\alpha_{x}^{n,k_{n}})$ for every variable $x$. Let $C=R(x_{1},\dots,x_{k})$ be a constraint of $\mathcal{I}$. For $j\in[S]$ and $i\in[k_{j}]$ let $M_{j,i}$ be the matrix whose columns are tuples from $R$ where each tuple $\mathbf{a}$ appears $s_{m_{j,i}}^{j}(C^{\mathbf{a}})$ times. Similarly, for $j\in\{S+1,S+2,\dots,n\}$ and $i\in[k_{j}]$ let $M_{j,i}$ be the matrix whose columns are tuples from $R$ where each tuple $\mathbf{a}$ appears $s_{m_{j,i}}^{1}(C^{\mathbf{a}})$ times. Our assumptions about $f$ imply that $f(M_{1,1},\dots,M_{1,k_{1}},\dots,M_{n,1},\dots,M_{n,k_{n}})=(s(x_{1}),\dots,s(x_{k}))$. Since $f\in\operatorname{Pol}(\mathbb{A},\mathbb{B})$, it follows that $(s(x_{1}),\dots,s(x_{k}))\in R^{\mathbb{B}}$. Thus, $s$ is a solution of $\mathcal{I}$ in $\mathbb{B}$. ∎
Theorem 3.6.
Let $(\mathbb{A},\mathbb{B})$ be a PCSP template. The following conditions are equivalent:
  1. (1)
    $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$;
  2. (2)
    $\mathrm{CSingl}(\mathrm{BLP}+\mathrm{AIP})$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$;
  3. (3)
    $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette symmetric function for every $\ell,n\in\mathbb{N}$;
  4. (4)
    for every $N\in\mathbb{N}$ there exist $n\geq N$, $k_{1},\dots,k_{n}\in\mathbb{N}$, and $m_{1,1},\dots,m_{1,k_{1}},\dots,m_{n,1},\dots,m_{n,k_{n}}\geq N$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ contains an $(\overline{m_{1,1},\dots,m_{1,k_{1}}},\overline{m_{2,1},\dots,m_{2,k_{2}}},\dots,\overline{m_{n,1},\dots,m_{n,k_{n}}})$-palette symmetric function.
Proof.
[…]
6 Temporal constraint satisfaction problem
As we show in this section, palette functions can also be useful for solving CSPs over infinite domain. Notice that such CSPs can be very complicated. In fact, as was shown in [8], every computational problem is equivalent (under polynomial-time Turing reductions) to a problem of the form $\operatorname{CSP}(\mathbb{A})$ for some infinite structure $\mathbb{A}$. To bring the problem back to the class NP, mathematicians usually require the constraint language to satisfy additional assumptions [11, 6].
6.1 Tractable temporal CSPs
One of the main results on the complexity of CSPs over infinite domain is the following classification of temporal CSPs. A relational structure $\mathbb{A}$ is temporal if its domain is $\mathbb{Q}$ and its relations are definable by Boolean combinations of atomic formulas of the form $x<y$.
Theorem 6.1.
[9] Suppose $\mathbb{A}$ is a temporal relational structure. Then $\operatorname{CSP}(\Gamma)$ is solvable in polynomial time, or NP-complete.
Even though this classification has been known for many years, only recently Mottet [36] discovered a new uniform algorithm for the tractable cases, based on the following theorem. For a temporal relational structure $\mathbb{A}=(\mathbb{Q};R_{1},\dots,R_{s})$ we denote by $\mathbb{A}|_{[N]}$ its restriction to the set $[N]$, that is, the structure $([N];R_{1}\cap[N]^{\operatorname{ar}(R_{1})},\dots,R_{s}\cap[N]^{\operatorname{ar}(R_{s})})$.
Theorem 6.2 (Theorem 3 in [36]).
Let $\mathbb{A}$ be a temporal structure, $N\in\mathbb{N}$. Then one of the following holds:
  • $\operatorname{PCSP}(\mathbb{A}|_{[N]},\mathbb{A})$ is solvable by local consistency or $\mathrm{Singl}\mathrm{AIP}$,
  • $\operatorname{CSP}(\mathbb{A})$ is NP-hard.
This gives a very simple algorithm for $\operatorname{CSP}(\mathbb{A})$ whenever it is tractable. In fact, for any instance $\mathbb{X}$ of $\operatorname{CSP}(\mathbb{A})$ with $|X|=N$, the existence of a homomorphism $\mathbb{X}\to\mathbb{A}|_{[N]}$ is equivalent to existence of a homomorphism $\mathbb{X}\to\mathbb{A}$. The idea of going from a temporal CSP instance to a finite one by restricting its domain first appeared in [10] under the name “sampling”. Hence, by Theorem 6.2, it is sufficient to run the consistency algorithm and $\mathrm{Singl}\mathrm{AIP}$ on $\mathbb{X}$ as an instance of $\operatorname{PCSP}(\mathbb{A}|_{[N]},\mathbb{A})$.
The proof in [36] still uses a lot of knowledge about temporal CSPs. We show that the soundness of the corresponding algorithms can be witnessed by concrete (palette) symmetric functions from $\operatorname{Pol}(\mathbb{A}|_{[N]},\mathbb{A})$.
Let us define several relational structures that are the most expressible among all temporal structures whose CSP is tractable. Put
$\mathbb{D}$ $=(\mathbb{Q};x\neq y,y\leq x\vee z<x,y\neq x\vee z\leq x),$
$\mathbb{E}$ $=(\mathbb{Q};x=y<z\vee x=z<y\vee y=z<x),$
$\mathbb{F}$ $=(\mathbb{Q};x\neq y,x=y\Rightarrow u=v,x>y\vee x>z\vee x=y=z).$
Thus, $\mathbb{D}$ has one binary and two ternary relations, and $\mathbb{E}$ has only one ternary. For a relation $R$ by $R^{dual}$ we denote the relation $\{(a_{1},\dots,a_{\operatorname{ar}(R)})\mid(-a_{1},\dots,-a_{\operatorname{ar}(R)})\in R\}$. For a temporal structure $\mathbb{A}$ by $\mathbb{A}^{dual}$ we denote the structure obtained by replacing each relation $R$ with its dual relation $R^{dual}$. We have the following classification of tractable temporal CSPs:
Theorem 6.3 ([12]).
Suppose $\mathbb{A}$ is a temporal relational structure such that $\operatorname{CSP}(\mathbb{A})$ is not NP-hard. Then one of the following holds:
  • (c)
    $\operatorname{Pol}(\mathbb{A})$ contains a constant operation;
  • (min)
    $\min\in\operatorname{Pol}(\mathbb{A})$ or $\max\in\operatorname{Pol}(\mathbb{A})$;
  • (mi)
    $\mathbb{D}$ or $\mathbb{D}^{dual}$ $p$-defines $\mathbb{A}$;
  • (mx)
    $\mathbb{E}$ or $\mathbb{E}^{dual}$ $p$-defines $\mathbb{A}$;
  • ($l$)
    $\mathbb{F}$ or $\mathbb{F}^{dual}$ $p$-defines $\mathbb{A}$.
6.2 Palette functions
In this subsection we build concrete palette polymorphisms for $\operatorname{Pol}(\mathbb{D}|_{[N]},\mathbb{D})$ and $\operatorname{Pol}(\mathbb{E}|_{[N]},\mathbb{E})$, and prove that $\operatorname{Pol}(\mathbb{F}|_{[N]},\mathbb{F})$ does not contain one.
Lemma 6.4.
$\operatorname{Pol}(\mathbb{D}|_{[N]},\mathbb{D})$ contains an $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette totally symmetric function for all $N,\ell_{1},\dots, n\ell_{n}\in\mathbb{N}$.
Proof.
Put $M=\ell_{1}+\dots+\ell_{n}$. For $\mathbf{a}\in\mathbb{Q}^{M}$ by $\phi(\mathbf{a})$ denote the tuple $\mathbf{b}\in\{0,1\}^{M}$ such that $\mathbf{b}^{i}=\begin{cases}0,&\text{if $\mathbf{a}^{i}=\min_{j}\mathbf{a}^{j}$}\\ 1,&\text{otherwise}\end{cases}.$ Choose any injective monotone mapping $h\colon\{0,1\}^{M}\to\mathbb{Q}\cap(0,1)$. For a block tuple $(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$ by $\mathcal{F}(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$ we denote the minimal $k\in[n]$ such that $\mathbf{a}_{k}=(\min_{i,j}\mathbf{a}_{i}^{j},\dots,\min_{i,j}\mathbf{a}_{i}^{j})$. If no such $k$ exists, set $\mathcal{F}(\mathbf{a}_{1},\dots,\mathbf{a}_{n})=\infty$. Now define an $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette totally symmetric function by
[…]
Notice that any $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette tuple $(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$ contains a block $\mathbf{a}_{i}$ consisting only of the minimal element. Hence $\mathcal{F}(\mathbf{a}_{1},\dots,\mathbf{a}_{n})<\infty$, which implies that $f$ is $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette totally symmetric.
Let us show that $f\in\operatorname{Pol}(\mathbb{D}|_{[N]},\mathbb{D})$:
$x\neq y$. Consider two block tuples $\mathbf{a}=(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$ and $\mathbf{b}=(\mathbf{b}_{1},\dots,\mathbf{b}_{n})$ such that $\mathbf{a}_{i}^{j}\neq\mathbf{b}_{i}^{j}$ for all $i,j$. We need to show that $f(\mathbf{a})\neq f(\mathbf{b})$. If $\min_{i,j}\mathbf{a}_{i}^{j}\neq\min_{i,j}\mathbf{b}_{i}^{j}$ or $\mathcal{F}(\mathbf{a})\neq\mathcal{F}(\mathbf{b})$, then it follows immediately from the definition of $f$. If $\mathcal{F}(\mathbf{a})=\mathcal{F}(\mathbf{b})=\infty$, then we use injectivity of $h$ and the fact that $\phi(\mathbf{a})\neq\phi(\mathbf{b})$. It remains to notice that the case $\mathcal{F}(\mathbf{a})=\mathcal{F}(\mathbf{b})<\infty$ cannot happen.
$y\leq x\vee z<x$. Consider three block tuples $\mathbf{a}=(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$, $\mathbf{b}=(\mathbf{b}_{1},\dots,\mathbf{b}_{n})$, and $\mathbf{c}=(\mathbf{c}_{1},\dots,\mathbf{c}_{n})$ such that $\mathbf{b}_{i}^{j}\leq\mathbf{a}_{i}^{j}\vee\mathbf{c}_{i}^{j}<\mathbf{a}_{i}^{j}$ for all $i,j$. We need to prove that $f(\mathbf{b})\leq f(\mathbf{a})\vee f(\mathbf{c})<f(\mathbf{a})$. If $\mathcal{F}(\mathbf{a})<\infty$, then $\min_{i,j}\mathbf{b}_{i}^{j}<\min_{i,j}\mathbf{a}_{i}^{j}$, or $\min_{i,j}\mathbf{c}_{i}^{j}<\min_{i,j}\mathbf{a}_{i}^{j}$, or $\mathcal{F}(\mathbf{b})\leq\mathcal{F}(\mathbf{a})$ and $\min_{i,j}\mathbf{b}_{i}^{j}=\min_{i,j}\mathbf{a}_{i}^{j}$. In each case the required condition follows. If $\mathcal{F}(\mathbf{a})=\infty$, then $\min_{i,j}\mathbf{b}_{i}^{j}<\min_{i,j}\mathbf{a}_{i}^{j}$, or $\min_{i,j}\mathbf{c}_{i}^{j}<\min_{i,j}\mathbf{a}_{i}^{j}$, or $\phi(\mathbf{b})\leq\phi(\mathbf{a})$ and $\min_{i,j}\mathbf{b}_{i}^{j}=\min_{i,j}\mathbf{a}_{i}^{j}$. It remains to use the fact that $h$ is monotone.
$y\neq x\vee z\leq x$. Consider three block tuples $\mathbf{a}=(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$, $\mathbf{b}=(\mathbf{b}_{1},\dots,\mathbf{b}_{n})$, and $\mathbf{c}=(\mathbf{c}_{1},\dots,\mathbf{c}_{n})$ such that $\mathbf{b}_{i}^{j}\neq\mathbf{a}_{i}^{j}\vee\mathbf{c}_{i}^{j}\leq\mathbf{a}_{i}^{j}$ for all $i,j$. We need to prove that $f(\mathbf{b})\neq f(\mathbf{a})\vee f(\mathbf{c})\leq f(\mathbf{a})$. If $\min_{i,j}\mathbf{c}_{i}^{j}<\min_{i,j}\mathbf{a}_{i}^{j}$, then $f(\mathbf{c})<f(\mathbf{a})$, so we are done. Thus, we assume that $\min_{i,j}\mathbf{c}_{i}^{j}\geq\min_{i,j}\mathbf{a}_{i}^{j}$. Additionally we assume that $f(\mathbf{a})=f(\mathbf{b})$. This implies that $\min_{i,j}\mathbf{a}_{i}^{j}=\min_{i,j}\mathbf{b}_{i}^{j}$ and $\mathcal{F}(\mathbf{a})=\mathcal{F}(\mathbf{b})$. If $\mathcal{F}(\mathbf{a})=\mathcal{F}(\mathbf{b})<\infty$, then $\min_{i,j}\mathbf{c}_{i}^{j}=\min_{i,j}\mathbf{a}_{i}^{j}$ and $\mathcal{F}(\mathbf{c})\leq\mathcal{F}(\mathbf{a})$, which implies the required condition. Otherwise, we have $\phi(\mathbf{a})=\phi(\mathbf{b})$. Therefore $\min_{i,j}\mathbf{c}_{i}^{j}=\min_{i,j}\mathbf{a}_{i}^{j}$, and $\mathcal{F}(\mathbf{c})<\infty$ or $\phi(\mathbf{c})\leq\phi(\mathbf{a})$. Since $h$ is monotone, we derive $f(\mathbf{c})\leq f(\mathbf{a})$. ∎
Using compactness argument we can prove the following corollary.
Corollary 6.5.
$\operatorname{Pol}(\mathbb{D})$ contains an $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette totally symmetric function for all $n,\ell_{1},\dots,\ell_{n}\in\mathbb{N}$.
Lemma 6.6.
Suppose $\mathbb{D}$ $p$-defines $\mathbb{A}$. Then $\operatorname{PCSP}(\mathbb{A}|_{[N]},\mathbb{A})$ can be solved by $\mathrm{Singl}\mathrm{AC}$ for all $N\in\mathbb{N}$.
Proof.
By Corollary 6.5, $\operatorname{Pol}(\mathbb{D})$ contains an $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette totally symmetric function for all $n,\ell_{1},\dots,\ell_{n}\in\mathbb{N}$. Therefore, $\operatorname{Pol}(\mathbb{A})$ and $\operatorname{Pol}(\mathbb{A}|_{[N]},\mathbb{A})$ also contain them. By Lemma 5.8, $\operatorname{PCSP}(\mathbb{A}|_{[N]},\mathbb{A})$ can be solved by $\mathrm{Singl}\mathrm{AC}$. ∎
Lemma 6.7.
$\operatorname{Pol}(\mathbb{E}|_{[N]},\mathbb{E})$ contains a $(\overline{2\ell_{1}+1},\dots,\overline{2\ell_{n}+1})$-palette alternating function for all $N,\ell_{1}, \dots, n\ell_{n}\in\mathbb{N}$.
Proof.
Put $M=(2\ell_{1}+1)+\dots+(2\ell_{n}+1)$. For a block tuple $(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$ by $\mathcal{S}(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$ we denote the minimal $k\in[n]$ such that $\mathbf{a}_{k}$ contains an odd number of elements equal to $\min_{i,j}\mathbf{a}_{i}^{j}$. If no such $k$ exists, set $\mathcal{S}(\mathbf{a}_{1},\dots,\mathbf{a}_{n})=\infty$. Similarly, by $\mathcal{P}(\mathbf{a})$ we denote the minimal $k$ such that $\mathbf{a}^{k}=\min_{i}\mathbf{a}^{i}$. Thus, $\mathcal{P}(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$ is the position $k\in[M]$ of the first element in the tuple equal to the minimal element.
A $(\overline{2\ell_{1}+1},\dots,\overline{2\ell_{n}+1})$-palette alternating function can be defined by
[…]
Notice that any $(\overline{2\ell_{1}+1},\dots,\overline{2\ell_{n}+1})$-palette tuple $(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$ contains a block $\mathbf{a}_{i}$ consisting only of the minimal element. Hence $\mathcal{S}(\mathbf{a}_{1},\dots,\mathbf{a}_{n})<\infty$, which implies that $f$ is $(\overline{2\ell_{1}+1},\dots,\overline{2\ell_{n}+1})$-palette alternating. Also notice that if $\min_{i,j}\mathbf{a}_{i}^{j}<\min_{i,j}\mathbf{b}_{i}^{j}$ then $f(\mathbf{a}_{1},\dots,\mathbf{a}_{n})<(\mathbf{b}_{1},\dots,\mathbf{b}_{n})$.
Let us show that $f\in\operatorname{Pol}(\mathbb{E}|_{[N]},\mathbb{E})$. Denote the only ternary relation of $\mathbb{E}$ by $R$. Consider three block tuples $\mathbf{a}=(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$, $\mathbf{b}=(\mathbf{b}_{1},\dots,\mathbf{b}_{n})$, and $\mathbf{c}=(\mathbf{c}_{1},\dots,\mathbf{c}_{n})$ such that $(\mathbf{a}_{i}^{j},\mathbf{b}_{i}^{j},\mathbf{c}_{i}^{j})\in R$ for all $i,j$. We need to prove that $(f(\mathbf{a}),f(\mathbf{b}),f(\mathbf{c}))\in R$. Let $m=\min\{\min_{i,j}\mathbf{a}_{i}^{j},\min_{i,j}\mathbf{b}_{i}^{j},\min_{i,j}\mathbf{c}_{i}^{j}\}$. If $m$ appears only in two tuples, say $\mathbf{a}$ and $\mathbf{b}$, then $\mathcal{S}(\mathbf{a})=\mathcal{S}(\mathbf{b})$ and $\mathcal{P}(\mathbf{a})=\mathcal{P}(\mathbf{b})$, which implies $f(\mathbf{a})=f(\mathbf{b})<f(\mathbf{c})$. Thus, assume that $m$ appears in each of $\mathbf{a},\mathbf{b},\mathbf{c}$.
Let $q=\min\{\mathcal{S}(\mathbf{a}),\mathcal{S}(\mathbf{b}),\mathcal{S}(\mathbf{c})\}$. If $q=\infty$ then exactly two of the numbers $\mathcal{P}(\mathbf{a}),\mathcal{P}(\mathbf{b}),\mathcal{P}(\mathbf{c})$ are equal and smaller then the remaining one, which implies $(f(\mathbf{a}),f(\mathbf{b}),f(\mathbf{c}))\in R$. If $q<\infty$, we may assume (by symmetry of $R$) that $\mathcal{S}(\mathbf{a})=q$. Thus, $\mathbf{a}_{q}$ contains an odd number of $m$. Since $(\mathbf{a}_{i}^{j},\mathbf{b}_{i}^{j},\mathbf{c}_{i}^{j})$ contains an even number of $m$ for all $i,j$, it follows that either $\mathbf{b}_{q}$ or $\mathbf{c}_{q}$ also contains an odd number of $m$. Hence one of $\mathcal{S}(\mathbf{b})$ or $\mathcal{S}(\mathbf{c})$ is equal to $q$, while the other is greater. This implies $(f(\mathbf{a}),f(\mathbf{b}),f(\mathbf{c}))\in R$. ∎
Corollary 6.8.
$\operatorname{Pol}(\mathbb{E})$ contains a $(\overline{2\ell_{1}+1},\dots,\overline{2\ell_{n}+1})$-palette alternating function for all $n,\ell_{1},\dots,\ell_{n}\in\mathbb{N}$.
Lemma 6.9.
Suppose $\mathbb{E}$ $p$-defines $\mathbb{A}$. Then $\operatorname{PCSP}(\mathbb{A}|_{[N]},\mathbb{A})$ can be solved by $\mathrm{Singl}\mathrm{AIP}$ for all $N\in\mathbb{N}$.
Proof.
By Corollary 6.8, $\operatorname{Pol}(\mathbb{E})$ contains a $(\overline{2\ell_{1}+1},\dots,\overline{2\ell_{n}+1})$-palette alternating function for all $n,\ell_{1},\dots,\ell_{n}\in\mathbb{N}$. Therefore, $\operatorname{Pol}(\mathbb{A})$ and $\operatorname{Pol}(\mathbb{A}|_{[N]},\mathbb{A})$ also contain them. Following the argument in Section 3.4 we can show that these functions can be used to get a (true) solution from the output of $\mathrm{Singl}\mathrm{AIP}$, which completes the proof. ∎
Lemma 6.10.
$\operatorname{Pol}(\mathbb{F}|_{[N]},\mathbb{F})$ does not contain an $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette symmetric function for any $n,\ell_{1}$, $\dots$, $\ell_{n},N\geq 4$.
Proof.
Assume the contrary. Suppose $f\in\operatorname{Pol}(\mathbb{F}|_{[N]},\mathbb{F})$ is an $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette symmetric function. For $a\in[N]$ and $\ell\in\mathbb{N}$, by $a^{\ell}$ we denote the $\ell$-tuple consisting of an element $a$. Choose 3 different elements $a,b,c\in[N]$, and tuples $\mathbf{d}_{i},\mathbf{e}_{i}\in\{a,b,c\}^{\ell_{i}}$ for each $i\in\{3,4,\dots,n\}$ such that $\mathbf{d}_{i}$ and $\mathbf{e}_{i}$ contain the same multiset of elements but $\mathbf{d}_{i}^{j}\neq\mathbf{e}_{i}^{j}$ for every $j\in[\ell_{i}]$. Since $f$ is $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette symmetric, $f(a^{\ell_{1}},b^{\ell_{2}},c^{\ell_{3}},\mathbf{d}_{3},\dots,\mathbf{d}_{n})=f(a^{\ell_{1}},b^{\ell_{2}},c^{\ell_{3}},\mathbf{e}_{3},\dots,\mathbf{e}_{n}).$ Since $f$ preserves the relation $x=y\Rightarrow u=v$, we obtain
$f(\mathbf{u}_{1},\mathbf{u}_{2},\mathbf{u}_{3},\mathbf{u}_{4},\dots,\mathbf{u}_{n})=f(\mathbf{u}_{1},\mathbf{u}_{2},\mathbf{u}_{3},\mathbf{v}_{4},\dots,\mathbf{v}_{n})$
for all $\mathbf{u}_{1},\dots,\mathbf{u}_{n},\mathbf{v}_{4},\dots,\mathbf{v}_{n}$. Since $n\geq 4$, the last $\ell_{n}$ coordinates of $f$ are dummy. By permuting the arguments, we conclude that all the coordinates are dummy, so $f$ must be a constant function. But a constant function does not preserve the relation $x\neq y$. This contradiction completes the proof. ∎
6.3 Singleton vs constraint-singleton
Here we show that for infinite relational structures, the algorithm $\mathrm{CSingl}\mathrm{AC}$ can be stronger than $\mathrm{Singl}\mathrm{AC}$, as witnessed by the structure $\mathbb{F}$.
Lemma 6.11.
$\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ does not solve $\operatorname{PCSP}(\mathbb{F}|_{[N]},\mathbb{F})$ for $N\geq 3$.
Proof.
Consider an instance $\mathcal{I}$ of $\operatorname{Pol}(\mathbb{F}|_{[N]},\mathbb{F})$ given by $(x=x\Rightarrow y=z)\wedge(y\neq z)$, which clearly has no solution. Let us show that $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ nevertheless returns “Yes” on this instance. If we restrict $x$ to some $c\in[N]$, then for a BLP solution of the first constraint we take a uniform distribution over all tuples $(c,c,a,a)$ where $a\in[N]$, and uniform distribution over tuples $(a,b)$ with $a\neq b$ for the second constraint. By Claim 4.2, AIP gives the same result if we replace the constraint $y\neq z$ by its parallelogram closure, which is the full relation. Therefore, $\mathrm{BLP}+\mathrm{AIP}$ returns “Yes” after this restriction.
If we restrict $y$ to some $c\in[N]$, then for the first constraint we take a uniform distribution over tuples $(a,b,c,d)$ with $a\neq b$, $c\neq d$, $a,b,d\in[N]$, and for the second constraint we take a uniform distribution over tuples $(c,d)$ with $c\neq d$ and $d\in[N]$. Again, by Claim 4.2, AIP gives the same result if we replace the first constraint by its parallelogram closure, which includes all tuples $(a,a,c,c)$ where $a\in[N]$. Hence, $\mathrm{BLP}+\mathrm{AIP}$ also returns “Yes” after this restriction.
Thus, $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ never restricts any domain and therefore returns “Yes” on this instance. ∎
To show that $\operatorname{PCSP}(\mathbb{F}|_{[N]},\mathbb{F})$ can be solved by $\mathrm{CSingl}\mathrm{AC}$, we need a palette version of a polymorphism. Let $n,k_{1},\dots,k_{n}\in\mathbb{N}$ with $k_{1}+\dots+k_{n}=m$. A function $f\colon A^{m}\to B$ is called a $(\overline{k_{1}},\dots,\overline{k_{n}})$-palette polymorphism from a $\sigma$-structure $\mathbb{A}$ to a $\sigma$-structure $\mathbb{B}$ if, for each symbol $R\in\sigma$ and each $(\overline{k_{1}},\dots,\overline{k_{n}})$-palette tuple $(\mathbf{a}_{1},\dots,\mathbf{a}_{m})\in(R^{\mathbb{A}})^{m}$, we have $f(\mathbf{a}_{1},\dots,\mathbf{a}_{m})\in R^{\mathbb{B}}$. By $(\overline{k_{1}},\dots,\overline{k_{n}})$-$\operatorname{Pol}(\mathbb{A},\mathbb{B})$ we denote the set of all $(\overline{k_{1}},\dots,\overline{k_{n}})$-palette polymorphism from $\mathbb{A}$ to $\mathbb{B}$. Informally, for functions from $(\overline{k_{1}},\dots,\overline{k_{n}})$-$\operatorname{Pol}(\mathbb{A},\mathbb{B})$ we only require the resulting tuple to belong to the relation if each tuple that appears fills at least one block completely.
Lemma 6.12.
Suppose $\mathbb{A}$ $p$-defines $\mathbb{B}$ and $f\in(\overline{k_{1}},\dots,\overline{k_{n}})$-$\operatorname{Pol}(\mathbb{A})$. Then $f\in(\overline{k_{1}},\dots,\overline{k_{n}})$-$\operatorname{Pol}(\mathbb{B})$.
Proof.
Let $m=k_{1}+\dots+k_{n}$. The proof follows almost word for word the standard argument showing that $f\in\operatorname{Pol}(\mathbb{A})\Rightarrow f\in\operatorname{Pol}(\mathbb{B})$. Here, we illustrate why this holds in the case where $\mathbb{A}$ has two binary relations $R_{1}$ and $R_{2}$, $\mathbb{B}$ has one relation $R$, and $R(x,y)=\exists zR_{1}(x,z)\wedge R_{2}(y,z)$. Choose a mapping $\phi\colon R\to A$ such that $(a_{1},\phi(a_{1},a_{2}))\in R_{1}$ and $(a_{2},\phi(a_{1},a_{2}))\in R_{2}$ for all $(a_{1},a_{2})\in R$. Let $(\mathbf{a}_{1},\dots,\mathbf{a}_{m})\in(R)^{m}$ be a $(\overline{k_{1}},\dots,\overline{k_{n}})$-palette tuple. Set $\mathbf{b}_{i}=(\mathbf{a}_{i}^{1},\phi(\mathbf{a}))$, $\mathbf{c}_{i}=(\mathbf{a}_{i}^{2},\phi(\mathbf{a}))$ for every $i\in[m]$. Then $(\mathbf{b}_{1},\dots,\mathbf{b}_{m})\in(R_{1})^{m}$ and $(\mathbf{c}_{1},\dots,\mathbf{c}_{m})\in(R_{2})^{m}$ are $(\overline{k_{1}},\dots,\overline{k_{n}})$-palette tuples. Let $f(\mathbf{b}_{1},\dots,\mathbf{b}_{m})=(b,d)$ and $f(\mathbf{c}_{1},\dots,\mathbf{c}_{m})=(c,d)$. Since $f\in(\overline{k_{1}},\dots,\overline{k_{n}})$-$\operatorname{Pol}(\mathbb{A})$, $(b,d)\in R_{1}$ and $(c,d)\in R_{2}$. Therefore $f(\mathbf{a}_{1},\dots,\mathbf{a}_{m})=(b,c)\in R$. ∎
Lemma 6.13.
Suppose $\mathbb{A}$ and $\mathbb{B}$ are relational structures over the same signature $\sigma$, with $A$ finite, and suppose that for all $n,k_{1},\dots,k_{n}\in\mathbb{N}$ there exists a $(\overline{k_{1}},\dots,\overline{k_{n}})$-palette totally symmetric function $f\in(\overline{k_{1}},\dots,\overline{k_{n}})$-$\operatorname{Pol}(\mathbb{A},\mathbb{B})$. Then $\mathrm{CSingl}\mathrm{AC}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$.
Proof.
The argument follows the proof of Lemma 5.8 with one modification. As before, notice that by identifying variables within the blocks of a $(\overline{k_{1}},\dots,\overline{k_{n}})$-palette totally symmetric (PTS) function we obtain a PTS function with smaller block sizes. Therefore, for all $L_{1},\dots,L_{n}\in\mathbb{N}$ we can choose an $(\overline{L_{1}},\dots,\overline{L_{n}})$-PTS function $f_{L_{1},\dots,L_{n}}\in(\overline{L_{1}},\dots,\overline{L_{n}})$-$\operatorname{Pol}(\mathbb{A},\mathbb{B})$ such that identifying variables within blocks of $f_{L_{1},\dots,L_{n}}$ yields $f_{\ell_{1},\dots,\ell_{n}}$, where $\ell_{i}\leq L_{i}$ for every $i$. We write $f(S_{1},\dots,S_{n})$ meaning $f_{k_{1},\dots,k_{n}}(s_{1}^{1},\dots,s_{1}^{k_{1}},\dots,s_{n}^{1},\dots,s_{n}^{k_{n}})$, where $S_{i}=\{s_{i}^{1},\dots,s_{i}^{k_{i}}\}$ for every $i\in[n]$. This notation is justified because we only apply $f_{k_{1},\dots,k_{n}}$ to palette tuples, so the order of elements in the sets $S_{i}$ does not matter.
We need to check that if $\mathrm{CSingl}\mathrm{AC}$ returns “Yes” on an instance $\mathcal{I}$, then the instance has a solution in $\mathbb{B}$. Let $P_{1},\dots,P_{S}$ be all pairs $(C,\mathbf{a})$, where $C$ is a constraint of $\mathcal{I}$ and $\mathbf{a}$ is a tuple from the reduced constraint relation at the end of the algorithm. For each $P_{j}=(C,\mathbf{a})$ and each constraint $C_{0}=R_{0}(x_{1},\dots,x_{k})$, let $R_{0}^{j}$ be the reduced constraint relation of $C_{0}$ after restricting $C$ to $\mathbf{a}$ and running arc-consistency. Then every constraint $C_{0}$ can be consistently evaluated to the tuple $f(R_{0}^{1},\dots,R_{0}^{S})\in R_{0}^{\mathbb{B}}$. Moreover, these evaluations agree on shared variables across all constraints because the function $f$ is palette totally symmetric. This confirms that the instance has a solution in $\mathbb{B}$. ∎
Lemma 6.14.
$(\overline{k_{1}},\dots,\overline{k_{n}})$-$\operatorname{Pol}(\mathbb{F}|_{[N]},\mathbb{F})$ contains a block totally symmetric function for all $n,N,k_{1},\dots$, $k_{n}\in\mathbb{N}$.
Proof.
Let us define a block totally symmetric function $f\in(\overline{k_{1}},\dots,\overline{k_{n}})$-$\operatorname{Pol}(\mathbb{F}|_{[N]},\mathbb{F})$ by
$f(\mathbf{a}_{1},\dots,\mathbf{a}_{n})=\sum\limits_{i\in[n]}(\min_{j}\mathbf{a}_{i}^{j})/(N+1)^{i}.$
Notice that the function $g\colon[N]^{n}\to\mathbb{Q}$ defined by $g(x_{1},\dots,x_{n})=\sum\limits_{i\in[n]}x_{i}/(N+1)^{i}$ preserves the lexicographic order, that is, if $a_{1}=b_{1},\dots,a_{j-1}=b_{j-1},a_{j}<b_{j}$, then $g(a_{1},\dots,a_{n})<g(b_{1},\dots,b_{n})$. This follows from
$(b_{j}-a_{j})\geq 1=N\cdot\sum\limits_{i=1}^{\infty}\frac{1}{(N+1)^{i}}>N\cdot\sum\limits_{i=1}^{n-j-1}\frac{1}{(N+1)^{i}}>(b_{j+i}-a_{j+i})\cdot\sum\limits_{i=1}^{n-j-1}\frac{1}{(N+1)^{i}}.$
It follows from the definition that $f$ depends only on the set of elements in each block. Hence, $f$ is block totally symmetric. Set $M=k_{1}+\dots+k_{n}$. Let us show that $f\in(\overline{k_{1}},\dots,\overline{k_{n}})$-$\operatorname{Pol}(\mathbb{F}|_{[N]},\mathbb{F})$.
$x\neq y$. Consider two block tuples $\mathbf{a}=(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$ and $\mathbf{b}=(\mathbf{b}_{1},\dots,\mathbf{b}_{n})$ such that $\mathbf{a}_{i}^{j}\neq\mathbf{b}_{i}^{j}$ for all $i,j$. Assume moreover that when we form the $M$-tuple whose entries (from $N]^{2}$) are taken from $\mathbf{a}$ and $\mathbf{b}$, we obtain a $(\overline{k_{1}},\dots,\overline{k_{n}})$-palette tuple. We need to show that $f(\mathbf{a})\neq f(\mathbf{b})$. Because of the palette property there exists $i\in[n]$ such that $\mathbf{a}_{i}^{j}=\mathbf{a}_{1}^{1}$ and $\mathbf{b}_{i}^{j}=\mathbf{b}_{1}^{1}$ for every $j\in[k_{i}]$. Since $\mathbf{a}_{1}^{1}\neq\mathbf{b}_{1}^{1}$, $\mathbf{a}$ and $\mathbf{b}$ have two different constant blocks. Hence $f(\mathbf{a})\neq f(\mathbf{b})$.
$x=y\Rightarrow u=v$. Consider four block tuples $\mathbf{a}=(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$, $\mathbf{b}=(\mathbf{b}_{1},\dots,\mathbf{b}_{n})$, $\mathbf{c}=(\mathbf{c}_{1},\dots,\mathbf{c}_{n})$, and $\mathbf{d}=(\mathbf{d}_{1},\dots,\mathbf{d}_{n})$ such that $\mathbf{a}_{i}^{j}=\mathbf{b}_{i}^{j}\Rightarrow\mathbf{c}_{i}^{j}=\mathbf{d}_{i}^{j}$ for all $i,j$. Assume moreover that when we form an $M$-tuple whose entries (from $N]^{4}$) are taken from $\mathbf{a}$, $\mathbf{b}$$\mathbf{c}$, and $\mathbf{d}$, we obtain a $(\overline{k_{1}},\dots,\overline{k_{n}})$-palette tuple. We need to prove that $f(\mathbf{a})=f(\mathbf{b})\Rightarrow f(\mathbf{c})=f(\mathbf{d})$. Consider two cases. Assume that $\mathbf{a}=\mathbf{b}$. Then $\mathbf{c}=\mathbf{d}$, and, therefore, $f(\mathbf{a})=f(\mathbf{b})$ and $f(\mathbf{c})=f(\mathbf{d})$, which completes this case. Assume that $\mathbf{a}_{i}^{j}\neq\mathbf{b}_{i}^{j}$ for some $i,j$. By the palette property, there exists $m\in[n]$ such that $\mathbf{a}_{m}^{\ell}=\mathbf{a}_{i}^{j}$ and $\mathbf{b}_{m}^{\ell}=\mathbf{b}_{i}^{j}$ for all $\ell\in[k_{m}]$. Thus, $\mathbf{a}$ and $\mathbf{b}$ have two different constant blocks. Hence, $f(\mathbf{a})\neq f(\mathbf{b})$.
$x>y\vee x>z\vee x=y=z$. Consider three block tuples $\mathbf{a}=(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$, $\mathbf{b}=(\mathbf{b}_{1},\dots,\mathbf{b}_{n})$, and $\mathbf{c}=(\mathbf{c}_{1},\dots,\mathbf{c}_{n})$ such that $\mathbf{a}_{i}^{j}>\mathbf{b}_{i}^{j}\vee\mathbf{a}_{i}^{j}>\mathbf{c}_{i}^{j}\vee\mathbf{a}_{i}^{j}=\mathbf{b}_{i}^{j}=\mathbf{c}_{i}^{j}$ for all $i,j$. It follows from the definition that for every $i\in[n]$ either $\min_{j}\mathbf{a}_{i}^{j}=\min_{j}\mathbf{b}_{i}^{j}=\min_{j}\mathbf{c}_{i}^{j}$, or $\min_{j}\mathbf{a}_{i}^{j}>\min_{j}\mathbf{b}_{i}^{j}$, or $\min_{j}\mathbf{a}_{i}^{j}>\min_{j}\mathbf{c}_{i}^{j}$. Since $g$ preserves the lexicographic order, we obtain $f(\mathbf{a})>f(\mathbf{b})\vee f(\mathbf{a})>f(\mathbf{c})\vee f(\mathbf{a})=f(\mathbf{b})=f(\mathbf{c})$. ∎
Corollary 6.15.
$(\overline{k_{1}},\dots,\overline{k_{n}})$-$\operatorname{Pol}(\mathbb{F})$ contains a block totally symmetric function for all $n,k_{1},\dots,k_{n}\in\mathbb{N}$.
Lemma 6.16.
Suppose $\mathbb{F}$ $p$-defines $\mathbb{A}$. Then $\operatorname{PCSP}(\mathbb{A}|_{[N]},\mathbb{A})$ can be solved by $\mathrm{CSingl}\mathrm{AC}$ for all $N\in\mathbb{N}$.
Proof.
By Corollary 6.15, $(\overline{k_{1}},\dots,\overline{k_{n}})$-$\operatorname{Pol}(\mathbb{F})$ contains a block totally symmetric function $f$ for every $n\in\mathbb{N}$ and every $k_{1},\dots,k_{n}\in\mathbb{N}$. By Lemma 6.12, each $(\overline{k_{1}},\dots,\overline{k_{n}})$-$\operatorname{Pol}(\mathbb{A})$, also contains it. Then Lemma 6.13 implies that $\operatorname{PCSP}(\mathbb{A}|_{[N]},\mathbb{A})$ can be solved by $\mathrm{CSingl}\mathrm{AC}$. ∎
6.4 Conclusions
Following the argument in Section 3.1 we conclude that if $\min\in\operatorname{Pol}(\mathbb{A})$, then $\operatorname{PCSP}(\mathbb{A}|_{[N]},\mathbb{A})$ can be solved by the arc-consistency algorithm, and a (true) solution can be obtained from its output by applying the totally symmetric function $\min$. Similarly, following the argument in Section 3.4 we can use an $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette totally symmetric function and a $(\overline{2\ell_{1}+1},\dots,\overline{2\ell_{n}+1})$-palette alternating function to obtain a (true) solution from the output of the algorithms $\mathrm{Singl}\mathrm{AC}$ and $\mathrm{Singl}\mathrm{AIP}$, respectively. Therefore, $\operatorname{PCSP}(\mathbb{D}|_{[N]},\mathbb{D})$ and $\operatorname{PCSP}(\mathbb{E}|_{[N]},\mathbb{E})$ can be solved by $\mathrm{Singl}\mathrm{AC}$ and $\mathrm{Singl}\mathrm{AIP}$, respectively.
On the other hand, $\operatorname{Pol}(\mathbb{F}|_{[N]},\mathbb{F})$ does not have palette totally symmetric functions, and even $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ fails to solve $\operatorname{PCSP}(\mathbb{F}|_{[N]},\mathbb{F})$. Nevertheless, one can define a palette polymorphism that witnesses that $\mathrm{CSingl}\mathrm{AC}$ solves $\operatorname{PCSP}(\mathbb{F}|_{[N]},\mathbb{F})$. This result confirms that $\mathrm{Singl}$- and $\mathrm{CSingl}$- algorithms have the same power only for finite relational structures, and highlights that the Hales-Jewett theorem is essential for proving the characterization in terms of palette symmetric functions.
We now summarize the results of this section and present the algorithm for each case from Theorem 6.3.
  • (1)
    $\operatorname{Pol}(\mathbb{A})$ contains a constant operation;
    $\operatorname{CSP}(\mathbb{A})$ is trivial: every instance has a solution.
  • (2)
    $\min\in\operatorname{Pol}(\mathbb{A})$ or $\max\in\operatorname{Pol}(\mathbb{A})$;
    $\operatorname{Pol}(\mathbb{A})$ has a totally symmetric function of any arity;
    $\operatorname{PCSP}(\mathbb{A}|_{[N]},\mathbb{A})$ can be solved by $\mathrm{AC}$.
  • (3)
    $\mathbb{D}$ or $\mathbb{D}^{dual}$ $p$-defines $\mathbb{A}$;
    $\operatorname{Pol}(\mathbb{D})$ contains an $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette totally symmetric function for all $n,\ell_{1},\dots,\ell_{n}\in\mathbb{N}$;
    $\operatorname{PCSP}(\mathbb{A}|_{[N]},\mathbb{A})$ can be solved by $\mathrm{Singl}\mathrm{AC}$.
  • (4)
    $\mathbb{E}$ or $\mathbb{E}^{dual}$ $p$-defines $\mathbb{A}$;
    $\operatorname{Pol}(\mathbb{E})$ contains a $(\overline{2\ell_{1}+1},\dots,\overline{2\ell_{n}+1})$-palette alternating function for all $n,\ell_{1},\dots,\ell_{n}\in\mathbb{N}$;
    $\operatorname{PCSP}(\mathbb{A}|_{[N]},\mathbb{A})$ can be solved by $\mathrm{Singl}\mathrm{AIP}$.
  • (5)
    $\mathbb{F}$ or $\mathbb{F}^{dual}$ $p$-defines $\mathbb{A}$.
    $\operatorname{Pol}(\mathbb{F}|_{[N]},\mathbb{F})$ does not contain an $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette symmetric function for any $n,\ell_{1},\dots,\ell_{n},N\geq 4$;
    $(\overline{k_{1}},\dots,\overline{k_{n}})$-$\operatorname{Pol}(\mathbb{F})$ contains a block totally symmetric function for all $n,k_{1},\dots,k_{n}\in\mathbb{N}$;
    $\operatorname{PCSP}(\mathbb{A}|_{[N]},\mathbb{A})$ can be solved by $\mathrm{CSingl}\mathrm{AC}$, but $\operatorname{PCSP}(\mathbb{F}|_{[N]},\mathbb{F})$ cannot be solved by $\mathrm{Singl}\mathrm{AC}$ or even $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$.
Combining the above results with Theorem 6.3 we obtain the following theorem from Section 1.3.
Theorem 1.3.
Let $\mathbb{A}$ be a temporal structure, $N\in\mathbb{N}$. Then one of the following holds:
  • $\operatorname{PCSP}(\mathbb{A}|_{[N]},\mathbb{A})$ is solvable by $\mathrm{CSingl}\mathrm{AC}\wedge\mathrm{Singl}\mathrm{AIP}$;
  • $\operatorname{CSP}(\mathbb{A})$ is NP-hard.
Remark 6.1.
Unlike polymorphisms of $\mathbb{D}$ and $\mathbb{E}$, which come from compactness arguments and rely on the axiom of choice, the polymorphisms constructed in Lemmas 6.4, 6.7, and 6.14 are concrete and can be used to obtain a (true) solution from the outputs of $\mathrm{CSingl}\mathrm{AC}$ and $\mathrm{Singl}\mathrm{AIP}$.
7 Dihedral group $\mathbf{D}_{4}$.
In this section, we prove two properties of the dihedral group $\mathbf{D}_{4}$ that were stated in Section 4.4. Recall that $\mathbb{D}_{4}=(\mathbb{Z}_{2}^{3};L_{3}^{1,2},L_{2,3}^{1},L_{1,3}^{2},E_{1-2},E_{2-3}^{0},E_{1-3}^{0},R,\{(0,0,0)\})$. In this section addition and multiplication are always modulo 2, but $x_{i}^{2}$ denotes the second component of the vector $x_{i}$, not the second power. The following claims explain the meaning of some of the relations in $\mathbb{D}_{4}$. We assume that $f$ is an $n$-ary operation on $\mathbb{Z}_{2}^{3}$.
Claim 7.1.
$f\in\operatorname{Pol}(\{L_{3}^{1,2}\})$ if and only if there exist $g_{1},g_{2},g_{3}:\mathbb{Z}_{2}^{2n}\to\mathbb{Z}_{2}$ and $a_{1},\dots,a_{n}\in\mathbb{Z}_{2}$ such that
$f^{1}(x_{1},\dots,x_{n})$ $=g_{1}(x_{1}^{1},\dots,x_{n}^{1},x_{1}^{2},\dots,x_{n}^{2}),$
$f^{2}(x_{1},\dots,x_{n})$ $=g_{2}(x_{1}^{1},\dots,x_{n}^{1},x_{1}^{2},\dots,x_{n}^{2}),$
$f^{3}(x_{1},\dots,x_{n})$ $=a_{1}x_{1}^{3}+\dots+a_{n}x_{n}^{3}+g_{3}(x_{1}^{1},\dots,x_{n}^{1},x_{1}^{2},\dots,x_{n}^{2}).$
Proof.
Let $E=\operatorname{pr}_{1,2}(L_{3}^{1,2})$. Then $E$ is an equivalence relation and it is preserved by $f$. This implies that the first two components of $f$ depend only on the first two components of the inputs, so we can choose the corresponding $g_{1}$ and $g_{2}$. Applying $f$ to the tuples $((x_{i}^{1},x_{i}^{2},x_{i}^{3}),(x_{i}^{1},x_{i}^{2},0),(0,0,x_{i}^{3}),(0,0,0))\in L_{3}^{1,2}$ for $i=1,\dots,n$ and looking at the third component, we obtain
$f^{3}(x_{1},\dots,x_{n})=f^{3}((x_{1}^{1},x_{1}^{2},0),\dots,(x_{n}^{1},x_{n}^{2},0))+f^{3}((0,0,x_{1}^{3}),\dots,(0,0,x_{n}^{3}))+f^{3}((0,0,0),\dots,(0,0,0)).$ (1)
Similarly, we derive
$f^{3}((0,0,x_{1}^{3}),\dots,(0,0,x_{n}^{3}))=$
$f^{3}((0,0,0),(0,0,x_{2}^{3}),\dots,$ $(0,0,x_{n}^{3}))+f^{3}((0,0,x_{1}^{3}),(0,0,0),\dots,(0,0,0))+f^{3}((0,0,0),\dots,(0,0,0)).$
This holds analogously when considering the $i$-th coordinate instead of the first. Choose $a_{1},\dots,a_{n},b_{1},\dots,b_{n}\in\mathbb{Z}_{2}$ so that
$f^{3}(\underbrace{(0,0,0),\dots,(0,0,0),(0,0,z)}_{i},(0,0,0),\dots,(0,0,0))=a_{i}z+b_{i}\text{\quad\quad\quad for every $i\in[n]$.}$
Using (1), we derive that $f^{3}(x_{1},\dots,x_{n})=a_{1}x_{1}^{3}+\dots+a_{n}x_{n}^{3}+g_{3}(x_{1}^{1},\dots,x_{n}^{1},x_{1}^{2},\dots,x_{n}^{2})$ for some $g_{3}$. ∎
Claim 7.2.
$f\in\operatorname{Pol}(\{E_{1-2}\})$ if and only if there exists $g:\mathbb{Z}_{2}^{n}\to Z_{2}$ such that
$f^{1}(x_{1},\dots,x_{n})$ $=g(x_{1}^{1},\dots,x_{n}^{1}),$
$f^{2}(x_{1},\dots,x_{n})$ $=g(x_{1}^{2},\dots,x_{n}^{2}).$
Proof.
Applying $f$ to tuples $((a_{i},b_{i},c_{i}),(d_{i},a_{i},e_{i})\in E_{1-2}$ for $i=1,\dots,n$, and using the fact that the result must be in $E_{1-2}$, we obtain
$f^{1}((a_{1},b_{1},c_{1}),\dots,(a_{n},b_{n},c_{n}))=f^{2}((d_{1},a_{1},e_{1}),\dots,(d_{n},a_{n},e_{n})).$
This implies that $f^{1}$ depends only on the first components, $f^{2}$ depends only on the second components, and the operations defining $f^{1}$ and $f^{2}$ are identical. ∎
Claim 7.3.
Suppose $f\in\operatorname{Pol}(\{L_{3}^{1,2},L_{2,3}^{1}\})$. Then every monomial in the polynomial representation of $g_{3}$ (see Claim 7.1) contains at most one $x_{i}^{2}$-variable (i.e, terms like $x_{1}^{2}\cdot x_{2}^{2}$ are not allowed).
Proof.
Assume the contrary. Choose monomials of the form $x_{i_{1}}^{1}\dots x_{i_{k}}^{1}x_{j_{1}}^{2}\dots x_{j_{m}}^{2}$ with minimal $m\geq 2$; among them choose one with minimal $k\geq 0$. Set $b_{i}=1$ if $i\in\{i_{1},\dots,i_{k}\}$ and $b_{i}=0$ otherwise; $c_{i}=1$ if $i\in\{j_{1},\dots,j_{m}\}$ and $c_{i}=0$ otherwise; $d_{i}=1$ if $i\in\{j_{2},\dots,j_{m}\}$ and $d_{i}=0$ otherwise; $e_{j_{1}}=1$ and $e_{i}=0$ for $i\neq j_{1}$. Applying $f$ to the tuples $((b_{i},c_{i},0),(b_{i},d_{i},0),(b_{i},e_{i},0),(b_{i},0,0))\in L_{2,3}^{1}$ for $i=1,\dots,n$, we get a tuple from $L_{2,3}^{1}$. Examining the third components of the obtained tuple, we derive:
$g_{3}(b_{1},\dots,b_{n},c_{1},\dots,c_{n})+g_{3}(b_{1},\dots,b_{n},d_{1},\dots,d_{n})+g_{3}(b_{1},\dots,b_{n},e_{1},\dots,e_{n})+g_{3}(b_{1},\dots,b_{n},0,\dots,0)=0.$
All the monomials of $g_{3}$ involving only the first components vanish, as do all linear monomials. The only monomial contributing 1 to the above sum is $x_{i_{1}}^{1}\dots x_{i_{k}}^{1}x_{j_{1}}^{2}\dots x_{j_{m}}^{2}$, which gives us a contradiction. ∎
The next claim differs from the previous one only by swapping 1 and 2, so we omit the proof.
Claim 7.4.
Suppose $f\in\operatorname{Pol}(\{L_{3}^{1,2},L_{1,3}^{2}\})$. Then every monomial in the polynomial representation of $g_{3}$ (see Claim 7.1) contains at most one $x_{i}^{1}$-variable (i.e, terms like $x_{1}^{1}\cdot x_{2}^{1}$ are not allowed).
Let us describe a set of operations $\mathcal{D}_{4}$, which we will later prove to be the clone generated by the dihedral group. Suppose $n\in\mathbb{N}$, $a:[n]\to\mathbb{Z}_{2}$, and $c:[n]\times[n]\to\mathbb{Z}_{2}$ are such that $c_{i,j}+c_{j,i}=a_{i}\cdot a_{j}$ for every $\in[n], iji\neq j$. Then we define the following $n$-ary operation $f$ on $\mathbb{Z}_{2}^{3}$:
$f^{1}(x_{1},\dots,x_{n})$ $=a_{1}x_{1}^{1}+\dots+a_{n}x_{n}^{1},$
$f^{2}(x_{1},\dots,x_{n})$ $=a_{1}x_{1}^{2}+\dots+a_{n}x_{n}^{2},$
$f^{3}(x_{1},\dots,x_{n})$ $=a_{1}x_{1}^{3}+\dots+a_{n}x_{n}^{3}+\sum\limits_{i,j\in[n]}c_{i,j}x_{i}^{1}x_{j}^{2}.$
By $\mathcal{D}_{4}$ we denote the set of all such operations $f$ for all $n$, $a$ and $c$.
Claim 7.5.
$\mathcal{D}_{4}\subseteq\operatorname{Clo}(x\circ y)$.
Proof.
Let $f$ be an $n$-ary operation from $\mathcal{D}_{4}$ defined by some mappings $a$ and $c$. Let $\{i_{1},\dots,i_{s}\}=\{i\mid a_{i}=1\}$ with $i_{1}<i_{2}<\dots<i_{s}$. Let us show that $f$ can be defined by
$f(x_{1},\dots,x_{n})=x_{i_{1}}\circ\dots\circ x_{i_{s}}\circ\prod_{i<j\colon c_{j,i}=1}(x_{i}\circ x_{j}\circ x_{i}^{-1}\circ x_{j}^{-1})\circ\prod_{i\colon c_{i,i}=1}(x_{i}\circ x_{i}).$
At the first two components the operation $\circ$ is just an addition; hence the last two products do not affect them, and the result matches the definition. We claim that the third component of $x_{i_{1}}\circ\dots\circ x_{i_{s}}$ is $x_{i_{1}}^{3}+\dots+x_{i_{s}}^{3}+\sum\limits_{j<k}x_{i_{j}}^{1}x_{i_{k}}^{2}$. This can be shown by induction on $s$, with the case $s=2$ following directly from the definition and the inductive step being straightforward. Then we check that $(x_{i}\circ x_{j}\circ x_{i}^{-1}\circ x_{j}^{-1})=(0,0,x_{i}^{1}x_{j}^{2}+x_{i}^{2}x_{j}^{1})$ and $(x_{i}\circ x_{i})=(0,0,x_{i}^{1}x_{i}^{2})$. It remains to apply the definition of $\circ$ to verify that the third component also matches the definition. ∎
Claim 7.6.
$\operatorname{Pol}(\mathbb{D}_{4})\subseteq\mathcal{D}_{4}$.
Proof.
Let $f\in\operatorname{Pol}(\mathbb{D}_{4})$. By Claim 7.1 we choose the corresponding $g_{1},g_{2},g_{3}$ and $a_{1},\dots,a_{n}$. By Claim 7.2 we can replace $g_{1}$ and $g_{2}$ with a single operation $g$. Since the relation $\{(0,0,0)\}$ is preserved by $f$, $g_{3}(0,0,\dots,0)=0$. Applying $f$ to tuples $((0,b_{1},c_{1}),(0,d_{1},b_{1})),\dots,((0,b_{n},c_{n}),(0,d_{n},b_{n}))\in E_{2-3}^{0}$, we obtain
$g(b_{1},\dots,b_{n})=a_{1}b_{1}+\dots+a_{n}b_{n}+g_{3}(0,0,\dots,0,d_{1},\dots,d_{n})$
for all $b_{1},\dots,b_{n},d_{1},\dots,d_{n}\in\mathbb{Z}_{2}$. Similarly, using the relation $E_{1-3}^{0}$, we obtain
$g(b_{1},\dots,b_{n})=a_{1}b_{1}+\dots+a_{n}b_{n}+g_{3}(d_{1},\dots,d_{n},0,\dots,0)$
for all $b_{1},\dots,b_{n},d_{1},\dots,d_{n}\in\mathbb{Z}_{2}$. These two conditions imply that $g(z_{1},\dots,z_{n})=a_{1}z_{1}+\dots+a_{n}z_{n}$ and $g_{3}$ returns 0 whenever the first $n$ or the last $n$ arguments equal 0. Therefore, the polynomial representation of $g_{3}$ cannot have a monomial $x_{i}^{1}$ or $x_{i}^{2}$. By Claims 7.3 and 7.4, the only possible monomials are of the form $x_{i}^{1}\cdot x_{j}^{2}$. Let $g_{3}(x_{1}^{1},\dots,x_{n}^{1},x_{1}^{2},\dots,x_{n}^{2})=\sum\limits_{i,j\in[n]}c_{i,j}x_{i}^{1}x_{j}^{2}$. It remains to check that the relation $R$ enforces the property $c_{i,j}+c_{j,i}=a_{i}\cdot a_{j}$. Substituting the tuple $((1,0,0),(0,1,0))$ into the $i$-th position, $((0,1,0),(1,0,0))$ into the $j$-th position, and $((0,0,0),(0,0,0))$ into the remaining positions of $f$, we obtain the tuple $((a_{i},a_{j},c_{i,j}),(a_{j},a_{i},c_{j,i}))$, which must belong to $R$. Hence, the definition of $R$ gives $a_{i}\cdot a_{j}+c_{i,j}+c_{j,i}=0$. ∎
Lemma 4.1.
$\operatorname{Pol}(\mathbb{D}_{4})=\operatorname{Clo}(x\circ y)=\mathcal{D}_{4}$.
Proof.
We prove the inclusions $\operatorname{Pol}(\mathbb{D}_{4})\subseteq\mathcal{D}_{4}\subseteq\operatorname{Clo}(x\circ y)\subseteq\operatorname{Pol}(\mathbb{D}_{4})$. The first inclusion follows from Claim 7.6, the second follows from Claim 7.5, and the last follows from the fact that $x\circ y$ preserves all the relations from $\mathbb{D}_{4}$. ∎
Lemma 4.3.
Suppose $n,\ell\geq 4$. Then $\operatorname{Clo}(x\circ y)$ does not contain an idempotent $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette symmetric operation.
Proof.
Assume the contrary. Let $f\in\operatorname{Clo}(x\circ y)$ be such an operation. By Lemma 4.1, $f\in\mathcal{D}_{4}$, so we can choose $a:[n]\to\mathbb{Z}_{2}$ and $c:[n]\times[n]\to\mathbb{Z}_{2}$ defining $f$. Take $i,j\in[\ell+1]$ with $i\neq j$. Substitute $(1,0,0)$ for the $i$-th coordinate and into the blocks 3 and 4 (the second overlined part), substitute $(0,1,0)$ for the $j$-th coordinate and into the blocks 5 and 6 (the third overlined part), and $(0,0,0)$ for the remaining coordinates. Since $f$ is $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette symmetric, the result must be the same for all $i$ and $j$. The first coordinate of the result is $a_{i}+a_{2\ell+2}+a_{2\ell+3}+\dots+a_{4\ell+2}$, and the third one is
$c_{i,j}+\sum\limits_{m=1}^{2\ell+1}c_{i,4\ell+2+m}+\sum\limits_{m=1}^{2\ell+1}c_{2\ell+1+m,j}+\sum\limits_{m_{1},m_{2}\in[2\ell+1]}c_{2\ell+1+m_{1},4\ell+2+m_{1}}.$
Denote $s_{i}=\sum\limits_{m=1}^{2\ell+1}c_{i,4\ell+2+m}$, $t_{j}=\sum\limits_{m=1}^{2\ell+1}c_{2\ell+1+m,j}$. Hence, $a_{1}=a_{2}=\dots=a_{\ell+1}$, and the value $c_{i,j}+s_{i}+t_{j}$ does not depend on $i$ and $j$. Since $\ell\geq 2$, there exist $i,j\in[\ell+1],i\neq j$, such that $s_{i}+t_{i}=s_{j}+t_{j}$. This implies $c_{i,j}+c_{j,i}=0$, which, by the definition of $\mathcal{D}_{4}$, means that $a_{i}\cdot a_{j}=0$, and therefore $a_{i}=a_{j}=0$. Thus, we have showed that $a_{1}=\dots=a_{\ell+1}=0$. Repeating the argument word to word, we obtain $a_{\ell+2}=\dots=a_{2\ell+1}=0$. Finally, by swapping the blocks and repeating the same argument, we prove that $a_{1}=\dots=a_{n(2l+1)}=0$. This contradicts the fact that $f$ is idempotent, since $f((1,0,0),\dots,(1,0,0))=(1,0,0)$ implies $\sum_{i=1}^{n(2\ell+1)}a_{i}=1$. ∎
8 Palette symmetric terms for small algebras
In this section we prove the existence of palette symmetric (PS) term operations in a product of small Taylor algebras. To do so, we develop machinery that is considerably more general than what is needed for this particular claim. In most cases, we show that palette-symmetric terms for an algebra can be constructed from palette-symmetric terms for smaller algebras. The situation in which this construction fails is very specific: the algebra must be Maltsev and supernilpotent (see Theorem 8.46). Attempts to strengthen the result by replacing the bound on the size of the domain with a purely algebraic property fail because the smaller algebras appearing in the construction are not always subalgebras; sometimes it is a retract algebra $h(\mathbf{A})$ induced by a mapping $h\colon A\to A$, sometimes it is minimal Taylor reduct.
We begin with the necessary algebraic definitions. In Section 8.3 we define strong subalgebras and state several of their properties from [45]. Some of these properties are stronger for minimal Taylor algebras, which we introduce in Section 8.4. In Section 8.5 we show how to prove the existence of a term operation satisfying certain identities by considering a freely generated relation. In Sections 8.6, 8.7, and 8.8 we explain how to construct a suitable family of terms, how to find squeezing terms within this family, and what exactly can be squeezed. In Section 8.9 we present three methods for constructing a palette symmetric term from palette symmetric terms on smaller algebras. Finally, in Section 8.10 we derive the main results: Theorems 8.46 and 3.7.
8.1 Necessary definitions
Algebras. We denote algebras by bold letters $\mathbf{A},\mathbf{B},\mathbf{C},\dots$, their domains by $A,B,C,\dots$, and the basic operations by $f^{\mathbf{A}},f^{\mathbf{B}},g^{\mathbf{C}},\dots$. We use standard universal algebraic notions of term operation, subalgebra, factor algebra, product of algebras, see [7]. Algebras are called similar if they have the same signature. We write $\mathbf{B}\leq\mathbf{A}$ if $\mathbf{B}$ is a subalgebra of $\mathbf{A}$. By $\operatorname{Clo}(\mathbf{A})$ we denote the set of all term operations in $\mathbf{A}$, i.e. the clone generated by the basic operations of $\mathbf{A}$. An algebra is called idempotent if all of its operations are idempotent. In this paper we are only interested in finite idempotent Taylor algebras. An algebra is called Taylor if it satisfies a nontrivial idempotent identity. To distinguish identities and equations, we use $\approx$ instead of == for identities. Thus, whenever we use the symbol $\approx$, we assume that all variable symbols (usually $x,y,z$) are universally quantified. For example, an operation $m$ is called Maltsev if $m(x,x,y)\approx m(y,x,x)\approx y$. There are many equivalent definitions of Taylor algebras [40, 39, 30]. We will need only the following.
Theorem 8.1.
Suppose $\mathbf{A}$ is a finite idempotent algebra. Then the following conditions are equivalent:
  1. 1.
    $\mathbf{A}$ is a Taylor algebra;
  2. 2.
    $\mathbf{A}$ has a WNU term operation of any prime arity $p>|A|$ [35];
  3. 3.
    $\mathbf{A}$ has a cyclic term operation of any prime arity $p>|A|$, i.e. an operation $f$ satisfying
[…]
  4. 4.
    $\mathbf{A}$ has an XY-symmetric term operation of any prime arity $p>|A|$, i.e. an operation $f$ such that $f(a_{1},\dots,a_{p})=f(a_{\sigma(1)},\dots,a_{\sigma(p)})$ for any $a_{1},\dots,a_{p}\in A$ and any permutation $\sigma$ on $[p]$ such that $|\{a_{1},\dots,a_{p}\}|=2$ [45].
In this paper we assume that every algebra is finite, idempotent, and Taylor. By $0_{\mathbf{A}}$ we denote the equality relation on $A$, which is the 0-congruence on $\mathbf{A}$. An algebra $(A;F_{A})$ is called polynomially complete (PC) if the clone generated by $F_{A}$ and all constants on $A$ is the clone of all operations on $A$ (see [28, 33]).
By $\mathcal{V}_{k}$ we denote the class of finite algebras $\mathbf{A}=(A;w^{\mathbf{A}})$ whose basic operation $w^{\mathbf{A}}$ is a $k$-ary idempotent WNU operation. So the symbol $w$ always means the basic operation of an algebra from $\mathcal{V}_{k}$. Since we only consider finite algebras, $\mathcal{V}_{k}$ is not a variety. For a prime $p$ by $\mathbf{Z}_{p}$ we denote the algebra whose domain is $\{0,1,\dots,p-1\}$ and whose basic operation $w^{\mathbf{Z}_{p}}$ is $a\cdot x_{1}+\dots+a\cdot x_{k}\pmod{p}$, where $a\cdot k=1\pmod{p}$ and $k$ is specified by the membership in $\mathcal{V}_{k}$.
For an algebra $\mathbf{A}\in\mathcal{V}_{k}$ and $h\colon A\to A$ such that $h\circ h=h$ by $h(\mathbf{A})$ we denote the retract algebra whose domain is the image of $h$ and whose only operation $w$ is defined by $w^{h(\mathbf{A})}(x_{1},\dots,x_{k}):=h(w^{\mathbf{A}}(x_{1},\dots,x_{k}))$. For a function $h$ the image of $h$ we denote by $\operatorname{Im}(h)$. It is not hard to verify that $h(\mathbf{A})\in\mathcal{V}_{k}$.
Notations. A relation $R\subseteq A_{1}\times\dots\times A_{s}$ is called subdirect if $\operatorname{pr}_{i}(R)=A_{i}$ for every $i\in[s]$. We write $\mathbf{R}\leq_{sd}\mathbf{A}_{1}\times\dots\times\mathbf{A}_{m}$ if $R$ is a subdirect relation and $\mathbf{R}\leq\mathbf{A}_{1}\times\dots\times\mathbf{A}_{m}$. For an algebra $\mathbf{A}$ and $R\subseteq A^{n}$ by $\operatorname{Sg}_{\mathbf{A}}(R)$ we denote the minimal subuniverse of $\mathbf{A}^{n}$ containing $R$, that is the subuniverse of $\mathbf{A}^{n}$ generated from $R$. Notice that if $R=\{\alpha_{1},\dots,\alpha_{n}\}$ then $\operatorname{Sg}_{\mathbf{A}}(R)=\{t^{\mathbf{A}}(\alpha_{1},\dots,\alpha_{n})\mid t\text{ is a term over }x_{1},\dots,x_{n}\}$. We say that a tuple $(a_{1},\dots,a_{n})$ generates an algebra $\mathbf{A}$ if $\operatorname{Sg}_{\mathbf{A}}(\{a_{1},\dots,a_{n}\})=A$. For an equivalence relation $\sigma$ on $A$ and $a\in A$ by $a/\sigma$ we denote the equivalence class containing $a$. For an equivalence relation $\sigma$ on $A$ and $B\subseteq A$ denote $B/\sigma=\{b/\sigma\mid b\in B\}$.
For two binary relations $\delta_{1}\subseteq A_{1}\times A_{2}$ and $\delta_{2}\subseteq A_{2}\times A_{3}$ by $\delta_{1}\circ\delta_{2}$ we denote the composition, i.e. the binary relation $\{(a,b)\mid\exists c\colon(a,c)\in\delta_{1}\wedge(c,b)\in\delta_{2}\}$. A binary subdirect relation $\delta\subseteq A\times B$ is called linked if the bipartite graph corresponding to $\delta$ is connected.
Irreducible congruences. A congruence $\sigma$ on $\mathbf{A}$ is called irreducible if there are no subalgebras $\mathbf{S}_{1},\mathbf{S}_{2},\dots,\mathbf{S}_{k}\leq\mathbf{A}/\sigma\times\mathbf{A}/\sigma$ such that $0_{\mathbf{A}/\sigma}=S_{1}\cap S_{2}\cap\dots\cap S_{k}$ and $0_{\mathbf{A}/\sigma}\neq S_{i}$ for every $i\in[k]$. Similarly, we say that a congruence $\sigma$ is $\wedge$-irreducible if it cannot be represented as an intersection of other congruences $\sigma_{1},\dots,\sigma_{n}$ on $\mathbf{A}$. Thus, the only difference between these two notions is that in one of them we allow any binary subalgebras and in another only congruences. For a $\wedge$-irreducible congruence by $\sigma^{+}$ we denote the minimal congruence above $\sigma$.
Bridges. Suppose $\sigma_{1}$ and $\sigma_{2}$ are congruences on $\mathbf{D}_{1}$ and $\mathbf{D}_{2}$, respectively. A relation $\delta\leq\mathbf{D}_{1}^{2}\times\mathbf{D}_{2}^{2}$ is called a bridge from $\sigma_{1}$ to $\sigma_{2}$ if the following conditions hold:
  1. 1.
    $(a_{1},a_{2},a_{3},a_{4})\in\delta\wedge(a_{1},b_{1})\in\sigma_{1}\wedge(a_{2},b_{2})\in\sigma_{1}\wedge(a_{3},b_{3})\in\sigma_{2}\wedge(a_{4},b_{4})\in\sigma_{2}\Longrightarrow(b_{1},b_{2},b_{3},b_{4})\in\delta$;
  2. 2.
    $\operatorname{pr}_{1,2}(\delta)\supsetneq\sigma_{1}$, $\operatorname{pr}_{3,4}(\delta)\supsetneq\sigma_{2}$,
  3. 3.
    $(a_{1},a_{2},a_{3},a_{4})\in\delta$ implies $(a_{1},a_{2})\in\sigma_{1}\Leftrightarrow(a_{3},a_{4})\in\sigma_{2}.$
An example of a bridge is the relation $\delta=\{(a_{1},a_{2},a_{3},a_{4})\mid a_{1},a_{2},a_{3},a_{4}\in\mathbb{Z}_{4}:a_{1}-a_{2}=2a_{3}-2a_{4}\}$. We can check that $\delta$ is a bridge from the equality relation (0-congruence) on $\mathbf{Z}_{4}$ and the (mod 2)(mod\;2)-equivalence relation on $\mathbf{Z}_{4}$. The notion of a bridge is strongly related to other notions in Universal Algebra and Tame Congruence Theory such as similarity and centralizers (see [41] for the detailed comparison). For a bridge $\delta$ by $\operatorname{tr}(\delta)$ we denote the binary relation defined by $\operatorname{tr}(\delta)(x,y)=\delta(x,x,y,y)$.
We can compose a bridge $\delta_{1}$ from $\sigma_{0}$ to $\sigma_{1}$ and a bridge $\delta_{2}$ from $\sigma_{1}$ to $\sigma_{2}$ using the following formula:
$\delta(x_{1},x_{2},z_{1},z_{2})=\exists y_{1}\exists y_{2}\;\delta_{1}(x_{1},x_{2},y_{1},y_{2})\wedge\delta_{2}(y_{1},y_{2},z_{1},z_{2}).$
We can prove (Lemma 28 in [45]) that $\delta$ is a bridge from $\sigma_{0}$ to $\sigma_{2}$ whenever the congruences $\sigma_{1}$, $\sigma_{2}$, and $\sigma_{3}$ are irreducible. Moreover, $\operatorname{tr}(\delta)=\operatorname{tr}(\delta_{1})\circ\operatorname{tr}(\delta_{2})$.
8.2 Two types of irreducible congruences
An irreducible congruence $\sigma$ on $\mathbf{A}$ is called linear if there exists a bridge $\delta$ from $\sigma$ to $\sigma$ such that $\operatorname{tr}(\delta)\supsetneq\sigma$. An irreducible congruence that is not linear is called a PC congruence. Note that a different definition is provided in [45]; however, as shown in Lemma 7 of [45], it is equivalent to the one presented here.
Notice that a congruence $\sigma$ is an irreducible/PC/linear congruence if and only if $0_{\mathbf{A}/\sigma}$ is an irreducible/PC/linear congruence. A congruence $\sigma$ on $\mathbf{A}\in\mathcal{V}_{k}$ is called perfect linear if there exists a bridge $\delta$ from $\sigma$ to $\sigma$ such that $\operatorname{tr}(\delta)=A^{2}$. This definition differs from the original one, but it is equivalent by Lemma 27 of [45]. An alternative equivalent characterization follows from commutator theory: $\sigma$ is perfect linear iff $\sigma^{+}$ is abelian and the centralizer of $\sigma^{+}$ modulo $\sigma$ is the 1-congruence [41].
Another important fact is that a bridge can only connect congruences of the same type.
Lemma 8.2 (Lemmas 8 and 10 in [45]).
Suppose $\sigma_{1}$ is a linear/PC/perfect linear congruence on $\mathbf{A}_{1}$, $\sigma_{2}$ is an irreducible congruence on $\mathbf{A}_{2}$, there is a bridge from $\sigma_{1}$ to $\sigma_{2}$. Then $\sigma_{2}$ is a linear/PC/perfect linear congruence. Moreover, if $\sigma_{1}$ is a PC congruence, then $\mathbf{A}_{1}/\sigma_{1}\cong\mathbf{A}_{2}/\sigma_{2}$.
The following claims justify the names chosen for the two types:
Lemma 8.3 (Lemma 11 in [45]).
Suppose $\sigma$ is a maximal linear congruence on $\mathbf{A}\in\mathcal{V}_{k}$. Then $\mathbf{A}/\sigma\cong\mathbf{Z}_{p}$ for some prime $p$.
Notice that $0_{\mathbf{Z}_{p}}$ is a perfect linear congruence on $\mathbf{Z}_{p}$.
Lemma 8.4 (Lemma 12 in [45]).
Suppose $\sigma$ is a maximal PC congruence on a BA and center free algebra $\mathbf{A}\in\mathcal{V}_{k}$. Then $\mathbf{A}/\sigma$ is a PC algebra.
$\boxtimes$-product of $\mathbf{B}$ and $\mathbf{Z}_{p}$. For $x=(a,b)$ by $x^{(1)}$ and $x^{(2)}$ we denote $a$ and $b$ respectively. For an algebra $\mathbf{B}=(B;g^{\mathbf{B}})$ by $\mathbf{B}\boxtimes\mathbf{Z}_{p}$ we denote the set of algebras $\mathbf{A}$ such that $A=B\times Z_{p}$, $(g^{\mathbf{A}}(x_{1},\dots,x_{n}))^{(1)}=g^{\mathbf{B}}(x_{1}^{(1)},\dots,x_{n}^{(1)})$ and $(g^{\mathbf{A}}(x_{1},\dots,x_{n}))^{(2)}=f(x_{1}^{(1)},\dots,x_{n}^{(1)})+a_{1}x_{1}^{(2)}+\dots+a_{n}x_{n}^{(2)}$ for some mapping $f\colon B^{n}\to\mathbb{Z}_{p}$ and $a_{1},\dots,a_{n}\in\mathbb{Z}_{p}$. Notice that any $m$-ary term operation of $\mathbf{C}\in\mathbf{B}\boxtimes\mathbf{Z}_{p}$ is also determined by their behavior on the first coordinate, the corresponding coefficients $a_{1},\dots,a_{m}$, and a mapping $f\colon B^{m}\to\mathbb{Z}_{p}$.
Lemma 8.5 ([45], Theorem 52).
Suppose $\mathbf{A}\in\mathcal{V}_{k}$, $0_{\mathbf{A}}$ is a perfect linear congruence. Then there exist a prime $p$, an algebra $\mathbf{C}\in(\mathbf{A}/0_{\mathbf{A}}^{+}\boxtimes\mathbf{Z}_{p})\cap\mathcal{V}_{k}$ and an injective homomorphism $h\colon\mathbf{A}\to\mathbf{C}$ such that $h(a)^{(1)}=a/0_{\mathbf{A}}^{+}$ for every $a\in A$.
Lemma 8.6.
Suppose $\mathbf{A}\in\mathcal{V}_{k}$, $0_{\mathbf{A}}$ is a perfect linear congruence. Then there exists a prime $p$ such that each equivalence class of $0_{\mathbf{A}}^{+}$ is either of size 1, or of size $p$, and the union of all the large equivalence classes binary absorbs $\mathbf{A}$ (see the definition of absorption in Section 8.3).
Proof.
By Lemma 8.5 there exists an algebra $\mathbf{C}\in(\mathbf{A}/0_{\mathbf{A}}^{+}\boxtimes\mathbf{Z}_{p})\cap\mathcal{V}_{k}$ and an injective homomorphism $h\colon\mathbf{A}\to\mathbf{C}$. Then to every equivalence class of $0_{\mathbf{A}}^{+}$ we assign a subalgebra of $\mathbf{Z}_{p}$, which implies that there are two types of equivalence classes: of size $p$ and of size 1. We claim that a binary absorbing term can be defined by $g(x,y)=w(x,x,\dots,x,y)$. Let $D$ be a large equivalence class of $0_{\mathbf{A}}^{+}$ and $a\in A$. We need to prove that $w(D,\dots,D,a)$ and $w(a,\dots,a,D)$ belong to large equivalence classes. To see this, choose two different elements $d_{1},d_{2}\in D$ and notice that $w^{\mathbf{A}}(d_{2},\dots,d_{2},a)\neq w^{\mathbf{A}}(d_{1},d_{2},\dots,d_{2},a)$ and $w^{\mathbf{A}}(a,\dots,a,d_{1})\neq w^{\mathbf{A}}(a,\dots,a,d_{2})$. ∎
Lemma 8.7.
Suppose $\mathbf{C}\in(\mathbf{Z}_{2}\boxtimes\mathbf{Z}_{2})\cap\mathcal{V}_{k}$. Then any XY-symmetric term operation is also symmetric term operation.
Proof.
Let $t$ be an XY-symmetric term operation. Since $t$ is symmetric on elements $(0,0)$ and $(0,1)$, we derive that $a_{1}=\dots=a_{n}$ in the definition of $\boxtimes$-product. Since $t$ is symmetric on elements $(0,0)$ and $(1,0)$, we derive that both $g^{(1)}$ and $f$ in this definition are symmetric, which completes the proof. ∎
8.3 Strong subuniverses
(Binary) absorbing subuniverse. We say $B$ is an absorbing subuniverse of an algebra $\mathbf{A}$ if there exists $t\in\operatorname{Clo}(\mathbf{A})$ such that $t(B,B,\dots,B,A,B,\dots,B)\subseteq B$ for any position of $A$. Also in this case we say that $B$ absorbs $\mathbf{A}$ with a term operation $t$. We write $B\trianglelefteq_{n}\mathbf{A}$, where $B$ absorbs $\mathbf{A}$ and the term operation can be chosen $n$-ary. Since we can always add a dummy variable $B\trianglelefteq_{n}\mathbf{A}\Rightarrow B\trianglelefteq_{n+1}\mathbf{A}$.
We say that $B$ binary absorbs $\mathbf{A}$ if $B\trianglelefteq_{2}\mathbf{A}$, and $B$ ternary absorbs $\mathbf{A}$ if $B\trianglelefteq_{3}\mathbf{A}$. To shorten sometimes we will write BA instead of binary absorbing.
Central subuniverse. A subuniverse $C$ of $\mathbf{A}$ is called central if it is an absorbing subuniverse and for every $a\in A\setminus C$ we have $(a,a)\notin\operatorname{Sg}_{\mathbf{A}}((\{a\}\times C)\cup(C\times\{a\}))$. Central subuniverses are strongly connected with ternary absorption.
Lemma 8.8 ([44], Corollary 6.11.1).
Suppose $B$ is a central subuniverse of $\mathbf{A}$, then $B\trianglelefteq_{3}\mathbf{A}$.
We say that an algebra $\mathbf{A}$ is BA and center free if $\mathbf{A}$ has no nontrivial binary absorbing subuniverse and nontrivial central subuniverse.
All types of subuniverses. Suppose $\varnothing\neq\mathbf{C}\lneq\mathbf{B}\leq\mathbf{A}$. We write
  • $C<_{\mathcal{BA}}^{\mathbf{A}}\mathbf{B}$ if $C$ is a BA subuniverse of $\mathbf{B}$.
  • $C<_{\mathcal{C}}^{\mathbf{A}}\mathbf{B}$ if $C$ is a central subuniverse of $\mathbf{B}$.
  • $C<_{\mathcal{D}(\sigma)}^{\mathbf{A}}\mathbf{B}$ if there exists an irreducible congruence $\sigma$ such that
    1. 1.
      $B^{2}\subseteq\sigma^{+}$;
    2. 2.
      $C=B\cap E$ for some block $E$ of $\sigma$;
    3. 3.
      $\mathbf{B}/\sigma$ is BA and center free.
  • $C<_{\mathcal{L}(\sigma)}^{\mathbf{A}}\mathbf{B}$ if $C<_{\mathcal{D}(\sigma)}^{\mathbf{A}}B$ and $\sigma$ is linear.
  • $C<_{\mathcal{PC}(\sigma)}^{\mathbf{A}}\mathbf{B}$ if $C<_{\mathcal{D}(\sigma)}^{\mathbf{A}}\mathbf{B}$ and $\sigma$ is a PC congruence.
  • $C<_{\mathcal{S}}^{\mathbf{A}}\mathbf{B}$ if there exists a BA and central (simultaneously) subuniverse $D$ in $\mathbf{B}$ such that $D\leq\mathbf{C}$.
Notice that the above definition of types differs from the original one in [45], since we write $\sigma^{+}$ instead of $\sigma^{*}$; that is, we use the minimal congruence above rather than the minimal compatible binary relation above. Nevertheless, this does not affect the definition. Indeed, if $\sigma^{*}\not\supseteq B^{2}$ while $\sigma^{+}\supseteq B^{2}$, then $\sigma^{*}\cap B^{2}$ defines a linked binary relation on $\mathbf{B}/\sigma$ and yields either a BA or central subuniverse, which is forbidden (see Theorem 3.11.1 in [13] or Lemma 70 in [45]).
If we do not want to specify the congruence we write $C<_{T}^{\mathbf{A}}\mathbf{B}$ instead of $C<_{T(\sigma)}^{\mathbf{A}}\mathbf{B}$. If $C<_{T}^{\mathbf{A}}\mathbf{B}$ then we say that $C$ is a subuniverse of $\mathbf{B}$ of type $T$. We write $C\leq_{T(\sigma)}^{\mathbf{A}}\mathbf{B}$ to mean that either $C<_{T(\sigma)}^{\mathbf{A}}\mathbf{B}$ or $C=B$. Often, instead of $_{\mathcal{BA}}^{\mathbf{A}}\mathbf{B}, C<𝒞$$_{\mathcal{C}}^{\mathbf{A}}\mathbf{B}$, and $C<_{\mathcal{S}}^{\mathbf{A}}\mathbf{B}$ we write $_{\mathcal{BA}}\mathbf{B}, C<𝒞$$_{\mathcal{C}}\mathbf{B}$, and $C<_{\mathcal{S}}\mathbf{B}$, respectively, since $\mathbf{A}$ is irrelevant to the definition. Also, we write $C<_{\mathcal{BA},\mathcal{C}}\mathbf{B}$ to mean that both $C<_{\mathcal{BA}}\mathbf{B}$ and $C<_{\mathcal{C}}\mathbf{B}$ hold.
We write $C\lll^{\mathbf{A}}\mathbf{B}$ if there exist $B_{0},B_{1},\dots,B_{n}\subseteq B$ and $T_{1},\dots,T_{n}\in\{\mathcal{BA},\mathcal{C},\mathcal{S},\mathcal{D}\}$ such that $\mathbf{C}=\mathbf{B}_{n}<_{T_{n}}^{\mathbf{A}}\mathbf{B}_{n-1}<_{T_{n-1}}^{\mathbf{A}}<\dots<_{T_{2}}^{\mathbf{A}}<\mathbf{B}_{1}<_{T_{1}}^{\mathbf{A}}\mathbf{B}_{0}=\mathbf{B}$. Note that $n$ may be 0 and the relation $\lll^{\mathbf{A}}$ is reflexive. We do not allow empty subuniverses and the condition $\varnothing\lll\mathbf{A}$ never holds. Nevertheless, it sometimes convenient to allow the empty set. In this case we place a dot above the symbol and write $B\dot{\lll}\mathbf{A}$ to mean that either $B\lll\mathbf{A}$ or $B=\varnothing$. With the same meaning we use dots in the notations $C\dot{<}_{T}^{\mathbf{A}}\mathbf{B}$ or $C\dot{\leq}_{T}^{\mathbf{A}}\mathbf{B}$.
Lemma 8.9 (Lemma 6.4 in [44]).
Suppose $\mathbf{C}_{1}\leq_{\mathcal{C}}\mathbf{A}$ and $\mathbf{C}_{2}\leq_{\mathcal{C}}\mathbf{A}$. Then $C_{1}\cap C_{2}\dot{\leq}_{\mathcal{C}}\mathbf{A}$.
Lemma 8.10 (Lemma 6.5 in [44]).
Suppose $\mathbf{C}\leq_{\mathcal{C}}\mathbf{A}$ and $\mathbf{B}\leq\mathbf{A}$. Then $C\cap B\dot{\leq}_{\mathcal{C}}\mathbf{A}$.
Lemma 8.11.
Suppose $B<_{\mathcal{BA}}\mathbf{A}$, $R\leq_{sd}\mathbf{A}^{n}$. Then $R\cap B^{n}\neq\varnothing$
Proof.
To get a tuple from $R\cap B^{n}$ we apply a binary absorbing operation to all the tuples from $R$. ∎
Lemma 8.12 (Proposition 2.4 in [4]).
Suppose $\mathbf{C}\trianglelefteq\mathbf{B}\trianglelefteq\mathbf{A}$. Then $\mathbf{C}\trianglelefteq\mathbf{A}$.
Lemma 8.13.
Suppose $\mathbf{C}\leq_{\mathcal{C}}\mathbf{B}\leq_{\mathcal{C}}\mathbf{A}$. Then $\mathbf{C}\leq_{\mathcal{C}}\mathbf{A}$.
Proof.
By Lemma 8.12 $\mathbf{C}\trianglelefteq\mathbf{A}$. If $a\in A\setminus B$ then, since $B<_{\mathcal{C}}\mathbf{A}$, we have
$(a,a)\notin\operatorname{Sg}_{\mathbf{A}}((\{a\}\times B)\cup(B\times\{a\}))\supseteq\operatorname{Sg}_{\mathbf{A}}((\{a\}\times C)\cup(C\times\{a\}))$
If $a\in B\setminus C$ then, since $C<_{\mathcal{C}}\mathbf{B}$, we have $(a,a)\notin\operatorname{Sg}_{\mathbf{A}}((\{a\}\times C)\cup(C\times\{a\}))$
Lemma 8.14 (Corollary 6.9.2 in [44]).
Suppose $R\leq\mathbf{A}_{1}\times\dots\times\mathbf{A}_{n}$, $\operatorname{pr}_{1}(R)=A_{1}$, $C_{i}\leq_{\mathcal{C}}\mathbf{A}_{i}$ for every $i\in[n]$. Then $\operatorname{pr}_{1}(R\cap(C_{1}\times\dots\times C_{N}))\dot{\leq}_{\mathcal{C}}\mathbf{A}_{1}$.
Lemma 8.15.
Suppose $B\lll\mathbf{A}$ and $|B|>1$. Then there exists $C<_{T}^{\mathbf{A}}\mathbf{B}$, where $T\in\{\mathcal{BA},\mathcal{C},\mathcal{L},\mathcal{PC}\}$.
Lemma 8.16 (Lemma 20(i) in [45]).
Suppose $B\lll\mathbf{A}$, $D\lll\mathbf{A}$. Then $B\cap D\dot{\lll}\mathbf{A}$.
Lemma 8.17 (Corollary 22 in [45]).
Suppose
  1. 1.
    $R\leq_{sd}\mathbf{A}_{1}\times\dots\times\mathbf{A}_{n}$, $n\geq 2$;
  2. 2.
    $C_{i}<_{T_{i}(\sigma_{i})}^{\mathbf{A}_{i}}\mathbf{B}_{i}\lll\mathbf{A}_{i}$, where $T_{i}\in\{\mathcal{BA},\mathcal{C},\mathcal{S},\mathcal{L},\mathcal{PC}\}$ for $i=1,2,\dots,n$;
  3. 3.
    $R\cap(C_{1}\times\dots\times C_{n})=\varnothing$;
  4. 4.
    $R\cap(C_{1}\times\dots\times C_{j-1}\times B_{j}\times C_{j+1}\times\dots\times C_{n})\neq\varnothing$ for every $j\in[n]$.
Then one of the following conditions holds:
  1. (ba)
    $T_{1}=\dots=T_{n}=\mathcal{BA}$;
  2. (l)
    $T_{1}=\dots=T_{n}=\mathcal{L}$ and for every $k,\ell\in[n]$ there exists a bridge $\delta$ from $\sigma_{k}$ and $\sigma_{\ell}$ such that $\operatorname{tr}(\delta)=\sigma_{k}\circ\operatorname{pr}_{k,\ell}(R)\circ\sigma_{\ell}$;
  3. (c)
    $n=2$ and $T_{1}=T_{2}=\mathcal{C}$;
  4. (pc)
    $n=2$, $T_{1}=T_{2}=\mathcal{PC}$, $\mathbf{A}_{1}/\sigma_{1}\cong\mathbf{A}_{2}/\sigma_{2}$, and the relation $\{(a/\sigma_{1},b/\sigma_{2})\mid(a,b)\in R\}$ is bijective.
Lemma 8.18 (Corollary 18(r1) in [45]).
Suppose $R\leq_{sd}\mathbf{A}_{1}\times\dots\times\mathbf{A}_{n}$, $B_{i}\lll\mathbf{A}_{i}$ for every $i\in[n]$. Then $\operatorname{pr}_{1}(R\cap(B_{1}\times\dots\times B_{n}))\dot{\lll}\mathbf{A}_{1}$.
8.4 Minimal Taylor algebras
An algebra $\mathbf{B}$ is a reduct of $\mathbf{A}$ if $A=B$ and $\operatorname{Clo}(\mathbf{B})\subseteq\operatorname{Clo}(\mathbf{A})$. An algebra $\mathbf{A}$ is called a minimal Taylor algebra if it is Taylor but no proper reduct of $\mathbf{A}$ is.
Lemma 8.19 (Propositions 5.2, 5.4 in [2]).
Every Taylor algebra has a minimal Taylor reduct. Any subalgebra, finite power, or quotient of a minimal Taylor algebra is again a minimal Taylor algebra.
Lemma 8.20 (Theorem 5.10 in [2] , Theorem 6.15 in [44]).
Let $\mathbf{A}$ be a minimal Taylor algebra and $B\subseteq A$. Then the following conditions are equivalent.
  1. (a)
    $B\trianglelefteq_{3}\mathbf{A}$.
  2. (b)
    $B\leq_{\mathcal{C}}\mathbf{A}$.
  3. (c)
    The relation $R(x,y)=B(x)\vee B(y)$ is a subuniverse of $\mathbf{A}^{2}$.
For an algebra $\mathbf{A}$ by $\mathcal{C}(\mathbf{A})$ we denote the set of all proper subuniverses $B\lneq\mathbf{A}$ such that the relation $R(x,y)=B(x)\vee B(y)$ is a subuniverse of $\mathbf{A}^{2}$. If $\mathbf{B}$ is a reduct of $\mathbf{A}$, then $\mathcal{C}(\mathbf{B})\supseteq\mathcal{C}(\mathbf{A})$; this property does not hold for central or ternary absorbing subuniverses, which is exactly why we use $\mathcal{C}(\mathbf{A})$ instead.
Lemma 8.21.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$ are Taylor algebras such that $|A_{i}|<p$ for a prime $p$. Then $\mathbf{A}_{1}\times\dots\times\mathbf{A}_{s}$ has a cyclic term operation of arity $p$.
Proof.
The proof is by induction on $s$. For $s=1$ it follows from Theorem 8.1. Let $w_{1}$ be a $p$-ary cyclic term in $\mathbf{A}_{1},\dots,\mathbf{A}_{s-1}$, $w_{2}$ be a $p$-ary cyclic term in $\mathbf{A}_{s}$. Then the term
$w(x_{1},\dots,x_{p})\approx w_{1}(w_{2}(x_{1},\dots,x_{p}),w_{2}(x_{2},\dots,x_{p},x_{1}),\dots,w_{2}(x_{p},x_{1}\dots,x_{p-1}))$
is a cyclic term for $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$. ∎
Lemma 8.22.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{p}$, such that $p>|A_{i}|$ for every $i\in[s]$. Then there exists a $p$-ary term $c$ such that $(A_{i};c^{\mathbf{A}_{i}})$ is a minimal Taylor algebra and $c^{\mathbf{A}_{i}}$ is cyclic for every $i\in[s]$.
Proof.
Using Lemma 8.19 choose a minimal Taylor reduct $\mathbf{B}$ of $\mathbf{A}_{1}\times\dots\times\mathbf{A}_{s}$. Notice that $\mathbf{B}\cong\mathbf{A}_{1}^{\prime}\times\dots\times\mathbf{A}_{s}^{\prime}$, where each $\mathbf{A}_{i}^{\prime}$ is a minimal Taylor reduct of $\mathbf{A}_{i}$. By Lemma 8.21, $\mathbf{B}$ has a cyclic term of arity $p$. Then the term defining $c$ from the basic operation $w$ is exactly the term we need. ∎
8.5 Free generated relations and symmetric reductions
The free generated relations. For algebras $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{k}$ by $R^{n}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ we denote the following relation. Coordinates of the relation are indexed by $(\mathbf{A}_{i},\alpha)$, where $\alpha\in A_{i}^{n}$ is a tuple generating $\mathbf{A}_{i}$. The set of all indices denote by $I^{n}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$. For $i\in[n]$ by $\gamma_{i}$ we denote the tuple whose elements are indexed by $I^{n}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ and whose $(\mathbf{A}_{i},\alpha)$-th element is equal to $\alpha(i)$ for every $(\mathbf{A}_{i},\alpha)\in I$. Then $R^{n}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ is the subalgebra generated by $\gamma_{1},\dots,\gamma_{n}$, and the tuples $\gamma_{1},\dots,\gamma_{n}$ are called the generators of $R^{n}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$. By $\mathcal{R}^{n}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ we denote the set of all relations $R$ whose coordinates are indexed by $I^{n}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$, and the projection of $R$ onto the $(\mathbf{A}_{i},\alpha)$-th coordinate is $A_{i}$ for every $(\mathbf{A}_{i},\alpha)\in I^{n}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$.
Similarly, for $k_{1},\dots,k_{n}\in\mathbb{N}$ and algebras $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{k}$ by $R^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ we denote the following relation. Coordinates of the relation are indexed by $(\mathbf{A}_{i},\alpha)$, where $\alpha\in A_{i}^{k_{1}+\dots+k_{n}}$ is a $(\overline{k_{1}},\dots,\overline{k_{n}})$-palette tuple generating $\mathbf{A}_{i}$. The set of all indices we denote by $I^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$. For $i\in[k_{1}+\dots+k_{n}]$ by $\gamma_{i}$ we denote the tuple whose elements are indexed by $I^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ and whose $(\mathbf{A}_{j};\alpha)$-th element is equal to $\alpha(i)$ for every $(\mathbf{A}_{j};\alpha)\in I^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$. Then $R^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{n}}$ is the subuniverse generated by $\gamma_{1},\dots,\gamma_{k_{1}+\dots+k_{n}}$, and the tuples $\gamma_{1},\dots,\gamma_{k_{1}+\dots+k_{n}}$ are called the generators of $R^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{n}}$. By $\mathcal{R}^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ we denote the set of all relations $R$ whose coordinates are indexed by $I^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$, and the projection of $R$ onto the $(\mathbf{A}_{i},\alpha)$-th coordinate is $A_{i}$ for every $(\mathbf{A}_{i},\alpha)\in I^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$. Note that we will use the same notations for $(\overline{k_{1}},\dots,\overline{k_{n}},\ell)$-palette tuples. For example, the coordinates of the free generated relation $R^{\overline{k_{1}},\dots,\overline{k_{n}},\ell}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ are indexed by $(\overline{k_{1}},\dots,\overline{k_{n}},\ell)$-palette tuples generating $\mathbf{A}_{i}$.
Reductions. To avoid repetitions, we define reductions for relations from $\mathcal{R}^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$, assuming that the same definitions can be applied to $\mathcal{R}^{n}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$. In our proof we reduce a relation $R\in\mathcal{R}^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ by reducing their coordinates. A reduction $D^{(\top)}$ for $R\in\mathcal{R}^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ is a mapping that assigns a subuniverse $D_{(\mathbf{A}_{i},\alpha)}^{(\top)}\leq\mathbf{A}_{i}$ to every $(\mathbf{A}_{i},\alpha)\in I^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$. We write $D^{(\bot)}\lll D^{(\top)}$ and $D^{(\bot)}\leq_{T}D^{(\top)}$ whenever $D_{i}^{(\bot)}\lll D_{i}^{(\top)}$ for every $i\in I^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ and $D_{i}^{(\bot)}\leq_{T}D_{i}^{(\top)}$ for every $i\in I^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$, respectively. Notice that any reduction $D^{(\top)}$ can be viewed as a relation from $\mathcal{R}^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$. Then for any $R\in\mathcal{R}^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ and a reduction $D^{(\bot)}$ by $R^{(\bot)}$ we denote $R\cap D^{(\bot)}$. A reduction $D^{(\bot)}$ is called 1-consistent for $R\in\mathcal{R}^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ if $\operatorname{pr}_{i}(R^{(\bot)})=D_{i}^{(\bot)}$ for every $i\in I^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$.
Permutations and symmetries. By $\operatorname{Perm}^{k_{1},\dots,k_{n}}$ we denote the set of all permutations on $k_{1}+\dots+k_{n}]$ such that the first $k_{1}$ elements remain in the first $k_{1}$ positions, the next $k_{2}$ elements remain in the next $k_{2}$ positions, and so on. For a tuple $\alpha\in A^{k_{1}+\dots+k_{n}}$ denote $\operatorname{Perm}^{k_{1},\dots,k_{n}}(\alpha)=\{\alpha^{\sigma}\mid\sigma\in\operatorname{Perm}^{k_{1},\dots,k_{n}}\}$. For an index $i=(\mathbf{A}_{j},\alpha)$ we denote by $\operatorname{Perm}^{k_{1},\dots,k_{n}}(i)$ the set of indices $(\mathbf{A}_{j},\beta)$ with $\beta\in\operatorname{Perm}^{k_{1},\dots,k_{n}}(\alpha)$. Let $\gamma$ be a tuple whose coordinates are indexed by elements from $I^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$, and let $\sigma\colon[k_{1}+\dots+k_{n}]\to[k_{1}+\dots+k_{n}]$ be a permutation. By $\gamma^{\langle{\sigma}\rangle}$ we denote the tuple $\gamma^{\prime}$ defined by $\gamma^{\prime}((\mathbf{A}_{i},\alpha))=\gamma((\mathbf{A}_{i},\alpha^{\sigma}))$ for every $(\mathbf{A}_{i},\alpha)\in I^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$. Similarly, for a relation $R\in\mathcal{R}^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ put $R^{\langle{\sigma}\rangle}=\{\gamma^{\langle{\sigma}\rangle}\mid\gamma\in R\}$. A relation $R$ is called $\sigma$-symmetric if $R^{\langle{\sigma}\rangle}=R$. A relation $R\in\mathcal{R}^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ is called symmetric if it is $\sigma$-symmetric for every $\sigma\in\operatorname{Perm}^{k_{1},\dots,k_{n}}$. Similarly, a reduction $D^{(\top)}$ is called symmetric if $D_{i}^{(\top)}=D_{j}^{(\top)}$ for any $j\in\operatorname{Perm}^{k_{1},\dots,k_{n}}(i)$.
We say that a finite idempotent algebra $\mathbf{A}$ has bounded width if there is no linear congruence on any subalgebra of $\mathbf{A}$ (see [5] for alternative definitions). In the theory of strong subalgebras this means that only types $\mathcal{BA},\mathcal{C},\mathcal{S},\mathcal{PC}$ may appear. We say that a term $t$ is local $(\overline{k_{1}},\dots,\overline{k_{n}})$-palette symmetric in $\mathbf{A}$ if we require symmetry only on $(\overline{k_{1}},\dots,\overline{k_{n}})$-palette tuples generating $\mathbf{A}$.
Lemma 8.23.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$ are algebras such that each $\mathbf{A}_{i}$ is a bounded width algebra or a BA and center free PC algebra, $k_{1},\dots,k_{n}\in\mathbb{N}$. Then there exists a local $(\overline{k_{1}},\dots,\overline{k_{n}})$-palette symmetric term operation for $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$.
Proof.
First, notice that if $\mathbf{A}$ is a BA and center free PC algebra, then $B\lll A$ implies that $|B|=1$ or $B=A$, which guaranties that the linear case cannot come from PC algebras. Let us consider an inclusion minimal symmetric 1-consistent reduction $D^{(\top)}$ for $R^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ such that $D_{(\mathbf{A}_{i},\alpha)}^{(\top)}\lll\mathbf{A}_{i}$ for every $(\mathbf{A}_{i},\alpha)\in I^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$. Such a reduction always exists as the trivial one satisfies the required properties. Consider two cases:
Case 1. $|D_{(\mathbf{A}_{i},\alpha)}^{(\top)}|=1$ for every $(\mathbf{A}_{i},\alpha)\in I^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$. Then $D^{(\top)}$ viewed as a relation contains only one symmetric tuple. Then there exists a term giving this tuple on the generators of $R^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$. Since the tuple is symmetric, the term operation is local $(\overline{k_{1}},\dots,\overline{k_{n}})$-PS for $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$.
Case 2. $|D_{(\mathbf{A}_{i},\alpha)}^{(\top)}|>1$ for some $(\mathbf{A}_{i},\alpha)\in I^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$. By Lemma 8.15 there exists $B<_{T}D_{(\mathbf{A}_{i},\alpha)}^{(\top)}$ for some $T\in\{\mathcal{BA},\mathcal{C},\mathcal{PC}\}$. Let us define a reduction $D^{(\triangle)}$ for $R^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ by $D^{(\triangle)}_{(\mathbf{A}_{i},\beta)}=B$ for any $\beta\in\operatorname{Perm}^{k_{1},\dots,k_{n}}(\alpha)$, and by the corresponding $A_{j}$ on the remaining coordinates. Since $\alpha$ is a palette tuple, $R^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\triangle)}\neq\varnothing$. If $R^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\top)}\cap D^{(\triangle)}=\varnothing$, then, by Lemma 8.11, $T\neq\mathcal{BA}$, and by Lemma 8.17 there should be 2 coordinates giving a contradiction. This cannot happen as the reduction $D^{(\top)}$ is 1-consistent.
Then the reduction $D^{(\bot)}$ defined by $D^{(\bot)}_{j}=\operatorname{pr}_{j}(R^{\overline{k_{1}},\dots,\overline{k_{n}}}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\top)}\cap D^{(\triangle)})$ is smaller than $D^{(\top)}$ and symmetric. By Lemma 8.18, $D^{(\bot)}\lll D^{(\top)}$, which contradicts our assumption about the minimality. ∎
Corollary 8.24.
Every finite bounded width algebra $\mathbf{A}$ has a $(\overline{k_{1}},\dots,\overline{k_{n}})$-palette symmetric term operation for every $k_{1},\dots,k_{n}\in\mathbb{N}$.
Proof.
Let $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$ be the set of all subalgebras of $\mathbf{A}$ including $\mathbf{A}$. Then Lemma 8.23 gives a local $(\overline{k_{1}},\dots,\overline{k_{n}})$-palette symmetric term operation for $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$, which is a $(\overline{k_{1}},\dots,\overline{k_{n}})$-palette symmetric term operation for $\mathbf{A}$. ∎
Lemma 8.25.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{k}$, then for every $n,k_{1},\dots,k_{n},\ell\in\mathbb{N}$ there exists a term $t$ of arity $k_{1}+\dots+k_{n}+\ell$ such that for any $i\in[s]$ and any $(\overline{k_{1}},\dots,\overline{k_{n}},\ell)$-palette tuple $(\mathbf{a}_{1},\dots,\mathbf{a}_{n},\mathbf{b})$ generating $\mathbf{A}_{i}$:
  1. 1.
    there exists a minimal central subuniverse $B$ of $\mathbf{A}_{i}$ such that $t^{\mathbf{A}}(\mathbf{a}_{1}^{\sigma_{1}},\dots,\mathbf{a}_{n}^{\sigma_{n}},\mathbf{b}^{\delta})\in B$ for any permutations $\sigma_{1},\dots,\sigma_{n},\delta$ on $k_{1}],\dots,[k_{n}]$, and $\ell$], respectively;
  2. 2.
    if $C<_{\mathcal{C}}\mathbf{A}_{i}$ and $\mathbf{b}\in C^{\ell}$ then $t^{\mathbf{A}_{i}}(\mathbf{a}_{1},\dots,\mathbf{a}_{n},\mathbf{b})\in C$.
Proof.
Let us consider the free-generated relation $R^{\overline{k_{1}},\dots,\overline{k_{n}},\ell}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$. For every $i\in[s]$ and every tuple $(\mathbf{a}_{1},\dots,\mathbf{a}_{n},\mathbf{b})$ put $D^{(\top)}_{(\mathbf{A}_{i},(\mathbf{a}_{1},\dots,\mathbf{a}_{n},\mathbf{b}))}=B$, where $B$ is the minimal central subuniverse of $\mathbf{A}_{i}$ containing all the elements of $\mathbf{b}$. It exists because the intersection of two central subuniverses is central (Lemma 8.9). Since $R^{\overline{k_{1}},\dots,\overline{k_{n}},\ell}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\top)}$ contains the last $\ell$ generators of $R^{\overline{k_{1}},\dots,\overline{k_{n}},\ell}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}},$ it is not empty. By Lemma 8.14, $\operatorname{pr}_{(\mathbf{A}_{i},\alpha)}(R^{\overline{k_{1}},\dots,\overline{k_{n}},\ell}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\top)})\leq_{\mathcal{C}}\mathbf{A}_{i}$ for every $(\mathbf{A}_{i},\alpha)\in I^{\overline{k_{1}},\dots,\overline{k_{n}},\ell}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$. Then, by the minimality of the central subuniverses the reduction $D^{(\top)}$ is 1-consistent.
Let us consider an inclusion minimal 1-consistent symmetric reduction $D^{(\bot)}\leq_{\mathcal{C}}D^{(\top)}$. Such a reduction exists as the reduction $D^{(\top)}$ satisfies this property. If $D^{(\bot)}_{(\mathbf{A}_{i},\alpha)}$ is a minimal central subuniverse for every $(\mathbf{A}_{i},\alpha)\in I^{\overline{k_{1}},\dots,\overline{k_{n}},\ell}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}},$ then we are done. Otherwise, fix $(\mathbf{A}_{i},\alpha)\in I^{\overline{k_{1}},\dots,\overline{k_{n}},\ell}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ and choose $B<_{\mathcal{C}}D^{(\bot)}_{(\mathbf{A}_{i},\alpha)}$. Let us define a reduction $D^{(\triangle)}$ for $R^{\overline{k_{1}},\dots,\overline{k_{n}},\ell}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ by $D^{(\triangle)}_{(\mathbf{A}_{i},\beta)}=B$ for any $\beta\in\operatorname{Perm}^{k_{1},\dots,k_{n},\ell}(\alpha)$, and by the corresponding $A_{j}$ on the remaining coordinates. Since $\alpha$ is a palette tuple, $R^{\overline{k_{1}},\dots,\overline{k_{n}},\ell}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\triangle)}\neq\varnothing$. If $R^{\overline{k_{1}},\dots,\overline{k_{n}},\ell}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\bot)}\cap D^{(\triangle)}=\varnothing$, then by Lemma 8.17 there should be 2 coordinates giving a contradiction. This cannot happen as the reduction $D^{(\bot)}$ is 1-consistent.
Then the reduction $D^{(\star)}$ defined by $D^{(\star)}_{j}=\operatorname{pr}_{j}(R^{\overline{k_{1}},\dots,\overline{k_{n}},\ell}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\bot)}\cap D^{(\triangle)})$ is smaller than $D^{(\bot)}$ and symmetric. By Lemma 8.14, $D^{(\star)}\leq_{\mathcal{C}}D^{(\bot)}$, by Lemma 8.13 $D^{(\star)}\leq_{\mathcal{C}}D^{(\top)}$, which contradicts our assumption about the minimality of $D^{(\bot)}$. ∎
8.6 Squeezing operations
For two terms $t_{1}$ and $t_{2}$ over the variables $x_{1},\dots,x_{n+1}$ we denote the term $t_{1}(x_{1},\dots,x_{n},t_{2}(x_{1},\dots,x_{n},x_{n+1}))$ by $t_{1}\circ t_{2}$. Similarly, by $t^{m}$ we denote $\underbrace{t\circ t\circ\dots\circ t}_{m}$. We say that a set $\mathcal{F}$ of terms over the variables $x_{1},\dots,x_{n+1}$ is closed under right composition if $f_{1}\circ f_{2}\in\mathcal{F}$ for all $f_{1},f_{2}\in\mathcal{F}$.
Let $\mathcal{F}$ be a set of terms over the variables $x_{1},\dots,x_{n+1}$. By $\operatorname{Squeezer}(\mathcal{F},\mathbf{A}_{1},\dots,\mathbf{A}_{s})$ we denote the set of all terms $f\in\mathcal{F}$ such that
  1. 1.
    for any term $t\in\mathcal{F}$, $j\in[s]$, and $a_{1},\dots,a_{n}\in A_{j}$
    $|\{f^{\mathbf{A}_{j}}(a_{1},\dots,a_{n},b)\mid b\in A_{j}\}|\leq|\{t^{\mathbf{A}_{j}}(a_{1},\dots,a_{n},b)\mid b\in A_{j}\}|;$
  2. 2.
    for all $j\in[s]$ we have $f^{\mathbf{A}_{j}}\circ f^{\mathbf{A}_{j}}=f^{\mathbf{A}_{j}}$.
Thus, $\operatorname{Squeezer}(\mathcal{F},\mathbf{A}_{1},\dots,\mathbf{A}_{s})$ consists of terms from $\mathcal{F}$ such that the corresponding images are the smallest ones among all such terms.
We say that a family of terms $\mathcal{F}$ is good for algebras $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{k}$ if for any $f\in\mathcal{F}$, any $i\in[s]$, any tuples $(a_{1},\dots,a_{n})$ and $(b_{1},\dots,b_{m})$ with $m\leq n$ generating $\mathbf{A}_{i}$, there exists $t\in\mathcal{F}$ such that $t(b_{1},\dots,b_{m},y_{1},\dots,y_{n-m},x)\approx f(a_{1},\dots,a_{n},x)$.
Lemma 8.26.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{k}$, $\mathcal{F}$ is a nonempty set of terms over the variables $x_{1},\dots,x_{n+1}$ closed under right composition. Then
  1. (1)
    $\operatorname{Squeezer}(\mathcal{F},\mathbf{A}_{1},\dots,\mathbf{A}_{s^{\prime}})\supseteq\operatorname{Squeezer}(\mathcal{F},\mathbf{A}_{1},\dots,\mathbf{A}_{s})$ if $s^{\prime}\leq s$;
  2. (2)
    $\operatorname{Squeezer}(\mathcal{F},\mathbf{A}_{1},\dots,\mathbf{A}_{s})=\operatorname{Squeezer}(\mathcal{F},\mathbf{A}_{1},\dots,\mathbf{A}_{s},\mathbf{A}_{i}/\sigma,\mathbf{B})$ for any congruence $\sigma$ on $\mathbf{A}_{i}$ and subalgebra $\mathbf{B}\leq\mathbf{A}_{i}$;
  3. (3)
    $(t_{1}\circ t_{2}\circ t_{3})^{N!}\in\operatorname{Squeezer}(\mathcal{F},\mathbf{A}_{1},\dots,\mathbf{A}_{s})$ for any $t_{1},t_{3}\in\mathcal{F}$ and $t_{2}\in\operatorname{Squeezer}(\mathcal{F},\mathbf{A}_{1},\dots,\mathbf{A}_{s})$, where $N=\max\{|A_{1}|,\dots,|A_{s}|\}$;
  4. (4)
    $\operatorname{Squeezer}(\mathcal{F},\mathbf{A}_{1},\dots,\mathbf{A}_{s})\neq\varnothing$.
Proof.
Properties (1) and (2) follow immediately from the definition. To prove property (3) notice that $(t_{1}\circ t_{2}\circ t_{3})$ is from $\mathcal{F}$ and satisfies property 1 of squeezers. Property 2 is guaranteed by the $N!$-th power.
Let us prove that $\operatorname{Squeezer}(\mathcal{F},\mathbf{A}_{1},\dots,\mathbf{A}_{s})$ is not empty (property (4)). Since we can compose terms, we can derive a term $t$ such that the set $\{t^{\mathbf{A}_{j}}(a_{1},\dots,a_{n},b)\mid b\in A_{j}\}$ is of minimal size for all $j\in[s]$ and $a_{1},\dots,a_{n}\in A_{j}$. It remains to obtain property 2 of squeezers by taking the $N!$-th power of the term, where $N=\max\{|A_{1}|,\dots,|A_{s}|\}$. ∎
Lemma 8.27.
Suppose $\mathbf{A}\in\mathcal{V}_{k}$, $\mathcal{F}$ is a set of terms over the variables $x_{1},\dots,x_{n+1}$ closed under right composition, $\mathcal{F}$ is good for $\mathbf{A}$, $\operatorname{Sg}_{\mathbf{A}}(\{a_{1},\dots,a_{n}\})=\operatorname{Sg}_{\mathbf{A}}(\{b_{1},\dots,b_{n}\})=A$. Then for all $f\in\operatorname{Squeezer}(\mathcal{F},\mathbf{A})$ and $t\in\mathcal{F}$ we have $|\{f(a_{1},\dots,a_{n},c)\mid c\in A\}|\leq|\{t(b_{1},\dots,b_{n},c)\mid c\in A\}|$.
Proof.
Since $\mathcal{F}$ is good for $\mathbf{A}$, there exists $t^{\prime}\in\mathcal{F}$ such that $t^{\prime}(a_{1},\dots,a_{n},x)\approx t(b_{1},\dots,b_{n},x)$. Then the required condition follows from the definition of a squeezer. ∎
Lemma 8.28.
Suppose $f,g_{1},g_{2},h$ are unary mappings on a finite set $A$ such that $|\operatorname{Im}(h)|=|\operatorname{Im}(f)|=|\operatorname{Im}(f\circ g_{1}\circ h)|=|\operatorname{Im}(f\circ g_{2}\circ h)|=k$. Then $(f\circ g_{1}\circ h)^{|A|!}=(f\circ g_{2}\circ h)^{|A|!}$.
Proof.
Let us denote the left-hand and the right-hand sides of the equality by $f_{1}$ and $f_{2}$ respectively. As $|A|!$ is divisible by the size of any cycle in $f\circ g_{1}\circ h$ and $f\circ g_{2}\circ h$, we have $f_{1}\circ f_{1}=f_{1}$ and $f_{2}\circ f_{2}=f_{2}$. Then the equality follows from $\operatorname{Im}(f_{1})=\operatorname{Im}(f_{2})=\operatorname{Im}(f)$ and $f_{1}(a)=f_{1}(b)\Leftrightarrow h(a)=h(b)\Leftrightarrow f_{2}(a)=f_{2}(b)$ for all $a,b\in A$. ∎
Lemma 8.29.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{k}$, $\mathcal{F}$ is a set of terms over the variables $x_{1},\dots,x_{n+1}$ closed under right composition, $\mathcal{F}$ is good for $\mathbf{A}_{1}$, $m\leq n$, $g\in\operatorname{Squeezer}(\mathcal{F},\mathbf{A}_{1},\dots,\mathbf{A}_{s})$, $\operatorname{Sg}_{\mathbf{A}_{1}}(\{a_{1},\dots,a_{n}\})=\operatorname{Sg}_{\mathbf{A}_{1}}(\{b_{1},\dots,b_{m}\})=A_{1}$. Then there exists a term $f\in\operatorname{Squeezer}(\mathcal{F},\mathbf{A}_{1},\dots,\mathbf{A}_{s})$ such that
$f^{\mathbf{A}_{1}}(b_{1},\dots,b_{m},y_{1},\dots,y_{n-m},x)\approx g^{\mathbf{A}_{1}}(a_{1},\dots,a_{n},x).$
Proof.
Since $\mathcal{F}$ is good for $\mathbf{A}_{1}$, there exists $g^{\prime}\in\mathcal{F}$ such that $g^{\prime}(b_{1},\dots,b_{m},y_{1},\dots,y_{n-m},x)\approx g(a_{1},\dots,a_{n},x)$. Choose any term $f^{\prime}\in\operatorname{Squeezer}(\mathcal{F},\mathbf{A}_{1},\dots,\mathbf{A}_{s}).$ Put $f=(g^{\prime}\circ f^{\prime}\circ g^{\prime})^{S!}$, where $S$ is the maximal size of an algebra in $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$. By Lemma 8.26(3), $f\in\operatorname{Squeezer}(\mathcal{F},\mathbf{A}_{1},\dots,\mathbf{A}_{s})$. Choose some $c_{1},\dots,c_{n-m}\in A_{j}$. Denote
$g_{0}^{\prime}(x)$ $:=g^{\prime}(b_{1},\dots,b_{m},c_{1},\dots,c_{n-m},x)=g(a_{1},\dots,a_{n},x),$
$f_{0}^{\prime}(x)$ $:=f^{\prime}(b_{1},\dots,b_{m},c_{1},\dots,c_{n-m},x).$
By Lemma 8.27, we have $|\operatorname{Im}(g_{0}^{\prime})|=|\operatorname{Im}(f_{0}^{\prime})|=|\operatorname{Im}(g_{0}^{\prime}\circ f_{0}^{\prime}\circ g_{0}^{\prime})|$. By property 2 of squeezers, we obtain $(g_{0}^{\prime}\circ g_{0}^{\prime}\circ g_{0}^{\prime})^{S!}=g_{0}^{\prime}$. Then Lemma 8.28 implies $(g_{0}^{\prime}\circ f_{0}^{\prime}\circ g_{0}^{\prime})^{S!}=(g_{0}^{\prime}\circ g_{0}^{\prime}\circ g_{0}^{\prime})^{S!}=g_{0}^{\prime}$, which is exactly the equality we need. ∎
Lemma 8.30.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{k}$, $\mathcal{F}$ is a set of terms over the variables $x_{1},\dots,x_{n+1}$ closed under right composition, $\mathcal{F}$ is good for $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$. Then there exists a finite set $\mathcal{F}_{0}\subseteq\operatorname{Squeezer}(\mathcal{F},\mathbf{A}_{1},\dots,\mathbf{A}_{s})$ such that for every $j\in[s]$, every $(\overline{k_{1}},\dots,\overline{k_{n}})$-palette tuple $(\mathbf{a}_{1},\dots,\mathbf{a}_{n})\in A_{j}^{k_{1}+\dots+k_{n}}$ generating $\mathbf{A}_{j}$, and every mapping $h:A_{j}\to A_{j}$ the following conditions are equivalent:
  1. (1)
    $\exists f\in\mathcal{F}_{0}\;\exists i_{1}\in[k_{1}]\dots\exists i_{n}\in[k_{n}]\colon f^{\mathbf{A}_{j}}(\mathbf{a}_{1}^{i_{1}},\dots,\mathbf{a}_{n}^{i_{n}},x)\approx h(x)$;
  2. (2)
    $\exists f\in\mathcal{F}_{0}\;\forall i_{1}\in[k_{1}]\dots\forall i_{n}\in[k_{n}]\colon f^{\mathbf{A}_{j}}(\mathbf{a}_{1}^{i_{1}},\dots,\mathbf{a}_{n}^{i_{n}},x)\approx h(x)$.
Proof.
For every $j\in[s]$ we build a set $\mathcal{F}_{j}$ as follows. For every $m\leq n$, every $g\in\operatorname{Squeezer}(\mathcal{F},\mathbf{A}_{1},\dots,\mathbf{A}_{s})$, and every $a_{1},\dots,a_{n},b_{1},\dots,b_{m}\in A_{j}$ such that $\operatorname{Sg}_{\mathbf{A}_{j}}(\{a_{1},\dots,a_{n}\})=\operatorname{Sg}_{\mathbf{A}_{j}}(\{b_{1},\dots,b_{m}\})=A_{j}$ we add to $\mathcal{F}_{j}$ a term $f\in\operatorname{Squeezer}(\mathcal{F},\mathbf{A}_{1},\dots,\mathbf{A}_{s})$ such that for all $c_{1},\dots,c_{n-m}\in A_{j}$
$f^{\mathbf{A}_{j}}(b_{1},\dots,b_{m},c_{1},\dots,c_{n-m},x)\approx g^{\mathbf{A}_{j}}(a_{1},\dots,a_{n},x),$
which exists by Lemma 8.29. Let $S_{n}$ be the set of all permutations on $[n]$. Put
$\mathcal{F}_{0}=\{f\mid\exists f^{\prime}\in\mathcal{F}_{1}\cup\dots\cup\mathcal{F}_{s},\sigma\in S_{n},f(x_{1},\dots,x_{n},z)\approx f^{\prime}(x_{\sigma(1)},\dots,x_{\sigma(n)},z)\}.$
Let us show that $\mathcal{F}_{0}$ satisfies the required properties. Assume that condition (1) is satisfied, that is, for some $f\in\mathcal{F}_{0}$ and $i_{1},\dots,i_{n}$ we have $f^{\mathbf{A}_{j}}(\mathbf{a}_{1}^{i_{1}},\dots,\mathbf{a}_{n}^{i_{n}},x)\approx h(x)$. Since $(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$ is a $(\overline{k_{1}},\dots,\overline{k_{n}})$-palette tuple, there should be constant blocks $\mathbf{a}_{j_{1}},\dots,\mathbf{a}_{j_{m}}$ such that elements of $\mathbf{a}_{j_{1}},\dots,\mathbf{a}_{j_{m}}$ generate $A_{j}$. Let each $\mathbf{a}_{j_{\ell}}$ consist of an element $b_{\ell}$ for every $\ell\in[m]$. Then the required term satisfying condition (2) comes from the term chosen for $\mathbf{a}_{1}^{i_{1}},\dots,\mathbf{a}_{n}^{i_{n}}$, $b_{1},\dots,b_{m}$ and a permutation $\sigma$ that maps $\ell$ to $j_{\ell}$ for every $\ell\in[m]$. ∎
We introduce an order on finite sets of finite algebras by comparing lexicographically the sizes of algebras. Formally, we say that $\{\mathbf{A}_{1},\dots,\mathbf{A}_{s}\}<\{\mathbf{B}_{1},\dots,\mathbf{B}_{m}\}$ if there exists $n_{0}$ such that
  1. 1.
    for every $n>n_{0}$ both $\{\mathbf{B}_{1},\dots,\mathbf{B}_{m}\}$ and $\{\mathbf{A}_{1},\dots,\mathbf{A}_{s}\}$ have the same number of algebras of size $n$;
  2. 2.
    the set $\{\mathbf{B}_{1},\dots,\mathbf{B}_{m}\}$ has more algebras of size $n_{0}$ than $\{\mathbf{A}_{1},\dots,\mathbf{A}_{s}\}$.
8.7 The existence of good sets of terms
In this section we show how to build good sets of terms that will help us to squeeze the domains later. The idea is to choose terms that preserve a congruence or unary relations in the last coordinate. For the linear and central case it is enough to consider all such terms to get a good family satisfying all the necessary properties. However, for the PC case we have to use more involved definition, and this is precisely the reason why we introduce good sets of terms.
Lemma 8.31.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{k}$, $\mathcal{B}_{i}$ is a set of subsets of $\mathbf{A}_{i}$ for every $i$, $\mathcal{T}$ is the set of all terms $t$ over the variables $x_{1},\dots,x_{n+1}$ such that $t^{\mathbf{A}_{i}}(a_{1},\dots,a_{n+1})\in B$ for any $i\in[s]$, $B\in\mathcal{B}_{i}$, $a_{1},\dots,a_{n}\in A_{i}$, $a_{n+1}\in B$. Then
  • (n)
    $\mathcal{T}$ is not empty;
  • (c)
    $\mathcal{T}$ is closed under right composition;
  • (g)
    $\mathcal{T}$ is good for $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$;
  • (s)
    for every $i\in[s]$ one of the following conditions holds:
    1. 1.
      there exists $t\in\mathcal{T}$ and a tuple $(a_{1},\dots,a_{n})$ generating $\mathbf{A}_{i}$ such that $|\{t^{\mathbf{A}_{i}}(a_{1},\dots,a_{n},b)\mid b\in A_{i}\}|<|A_{i}|$;
    2. 2.
      there is no tuple $(a_{1},\dots,a_{n})$ generating $\mathbf{A}_{i}$;
    3. 3.
      $\{t^{\mathbf{A}_{i}}(a_{1},\dots,a_{n},b)\mid b\in A_{i}\}=A_{i}$ for every $t\in\mathcal{T}$ and $a_{1},\dots,a_{n}\in A_{i}$.
Proof.
$\mathcal{T}$ is nonempty because the projection onto the variable $x_{n+1}$ belongs to $\mathcal{T}$. Property (c) follows immediately from the definition. To prove property (g) for $i\in[s]$, tuples $(a_{1},\dots,a_{n})$ and $(b_{1},\dots,b_{m})$ generating $\mathbf{A}_{i}$, and $f\in\mathcal{T}$, we can define a required term as follows. For every $j\in[n]$ choose a term $t_{j}$ such that $t_{j}^{\mathbf{A}_{i}}(b_{1},\dots,b_{m})=a_{j}$ Then the term $f(t_{1}(x_{1},\dots,x_{m}),\dots,t_{n}(x_{1},\dots,x_{m}),x_{n+1})$ with dummy variables $x_{m+1},\dots,x_{n}$ belongs to $\mathcal{T}$ and satisfies the required condition.
Let us prove property (s). If conditions 2 or 3 are satisfied then we are done. Otherwise, consider a tuple $(b_{1},\dots,b_{n})$ generating $\mathbf{A}_{i}$, a tuple $(a_{1},\dots,a_{n})$ and a term $t\in\mathcal{T}$ such that $\{t^{\mathbf{A}_{i}}(a_{1},\dots,a_{n},b)\mid b\in A_{i}\}\neq A_{i}$. Again, for every $j\in[n]$ choose a term $t_{j}$ such that $t_{j}^{\mathbf{A}_{i}}(b_{1},\dots,b_{n})=a_{j}$. Then the term $t^{\prime}(x_{1},\dots,x_{n+1})\approx t(t_{1}(x_{1},\dots,x_{n}),\dots,t_{n}(x_{1},\dots,x_{n}),x_{n+1})$ and a tuple $(b_{1},\dots,b_{n})$ witness condition 1. ∎
Corollary 8.32.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{k}$, $\sigma$ is a maximal linear congruence on $\mathbf{A}_{1}$, $\mathcal{T}$ is the set of all terms $t$ over the variables $x_{1},\dots,x_{n+1}$ such that $t^{\mathbf{A}_{1}/\sigma}(x_{1},\dots,\dots,x_{n+1})\approx x_{n+1}$. Then $\mathcal{T}$ satisfies properties (n),(c),(g), and (s) from Lemma 8.31.
Proof.
It is sufficient to set $\mathcal{B}_{1}$ to be the set of all equivalence classes of $\sigma$, $\mathcal{B}_{i}=\varnothing$ for $i\geq 2$, and apply Lemma 8.31. ∎
Corollary 8.33.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{k}$, $\mathcal{T}$ is the set of all terms $t$ over the variables $x_{1},\dots,x_{n+1}$ such that $t^{\mathbf{A}_{i}}(a_{1},\dots,a_{n+1})\in B$ for any $i\in[s]$, $B<_{\mathcal{C}}\mathbf{A}_{i}$, $a_{1},\dots,a_{n}\in A_{i}$, $a_{n+1}\in B$. Then $\mathcal{T}$ satisfies properties (n),(c),(g), and (s) from Lemma 8.31.
Proof.
It is sufficient to set $\mathcal{B}_{i}$ to be the set of all proper central subuniverses of $\mathbf{A}_{i}$ for every $i\in[s]$ and apply Lemma 8.31. ∎
A term $t$ is called strongly stable for an algebra $\mathbf{A}$ if $\{t^{\mathbf{A}}(a_{1},\dots,a_{n})\}\lll\mathbf{A}$ for any tuple $(a_{1},\dots,a_{n})$ generating $\mathbf{A}$.
Lemma 8.34.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{k}$ are BA and center free algebras, $\mathcal{T}$ is the set of all terms $t$ over the variables $x_{1},\dots,x_{n+1}$ such that
  • $t$ is strongly stable in $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$;
  • for every $i\in[s]$, every maximal PC congruence $\sigma$ on $\mathbf{A}_{i}$, and every tuple $(a_{1},\dots,a_{n})$ generating $\mathbf{A}_{i}$ we have $t^{\mathbf{A}_{i}/\sigma}(a_{1}/\sigma,\dots,a_{n}/\sigma,x)\approx x$.
Then
  • (n)
    $\mathcal{T}$ is not empty;
  • (c)
    $\mathcal{T}$ is closed under right composition;
  • (g)
    $\mathcal{T}$ is good for $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$;
  • (s)
    for every $i\in[s]$ one of the following conditions holds:
    1. 1.
      there exists $t\in\mathcal{T}$ and a tuple $(a_{1},\dots,a_{n})$ generating $\mathbf{A}_{i}$ such that $|\{t^{\mathbf{A}_{i}}(a_{1},\dots,a_{n},b)\mid b\in A_{i}\}|<|A_{i}|$;
    2. 2.
      there is no tuple $(a_{1},\dots,a_{n})$ generating $\mathbf{A}_{i}$;
    3. 3.
      $\mathbf{A}_{i}\cong\mathbf{B}_{1}\times\dots\times\mathbf{B}_{\ell}$ for some PC algebras $\mathbf{B}_{1},\dots,\mathbf{B}_{\ell}$.
Proof.
(n). Consider the free generated relation $R^{n+1}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$. Let $\sigma_{i}$ be the intersection of all the maximal PC congruences on $\mathbf{A}_{i}$ for each $i\in[s]$. Let $D^{(\top)}$ be the reduction defined by $D^{(\top)}_{(\mathbf{A}_{i},(a_{1},\dots,a_{n+1}))}=a_{n+1}/\sigma_{i}$, where $a_{n+1}/\sigma_{i}$ is the equivalence class of $\sigma_{i}$ containing $a_{n+1}$. By Lemma 8.16, $D^{(\top)}_{(\mathbf{A}_{i},(a_{1},\dots,a_{n+1}))}\lll\mathbf{A}_{i}$. Note that $R^{n+1}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\top)}$ is not empty because the intersection contains the $(n+1)$-th generator. Using Lemmas 8.15 and 8.18, we can find a reduction $D^{(\bot)}\lll D^{(\top)}$ such that $|R^{n+1}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\bot)}|=1$. The term defining the only tuple in this intersection belongs to $\mathcal{T}$ and witnesses that it is not empty.
(c). Checking this property is straightforward.
(g). Let $f\in\mathcal{T}$, $i\in[s]$, $(a_{1},\dots,a_{n})$ and $(b_{1},\dots,b_{m})$ generate $\mathbf{A}_{i}$. Choose terms $t_{1},\dots,t_{j}$ such that $t_{j}^{\mathbf{A}_{i}}(b_{1},\dots,b_{m})=a_{i}$ for every $j\in[n]$. Let $t^{\prime}(x_{1},\dots,x_{n})\approx f(t_{1}(x_{1},\dots,x_{m}),\dots,t_{n}(x_{1},\dots,x_{m}),x_{n+1})$, where variables $x_{m+1},\dots,x_{n}$ are dummy. To define the required term consider the free generated relation $R^{n+1}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$. Let $D^{(\top)}$ be the reduction defined by $D^{(\top)}_{(\mathbf{A}_{j},(c_{1},\dots,c_{n+1}))}=c_{n+1}/\sigma_{j}$ for any $j\in[s]$ and $c_{1},\dots,c_{n+1}\in A_{j}$. Since the $(n+1)$-th generator of $R^{n+1}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ is in $D^{(\top)}$, $R^{n+1}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\top)}\neq\varnothing$. Let $D^{(\triangle)}$ be the reduction defined by $D^{(\triangle)}_{(\mathbf{A}_{i},(b_{1},\dots,b_{n+1}))}=\{f(a_{1},\dots,a_{n},b_{n+1})\}$ for any $b_{m+1},\dots,b_{n+1}\in A_{i}$. Since $f$ is strongly stable in $\mathbf{A}_{i}$, $D^{(\triangle)}_{(\mathbf{A}_{i},(b_{1},\dots,b_{n+1}))}\lll\mathbf{A}_{i}$. Since $t^{\prime}$ applied to $R^{n+1}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ gives a tuple from $D^{(\triangle)}$, the intersection $R^{n+1}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\triangle)}$ is not empty. Assume that $R^{n+1}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\top)}\cap D^{(\triangle)}=\varnothing$. By Lemma 8.17, there exists an equivalence class $C$ of a PC congruence on $\mathbf{A}_{i}$ and $(c_{1},\dots,c_{n+1})$ such that $c_{n+1}\in C$ and $R^{n+1}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\top)}$ has no tuples whose $(\mathbf{A}_{i},(c_{1},\dots,c_{n+1}))$-coordinate is from $C$. This cannot happen as $D^{(\top)}_{(\mathbf{A}_{i},(c_{1},\dots,c_{n+1}))}\subseteq C$. Thus, $R^{n+1}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\top)}\cap D^{(\triangle)}\neq\varnothing$. It remains to use Lemmas 8.15 and 8.18, to find a reduction $D^{(\bot)}\lll D^{(\top)}\cap D^{(\triangle)}$ such that $|R^{n+1}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\bot)}|=1$. The term defining the only tuple in this intersection satisfies the required conditions.
(s). If conditions 2 or 3 are satisfied then we are done. Otherwise, choose a tuple $(a_{1},\dots,a_{n})$ generating $\mathbf{A}_{i}$ and consider the free generated relation $R^{n+1}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$. As before, we define a reduction $D^{(\top)}$ by $D^{(\top)}_{(\mathbf{A}_{j},(c_{1},\dots,c_{n+1}))}=c_{n+1}/\sigma_{j}$ for any $j\in[s]$ and $c_{1},\dots,c_{n+1}\in A_{j}$. Since $\mathbf{A}_{i}$ is not a product of PC algebras, $\sigma_{i}\neq 0_{\mathbf{A}_{i}}$. Choose an equivalence class $C$ of $\sigma_{i}$ containing more than one element. Using Lemma 8.15 choose an element $e$ such that $\{e\}\lll^{\mathbf{A}_{i}}C$. Let us define a reduction $D^{(\triangle)}$ by $D^{(\triangle)}_{(\mathbf{A}_{i},(a_{1},\dots,a_{n},d))}=\{e\}$ for any $d\in C$. Notice that $R^{n+1}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\triangle)}$ is not empty because the elements $a_{1},\dots,a_{n}$ generate the whole $\mathbf{A}_{i}$ including the element $e$. Let us show that $R^{n+1}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\top)}\cap D^{(\triangle)}$ is also not empty. This follows from Lemma 8.17 and the fact that there is no maximal PC congruence that cuts the class $C$. Again, it remains to use Lemmas 8.15 and 8.18 to find a reduction $D^{(\bot)}\lll D^{(\top)}\cap D^{(\triangle)}$ such that $|R^{n+1}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\bot)}|=1$. The term defining the only tuple in this intersection satisfies the required conditions. ∎
8.8 Divide and conquer
First, we will show that the families of terms we defined earlier actually squeeze the domains unless the algebras satisfy very strong property. Second, we prove a technical claim that allows us to combine freely a term operation on one PC algebra and a strongly stable term on another algebra.
Lemma 8.35.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{k}$, $j\in[s]$, $\{c\}$ is not a central subuniverse of $\mathbf{A}_{j}$ for some $c\in A_{j}$, $(c_{1},\dots,c_{n})$ is a tuple generating $\mathbf{A}_{j}$. Then there exists a term $t$ such that
  1. 1.
    $t^{\mathbf{A}_{i}}(a_{1},\dots,a_{n},a_{n+1})\in C$ for all $i\in[s]$, $C<_{\mathcal{C}}\mathbf{A}_{i}$, $a_{1},\dots,a_{n}\in A_{i}$, and $a_{n+1}\in C$;
  2. 2.
    $\{t^{\mathbf{A}_{j}}(c_{1},\dots,c_{n},b)\mid b\in A_{j}\}\neq A_{j}$.
Proof.
First, let us add all subalgebras $\mathbf{A}_{s+1},\dots,\mathbf{A}_{s^{\prime}}$ of the algebras $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$ to the list. By Lemma 8.10, $C\leq_{\mathcal{C}}\mathbf{A}$ and $B\leq\mathbf{A}$ imply $C\cap B\leq_{\mathcal{C}}\mathbf{B}$. Then, having all subalgebras in the list, it is sufficient to prove condition 1 only for tuples $(a_{1},\dots,a_{n},a_{n+1})$ generating $\mathbf{A}_{i}$ for each $i\in[s^{\prime}]$.
Let us consider the free generated relation $R^{n+1}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s^{\prime}}}$. Let $\gamma_{1},\dots,\gamma_{n+1}$ be the generators of this relation. For every $(\mathbf{A}_{i},\alpha)\in I_{\mathbf{A}_{1},\dots,\mathbf{A}_{s^{\prime}}}^{n+1}$ by $D^{(\top)}_{(\mathbf{A}_{i},(a_{1},\dots,a_{n+1}))}$ we denote the minimal central subuniverse of $\mathbf{A}_{i}$ containing $a_{n+1}$, which can be equal to $A_{i}$. It exists because the intersection of two central subuniverses is central (Lemma 8.9). Let $C$ be the minimal central subuniverse of $\mathbf{A}_{j}$ containing $c$. By Lemma 8.14, $\operatorname{pr}_{(\mathbf{A}_{i},(c_{1},\dots,c_{n},c))}(R^{n+1}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s^{\prime}}}\cap D^{(\top)})\dot{\leq}_{\mathcal{C}}\mathbf{A}_{i}$. Since $R^{n+1}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s^{\prime}}}\cap D^{(\top)}$ contains the tuple $\gamma_{n+1}$, it is not empty. Then, by the minimality of the central subuniverses we have $\operatorname{pr}_{(\mathbf{A}_{i},(c_{1},\dots,c_{n},c))}(R^{n+1}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s^{\prime}}}\cap D^{(\top)})=C$. By Lemma 8.15 we have one of the two cases:
Case 1. There exists $B<_{\mathcal{C}}^{A_{j}}C$. Choose a tuple $\gamma\in R^{n+1}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s^{\prime}}}\cap D^{(\top)}$ such that $\gamma(\mathbf{A}_{j},(c_{1},\dots,c_{n},c))\in B$ and the corresponding term $t$ generating this tuple from $\gamma_{1},\dots,\gamma_{n+1}$. Since $\gamma\in D^{(\top)}$, the term $t$ satisfies the required condition 1. It remains to show that $h(x)\approx t^{\mathbf{A}_{j}}(c_{1},\dots,c_{n},x)$ is not a bijection. By Lemma 8.13, $B<_{\mathcal{C}}\mathbf{A}_{j}$, then condition 1 implies that $h(B)\subseteq B$. Since $h(c)\in B$ and $c\notin B$, the operation $h$ cannot be bijective.
Case 2. There exists $B<_{T}^{A_{j}}C$ for some $T\in\{\mathcal{BA},\mathcal{PC},\mathcal{L}\}$. Let $D^{(\bot)}$ be defined by $D^{(\bot)}_{(\mathbf{A}_{j},(a_{1},\dots,a_{n},d))}=B$ for every $d\in C$ and $D^{(\bot)}_{(\mathbf{A}_{i},\alpha)}=A_{i}$ for all other coordinates $(\mathbf{A}_{i},\alpha)\in I^{n+1}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s^{\prime}}}$. Since the tuple $(c_{1},\dots,c_{n})$ generates $\mathbf{A}_{j}$, $R^{n+1}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s^{\prime}}}\cap D^{(\bot)}\neq\varnothing$. Combining this with $R^{n+1}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s^{\prime}}}\cap D^{(\top)}\neq\varnothing$ and Lemma 8.17, we derive that $R^{n+1}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s^{\prime}}}\cap D^{(\top)}\cap D^{(\bot)}\neq\varnothing$. Then any term generating a tuple from this nonempty intersection satisfies the required conditions 1 and 2. ∎
Lemma 8.36.
Suppose $p$ is a prime, $\mathbf{A}\in\mathcal{V}_{k}$, $0_{\mathbf{A}}$ is $\wedge$-irreducible, there is no bridge from $0_{\mathbf{A}}$ to $0_{\mathbf{Z}_{p}}$, $(a_{1},\dots,a_{n})$ is a tuple generating $\mathbf{A}$. Then there exists a term $t$ such that
  1. 1.
    $t^{\mathbf{Z}_{p}}(x_{1},\dots,x_{n+1})\approx x_{n+1}$;
  2. 2.
    $\{t^{\mathbf{A}}(a_{1},\dots,a_{n},b)\mid b\in A\}\neq A$.
Proof.
Let us consider the free generated relation $R^{n+1}_{\mathbf{Z}_{p},\mathbf{A}}$. Let $\gamma_{1},\dots,\gamma_{n+1}$ be the generators for this relation. Choose two different elements $c$ and $c^{\prime}$ such that $(c,c^{\prime})\in 0_{\mathbf{A}}^{+}$. Let us define a reduction $D^{(\top)}$ as follows. For $(\mathbf{Z}_{p},\alpha)\in I^{n+1}_{\mathbf{Z}_{p},\mathbf{A}}$ let $D^{(\top)}_{(\mathbf{Z}_{p},\alpha)}=\{\alpha(n+1)\}$, and $D^{(\top)}_{(\mathbf{A},\alpha)}=A$ for $(\mathbf{A},\alpha)\in I^{n+1}_{\mathbf{Z}_{p},\mathbf{A}}$. We define a reduction $D^{(\bot)}$ by $D_{(\mathbf{A},\alpha)}^{(\bot)}=\operatorname{pr}_{(\mathbf{A},\alpha)}(R^{n+1}_{\mathbf{Z}_{p},\mathbf{A}}\cap D^{(\top)})$ for $(\mathbf{A},\alpha)\in I^{n+1}_{\mathbf{Z}_{p},\mathbf{A}}$, and $D_{(\mathbf{Z}_{p},\alpha)}^{(\bot)}=D_{(\mathbf{Z}_{p},\alpha)}^{(\top)}$ for $(\mathbf{Z}_{p},\alpha)\in I^{n+1}_{\mathbf{Z}_{p},\mathbf{A}}$. By Lemma 8.18, $D^{(\bot)}\lll D^{(\top)}$, and $R^{n+1}_{\mathbf{Z}_{p},\mathbf{A}}\cap D^{(\bot)}$ is not empty as it contains $\gamma_{n+1}$.
By Lemma 8.15 choose $\{b\}\lll^{A}D^{(\bot)}_{(\mathbf{A},(a_{1},\dots,a_{n},c))}$. If there exists a tuple in $R^{n+1}_{\mathbf{Z}_{p},\mathbf{A}}\cap D^{(\top)}$ such that its coordinates $(\mathbf{A},(a_{1},\dots,a_{n},c))$ and $(\mathbf{A},(a_{1},\dots,a_{n},c^{\prime}))$ are equal to $b$, then the term corresponding to this tuple is a required term. Otherwise, choose a minimal $C$ such that $\{b\}\lll^{A}C\lll^{A}A$ and $R^{n+1}_{\mathbf{Z}_{p},\mathbf{A}}\cap D^{(\top)}$ contains a tuple $\gamma$ whose $(\mathbf{A},(a_{1},\dots,a_{n},c))$-th coordinate equals $b$ and $(\mathbf{A},(a_{1},\dots,a_{n},c^{\prime}))$-th coordinate belongs to $C$. Choose $B$ such that $\{b\}\lll^{A}B<_{T(\delta)}^{\mathbf{A}}C$. Since $b\in\operatorname{Sg}_{\mathbf{A}}(\{a_{1},\dots,a_{n}\})$, $R^{n+1}_{\mathbf{Z}_{p},\mathbf{A}}$ contains a tuple whose coordinates $(\mathbf{A},(a_{1},\dots,a_{n},c))$ and $(\mathbf{A},(a_{1},\dots,a_{n},c^{\prime}))$ are equal to $b$. Then by Lemma 8.17, $T=\mathcal{L}$ and there is a bridge from $\delta$ to $0_{\mathbf{Z}_{p}}$. Since $(c,c^{\prime})\in 0_{\mathbf{A}}^{+}$, we have
$(\gamma(\mathbf{A},(a_{1},\dots,a_{n},c)),\gamma(\mathbf{A},(a_{1},\dots,a_{n},c^{\prime})))\in 0_{\mathbf{A}}^{+}\setminus\delta.$
Hence, $\delta\not\supseteq 0_{\mathbf{A}}^{+}$ and by the $\wedge$-irreducibility of $0_{\mathbf{A}}$ we get $\delta=0_{\mathbf{A}}$. The existence of a bridge between $0_{\mathbf{A}}$ and $0_{\mathbf{Z}_{p}}$ contradicts our assumptions and completes the proof.
Corollary 8.37.
Suppose $\mathbf{A}_{1},\mathbf{A}_{2}\in\mathcal{V}_{k}$, $\sigma$ is a maximal congruence on $\mathbf{A}_{1}$ of linear type, $\omega$ is a $\wedge$-irreducible congruence on $\mathbf{A}_{2}$, there is no bridge from $\omega$ to $\sigma$, $(a_{1},\dots,a_{n})$ is a tuple generating $\mathbf{A}_{2}$. Then there exists a term $t$ such that
  1. 1.
    $t^{\mathbf{A}_{1}/\sigma}(x_{1},\dots,x_{n+1})\approx x_{n+1}$;
  2. 2.
    $\{t^{\mathbf{A}_{2}}(a_{1},\dots,a_{n},b\mid b\in A_{2}\}\neq A_{2}$.
Proof.
By Lemma 8.3, $\mathbf{A}_{1}/\sigma\cong\mathbf{Z}_{p}$ for some prime $p$. It remains to apply Lemma 8.36 to the algebras $\mathbf{A}_{1}/\sigma$ and $\mathbf{A}_{2}/\omega$. ∎
Lemma 8.38.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{q}$, $\sigma$ is a maximal congruence on $\mathbf{A}_{1}$ of linear type, there does not exist a term $m$ such that $m^{\mathbf{A}_{i}}$ is a Maltsev operation for every $i\in[s]$. Then there exists a term $t$ such that
  1. 1.
    $t^{\mathbf{A}_{i}/\sigma}(x,y)\approx y$;
  2. 2.
    $\{t^{\mathbf{A}_{j}}(a,b\mid b\in A_{j}\}\neq A_{j}$ for some $a\in A_{j}$.
Proof.
By Lemma 8.3, $\mathbf{A}_{1}/\sigma\cong\mathbf{Z}_{p}$ for some prime $p$. Let us consider a term $g_{1}$ such that $g_{1}^{\mathbf{A}_{1}/\sigma}(x_{1},\dots,x_{p+1})\approx x_{1}+\dots+x_{p+1}$. For every $i\in\mathbb{N}$ we define $g_{i+1}$ as a composition of $g_{1}$ with $p+1$ copies of $g_{i}$ with different variables. So, $g_{i}$ is of arity $(p+1)^{i}$ for every $i$. Let $N=\max\limits_{i\in[s]}(|A_{i}|)$. Then $g_{N!}^{\mathbf{A}_{i}}(x,\dots,x,g_{N!}^{\mathbf{A}_{i}}(x,\dots,x,y))=g_{N!}^{\mathbf{A}_{i}}(x,\dots,x,y)$ for every $i$ and $g_{N!}^{\mathbf{A}_{1}/\sigma}(x,x,\dots,x,y)=g_{N!}^{\mathbf{A}_{1}/\sigma}(y,x,\dots,x,x)=y$. Put $t_{1}(x,y)=g_{N!}(x,\dots,x,y)$, $t_{2}(x,y)=g_{N!}(y,x,\dots,x)$, and $m(x,y,z)=g_{N!}(x,y,\dots,y,z)$. Then $m^{\mathbf{A}_{i}}$ is a Maltsev operation for every $i\in[s]$, or one of the $t_{1}$ or $t_{2}$ satisfies the required conditions 1 and 2. ∎
Lemma 8.39.
Suppose $\mathbf{B},\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{k}$, $\mathbf{B}$ is a BA and center free PC algebra, and
  1. (1)
    $\mathbf{B}$ is not isomorphic to $\mathbf{A}_{i}/\sigma$ for any $i\in[s]$ and any congruence $\sigma$;
  2. (2)
    $t_{1}$ is an $n$-ary term that is strongly stable in $\mathbf{A}_{i}$ for every $i\in[s]$;
  3. (3)
    $t_{2}$ is an $m$-ary term.
Then there exists an $(m+n)$-ary term $t$ such that
  • (a)
    $t^{\mathbf{A}_{i}}(b_{1},\dots,b_{m},a_{1},\dots,a_{n})=t_{1}^{\mathbf{A}_{i}}(a_{1},\dots,a_{n})$ for every $i\in[s]$ and any tuples $(b_{1},\dots,b_{m})$ and $(a_{1},\dots,a_{n})$ generating $\mathbf{A}_{i}$,
  • (b)
    $t^{\mathbf{B}}(b_{1},\dots,b_{m},a_{1},\dots,a_{n})=t_{2}^{\mathbf{B}}(b_{1},\dots,b_{m})$ for any tuples $(b_{1},\dots,b_{m})$ and $(a_{1},\dots,a_{n})$ generating $\mathbf{B}$.
Proof.
Let $R^{m+n}_{\mathbf{B},\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$ be the free generated relation. Let us define a reduction $D^{(\top)}$ by $D^{(\top)}_{(\mathbf{B},(b_{1},\dots,b_{m},a_{1},\dots,a_{n}))}=\{t_{2}^{\mathbf{B}}(b_{1},\dots,b_{m})\}$ when $(a_{1},\dots,a_{n})$ and $(b_{1},\dots,b_{m})$ generate $\mathbf{B}$, and $D^{(\top)}_{(\mathbf{C},\alpha)}=C$ otherwise. Since $\mathbf{B}$ is a BA and center free PC algebra, $D_{(\mathbf{B},(b_{1},\dots,b_{m},a_{1},\dots,a_{n}))}^{(\top)}<_{\mathcal{PC}}\mathbf{B}$. Similarly, define a reduction $D^{(\bot)}$ by $D^{(\bot)}_{(\mathbf{A}_{i},(b_{1},\dots,b_{m},a_{1},\dots,a_{n}))}=\{t_{1}^{\mathbf{A}_{i}}(a_{1},\dots,a_{n})\}$ when $(a_{1},\dots,a_{n})$ and $(b_{1},\dots,b_{m})$ generate $\mathbf{A}_{i}$, and $D^{(\bot)}_{(\mathbf{C},\alpha)}=C$ otherwise. The terms $t_{1}$ and $t_{2}$ witness that $R^{m+n}_{\mathbf{B},\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\top)}$ and $R^{m+n}_{\mathbf{B},\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\bot)}$ are not empty. If $R^{m+n}_{\mathbf{B},\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\top)}\cap D^{(\bot)}=\varnothing$ then Lemma 8.17 implies that there exists a congruence $\sigma$ on some $\mathbf{A}_{i}$ such that $\mathbf{A}_{i}/\sigma\cong\mathbf{B}$, which contradicts condition (1). Thus, $R^{m+n}_{\mathbf{B},\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\top)}\cap D^{(\bot)}\neq\varnothing$. It remains to choose any tuple from this nonempty intersection, then a term $t$ generating this tuple satisfies the required conditions. ∎
Lemma 8.40.
Suppose $t_{1}$ is an $n$-ary strongly stable term in algebras $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{k}$, then for any $i\in[s]$, $m\leq n$, and tuples $(b_{1},\dots,b_{m})$ and $(a_{1},\dots,a_{n})$ generating $\mathbf{A}_{i}$ there exists a term $t_{2}$ such that
  • (s)
    $t_{2}$ is strongly stable in $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$,
  • (g)
    $t_{2}^{\mathbf{A}_{i}}(b_{1},\dots,b_{m},c_{1},\dots,c_{n-m})=t_{1}^{\mathbf{A}_{i}}(a_{1},\dots,a_{n})$ for any $c_{1},\dots,c_{n-m}\in A_{i}$.
Proof.
Since $a_{1},\dots,a_{n}\in\operatorname{Sg}_{\mathbf{A}_{i}}(\{b_{1},\dots,b_{m}\})$, there are $m$-ary terms $g_{1},\dots,g_{n}$ such that $g_{j}^{\mathbf{A}_{i}}(b_{1},\dots,b_{m})=a_{j}$ for every $j\in[n]$. Define $t_{3}(x_{1},\dots,x_{m})=t_{1}(g_{1}(x_{1},\dots,x_{m}),\dots,g_{n}(x_{1},\dots,x_{m}))$. Consider the free generated relation $R^{n}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}$. Let us define a reduction $D^{(\top)}$ by $D^{(\top)}_{(\mathbf{A}_{i},(b_{1},\dots,b_{m},c_{1},\dots,c_{n-m}))}=\{t_{3}^{\mathbf{B}}(b_{1},\dots,b_{m})\}$ for all $c_{1},\dots,c_{n-m}\in A_{i}$, and $D^{(\top)}_{(\mathbf{C},\alpha)}=C$ otherwise. Notice that $\{t_{3}^{\mathbf{A}_{i}}(b_{1},\dots,b_{m})\}=\{t_{1}^{\mathbf{A}_{i}}(a_{1},\dots,a_{n})\}\lll\mathbf{A}_{i}$. The term $t_{3}$ witnesses that $R^{n}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\top)}\neq\varnothing$. Using Lemmas 8.15 and 8.18, we can find a reduction $D^{(\bot)}\lll D^{(\top)}$ such that $|R^{n}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\bot)}|=1$. It remains to choose a term $t_{2}$ generating the only tuple in $R^{n}_{\mathbf{A}_{1},\dots,\mathbf{A}_{s}}\cap D^{(\bot)}$. ∎
8.9 Three ways to reduce the domain
In this section we show how squeezers can be used to reduce domains and define a PS operation using PS operations on smaller domains. First, we will show how to do this for a very special PC case even without squeezers. Then, using squeezers we do this for the linear case and remaining PC cases. Finally, we present the construction for the central case, which is the most complicated one.
Theorem 8.41.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{p}$, $|\mathbf{A}_{1}|\geq|\mathbf{A}_{j}|$ for all $j\in[s]$, $\mathbf{A}_{1}$ is a BA and center free PC algebra, and
  • (i)
    for every $\{\mathbf{B}_{1},\dots,\mathbf{B}_{u}\}<\{\mathbf{A}_{1},\dots,\mathbf{A}_{s}\}$ and every $m,n\in\mathbb{N}$ there exists a $(\underbrace{\overline{k^{n}},\dots,\overline{k^{n}}}_{m})$-PS term for $\mathbf{B}_{1},\dots,\mathbf{B}_{u}$.
Then $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$ admit a $(\underbrace{\overline{k},\dots,\overline{k}}_{n})$-PS term for every $n\in\mathbb{N}$.
Proof.
First, assume that $\mathbf{A}_{i}$ is not isomorphic to $\mathbf{A}_{1}$ for any $i\geq 2$ as otherwise we could exclude $\mathbf{A}_{i}$ from the list and use assumption (i). Let $\mathbf{A}_{s+1},\dots,\mathbf{A}_{s^{\prime}}$ be the set of all subalgebras of $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$.
Let $f_{1},\dots,f_{m}$ be a set of $n$-ary terms that are strongly stable in $\mathbf{A}_{2},\dots,\mathbf{A}_{s^{\prime}}$ such that for any other strongly stable term $f$ there exists $j\in[m]$ such that $f_{j}^{\mathbf{A}_{i}}=f^{\mathbf{A}_{i}}$ for all $i\in[s^{\prime}]$. Thus, these terms should cover all possible (strongly stable) term operations on each $\mathbf{A}_{i}$.
Recall that local palette symmetric terms in $\mathbf{A}$ differ from ordinary palette symmetric terms in that the symmetry is required only on tuples that generate $\mathbf{A}$. By Lemma 8.23, there exists a local $(\underbrace{\overline{k},\dots,\overline{k}}_{n})$-PS term $t_{1}$ for $\mathbf{B}$. Using Lemma 8.39 we build terms $g_{1},\dots,g_{m}$ of arity $k\cdot n+n$ such that for every $j\in[m]$
  • (a)
    $g_{j}^{\mathbf{A}_{i}}(b_{1},\dots,b_{k\cdot n},a_{1},\dots,a_{n})=f_{j}^{\mathbf{A}_{i}}(a_{1},\dots,a_{n})$ for every $i\in[s^{\prime}]$ and any tuples $(b_{1},\dots,b_{k\cdot n})$ and $(a_{1},\dots,a_{n})$ generating $\mathbf{A}_{i}$,
  • (b)
    $g_{j}^{\mathbf{A}_{1}}(b_{1},\dots,b_{k\cdot n},a_{1},\dots,a_{n})=t_{1}^{\mathbf{A}_{1}}(b_{1},\dots,b_{k\cdot n})$ for any tuples $(b_{1},\dots,b_{k\cdot n})$ and $(a_{1},\dots,a_{n})$ generating $\mathbf{A}_{1}$.
By condition (i), we can find a $(\underbrace{\overline{k^{n}},\dots,\overline{k^{n}}}_{m})$-PS term $t_{2}$ for $\mathbf{A}_{2},\dots,\mathbf{A}_{s^{\prime}}$. Then the required $(\underbrace{\overline{k},\dots,\overline{k}}_{n})$-PS term can be defined by
$t(\mathbf{x}_{1},\dots,\mathbf{x}_{n}):=t_{2}((g_{1}(\mathbf{x}_{1},\dots,\mathbf{x}_{n},\mathbf{x}_{1}^{j_{1}},\dots,\mathbf{x}_{n}^{j_{n}}))_{j_{1},\dots,j_{n}\in[k]},\dots,$
$\dots,(g_{m}(\mathbf{x}_{1},\dots,\mathbf{x}_{n},\mathbf{x}_{1}^{j_{1}},\dots,$ $\mathbf{x}_{n}^{j_{n}}))_{j_{1},\dots,j_{n}\in[k]}).$
Here, the first $k^{n}$ arguments of $t_{2}$ are filled with $g_{1}$, where $\mathbf{x}_{1}^{j_{1}},\dots,\mathbf{x}_{n}^{j_{n}}$ are substituted into the last $n$ coordinates for $j_{1},\dots,j_{n}\in[k]$. The next $k^{n}$ arguments are filled with $g_{2}$ and so on. Let us show that $t$ is in fact a $(\overline{k},\dots,\overline{k})$-PS term for $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$. Since we added all the subalgebras to the list, it is sufficient to prove that $t$ is local $(\overline{k},\dots,\overline{k})$-PS term for each $i\in[s^{\prime}]$. Let $(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$ be a $(\overline{k},\dots,\overline{k})$-palette tuple generating $A_{i}$. Notice $(\mathbf{a}_{1}^{j_{1}},\dots,\mathbf{a}_{n}^{j_{n}})$ also generates $\mathbf{A}_{i}$ for all $j_{1},\dots,j_{n}\in[k]$. First, we consider the case when $i=1$. By condition (b), $g_{i}^{\mathbf{A}_{1}}(\mathbf{a}_{1},\dots,\mathbf{a}_{n},\mathbf{a}_{1}^{j_{1}},\dots,\mathbf{a}_{n}^{j_{n}})=t_{1}^{\mathbf{A}_{1}}(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$. Hence, the result does not depend on $j_{1},\dots,j_{n}$ and we substitute only one value in $t_{2}$. Since, $t_{1}$ is local $(\overline{k},\dots,\overline{k})$-PS in $\mathbf{A}_{1}$, permutations in $\mathbf{a}_{1},\dots,\mathbf{a}_{n}$ do not change this one value. Thus, we proved the required condition for such tuples.
Consider the case when $i\geq 2$. Property (a) implies that each $g_{j}^{\mathbf{A}_{i}}$ on such tuples only depends on the last $n$ coordinates, hence permutation of elements inside blocks of $(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$ only imply permutations in the corresponding blocks of size $k^{n}$ of $t_{2}$ but the content of each block does not depend on the permutations. Since $t_{2}$ is $(\overline{k^{n}},\dots,\overline{k^{n}})$-PS in $\mathbf{A}_{i}$, it remains to prove that the tuple we substitute in $t_{2}^{\mathbf{A}_{i}}$ is $(\overline{k^{n}},\dots,\overline{k^{n}})$-palette.
If a value $c\in A_{i}$ appears in the tuple we substitute in $t_{2}^{\mathbf{A}_{i}}$, then there should be $j\in[m]$ and a tuple $(c_{1},\dots,c_{n})$ generating $\mathbf{A}_{i}$ such that $f_{j}^{\mathbf{A}_{i}}(c_{1},\dots,c_{n})=c$. Since the tuple $(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$ is palette, there exist constant blocks $\mathbf{a}_{\ell_{1}},\dots,\mathbf{a}_{\ell_{v}}$ containing all the elements $c_{1},\dots,c_{n}$. By Lemma 8.40 there exists $j^{\prime}\in[m]$ such that $f_{j^{\prime}}(\mathbf{a}_{1}^{j_{1}},\dots,\mathbf{a}_{n}^{j_{n}})=f_{j}(b_{1},\dots,b_{n})$ for all $j_{1},\dots,j_{n}\in[k]$. Hence, the $j^{\prime}$-th block of the tuple we substitute in $t_{2}$ consists of the element $c$. Thus, this tuple is palette. ∎
Theorem 8.42.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{p}$, $\sigma$ is a congruence on $\mathbf{A}_{1}$, $\mathcal{F}$ is a set of terms over the variables $x_{1},\dots,x_{n+1}$ closed under right composition, $\mathcal{F}$ is good for algebras $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$ and their subalgebras, $\varnothing\neq J\subseteq[s]$, $I\subseteq[s]$,
  1. (p)
    $f^{\mathbf{A}_{1}/\sigma}(a_{1}/\sigma,\dots,a_{n}/\sigma,x)\approx x$ for any $f\in\mathcal{F}$ and any tuple $(a_{1},\dots,a_{n})$ generating $\mathbf{A}_{1}$;
  2. (s)
    for every $j\in J$ there exist $t\in\mathcal{F}$ and a tuple $(a_{1},\dots,a_{n})$ generating $\mathbf{A}_{j}$ such that $|\{t^{\mathbf{A}_{j}}(a_{1},\dots,a_{n},b)\mid b\in A_{j}\}|<|A_{j}|$;
  3. (k)
    $\{t^{\mathbf{A}_{i}}(a_{1},\dots,a_{n},b)\mid b\in A_{i}\}=A_{i}$ for every $i\in I$, $t\in\mathcal{F}$, and $a_{1},\dots,a_{n}\in A_{i}$;
  4. ($\ell$)
    $|A_{i}|\leq\max\limits_{j\in J}|A_{j}|$ for any $i\notin J\cup I$;
  5. (i)
    for every $\{\mathbf{B}_{1},\dots,\mathbf{B}_{u}\}<\{\mathbf{A}_{1},\dots,\mathbf{A}_{s}\}$ and every $m,\ell\in\mathbb{N}$ there exists a $(\underbrace{\overline{k^{\ell}},\dots,\overline{k^{\ell}}}_{m})$-PS term for $\mathbf{B}_{1},\dots,\mathbf{B}_{u}$.
Then $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$ admit a $(\underbrace{\overline{k},\dots,\overline{k}}_{n})$-PS term.
Before proving the theorem, let us comment some of the above conditions. Condition (s) describes the nonempty set $J$ of algebras we can squeeze with a tuple generating the whole algebra, condition (k) describes the set $I$ of algebras that we cannot squeeze at all. Thus, $I\cap J=\varnothing$ but it is possible that $I\cup J\neq[s]$. Condition ($\ell$) says that any algebra outside of $I$ and $J$ is not larger than one of the algebras in $J$.
Proof.
Let $\mathbf{A}_{s+1},\dots,\mathbf{A}_{s^{\prime}}$ be the set of all subalgebras of $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$. Using Lemma 8.30 we choose a finite set $\{f_{1},\dots,f_{m}\}\subseteq\operatorname{Squeezer}(\mathcal{F},\mathbf{A}_{1},\dots,\mathbf{A}_{s^{\prime}})$. Let $\mathcal{U}$ be the set of all algebras $\mathbf{B}\in\mathcal{V}_{p}$ such that $B\subsetneq A_{i}$ for some $i\in J$. Then $\mathcal{U}\cup\{\mathbf{A}_{i}\mid i\in[s]\setminus J\}<\{\mathbf{A}_{1},\dots,\mathbf{A}_{s}\}$, and, by condition (i), $\mathcal{U}\cup\{\mathbf{A}_{i}\mid i\in[s]\setminus J\}$ admits a $(\underbrace{\overline{k},\dots,\overline{k}}_{m})$-PS term $t_{0}$. Let $\mathcal{U}_{0}$ be the set of all proper subalgebras of $\mathbf{A}_{1}$. Since $\mathcal{U}_{0}\cup\{\mathbf{A}_{2},\dots,\mathbf{A}_{s}\}<\{\mathbf{A}_{1},\dots,\mathbf{A}_{s}\}$, by assumption (i) there exists a $(\underbrace{\overline{k^{n}},\dots,\overline{k^{n}}}_{m})$-PS term $t_{1}$ for $\mathcal{U}_{0}$ and $\mathbf{A}_{2},\dots,\mathbf{A}_{s}$. Then a $(\underbrace{\overline{k},\dots,\overline{k}}_{n})$-PS term $t$ for $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$ can be obtained as follows.
First, for every $i\in[m]$ let the term $g_{i}(\mathbf{x}_{1},\dots,\mathbf{x}_{n},y_{1},\dots,y_{n})$ be obtained from $t_{0}(\mathbf{x}_{1},\dots,\mathbf{x}_{n})$ by replacing each $w(z_{1},\dots,z_{p})$ by
$f_{i}(y_{1},\dots,y_{n},w(f_{i}(y_{1},\dots,y_{n},z_{1}),\dots,f_{i}(y_{1},\dots,y_{n},z_{p}))).$
Thus, we add $f_{i}(y_{1},\dots,y_{n},z)$ to every variable $z$ to reduce the corresponding domain. Then put
$t(\mathbf{x}_{1},\dots,\mathbf{x}_{n}):=t_{1}((g_{1}(\mathbf{x}_{1},\dots,\mathbf{x}_{n},\mathbf{x}_{1}^{j_{1}},\dots,\mathbf{x}_{n}^{j_{n}}))_{j_{1},\dots,j_{n}\in[k]},\dots,$
$\dots,(g_{m}(\mathbf{x}_{1},\dots,\mathbf{x}_{n},\mathbf{x}_{1}^{j_{1}},\dots,$ $\mathbf{x}_{n}^{j_{n}}))_{j_{1},\dots,j_{n}\in[k]}).$
Here, the first $k^{n}$ arguments of $t_{1}$ are filled with $g_{1}$ where we replace the variables $y_{1},\dots,y_{n}$ by $\mathbf{x}_{1}^{j_{1}},\dots,\mathbf{x}_{n}^{j_{n}}$ for $j_{1},\dots,j_{n}\in[k]$. The next $k^{n}$ arguments are filled with $g_{2}$ and so on. Let us show that $t$ is in fact a $(\overline{k},\dots,\overline{k})$-PS term for $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$.
First, let us prove this for $\mathbf{A}_{i}$, where $i\in I$. Let $(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$ be a $(\overline{k},\dots,\overline{k})$-palette tuple from $A_{i}^{nk}$. Condition $(k)$ together with property 2 of squeezers imply that $f_{\ell}(a_{1},\dots,a_{n},x)\approx x$ for any $\ell\in[m]$ and $a_{1},\dots,a_{n}\in A_{i}$. Hence $g_{\ell}^{\mathbf{A}_{i}}(\mathbf{a}_{1},\dots,\mathbf{a}_{n},\mathbf{a}_{1}^{j_{1}},\dots,\mathbf{a}_{n}^{j_{n}})=t_{0}^{\mathbf{A}_{i}}(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$ for all $\ell\in[m]$ and $j_{1},\dots,j_{n}\in[k]$. Since $t_{0}$ is PS in $\mathbf{A}_{i}$, permutations in $\mathbf{a}_{1},\dots,\mathbf{a}_{n}$ does not change the result. Then the idempotency of $t_{1}^{\mathbf{A}_{i}}$ implies that $t$ is $(\overline{k},\dots,\overline{k})$-PS in $\mathbf{A}_{i}$.
Second, let us prove that $\mathbf{A}_{i}$ is $(\overline{k},\dots,\overline{k})$-PS for $i\in[s]\setminus I$. Again, let $(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$ be a $(\overline{k_{1}},\dots,\overline{k_{n}})$-palette tuple from $A_{i}^{nk}$. Let us fix $j_{1},\dots,j_{n}\in[k]$ and define $h(x)\approx f_{\ell}(\mathbf{a}_{1}^{j_{1}},\dots,\mathbf{a}_{n}^{j_{n}},x)$. Let $\mathbf{B}=\operatorname{Sg}_{\mathbf{A}_{i}}(\{\mathbf{a}_{1}^{j_{1}},\dots,\mathbf{a}_{n}^{j_{n}}\})$ and $\mathbf{B}^{\prime}=h(\mathbf{B})$. Let us show that $\mathbf{B}^{\prime}\in\mathcal{U}\cup\{\mathbf{A}_{j}\mid j\in[s]\setminus J\}$. In fact, if $i\in J$ and $(\mathbf{a}_{1}^{j_{1}},\dots,\mathbf{a}_{n}^{j_{n}})$ generates $\mathbf{A}_{i}$ then it follows from condition (s) and Lemma 8.27 that $\mathbf{B}^{\prime}\lneq\mathbf{B}=\mathbf{A}_{i}$ and $\mathbf{B}^{\prime}\in\mathcal{U}$. If $i\in J$ and $(\mathbf{a}_{1}^{j_{1}},\dots,\mathbf{a}_{n}^{j_{n}})$ does not generate $\mathbf{A}_{i}$, then $\mathbf{B},\mathbf{B}^{\prime}\in\mathcal{U}$. If $i\notin I\cup J$, then either $h(x)\approx x$ and $\mathbf{B}^{\prime}=\mathbf{B}=\mathbf{A}_{i}$, or by condition ($\ell$) $|B^{\prime}|<|A_{i}|\leq\max\limits_{j\in J}|A_{j}|$ and $\mathbf{B}^{\prime}\in\mathcal{U}$. Thus, the term $t_{0}$ is PS in $\mathbf{B}^{\prime}$. Recall that the choice of $j_{1},\dots,j_{n}\in[k]$ determines the function $h$, and the function $h$ determines the algebra $\mathbf{B}^{\prime}$. Since $t_{0}$ is $(\overline{k},\dots,\overline{k})$-PS in each $\mathbf{B}^{\prime}$, permutations in blocks $\mathbf{a}_{1},\dots,\mathbf{a}_{n}$ only imply permutations in the corresponding blocks of size $k^{n}$ of $t_{1}$ but the content of each block does not depend on the permutations.
Let us show that the tuple we substitute in $t_{1}^{\mathbf{A}_{i}}$ is $(\overline{k^{n}},\dots,\overline{k^{n}})$-palette. This follows from the fact that any value $c$ that ever appears in this tuple comes from some $g_{e}^{\mathbf{A}_{i}}(\mathbf{a}_{1},\dots,\mathbf{a}_{n},\mathbf{a}_{1}^{j_{1}},\dots,\mathbf{a}_{n}^{j_{n}})=c$. Since conditions (1) and (2) in Lemma 8.30 are equivalent, there exists $d\in[m]$ such that $g_{d}^{\mathbf{A}_{i}}(\mathbf{a}_{1},\dots,\mathbf{a}_{n},\mathbf{a}_{1}^{i_{1}},\dots,\mathbf{a}_{n}^{i_{n}})=c$ for all $i_{1},\dots,i_{n}\in[k]$. This implies that the $d$-th block of the tuple we substitute into $t_{1}^{\mathbf{A}_{i}}$ contains only value $c$, which means that it is palette.
Thus, we showed that the tuple we substitute in $t_{1}^{\mathbf{A}_{i}}$ is $(\overline{k^{n}},\dots,\overline{k^{n}})$-palette and permutations in $\mathbf{a}_{1},\dots,\mathbf{a}_{n}$ only imply permutations in the corresponding blocks of size $k^{n}$. Since $t_{1}$ is $(\overline{k^{n}},\dots,\overline{k^{n}})$-PS in $\mathbf{A}_{i}$ for $i\geq 2$, the term $t$ is $(\overline{k},\dots,\overline{k})$-PS in $\mathbf{A}_{i}$ for $i\geq 2$.
Assume that $i=1$ and $(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$ generates $\mathbf{A}_{1}$. Let us show that the values $g_{d}^{\mathbf{A}_{1}}(\mathbf{a}_{1},\dots,\mathbf{a}_{n},\mathbf{a}_{1}^{j_{1}},\dots,\mathbf{a}_{n}^{j_{n}})$ and $g_{e}^{\mathbf{A}_{1}}(\mathbf{a}_{1},\dots,\mathbf{a}_{n},\mathbf{a}_{1}^{i_{1}},\dots,\mathbf{a}_{n}^{i_{n}})$ are equivalent modulo $\sigma$ for all $d,e\in[m]$ and $j_{1},\dots,j_{n},i_{1},\dots,i_{n}\in[k]$. By condition (p) and since $\sigma$ is preserved by $w^{\mathbf{A}_{1}}$, the value $g_{e}^{\mathbf{A}_{1}}(\mathbf{a}_{1},\dots,\mathbf{a}_{n},\mathbf{a}_{1}^{i_{1}},\dots,\mathbf{a}_{n}^{i_{n}})$ is equivalent to $t_{0}^{\mathbf{A}_{1}}(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$ for all $e\in[m]$ and $i_{1},\dots,i_{n}\in[k]$. Therefore, we always substitute into $t_{1}^{\mathbf{A}_{1}}$ only values from one equivalence class of $\sigma$. Since $t_{1}$ is $(\overline{k^{n}},\dots,\overline{k^{n}})$-PS for all proper subalgebras of $\mathbf{A}_{1}$, permutations inside blocks $\mathbf{a}_{1},\dots,\mathbf{a}_{n}$ do not change the result and $t$ is $(\overline{k},\dots,\overline{k})$-PS in $\mathbf{A}_{1}$.
Finally, if $i=1$ and $(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$ does not generate $\mathbf{A}_{1}$, then it follows from the fact that $t_{1}$ is $(\overline{k^{n}},\dots,\overline{k^{n}})$-PS in the subalgebra generated by $(\mathbf{a}_{1},\dots,\mathbf{a}_{n})$. ∎
Recall that by $\mathcal{C}(\mathbf{A})$ we denote the set of all proper subuniverses $B\leq\mathbf{A}$ such that $(B\times A)\cup(A\times B)$ is a subuniverse of $\mathbf{A}^{2}$ (see Subsection 8.4).
Lemma 8.43.
Suppose $\mathbf{A}\in\mathcal{V}_{p}$, $\colon A\to A, hh$$\circ h=h$, $B\in\mathcal{C}(\mathbf{A})$, $h(B)\subseteq B$. Then $h(B)\in\mathcal{C}(h(\mathbf{A}))\cup\{h(A)\}$.
Proof.
It is sufficient to check that $w^{h(\mathbf{A})}(x_{1},\dots,x_{k}):=h(w^{\mathbf{A}}(x_{1},\dots,x_{k}))$ preserves the relation $(h(B)\times h(A))\cup(h(A)\times h(B))$, which is straightforward. ∎
We say that a $(\overline{k_{1}},\dots,\overline{k_{n}},\ell)$-PS term operation $f$ in $\mathbf{A}$ satisfies property (c) if for any $B\in\operatorname{\mathcal{C}}(\mathbf{A})$ and any $(\overline{k_{1}},\dots,\overline{k_{n}},\ell)$-palette tuple $(\mathbf{a}_{1},\dots,\mathbf{a}_{n},\mathbf{b})$ such that $\mathbf{b}\in B^{\ell}$ we have $f(\mathbf{a}_{1},\dots,\mathbf{a}_{n},\mathbf{b})\in B$.
For an algebra $\mathbf{A}$, a tuple $(a_{1},\dots,a_{n})\in A^{n}$ is called self-contained if $\{a_{1},\dots,a_{n}\}$ is a subuniverse of an algebra $\mathbf{A}$. We say that a function is weak $(\underbrace{\overline{k^{n}},\dots,\overline{k^{n}}}_{m},\ell)$-PS satisfying property (c), if we require both palette symmetry and property (c) only on self-contained tuples.
Lemma 8.44.
Suppose $k,n,\ell\in\mathbb{N}$, algebras $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$ admit a weak $(\underbrace{\overline{k^{n}},\dots,\overline{k^{n}}}_{m},\ell)$-PS term satisfying property (c) for all $m\in\mathbb{N}$. Then $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$ admit $(\underbrace{\overline{k},\dots,\overline{k}}_{n},\ell)$-PS term satisfying property (c).
Proof.
Choose terms $t_{1},\dots,t_{m}$ of arity $n$ so that for any $i\in[s]$ and any $n$-ary term operation $g$ in $\mathbf{A}_{i}$ there exists $j\in[m]$ such that $t_{j}^{\mathbf{A}_{i}}=g$. Let $f$ be a weak $(\underbrace{\overline{k^{n}},\dots,\overline{k^{n}}}_{m},\ell)$-PS term satisfying property (c) for $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$. Then a $(\underbrace{\overline{k},\dots,\overline{k}}_{n},\ell)$-PS term $t$ can be defined by
$t(\mathbf{x}_{1},\dots,\mathbf{x}_{n},\mathbf{x}_{n+1}):=f((t_{1}(\mathbf{x}_{1}^{j_{1}},\dots,\mathbf{x}_{n}^{j_{n}}))_{j_{1},\dots,j_{n}\in[k]},\dots,(t_{m}(\mathbf{x}_{1}^{j_{1}},\dots,\mathbf{x}_{n}^{j_{n}}))_{j_{1},\dots,j_{n}\in[k]},\mathbf{x}_{n+1}).$
Notice that permutations in $\mathbf{x}_{1},\dots,\mathbf{x}_{n}$ only imply permutations inside blocks of size $k^{n}$. Thus, we only need to check that the tuple we substitute in $f$ is $(\underbrace{\overline{k^{n}},\dots,\overline{k^{n}}}_{m},\ell)$-palette and self-contained. Fix $i\in[s]$ and a $(\underbrace{\overline{k},\dots,\overline{k}}_{n},\ell)$-palette tuple $(\mathbf{a}_{1},\dots,\mathbf{a}_{n+1})\in A_{i}^{kn+\ell}$. Let $b$ be an element of the subalgebra generated by elements of this tuple in $\mathbf{A}_{i}$. Let $\mathbf{a}_{j_{1}},\dots,\mathbf{a}_{j_{r}}$ be all the constant blocks. Let $g$ be an $n$-ary term operation in $\mathbf{A}_{i}$ such that only variables on positions $j_{1},\dots,j_{r}$ are not dummy, and $g(\mathbf{a}_{1}^{1},\dots,\mathbf{a}_{n}^{1})=b$. By our assumptions, $g=t_{j}^{\mathbf{A}_{i}}$ for some $j\in[m]$. Then, the $j$-th block of the tuple we substitute consists only of element $b$, which means that the corresponding tuple is palette and self-contained. ∎
Remark 8.1.
For any $(\overline{k_{1}},\dots,\overline{k_{n}},k_{i})$-PS function $f$ we can derive a $(\overline{k_{1}},\dots,\overline{k_{n}})$-PS function $g$ by
$g(\mathbf{x}_{1},\dots,\mathbf{x}_{n}):=f(\mathbf{x}_{1},\dots,\mathbf{x}_{n},\mathbf{x}_{i}).$
Theorem 8.45.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{p}$ are minimal Taylor, $\mathcal{F}$ is a set of terms over the variables $x_{1},\dots,x_{n+1}$ closed under right composition, $\mathcal{F}$ is good for algebras $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$ and their subalgebras, $\varnothing\neq I\subseteq[s]$,
  1. (p)
    $f^{\mathbf{A}_{i}}(a_{1},\dots,a_{n},b)\in B$ for any $f\in\mathcal{F}$, $i\in[s]$, $B\in\mathcal{C}(\mathbf{A}_{i})$, $a_{1},\dots,a_{n}\in A_{i}$, and $b\in B$;
  2. (s)
    for every $i\in I$ there exist $t\in\mathcal{F}$ and a tuple $(a_{1},\dots,a_{n})$ generating $\mathbf{A}_{i}$ such that $|\{t^{\mathbf{A}_{i}}(a_{1},\dots,a_{n},b)\mid b\in A_{i}\}|<|A_{i}|$;
  3. (k)
    $\{t^{\mathbf{A}_{i}}(a_{1},\dots,a_{n},b)\mid b\in A_{i}\}=A_{i}$ for every $i\in[s]\setminus I$, $t\in\mathcal{F}$, and $a_{1},\dots,a_{n}\in A_{i}$;
  4. (n)
    $\mathcal{C}(\mathbf{A}_{i})$ is not empty for some $i\in[s]$;
  5. (i)
    for every $\{\mathbf{B}_{1},\dots,\mathbf{B}_{u}\}<\{\mathbf{A}_{1},\dots,\mathbf{A}_{s}\}$ and every $m,\ell,n_{0}\in\mathbb{N}$ there exists a $(\underbrace{\overline{k^{n_{0}}},\dots,\overline{k^{n_{0}}}}_{m},\ell)$-PS term $t$ for $\mathbf{B}_{1},\dots,\mathbf{B}_{u}$ satisfying property (c).
Then for every $\ell\in\mathbb{N}$ the algebras $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$ admit a weak $(\underbrace{\overline{k},\dots,\overline{k}}_{n},\ell)$-PS term satisfying property (c).
Proof.
Let $\mathbf{A}_{s+1},\dots,\mathbf{A}_{s^{\prime}}$ be the set of all subalgebras of $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$. Using Lemma 8.30 we choose a finite set $\{f_{1},\dots,f_{m}\}\subseteq\operatorname{Squeezer}(\mathcal{F},\mathbf{A}_{1},\dots,\mathbf{A}_{s^{\prime}})$. Let us fix $\ell\in\mathbb{N}$. Put $M=(k!)^{n}\cdot\ell!$. Let $\mathcal{U}$ be the set of all algebras $\mathbf{B}\in\mathcal{V}_{p}$ such that $B\subsetneq A_{i}$ for some $i\in I$. Then $\mathcal{U}\cup\{\mathbf{A}_{i}\mid i\in[s]\setminus I\}<\{\mathbf{A}_{1},\dots,\mathbf{A}_{s}\}$, and, by condition (i), $\mathcal{U}\cup\{\mathbf{A}_{i}\mid i\in[s]\setminus I\}$ admits a $(\underbrace{\overline{k},\dots,\overline{k}}_{n},M)$-PS term $t_{0}$ satisfying property (c). Let $J$ be the set of all $j\in[s]$ such that $\mathcal{C}(\mathbf{A}_{j})$ is not empty. By condition (n), $J\neq\varnothing$. Let $\mathcal{U}_{0}$ be the set of all algebras $\mathbf{B}\in\mathcal{V}_{p}$ such that $B\subsetneq A_{i}$ for some $i\in J$. Then $\mathcal{U}_{0}\cup\{\mathbf{A}_{i}\mid i\in[s]\setminus J\}<\{\mathbf{A}_{1},\dots,\mathbf{A}_{s}\}$ and, by condition (i) and Remark 8.1, $\mathcal{U}_{0}\cup\{\mathbf{A}_{i}\mid i\in[s]\setminus J\}$ admits a $(\underbrace{\overline{k^{n}},\dots,\overline{k^{n}}}_{m})$-PS term $t_{1}$. Finally, by Lemma 8.25 there exists a $(\underbrace{\overline{k},\dots,\overline{k}}_{n},\ell)$-block term $t_{2}$ for $\mathbf{A}_{1},\dots,\mathbf{A}_{s^{\prime}}$ satisfying properties 1 and 2 in Lemma 8.25. Note that the only reason for requiring the algebras $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$ to be minimal Taylor is to ensure that every subalgebra $\mathbf{B}\in\mathcal{C}(\mathbf{A}_{i})$ is central, and conversely that every proper central subalgebra of $\mathbf{A}_{i}$ belongs to $\mathcal{C}(\mathbf{A}_{i})$ (by Lemma 8.20). Thus, properties 1 and 2 in Lemma 8.25 can be reformulated for subalgebras from $\mathcal{C}(\mathbf{A}_{i})$. Then a weak $(\underbrace{\overline{k},\dots,\overline{k}}_{n},\ell)$-PS term $t$ for $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$ satisfying property (c) can be obtained as follows.
First, for every $i\in[m]$ let the term $g_{i}(\mathbf{x}_{1},\dots,\mathbf{x}_{n},\mathbf{u},y_{1},\dots,y_{n})$ be obtained from $t_{0}(\mathbf{x}_{1},\dots,\mathbf{x}_{n},\mathbf{u})$ by replacing each $w(z_{1},\dots,z_{p})$ by
$f_{i}(y_{1},\dots,y_{n},w(f_{i}(y_{1},\dots,y_{n},z_{1}),\dots,f_{i}(y_{1},\dots,y_{n},z_{p}))).$
Thus, we add $f_{i}(y_{1},\dots,y_{n},z)$ to every variable $z$ to reduce the corresponding domain $\mathbf{A}_{j}$ for $j\in I$. Then put
$t_{3}(\mathbf{x}_{1},\dots,\mathbf{x}_{n},\mathbf{u}):=t_{1}((g_{1}(\mathbf{x}_{1},\dots,\mathbf{x}_{n},\mathbf{u},\mathbf{x}_{1}^{j_{1}},\dots,\mathbf{x}_{n}^{j_{n}}))_{j_{1},\dots,j_{n}\in[k]},\dots,$
$\dots,(g_{m}(\mathbf{x}_{1},\dots,\mathbf{x}_{n},\mathbf{u},\mathbf{x}_{1}^{j_{1}},\dots,$ $\mathbf{x}_{n}^{j_{n}}))_{j_{1},\dots,j_{n}\in[k]}).$
Here, the first $k^{n}$ arguments of $t_{1}$ are filled with $g_{1}$ where we replace the variables $y_{1},\dots,y_{n}$ by $\mathbf{x}_{1}^{j_{1}},\dots,\mathbf{x}_{n}^{j_{n}}$ for $j_{1},\dots,j_{n}\in[k]$. The next $k^{n}$ arguments are filled with $g_{2}$ and so on.
Finally, to define $t$ we replace $\mathbf{u}$ in $t_{3}$ by the term $t_{2}$ applied to all possible permutations of the corresponding arguments. Formally, let $\{(\sigma_{i,1},\dots,\sigma_{i,n+1})\mid i\in[M]\}$ be the set of all tuples of permutations, where $\sigma_{i,1},\dots,\sigma_{i,n}$ are permutations on $[k]$ and $\sigma_{i,n+1}$ is a permutation on $\ell$]. Then
$t(\mathbf{x}_{1},\dots,\mathbf{x}_{n+1}):=t_{3}(\mathbf{x}_{1},\dots,\mathbf{x}_{n},t_{2}(\mathbf{x}_{1}^{\sigma_{1,1}},\dots,\mathbf{x}_{n+1}^{\sigma_{1,n+1}}),\dots,t_{2}(\mathbf{x}_{1}^{\sigma_{M,1}},\dots,\mathbf{x}_{n+1}^{\sigma_{M,n+1}})).$
Let us show that $t$ is in fact a weak $(\overline{k},\dots,\overline{k},\ell)$-PS term for $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$ satisfying property (c).
First, let us show property (c). Suppose $(\mathbf{a}_{1},\dots,\mathbf{a}_{n},\mathbf{a}_{n+1})$ is a self-contained $(\overline{k},\dots,\overline{k},\ell)$-palette tuple on $\mathbf{A}_{i}$ such that $\mathbf{a}_{n+1}\in B^{\ell}$ and $B\in\mathcal{C}(\mathbf{A}_{i})$. By properties of $t_{2}$, we have $t_{2}(\mathbf{a}_{1}^{\sigma_{i^{\prime},1}},\dots,\mathbf{a}_{n+1}^{\sigma_{i^{\prime},n+1}})\in B$ for every $i^{\prime}\in[M]$. By property (p) and Lemma 8.43, $h(B)\in\mathcal{C}(h(\mathbf{A}_{i}))\cup\{h(A_{i})\}$ for any $c_{1},\dots,c_{n}\in A_{i}$, where $h(x)\approx f(c_{1},\dots,c_{n},x)$. By Lemma 8.27 and condition (s), $|h(\mathbf{A}_{i})|<|A_{i}|$ if $i\in I$, and by condition (k) we have $h(\mathbf{A}_{i})=A_{i}$ if $i\notin I$. Therefore, $h(\mathbf{A}_{i})\in\mathcal{U}\cup\{\mathbf{A}_{i}\mid i\in[s]\setminus I\}$. Since the original tuple is self-contained, the tuple we substitute in $t_{0}$ is palette, which is exactly the reason why we only prove the existence of a weak(!) palette term. Since $t_{0}$ satisfies property (c) on $\mathcal{U}\cup\{\mathbf{A}_{i}\mid i\in[s]\setminus I\}$, each $g_{q}^{\mathbf{A}_{i}}$ returns an element of $B$. Since $B$ is a subalgebra of $\mathbf{A}_{i}$, $t_{1}^{\mathbf{A}_{i}}$ also returns an element of $B$. Hence $t^{\mathbf{A}_{i}}(\mathbf{a}_{1},\dots,\mathbf{a}_{n},\mathbf{a}_{n+1})\in B$.
Let us show that $t$ is $(\overline{k},\dots,\overline{k},\ell)$-PS term for each $\mathbf{A}_{i}$. We start with the case when $i\in[s]\setminus I$. In this case $t_{0}$ is $(\underbrace{\overline{k},\dots,\overline{k}}_{n},M)$-PS in $\mathbf{A}_{i}$. By property (k) and property 2 of squeezers, $f_{q}(c_{1},\dots,c_{n},x)\approx x$ for any $q\in[m]$ and $c_{1},\dots,c_{n}\in A_{i}$, hence the value $g_{q}^{\mathbf{A}_{i}}(\mathbf{a}_{1},\dots,\mathbf{a}_{n},\mathbf{u},y_{1},\dots,y_{n})$ only depends on $\mathbf{a}_{1},\dots,\mathbf{a}_{n},\mathbf{u}$, where $\mathbf{u}$ depends only on $\mathbf{a}_{1},\dots,\mathbf{a}_{n+1}$. Since $t_{0}$ is $(\underbrace{\overline{k},\dots,\overline{k}}_{n},M)$-PS in $\mathbf{A}_{i}$, permutations in $\mathbf{a}_{1},\dots,\mathbf{a}_{n},\mathbf{a}_{n+1}$ does not change this value. Thus, all the values we substitute in $t_{1}^{\mathbf{A}_{i}}$ are the same, and the idempotency of $t_{1}^{\mathbf{A}_{i}}$ implies that $t$ is $(\overline{k},\dots,\overline{k},\ell)$-PS in $\mathbf{A}_{i}$.
Let us prove that $t$ is $(\overline{k},\dots,\overline{k},\ell)$-PS in $\mathbf{A}_{i}$ for $i\in I$. Since the tuple $(\mathbf{a}_{1},\dots,\mathbf{a}_{n},\mathbf{a}_{n+1})$ is $(\overline{k},\dots,\overline{k},\ell)$-palette, $(\mathbf{a}_{1}^{j_{1}},\dots,\mathbf{a}_{n}^{j_{n}})$ generates the same subalgebra of $\mathbf{A}_{i}$ as $(\mathbf{a}_{1},\dots,\mathbf{a}_{n},\mathbf{a}_{n+1})$ for any $j_{1},\dots,j_{n}\in[k]$. Let us fix $q\in[m]$ and $j_{1},\dots,j_{n}\in[k]$ and define $h(x)\approx f_{q}(\mathbf{a}_{1}^{j_{1}},\dots,\mathbf{a}_{n}^{j_{n}},x)$. Let $\mathbf{B}=\operatorname{Sg}_{h(\mathbf{A}_{i})}(\{h(\mathbf{a}_{1}^{j_{1}}),\dots,h(\mathbf{a}_{n}^{j_{n}})\})$. Since $i\in I$, property (s) and Lemma 8.27 imply that $B\subsetneq A_{i}$. Since $\mathbf{B}\in\mathcal{U}\cup\{\mathbf{A}_{i}\mid i\in[s]\setminus I\}$, the term $t_{0}$ is a $(\overline{k},\dots,\overline{k},M)$-PS term on $\mathbf{B}$ satisfying property (c). Thus, the choice of $q\in[m]$ and $j_{1},\dots,j_{n}\in[k]$ determines the function $h$ and the function $h$ determines the algebra $\mathbf{B}$. Since $t_{0}$ is a $(\overline{k},\dots,\overline{k},M)$-PS on each $\mathbf{B}$, permutations in blocks $\mathbf{a}_{1},\dots,\mathbf{a}_{n},\mathbf{a}_{n+1}$ only imply permutations in the corresponding blocks of size $k^{n}$ of $t_{1}$ but the content of each block does not depend on the permutations.
Let us show that the tuple we substitute in $t_{1}^{\mathbf{A}_{i}}$ is $(\overline{k^{n}},\dots,\overline{k^{n}})$-palette. Any value $c$ that ever appears in this tuple comes from some $g_{q}^{\mathbf{A}_{i}}(\mathbf{a}_{1},\dots,\mathbf{a}_{n},\mathbf{u},\mathbf{a}_{1}^{j_{1}},\dots,\mathbf{a}_{n}^{j_{n}})=c$. Since conditions (1) and (2) in Lemma 8.30 are equivalent, there exists $d\in[m]$ such that $g_{d}^{\mathbf{A}_{i}}(\mathbf{a}_{1},\dots,\mathbf{a}_{n},\mathbf{u},\mathbf{a}_{1}^{i_{1}},\dots,\mathbf{a}_{n}^{i_{n}})=c$ for all $i_{1},\dots,i_{n}\in[k]$. This implies that the $d$-th block of the tuple we substitute into $t_{1}^{\mathbf{A}_{i}}$ contains only value $c$, which means that it is palette.
Thus, the tuple we substitute in $t_{1}^{\mathbf{A}_{i}}$ is $(\overline{k^{n}},\dots,\overline{k^{n}})$-palette and permutations in $\mathbf{a}_{1},\dots,\mathbf{a}_{n},\mathbf{a}_{n+1}$ only imply permutations in the corresponding blocks of size $k^{n}$.
If $i\notin J$, then $t_{1}$ is $(\overline{k^{n}},\dots,\overline{k^{n}})$-PS in $\mathbf{A}_{i}$. Therefore, $t$ is $(\overline{k},\dots,\overline{k},\ell)$-PS in $\mathbf{A}_{i}$.
Assume that $i\in J$ and the tuple $(\mathbf{a}_{1},\dots,\mathbf{a}_{n+1})$ does not generate $\mathbf{A}_{i}$. Since $t_{1}$ is $(\overline{k^{n}},\dots,\overline{k^{n}})$-PS in any proper subalgebra of $\mathbf{A}_{i}$, the term $t$ is $(\overline{k},\dots,\overline{k},\ell)$-PS in $\mathbf{A}_{i}$. Finally, assume that $i\in J$ and the tuple $(\mathbf{a}_{1},\dots,\mathbf{a}_{n+1})$ generates $\mathbf{A}_{i}$. Then there exists $B\in\mathcal{C}(\mathbf{A}_{i})$ such that $t_{2}(\mathbf{a}_{1}^{\sigma_{1,i^{\prime}}},\dots,\mathbf{a}_{n+1}^{\sigma_{n+1,i^{\prime}}})\in B$ for any $i^{\prime}\in[M]$. Since each $f_{q}$ right-preserves $\mathcal{C}(\mathbf{A}_{i})$ in $\mathbf{A}_{i}$ and $t_{0}$ satisfies property (c), all the elements we substitute into $t_{1}$ are from $B$. Since $t_{1}$ is $(\overline{k^{n}},\dots,\overline{k^{n}})$-PS on $\mathbf{B}$, $(\overline{k},\dots,\overline{k},\ell)$-PS in $\mathbf{A}_{i}$. ∎
8.10 Main result
Here we combine the results of the previous section and show that we can recursively build PS term operations using PS term operations on smaller algebras unless our algebras are Maltsev, nilpotent, and large enough.
Theorem 8.46.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}\in\mathcal{V}_{p}$ are minimal Taylor, where $p$ is a prime and $|A_{i}|<p$ for every $i\in[s]$, and
  1. (i)
    for every $\{\mathbf{B}_{1},\dots,\mathbf{B}_{u}\}<\{\mathbf{A}_{1},\dots,\mathbf{A}_{s}\}$ and every $m,n\in\mathbb{N}$ there exists a $(\underbrace{\overline{p^{n}},\dots,\overline{p^{n}}}_{m})$-PS term for $\mathbf{B}_{1},\dots,\mathbf{B}_{u}$;
  2. (c)
    $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$ do not admit a $(\underbrace{\overline{p^{n}},\dots,\overline{p^{n}}}_{m})$-PS term for some $m,n\in\mathbb{N}$.
Then the following conditions hold:
  1. (1)
    there exists a term $m$ such that $m^{\mathbf{A}_{i}}$ is a Maltsev operation for every $i\in[s]$;
  2. (2)
    any $\wedge$-irreducible congruence of $\mathbf{A}_{i}$ is perfect linear for any $i\in[s]$;
  3. (3)
    for any $i,j\in[s]$ and any $\wedge$-irreducible congruences on $\mathbf{A}_{i}$ and $\mathbf{A}_{j}$, respectively, there exists a bridge between them;
  4. (4)
    $|A_{i}|\geq 8$ for some $i$.
Proof.
First, choose a minimal set of algebras $\{\mathbf{B}_{1},\dots,\mathbf{B}_{u}\}\leq\{\mathbf{A}_{1},\dots,\mathbf{A}_{s}\}$ such that for some $n,\ell,m\in\mathbb{N}$ there does not exist a $(\underbrace{\overline{p^{n}},\dots,\overline{p^{n}}}_{m},\ell)$-PS term in $\mathbf{B}_{1},\dots,\mathbf{B}_{u}$ satisfying property (c). To make it clear , we assume that either $\{\mathbf{B}_{1},\dots,\mathbf{B}_{u}\}<\{\mathbf{A}_{1},\dots,\mathbf{A}_{s}\}$ in our partial order, or $\{\mathbf{B}_{1},\dots,\mathbf{B}_{u}\}=\{\mathbf{A}_{1},\dots,\mathbf{A}_{s}\}$. Such a set of algebras exists because there is no such a term for $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$, for some $m,n\in\mathbb{N}$, and $\ell=p^{n}$.
By Lemma 8.22, choose a $p$-ary cyclic term $t$ such that $\mathbf{B}_{i}^{\prime}:=(B_{i};t^{\mathbf{B}_{i}})$ is a minimal Taylor algebra for every $i\in[u]$. Since $\mathbf{B}_{i}^{\prime}$ contains less term operations, the property (c) only becomes stronger for $\mathbf{B}_{1}^{\prime},\dots,\mathbf{B}_{u}^{\prime}$. In fact, since $\mathcal{C}(\mathbf{B}_{i}^{\prime})\supseteq\mathcal{C}(\mathbf{B}_{i})$ for every $i\in[u]$, there does not exist a $(\underbrace{\overline{p^{n}},\dots,\overline{p^{n}}}_{m},\ell)$-PS term in $\mathbf{B}_{1}^{\prime},\dots,\mathbf{B}_{u}^{\prime}$ satisfying property (c). Notice that for every $j\in[u]$ there should be a tuple $(a_{1},\dots,a_{n})$ generating $\mathbf{B}_{j}^{\prime}$ as otherwise we could replace $\mathbf{B}_{j}^{\prime}$ by all its subalgebras and get a smaller set of algebras without PS term operations. Consider three cases:
Case 1. $\{c\}<_{\mathcal{C}}\mathbf{B}_{j}^{\prime}$ for every $j\in[u]$ and $c\in B_{j}^{\prime}$. By Lemma 8.10, $\{c\}<_{\mathcal{C}}\mathbf{S}$ for any subalgebra $\mathbf{S}$ of $\mathbf{B}_{j}^{\prime}$ and $c\in s$. Then Lemma 8.25 applied to $\mathbf{B}_{1}^{\prime},\dots,\mathbf{B}_{u}^{\prime}$ and their subalgebras gives $(\underbrace{\overline{p^{n}},\dots,\overline{p^{n}}}_{m},\ell)$-PS term in $\mathbf{B}_{1}^{\prime},\dots,\mathbf{B}_{u}^{\prime}$ satisfying property (c), which contradicts our assumptions.
Case 2. $\mathcal{C}(B_{i}^{\prime})$ is not empty for some $i\in[u]$ but $\{c\}<_{\mathcal{C}}\mathbf{B}_{j}^{\prime}$ for some $j\in[u]$ and $c\in B_{j}$. Consider a family of terms $\mathcal{F}$ defined in Corollary 8.33 for $\mathbf{B}_{1}^{\prime},\dots,\mathbf{B}_{u}^{\prime}$. By Lemma 8.35 there exists $t\in\mathcal{F}$ such that $\{t^{\mathbf{A}_{i}}(c_{1},\dots,c_{n},b)\mid b\in A_{j}\}\neq A_{j}$ for a tuple $(c_{1},\dots,c_{n})$ generating $\mathbf{A}_{j}$. Then we can check that the algebras $\mathbf{B}_{1}^{\prime},\dots,\mathbf{B}_{u}^{\prime}$ together with the family of terms $\mathcal{F}$ satisfy all the conditions of Theorem 8.45. Then there exists a weak $(\underbrace{\overline{p^{n}},\dots,\overline{p^{n}}}_{m},\ell)$-PS term in $\mathbf{B}_{1}^{\prime},\dots,\mathbf{B}_{u}^{\prime}$ satisfying property (c). Lemma 8.44 implies that there exists a $(\underbrace{\overline{p^{n}},\dots,\overline{p^{n}}}_{m},\ell)$-PS term in $\mathbf{B}_{1}^{\prime},\dots,\mathbf{B}_{u}^{\prime}$ satisfying property (c), which contradicts our assumption that there does not exist one for $\mathbf{B}_{1}^{\prime},\dots,\mathbf{B}_{u}^{\prime}$.
Case 3. $\mathcal{C}(B_{i}^{\prime})$ is empty for all $i\in[u]$. Then property (c) is void, and there does not exist $(\underbrace{\overline{p^{n}},\dots,\overline{p^{n}}}_{m})$-PS term in $\mathbf{B}_{1}^{\prime},\dots,\mathbf{B}_{u}^{\prime}$ as adding $\ell$ dummy variables to it would give a $(\underbrace{\overline{p^{n}},\dots,\overline{p^{n}}}_{m},\ell)$-PS term satisfying property (c). By condition (i) $\{\mathbf{B}_{1}^{\prime},\dots,\mathbf{B}_{u}^{\prime}\}$ has to coincide with $\{\mathbf{A}_{1},\dots,\mathbf{A}_{s}\}$. Notice that if $0_{\mathbf{A}_{i}}$ is not $\wedge$-irreducible for some $i\in[s]$ then we can replace $\mathbf{A}_{i}$ by $\mathbf{A}_{i}/\omega_{1},\dots,\mathbf{A}_{i}/\omega_{m}$, where $\omega_{1}\cap\dots\cap\omega_{m}=0_{\mathbf{A}_{i}}$, and use the fact that we have a PS term for all smaller sets of algebras. Thus, we assume that $0_{\mathbf{A}_{i}}$ is $\wedge$-irreducible for every $i\in[s]$. Since $\mathcal{C}(\mathbf{A}_{j})$ is empty and $\mathbf{A}_{j}$ is a minimal Taylor algebra, it has no central or binary absorbing subuniverses. By Lemma 8.15 we have one of the two subcases:
Subcase 3A. There exists a maximal linear congruence $\sigma$ on some $\mathbf{A}_{i}$. By Lemma 8.3 $\mathbf{A}_{i}/\sigma\cong\mathbf{Z}_{q}$ for some prime $q$. Consider a family of terms $\mathcal{F}$ defined in Corollary 8.32 for the chosen $\sigma$. We want to get a contradiction with Theorem 8.42 for the case when $I\cup J=[s]$ and, therefore, a void condition ($\ell$). It is sufficient to check that $J\neq\varnothing$. By property (s) in Lemma 8.31, this is equivalent to the existence of $j\in[s]$, $t\in\mathcal{F}$, and $a_{1},\dots,a_{n}\in A_{j}$ such that $|\{t^{\mathbf{A}_{j}}(a_{1},\dots,a_{n},b)\mid b\in A_{j}\}|<|A_{j}|$. Then Lemma 8.38 implies that there exists a term $m$ such that $m^{\mathbf{A}_{j}}$ is Maltsev operation for all $j\in[s]$. Corollary 8.37 implies that there exists a bridge between $\sigma$ and any $\wedge$-irreducible congruence on any $\mathbf{A}_{j}$. To complete this case it is sufficient to show that $|A_{j}|\geq 8$ for some $j\in[s]$. Assume the contrary. Choose some $j\in[s]$. By Lemma 8.6 each equivalence class of $0_{\mathbf{A}_{j}}^{+}$ contains $q$ elements. Therefore $A_{j}/0_{\mathbf{A}_{j}}^{+}$ contains at most 3 elements. If it contains 2 elements then $q=2$ and $\mathbf{A}_{j}/0_{\mathbf{A}_{j}}^{+}\cong\mathbf{Z}_{2}$. If it contains 3 elements and $0_{\mathbf{A}_{j}}^{+}$ is $\wedge$-irreducible, then $\mathbf{A}_{j}/0_{\mathbf{A}_{j}}^{+}\cong\mathbf{Z}_{3}$, which implies that $A_{j}$ has at least $3\cdot 3=9$ elements. If it contains 3 elements but $0_{\mathbf{A}_{j}}^{+}$ is not $\wedge$-irreducible, then there are congruences $\sigma_{1}$ and $\sigma_{2}$ on $\mathbf{A}_{j}$ such that $0_{\mathbf{A}_{i}}^{+}=\sigma_{1}\cap\sigma_{2}$ and $\sigma_{1}\circ\sigma_{2}\neq\sigma_{2}\circ\sigma_{1}$, which contradicts the existence of a Maltsev term [19]. Thus, we showed that each $\mathbf{A}_{j}$ is either isomorphic to $\mathbf{Z}_{q}$ or isomorphic to an algebra from $\mathbf{Z}_{q}\boxtimes\mathbf{Z}_{q}$. Let us consider an XY-symmetric term of arity $k^{n}$, which exists by Theorem 8.1. By Lemma 8.7 this term is also symmetric on each $\mathbf{A}_{i}$. By adding dummy variables we get a $(\overline{k^{n}},\dots,\overline{k^{n}})$-PS term for $\mathbf{A}_{1},\dots,\mathbf{A}_{n}$, which contradicts our assumptions.
Subcase 3B. Every maximal congruence on each $\mathbf{A}_{i}$ is a PC congruence. Consider a family of terms $\mathcal{T}$ defined in Lemma 8.34. Since $0_{\mathbf{A}_{i}}$ is $\wedge$-irreducible for every $i$, property (s) in Lemma 8.34 implies that for every $i\in[s]$ either there exist $t\in\mathcal{T}$ and a tuple $(a_{1},\dots,a_{n})$ generating $\mathbf{A}_{i}$ such that $|\{t^{\mathbf{A}_{i}}(a_{1},\dots,a_{n},b)\mid b\in A_{i}\}|<|A_{i}|$, or $\mathbf{A}_{i}$ is a PC algebra. Let $\mathbf{A}_{v}$ be a maximal algebra in the set $\{\mathbf{A}_{1},\dots,\mathbf{A}_{s}\}$. If $\mathbf{A}_{v}$ is a PC algebra, then we get a contradiction using Theorem 8.41. Otherwise, choose some maximal congruence $\sigma$ on $\mathbf{A}_{v}$. Then we can get a contradiction using Theorem 8.42 applied to the family $\mathcal{T}$ and the congruence $\sigma$, because $v\in J$ and condition ($\ell$) is satisfied. ∎
The main result of this section is now a simple corollary.
Theorem 3.7.
Suppose $\mathbf{A}_{1},\dots,\mathbf{A}_{s}$ are similar algebras admitting WNU term operations, $|A_{i}|<8$ for every $i\in[s]$. Then $\mathbf{A}_{1}\times\dots\times\mathbf{A}_{s}$ has a $(\underbrace{p^{m},\dots,p^{m}}_{n})$-palette symmetric term operation for every prime $p>8$ and $m,n\in\mathbb{N}$.
Proof.
Assume the converse. Choose a minimal set of similar Taylor algebras $\{\mathbf{B}_{1},\dots,\mathbf{B}_{u}\}\leq\{\mathbf{A}_{1},\dots,\mathbf{A}_{s}\}$ such that they do not admit the corresponding palette symmetric term operation. By Lemma 8.22 we can choose a cyclic term $c$ such that $(B_{i};c^{\mathbf{B}_{i}})$ is a minimal Taylor algebra for every $i\in[u]$. Then Theorem 8.46 implies that $|B_{i}|\geq 8$ for some $i$, which contradicts our assumptions. ∎
Acknowledgments
I am very grateful to Michael Kompatscher for showing that the dihedral group $\mathbf{D}_{4}$ does not admit palette symmetric terms, which prevented me from attempting to prove their existence for another five years. I thank Antoine Mottet for pointing out that a gap in my proof could be filled using the famous Hales-Jewett theorem. I am also grateful to Standa Živný and Lorenzo Ciardo for explaining to me the minion characterization of singleton algorithms. Finally, I thank Jakub Rydval for motivating me to find palette polymorphisms for temporal relational structures; this simple exercise led us to discover a structure demonstrating the limitations of the Hales-Jewett argument and to better understand the relationship between singleton and constraint-singleton algorithms.